### v1

# Minimax Decision Trees via Martingale Approximations

Abstract
We develop a martingale-based approach to constructing decision trees that efficiently approximate a target variable through recursive conditioning. We introduce MinimaxSplit, a novel splitting criterion that minimizes the worst-case variance at each step, and analyze its cyclic variant, proving an exponential error decay rate under mild conditions. Our analysis builds upon partition-based martingale approximations, providing new insights into their convergence behavior. Unlike traditional variance-based methods, MinimaxSplit avoids end-cut preference and performs well in noisy settings. We derive empirical risk bounds and also explore its integration into random forests.
Keywords: martingale approximation; convergence rates; tree-based models; non-parametric regression
1 Introduction
Let $d\geqslant 1$ and $(\mathbf{X},Y)$ be a coupling (or a joint distribution) on $\mathbb{R}^{d+1}$, where $\mathbf{X}\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}$. Throughout, vectors are denoted by bold symbols. We study the general problem of efficiently approximating $Y$ by conditioning on the value of $\mathbf{X}$ using finite partitions. If we adopt the $L^{2}$ norm as the minimization criterion, the problem can be formulated as
$\min_{m\in\mathcal{M}}\mathbb{E}[(Y-m(\mathbf{X}))^{2}],$ (1)
where $\mathcal{M}$ denotes some class of candidate functions defined on $\mathbb{R}^{d}$. In this paper, we focus on the case where $\mathcal{M}$ consists of piecewise constant functions. Generally speaking, the motivation is algorithmic: in many applications, such as decision trees (Breiman et al. (1984)), the goal is to provide an efficient approximation of $Y$ by a function of $\mathbf{X}$. For instance, if the conditioning on $\mathbf{X}$ is based on a partition of cardinality at most $K$, we have
=$\mathcal{M}=\bigg{\{}m:\mathbb{R}^{d}\to\mathbb{R}\,\Big{|}\,m(\mathbf{x})=\sum_{A\in\pi}m_{A}{\mathbbm{1}}_{\{\mathbf{x}\in A\}},\,\pi\in\mathfrak{P}(\mathbb{R}^{d}),\,|\pi|\leqslant K,\,m_{A}\in\mathbb{R}\bigg{\}},$
where $\mathfrak{P}(\mathbb{R}^{d})$ denotes the set of all partitions of $\mathbb{R}^{d}$. Clearly, the choice $m_{A}=\mathbb{E}[Y\mid\mathbf{X}\in A]$ is always optimal in (1), so it remains to optimize the partition $\pi$. However, finding a direct solution to (1) is generally computationally infeasible, so recursive optimizers are favorable. The common approach is to construct a sequence of nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ in $\mathfrak{P}(\mathbb{R}^{d})$ that generate the conditioning on $\mathbf{X}$, where the construction is recursive by sequentially optimizing a certain decision criterion, starting from $\pi_{0}=\{\mathbb{R}^{d}\}$. It does not hurt to consider binary partitions, as the general case follows analogously. We then formulate our main problem as follows.
(P-general). Construct nested binary partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ such that $\mathbb{E}[(Y-M_{k})^{2}]$ is small for each $k$, where
$M_{k}(\omega):=\mathbb{E}[Y\mid\mathbf{X}\in A],\quad\text{for }\mathbf{X}(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k}.$ (2)
Clearly, if $Y$ is not a function of $\mathbf{X}$, the quantity $\mathbb{E}[(Y-M_{k})^{2}]$ has a strictly positive lower bound $\mathbb{E}[\mathrm{Var}(Y\mid\mathbf{X})]$. Another obstruction arises from atoms of $\mathbf{X}$, which prevents an effective split of the partitions. The relation (2) and its analogues will be essential to this paper.
The main theme of this paper is providing solutions to (P-general) in various settings. Roughly speaking, our main contribution can be summarized as follows.
(Solving P-general). If $\mathbb{E}[Y\mid\mathbf{X}]$ is well-behaved, then modulo the lower bound $\mathbb{E}[\mathrm{Var}(Y\mid\mathbf{X})]$ and the atoms of $\mathbf{X}$, there exists an explicit construction of $\{\pi_{k}\}_{k\geqslant 0}$, not depending on the law of $(\mathbf{X},Y)$, such that $\mathbb{E}[(Y-M_{k})^{2}]$ decays exponentially in $k$, where $M_{k}$ is defined in (2).
An instance of (Solving P-general) is Theorem 9 below, where the construction is given by the cyclic MinimaxSplit algorithm detailed in Section 3.1.
In the rest of the Introduction, we illustrate the motivations of (P-general) and summarize our contribution via two concrete applications: partition-based martingale approximations and regression trees.
1.1 Partition-based martingale approximations
Let $U$ be a real-valued atomless random variable. A partition-based martingale approximation of $U$ is a discrete-time martingale $\{M_{k}\}_{k\geqslant 0}$ such that
$M_{k}(\omega):=\mathbb{E}[U\mid U\in A],\quad\text{for }U(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k},$ (3)
for some nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}$, where we assume that the partitions are binary and $\pi_{0}=\{\mathbb{R}\}$. The martingale property follows from the tower property of conditional expectations.
The connection to (P-general) is as follows. Suppose that $d=1$ (i.e., $\mathbf{X}$ is real-valued, and will be denoted by $X$ in the following), and the coupling $(X,Y)$ satisfies that $X$ is atomless and $Y$ is a strictly increasing function of $X$. Since in this case $\mathbb{E}[\mathrm{Var}(Y\mid X)]=0$ and $X$ is atomless, we expect that the approximation error $\mathbb{E}[(U-M_{k})^{2}]$ vanishes as $k\to\infty$. Moreover, each partition in $X$ bijectively maps to a partition in $Y$, and hence (2) reduces to (3). In other words, in the current setting, (P-general) is equivalent to the following problem:
(P-martingale). Given a real atomless random variable $U$, construct a partition-based martingale approximation $\{M_{k}\}_{k\geqslant 0}$ of $U$ such that $\mathbb{E}[(U-M_{k})^{2}]\to 0$ exponentially in $k$.
Motivated by a martingale embedding problem, Simons (1970) first introduced the Simons martingale and established the a.s. convergence $M_{k}\to U$ (and hence also in $L^{2}$), but did not analyze the convergence rate. Another recent motivation for studying the convergence rate of the Simons martingale arises from the construction of powerful e-values in hypothesis testing (Ramdas and Wang, 2024). Zhang et al. (2024) (Lemma 5.6) proved that if $U\in L^{2+\delta}$ for some $\delta>0$ and $\{M_{k}\}$ is the Simons martingale, then there exist $C>0$ and $r\in(0,1)$ such that $\mathbb{E}[(U-M_{k})^{2}]\leqslant Cr^{k}$, and $r<0.827$ is feasible if $U$ is bounded. Our Theorem 3(ii) improves the rate to $r=1/2$ and hence provides a tighter theoretical bound. We also show by example that the rate $r=1/2$ is optimal.
The terminology martingale approximation has been used extensively in the probability literature with different meanings. In Rüschendorf (1985), it refers to the best approximation of a random vector by a (single) martingale based on ideas from optimal transport. In Borovskikh and Korolyuk (1997) and Hall and Heyde (2014), it refers to techniques from martingale theory (such as inequalities and CLT rates) with various applications in statistics. Similar techniques are also used in the study of stationary ergodic sequences (Wu and Woodroofe, 2004; Zhao and Woodroofe, 2008) and Markovian walks (Grama et al., 2018). Note that in our setting, $Y$ is a sum of martingale differences, but whose variance is bounded, and hence one cannot apply the martingale CLT and its convergence rates.
1.2 Regression trees
Consider a regression problem where we have a data set $(\mathbf{X}_{i},Y_{i})_{1\leqslant i\leqslant N}$, where $\mathbf{X}_{i}\in\mathbb{R}^{d}$ represents the covariates (or features) and $Y_{i}\in\mathbb{R}$ represents the responses. We follow van der Vaart and Wellner (2013) and assume the regression model:
$Y_{i}=g_{*}(\mathbf{X}_{i})+\varepsilon_{i},$ (4)
where $g_{*}:\mathbb{R}^{d}\to\mathbb{R}$ is the true signal function, $\{\mathbf{X}_{i}\}_{1\leqslant i\leqslant N}$ are i.i.d. sampled from a certain law, and $\varepsilon_{i}$’s are i.i.d. random (Gaussian) errors with zero mean. In other words, the data set is sampled from some coupling $(\mathbf{X}_{*},Y_{*})$ where $Y_{*}\mid\mathbf{X}_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon$ with $\mathbf{X}_{*}$ and $\varepsilon$ independent. The function $g_{*}$ can be estimated by minimizing the empirical $L^{2}$ risk (i.e., sum of squares; also called $L^{2}$ loss or mean-squared error (MSE) in this paper):
$\hat{m}_{N}=\operatorname*{arg\,min}_{m\in\mathcal{M}}\frac{1}{N}\sum_{i=1}^{N}\left(Y_{i}-m(\mathbf{X}_{i})\right)^{2},$ (5)
where $\mathcal{M}$ denotes the class of candidate functions for the mean $m$. The quantity (5) serves as an approximation to $\mathbb{E}[\left(Y_{*}-m(\mathbf{X}_{*})\right)^{2}]$. We usually establish the consistency of the estimator $\hat{m}_{N}$ by showing the convergence in probability:
[…]
Here, the underlying probability measure is the true data generating mechanism $\mathbb{P}$, the joint law of $(\mathbf{X}_{*},Y_{*})$.
A decision tree regression model for (4) is a non-parametric model for constructing $\hat{m}_{N}$ by partitioning the input space into distinct regions $A_{1},A_{2},\dots$ and fitting a simple model $\hat{m}_{N,i}$ to each region $A_{i}$ (Breiman et al., 1984). In this case, $\hat{m}_{N}$ takes the form of
[…]
and the $\mathcal{M}$ in (5) is the class of all piecewise constant functions. The construction procedures of $A_{i}$’ s and $\hat{m}_{N,i}$’ s are data dependent, where the regions are associated to a tree grown sequentially by maximizing the chosen decision criteria, and hence the parameters associated with each tree node are chosen recursively.
The recursive nature of the decision tree also indicates the connection to (P-general). Let $\mathbb{P}_{N}$ denote the empirical measure of the training data and $(\mathbf{X},Y)\buildrel\mathrm{law}\over{\sim}\mathbb{P}_{N}$. A (simplified variant of the) decision tree algorithm produces nested binary partitions $\{\pi_{k}\}_{k\geqslant 0}$ with axis-aligned borders of the input space $\mathbb{R}^{d}$ based on the law of $(\mathbf{X},Y)$, where $\pi_{0}=\{\mathbb{R}^{d}\}$ and the index $k$ is the depth of the tree (Breiman et al., 1984). The prediction $M_{k}$ given by a tree of depth $k$ is defined by (2):
$M_{k}(\omega)=M_{k(N)}(\omega):=\mathbb{E}[Y\mid\mathbf{X}\in A],\quad\text{for }\mathbf{X}(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k}.$
Note that this notation $M_{k}$ can be interpreted as substituting a random input $\mathbf{X}$ into (7). The estimator of $g_{*}$ is then given by $\hat{m}_{k}(\mathbf{x})=\mathbb{E}[Y\mid\mathbf{X}\in A]$ for $\mathbf{x}\in A,\leavevmode\nobreak\ A\in\pi_{k}$.
To facilitate our discussion, we assume that the number of samples in any $A_{i}\in\pi_{k}$ is nonzero and tends to $\infty$ as $k\rightarrow\infty$ uniformly, and moreover, $N/2^{k}$ is exponentially growing in $k$ [^1] Gordon and Olshen (1984) (Theorem 3.6) imposed a similar set of stronger assumptions that require specific growth rates. Such an assumption will not be required in any of the main results in this paper, but is convenient for the explanation here. In this case, we are interested in obtaining the fast convergence rates of (6), a problem formulated as follows.
(P-regression). Under the model (7) and the above assumption, construct a decision tree algorithm such that if the true signal function $g_{*}$ belongs to a certain wide function class, $\mathbb{E}[\left\lVert\hat{m}_{k}-g_{*}\right\rVert^{2}]$ decays exponentially in $k$.
The classic CART algorithm (Breiman et al. (1984)) for decision trees involves the following widely adopted greedy algorithm that recursively constructs partitions $\{\pi_{k}\}_{k\geqslant 0}$. To partition a set $A\in\pi_{k}$, the greedy algorithm selects a dimension and a threshold where the set $A$ is split into two subsets such that the total remaining risk is minimized. See (11) below for a formal definition. A simplified variant of CART without pruning, called VarianceSplit, is described in Section 3.1.
In the theoretical analysis of the convergence rate of CART in the regression setting, researchers often assume a lower bound on the variance decay when the algorithm splits each node in the decision tree (Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Mazumder and Wang, 2024; Cattaneo et al., 2024). For example, a prevalent variance decay assumption is known as the sufficient impurity decrease (SID) condition introduced by Chi et al. (2022), which requires that each split in the CART algorithm decreases the variance of the parent node by at least a constant factor, leading to exponential decay rates of the MSE of the form $r^{k}$ for some $r\in(0,1)$.
The first general result without the variance decay assumption was obtained by Klusowski and Tian (2024). Their main result asserts that modulo model mis-specification, the MSE of CART is upper bounded by $1/k$ at depth $k$; see (23) below. In a similar vein, Cattaneo et al. (2024) studied the convergence rates of the oblique CART and obtained a $1/k$ upper bound for MSE in the general setting (Theorem 1 therein) as well as an $r^{k}$ upper bound under further assumptions such as node sizes (Theorem 4 therein).
The unconditional rate guarantee of $1/k$ appears significantly slower than the exponential decay rates under variance decay assumptions. On the other hand, Cattaneo et al. (2022) discussed why CART may have slower-than-polynomial convergence. This phenomenon was known as the end-cut preference (ECP) in the literature (see Section 11.8 of Breiman et al. (1984)).
We have chosen to work on the model (7) in the simple regression setting (4), where sequential construction is prevalent (Luo and Pratola, 2023; Luo et al., 2024; Liu et al., 2023). Although there are multiple approaches to construct $\hat{m}_{N}$ in (7) sequentially (Loh, 2014), we focus on the CART-like decision tree model (Breiman et al., 1984) in the regression setting above. The classification setting has been well studied in the computational complexity and encoding literature (Blanc et al., 2020; O’Donnell et al., 2005) in terms of using uniform bounds on Boolean functions between partial trees and optimal trees.
1.3 Our contribution
Partition-based martingale approximations.
We show that various martingale constructions are feasible and enjoy exponential convergence rates. For instance, a classic construction by Simons (1970) is the following: start from $\pi_{0}=\{\mathbb{R}\}$ and for each $k\geqslant 0$ and $A=[a,b)\in\pi_{k}$ where $-\infty\leqslant a<b\leqslant\infty$, split $A$ into $[a,\mathbb{E}[U\mid U\in A])$ and $[\mathbb{E}[U\mid U\in A],b)$ to form the sets in $\pi_{k+1}$. We show that if $U$ is bounded, the convergence rate of the Simons martingale is $C2^{-k}$, where $C$ does not depend on the law of $U$. Moreover, the rate $2^{-k}$ can be shown to be asymptotically optimal.
Other constructions of partition-based martingale approximations involve replacing the conditional mean in the Simons martingale by other criteria such as the median. The explicit constructions are given formally by Definition 1, whose convergence rates are summarized in Theorem 3. All constructions yield explicit exponential rates of convergence in both the bounded and the unbounded cases under moment constraints (Theorem 4). Under further assumptions on the law of $U$, we also establish exponential rates of $Cr^{k}$ where $r$ can be made arbitrarily close to $1/4$ (a threshold that cannot be surpassed with binary partitions) while allowing $C$ to depend on the law of $U$ (Theorem 5).
Regression trees.
We introduce new splitting criteria, namely the MinimaxSplit algorithm and its multivariate variation, the cyclic MinimaxSplit algorithm (see (14) and (15) below for formal definitions). On one hand, the VarianceSplit algorithm is based on variance reduction; that is, for each split, the remaining total variance within the children nodes is minimized (among the splits over all dimensions). On the other hand, by definition, the MinimaxSplit algorithm selects the split that minimizes the maximum variance within the children nodes (among splits over all dimensions), hence the term minimax. The cyclic MinimaxSplit algorithm selects the same active dimension on each level of the tree and the active dimension cycles over all dimensions.
We show that for the cyclic MinimaxSplit algorithm, the MSE at depth $k$ decays exponentially with rate $2^{-2k/(3d)}$ (given sufficient samples), without any further assumption (Theorem 9). We further develop empirical risk bounds for the cyclic MinimaxSplit algorithm (Theorem 10).
Paper outline.
The respective solutions to (P-general) in the partition-based martingale approximation and the regression tree settings will be detailed in Sections 2 and 3. All technical proofs are given in Appendix A.
2 On partition-based martingale approximations
In this section, we develop the framework of partition-based martingale approximations and show that a number of constructions feature exponential convergence rates, thus answering (P-martingale).
2.1 Basic concepts
We say that a sequence of finite partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ is nested if for each $k\geqslant 0$ and $A\in\pi_{k+1}$, there exists (a unique) $B\in\pi_{k}$ such that $A\subseteq B$. Consider nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}$ and a real-valued random variable $U$ with a finite second moment. Recall (3):
$M_{k}(\omega):=\mathbb{E}[U\mid U\in A],\quad\text{for }U(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k},$ (8)
where the MSE $\mathbb{E}[(U-M_{k})^{2}]$ is non-increasing in $k$ by the nested property of $\{\pi_{k}\}_{k\geqslant 0}$ and the total variance formula.
A discrete-time real-valued stochastic process $\{M_{k}\}_{k\geqslant 0}$ is a martingale if for any $0\leqslant\ell<k$, $\mathbb{E}[M_{k}\mid\sigma(M_{0},\dots,M_{\ell})]=M_{\ell}$. We use the abbreviation $\sigma(\pi_{k})$ to denote the $\sigma$-algebra generated by events of the form $\{U\in A\}_{A\in\pi_{k}}$. If $\Pi:=\{\pi_{k}\}_{k\geqslant 0}$ is a nested sequence of partitions, $\{\sigma(\pi_{k})\}_{k\geqslant 0}$ is a filtration generated by indicator functions of sets in $\pi_{k}$’s (Doob, 1953). It follows from the tower property of conditional expectations that the sequence $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ is a martingale (in fact, a Doob martingale (Doob, 1940, 1953)). We call this martingale the $\Pi$-based martingale approximation of the random variable $U$, or in general a partition-based martingale approximation that approximates the random variable in terms of MSE.
Unless $U$ is a constant, there are different kinds of partition, so there are different partition-based martingale approximations depending on the distribution of $U$. Our goal in this section is to identify a few partition-based martingale approximations that efficiently approximate $U$, where the construction algorithm is universal; see (P-martingale). The efficiency criterion is given by the decay of the MSE $\mathbb{E}[(U-M_{k})^{2}]$.
Without loss of generality, we give our construction of a partition of a generic interval $[a,b)\subset\mathbb{R}$, where $a,b\in\mathbb{R}\cup\{\pm\infty\}$. [^2] Here we slightly abuse notation that $[a,b)=(-\infty,b)$ if $a=-\infty$. The sequence of partitions $\Pi$ in the response space will be constructed recursively, where $\pi_{0}=\{\mathbb{R}\}$ and for every $k\geqslant 0$, each interval $A\in\pi_{k}$ splits into two intervals, by following the same construction, forming the elements in $\pi_{k+1}$. In the following, we introduce four distinct splitting rules that define partition-based martingale approximations. For simplicity, we assume that $U$ is atomless, so that the endpoints of the intervals do not matter and that no trivial split occurs.
Definition 1.
Suppose we are given an atomless law of $U$ and a non-empty interval $I=[a,b)$, where $a,b\in\mathbb{R}\cup\{\pm\infty\}$.
  1. (i)
    Define
    $u_{\text{var}}=\operatorname*{arg\,min}_{u\in I}\big{(}\mathbb{P}(U\in[a,u))\mathrm{Var}(U\mid U\in[a,u))+\mathbb{P}(U\in[u,b))\mathrm{Var}(U\mid U\in[u,b))\big{)}.$ (9)
    If the minimizer is not unique, we pick the largest minimizer. The variance splitting rule (corresponding to the VarianceSplit algorithm in Section 3.1) splits $I$ into the two sets $[a,u_{\text{var}})$ and $[u_{\text{var}},b)$.
  2. (ii)
    Define $u_{\text{Simons}}=\mathbb{E}[U\mid U\in I]$. The Simons splitting rule splits $I$ into the two sets $[a,u_{\text{Simons}})$ and $[u_{\text{Simons}},b)$.
  3. (iii)
    Define
    $u_{\text{minimax}}=\operatorname*{arg\,min}_{u\in I}\max\big{\{}\mathbb{P}(U\in[a,u))\mathrm{Var}(U\mid U\in[a,u)),\,\mathbb{P}(U\in[u,b))\mathrm{Var}(U\mid U\in[u,b))\big{\}}.$
    If the minimizer is not unique, we pick the largest minimizer. The minimax splitting rule (corresponding to the MinimaxSplit algorithm in Section 3.1) splits $I$ into the two sets $[a,u_{\text{minimax}})$ and $[u_{\text{minimax}},b)$.
  4. (iv)
    Define
    $u_{\text{median}}=\sup\{u\in I:\mathbb{P}(a\leqslant U<u)=\mathbb{P}(u\leqslant U<b)\}.$
    The median splitting rule splits $I$ into two sets $[a,u_{\text{median}})$ and $[u_{\text{median}},b)$.
In turn, when applying recursively the variance (resp. Simons, minimax, median) splitting rule starting from $\pi_{0}=\{\mathbb{R}\}$, we obtain a nested sequence of partitions $\Pi=\{\pi_{k}\}_{k\geqslant 0}$ depending on the law of $U$. We call the resulting $\Pi$-based martingale approximation $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ the variance (resp. Simons, minimax, median) martingale (with respect to $U$).
Example 2.
If $U$ is uniformly distributed on a compact interval, all four martingales coincide. For example, if $U\sim\mathrm{Uni}[0,1]$, each of the four martingales from Definition 1 satisfy $M_{k}=\mathbb{E}[U\mid\sigma(\pi_{k})]$, where $\pi_{k}:=\{[j/2^{k},(j+1)/2^{k}):0\leqslant j<2^{k}\}$ and it follows that $\mathbb{E}[(U-M_{k})^{2}]\asymp 4^{-k}$ (where $\asymp$ means up to universal constants).
The intuition for the minimax and median martingales is that at each splitting step, we balance the “sizes” of $U{\mathbbm{1}}_{\{U\in I\}}$ on the two sets. The “size” corresponds to the (unconditional) variance for the minimax martingale and the total probability for the median martingale. Intending to minimize the MSE $\mathbb{E}[(U-M_{k})^{2}]$ at each step $k$, the variance martingale naturally arises as an algorithm that greedily reduces the remaining risk within $U$ in each iteration through layers.
In a partition-based martingale, the nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ can be naturally identified as the vertices of a binary tree. A binary tree is the unique infinite tree $(V,E)$ such that all vertices have degree $3$, except for a unique vertex $\emptyset$ called the root, which has degree $2$. Denote by $V_{k}$ the set of all vertices in $V$ with the graph distance from the root equal to $k$. Then, each vertex $v\in V_{k}$ can be identified as a set $A\in\pi_{k}$. With each edge $e=(A_{k},A_{k+1})\in E$ where $A_{k}\in\pi_{k}$ and $A_{k+1}\in\pi_{k+1}$, we may associate a coefficient $p_{e}:=\mathbb{P}(U\in A_{k+1})\in[0,1]$. With each vertex $v=A_{k}$, we may associate a location $\ell_{v}:=\mathbb{E}[U\mid U\in A_{k}]$. It follows that $M_{k}$ is supported on the discrete points $\{\ell_{v}\}_{v\in V_{k}}$ and furthermore,
$\mathbb{E}[(M_{k}-M_{k+1})^{2}]=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}.$ (10)
This binary tree representation (10) of the risk will be frequently used in this paper. In the next two sections, we discuss two types of results: uniform rates for bounded $U$ (Section 2.2) and non-uniform rates (Section 2.3) where the asymptotic constant may depend on the possibly unbounded law of $U$.
2.2 Uniform convergence rates
Theorem 3.
Let $U$ be a $[0,1]$-valued atomless random variable and $\{M_{k}\}_{k\geqslant 0}$ be one of the four martingale approximations with respect to $U$ given by Definition 1. The following statements hold.
  1. (i)
    If $\{M_{k}\}_{k\geqslant 0}$ is the variance martingale, $\mathbb{E}[(U-M_{k})^{2}]\leqslant 2.71\cdot 2^{-2k/3}$.
  2. (ii)
    If $\{M_{k}\}_{k\geqslant 0}$ is the Simons martingale, $\mathbb{E}[(U-M_{k})^{2}]\leqslant 2^{1-k}$.
  3. (iii)
    If $\{M_{k}\}_{k\geqslant 0}$ is the minimax martingale, $\mathbb{E}[(U-M_{k})^{2}]\leqslant 0.4\cdot 2^{-2k/3}$.
  4. (iv)
    If $\{M_{k}\}_{k\geqslant 0}$ is the median martingale, $\mathbb{E}[(U-M_{k})^{2}]\leqslant 2^{-k}$.
The general case of a bounded $U$ can be derived by a scaling argument, since the constructions in Definition 1 are scale-invariant. Let us sketch the arguments for the minimax martingale, as the same idea becomes crucial when applied to minimax decision trees we develop in Section 3.
Assuming that $U$ is atomless and supported in $[0,1]$, we apply the representation (10). By construction, $\sum p_{(v_{k},v_{k+1})}=1$ and $\sum|\ell_{v_{k}}-\ell_{v_{k+1}}|\leqslant\sup\operatorname{supp}U-\inf\operatorname{supp}U=1$. Moreover, by the minimax property and the law of total variance, it is straightforward to verify (with details in Section 3.2) that the variances decay at least geometrically by half at each split at $u_{\text{minimax}}$, i.e., for any $v\in V_{k}$ with $v^{\prime}\sim w$ for $v^{\prime}\in V_{k-1}$,
$\max_{w\in V_{k+1},\,v\sim w}p_{(v,w)}(\ell_{v}-\ell_{w})^{2}\leqslant\frac{1}{2}p_{(v^{\prime},v)}(\ell_{v^{\prime}}-\ell_{v})^{2}.$
Therefore, by induction,
$\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\leqslant L2^{-k}\mathrm{Var}(U)\leqslant L2^{-k},$
where $L>0$ is a universal constant that may not be the same on each occurrence. By Hölder’s inequality,
$\mathbb{E}[(M_{k}-M_{k+1})^{2}]$ $=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}$
$\leqslant\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}\bigg{)}^{1/3}$
×$\hskip 28.45274pt\times\bigg{(}\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}$
$\leqslant\bigg{(}\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\bigg{)}^{2/3}\leqslant L2^{-2k/3}.$
The rest then follows from the martingale property: $\mathbb{E}[(U-M_{k})^{2}]\leqslant\sum_{\ell\geqslant k}\mathbb{E}[(M_{\ell}-M_{\ell+1})^{2}]\leqslant L2^{-2k/3}$.
Finding the optimal rate $r\in(0,1)$ (where $\mathbb{E}[(U-M_{k})^{2}]\leqslant Lr^{k}$ for some universal constant $L$) is also an intriguing question. Observe that one cannot have $r<1/4$ by Example 2 above. In Examples 11 and 12 below, we show that the rate $r$ in Theorem 3 is indeed optimal for the Simons and median martingales.
2.3 Non-uniform convergence rates
The above Theorem 3 requires that the support of $U$ is bounded and the asymptotic constant depends only on the range of $U$, $\sup U-\inf U$. We show in the next result that, without the bounded range assumption, non-uniform convergence rates can still be guaranteed.
Theorem 4.
Let $U$ be an atomless random variable with a finite $(2+\varepsilon)$-th moment for some $\varepsilon>0$, and $\{M_{k}\}_{k\geqslant 0}$ be one of the four partition-based martingale approximations given by Definition 1. Then there exist constants $r\in(0,1)$ and $C>0$, both possibly depending on the law of $U$, such that
$\mathbb{E}[(U-M_{k})^{2}]\leqslant Cr^{k}.$
In particular, $M_{k}\to U$ both a.s. and in $L^{2}$.
If $U$ is bounded, Theorem 4 is a special case of Theorem 3. The case of the Simons martingale has been established by Zhang et al. (2024), and our proof of Theorem 4 will follow a similar route. Since the asymptotic constant is allowed to depend on the law of $U$, the optimal non-uniform rate $r$ might be smaller than the uniform ones (in Theorem 3).
Recall from Section 2.2 that the rates obtained in Theorem 3 are asymptotically optimal for the Simons and median martingales (see Examples 11 and 12). For the variance and minimax martingales, we show in the next result that the rate $r=1/4$ is optimal under certain regularity conditions on the law of $U$.
Theorem 5.
Let $\{M_{k}\}_{k\geqslant 0}$ be either the minimax or variance martingale converging to $U$. Suppose that $U$ is bounded and has a bounded and continuous density $f$ with $\inf f>0$ on $\operatorname{supp}U$, which we assume is a connected interval. Then for any $r>1/4$, there exists a constant $C>0$ (depending on $r$ and the law of $U$) such that
$\mathbb{E}[|U-M_{k}|^{2}]\leqslant Cr^{k}.$
In other words, under the assumptions in Theorem 5, the MSE has an asymptotic convergence rate of $1/4+\varepsilon$. Unfortunately, we are unable to remove the assumptions for the density of $U$ in Theorem 5 nor give counterexamples. We conjecture that the same conclusion of Theorem 5 holds without those assumptions.
3 The MinimaxSplit algorithm
The goal of this section is to provide a solution to (P-regression) by constructing decision tree algorithms with exponential convergence. Section 3.1 gives the constructions inspired by the minimax martingale construction in Section 2. In traditional regression decision tree settings that minimize (1), models greedily split in order to minimize the sum of variances over the responses of the children nodes (Liu et al., 2023; Breiman et al., 1984), yet we propose to greedily minimize the maximum variance among the children nodes.
In the following, we say that a law on $\mathbb{R}^{d}$ is marginally atomless if its projection onto any of the $d$ dimensions is atomless. We will start from a marginally atomless coupling (that may not arise from empirical measures $\mathbb{P}_{N}$) and prove exponential rates in Section 3.2. Section 3.3 builds upon this case, covering more generally the non-marginally atomless setting and, in particular, the empirical risk bounds. Further numerical analysis will be presented in Section 3.4. All proofs can be found in Appendix A.4.
3.1 Formulation of the algorithms
We begin by recapping the greedy splitting regime in the VarianceSplit algorithm to construct an efficient decision tree. Subsequently, we propose alternative CART algorithms, namely the MinimaxSplit algorithm, along with its variation, the cyclic MinimaxSplit algorithm, and make comparisons against the VarianceSplit algorithm. Roughly speaking, the (cyclic) MinimaxSplit algorithms are analogues of the minimax martingale from Definition 1. Hereafter, we denote by $[d]=\{1,\dots,d\}$ for $d\in\mathbb{N}$. We also need the following notion of a splittable set. If the coupling is defined on a continuous space, the split can always be performed; yet, for atomic measures like empirical measures, we may halt decision splits when there is no probability mass remained. Intuitively, in the following decision tree algorithms, the decision trees only split a node if it corresponds to a splittable set. A similar notion is also introduced by Definition 3.1 of Klusowski and Tian (2024).
Definition 3.1.
Consider a coupling $(\mathbf{X},Y)$ on $\mathbb{R}^{d+1}$. We say a set $A\subseteq\mathbb{R}^{d}$ is $(\mathbf{X},Y)$-non-splittable, or non-splittable, if any of the following occurs:
  • $\mathbb{P}(\mathbf{X}\in A)=0$;
  • $\mathbb{P}(\mathbf{X}\in A)>0$ and $Y\mid\mathbf{X}\in A$ is a constant;
  • $\mathbb{P}(\mathbf{X}\in A)>0$ and $\mathbf{X}\mid\mathbf{X}\in A$ is a constant.
Otherwise, we say that $A$ is $(\mathbf{X},Y)$-splittable, or splittable.
Revisiting the VarianceSplit algorithm.
The greedy VarianceSplit construction (Section 8.4 of Breiman et al. (1984) and Luo and Li (2024)) proceeds by introducing nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ with axis-aligned borders, in a way that sequentially and greedily minimizes the total risk $\mathbb{E}[(Y-M_{k})^{2}]$ at each step $k$ (while splitting from the fixed partition $\pi_{k-1}$), where $M_{k}$ is taken as the conditional mean of $Y$ given the $\sigma$-algebra generated by the collection ${\mathbbm{1}}_{\{\mathbf{X}\in A\}},\,A\in\pi_{k}$. So, the partitions $\{\pi_{k}\}_{k\geqslant 0}$ of the input space are nested and we require that each $\pi_{k}$ consists of sets that split each element $A\in\pi_{k-1}$ into two hyper-rectangles with axis-aligned borders, unless all covariates and/or all response values on the event $\{\mathbf{X}\in A\}$ are the same, in which case we do not perform split and $A\in\pi_{k}$.
Formally, suppose that a hyper-rectangle $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})$ belongs to the partition $\pi_{k-1}$ and is splittable. In particular, $M_{k-1}(\omega)=\mathbb{E}[Y\mid\mathbf{X}\in A]$ if $\mathbf{X}(\omega)\in A$. To find a split of $A$ at depth $k$, the VarianceSplit algorithm looks for a covariate $j\in[d]$ and $x_{j}\in[a_{j},b_{j})$ such that the remaining risk $\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]$ is the smallest after splitting the $A$ at $x_{j}$ in covariate $j$. In other words, one seeks for
$\begin{split}(j,x_{j})&=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big{(}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j})\\ &\hskip 56.9055pt+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{)}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big{(}\mathbb{E}\big{[}(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}\big{]}\\ &\hskip 56.9055pt+\mathbb{E}\big{[}(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}\big{]}\Big{)}\\ &=:\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}(V_{\text{left}}+V_{\text{right}}),\end{split}$ (11)
where we break ties arbitrarily. [^3] Strictly speaking, the $\operatorname*{arg\,min}$ in (11) may not always be attained for general couplings $(\mathbf{X},Y)$, unless we assume that $\mathbf{X}$ is marginally either atomless or has a finite support, which will always be the case in our analysis. However, in the general setting, if we allow splits in which atoms can be duplicated and assigned to both nodes (instead of only on the right node as in (11)), $\operatorname*{arg\,min}$ can be attained. In this paper, we will implicitly assume that $\operatorname*{arg\,min}$ is always attained. Consequently, the splittable set $A$ splits into the sets $A_{L}:=[a_{1},b_{1})\times\dots\times[a_{j},x_{j})\times\dots\times[a_{d},b_{d})$ and $A_{R}:=[a_{1},b_{1})\times\dots\times[x_{j},b_{j})\times\dots\times[a_{d},b_{d})$ to form its two descendants in the set $\pi_{k}$, and define
$M_{k}(\omega)=\begin{cases}\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{L};\\ \mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{R}.\end{cases}$ (12)
In other words, at depth $k\geqslant 0$ and after constructing the partition $\pi_{k}$, define $M_{k}$ by
$M_{k}(\omega)=\mathbb{E}[Y\mid\mathbf{X}\in A],\quad\text{if }\mathbf{X}(\omega)\in A,\quad\text{for }A\in\pi_{k}.$ (13)
This leads to the desired coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$, where $\{M_{k}\}_{k\geqslant 0}$ is a martingale. It is also common practice to weigh the terms $V_{\text{left}}$ and $V_{\text{right}}$ in (11) by the sample sizes of the left and right children nodes (resulting in a minimization of $n_{\text{left}}V_{\text{left}}+n_{\text{right}}V_{\text{right}}$), for which we call the Weighted Variance algorithm.
The MinimaxSplit algorithm.
A common feature of the MinimaxSplit and VarianceSplit algorithms is that they both start from a nested sequence of partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ consisting of (at-most) binary splits into hyper-rectangles with axis-aligned borders. The $k$-th approximation $M_{k}$ is defined by (13). The two algorithms differ in the way the partition is constructed: in the MinimaxSplit setting, the split is no longer the best split that reduces total variance but instead is the best split that minimizes the maximum variance within the two descendants. Formally, for a splittable set $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})\in\pi_{k-1}$, define the split location as
$\begin{split}(j,\hat{x}_{j})&=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}],\\ &\hskip 85.35826pt\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}]\Big{\}}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}),\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{\}}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}V_{\text{left}},V_{\text{right}}\Big{\}},\end{split}$ (14)
where we break ties arbitrarily. Consequently, the splittable set $A$ splits into the two hyper-rectangles $[a_{1},b_{1})\times\dots\times[a_{j},\hat{x}_{j})\times\dots\times[a_{d},b_{d})$ and $[a_{1},b_{1})\times\dots\times[\hat{x}_{j},b_{j})\times\dots\times[a_{d},b_{d})$ to form its two descendants in the set $\pi_{k}$. We call this the MinimaxSplit rule. After obtaining the partition $\pi_{k}$, define $M_{k}$ through (13). This leads to a coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$, where $\{M_{k}\}_{k\geqslant 0}$ is a martingale.
The cyclic MinimaxSplit algorithm.
The cyclic MinimaxSplit algorithm is a variation of the MinimaxSplit algorithm. Instead of optimizing over all covariates $j\in[d]$ in (14), we cycle through the $d$ covariates as the tree grows. That is, for $k\geqslant 1$, let $j=j_{k}=(k-1\text{ (mod }d))+1\in[d]$. For all splittable sets $A\in\pi_{k-1}$, we split $A$ in the $j$-th coordinate to form its two descendants at depth $k$. The split location is then defined as
$\begin{split}\hat{x}_{j}&=\operatorname*{arg\,min}_{x_{j}\in[a_{j},b_{j})}\max\Big{\{}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}],\\ &\hskip 85.35826pt\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}]\Big{\}}\\ &=\operatorname*{arg\,min}_{x_{j}\in[a_{j},b_{j})}\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}),\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{\}},\end{split}$ (15)
where we break ties arbitrarily. Note that if $\mathbf{X}$ is marginally atomless, the minimizer $\hat{x}_{j}$ is essentially unique, in the sense that $\mathbb{P}(\mathbf{X}\in[a_{1},b_{1})\times\dots\times(\hat{x}_{j},\hat{x}_{j}^{\prime})\times\dots\times[a_{d},b_{d}))=0$ between two minimizers $\hat{x}_{j}<\hat{x}_{j}^{\prime}$. Define $M_{k}$ by (13). This leads to the desired coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$, where again, $\{M_{k}\}_{k\geqslant 0}$ is a martingale aiming to approximate $Y$.
If $d=1$, the cyclic MinimaxSplit algorithm coincides with the MinimaxSplit algorithm. If $d>1$, compared to the MinimaxSplit algorithm, the cyclic MinimaxSplit algorithm is technically more tractable. Moreover, since all features are considered equally, the cyclic MinimaxSplit algorithm appears more robust in recovering the underlying geometry of the regression function in the setting of (4) in low dimensions, which we illustrate numerically in Section 3.4. On the other hand, the cyclic MinimaxSplit algorithm may be less effective in higher dimensions because the major dimensions may not be easily identified or reached in the first few splits.
(At-most-binary) tree representation.
In the context of regression trees, the same development of the tree representation (10) holds, except that the tree may not be a binary tree due to possible non-splittable sets. Instead, there exists a tree $(V,E)$ that is a sub-tree of a binary tree whose vertices are identified as the sets in the nested partitions. With the same notation leading to (10), we have
$\mathbb{E}[(M_{k}-M_{k+1})^{2}]=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}.$
The goal of the next section is to provide a partial answer to (P-general) under the cyclic MinimaxSplit rule. Assuming marginal non-atomicity, we show that up to universal constants and model mis-specification, the approximation error decays exponentially with rate $r^{k}$ for some explicit $r\in(0,1)$ that depends only on dimension $d$.
3.2 Cyclic MinimaxSplit risk bound
We first claim that if $\mathbf{X}$ is marginally atomless and $j\in[d]$, the $L^{2}$ loss
$\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j})$ (16)
is non-decreasing and continuous in $x_{j}$. To see this, we take $a<b$ and define events $E=E_{1}\cup E_{2}$, where $E_{1}=\{\mathbf{X}\in A,\,X_{j}<a\}$ and $E_{2}=E_{1}^{c}=\{\mathbf{X}\in A,\,a\leqslant X_{j}<b\}.$ By the total variance formula,
$\begin{split}\mathrm{Var}(Y\mid E)&=\mathbb{E}[\mathrm{Var}(Y\mid\sigma(E_{1},E_{2}))\mid E]+\mathrm{Var}(\mathbb{E}[Y\mid\sigma(E_{1},E_{2})]\mid E)\\ &\geqslant\mathbb{E}[\mathrm{Var}(Y\mid\sigma(E_{1},E_{2}))\mid E]\geqslant\mathbb{P}(E_{1}\mid E)\mathrm{Var}(Y\mid E_{1}).\end{split}$ (17)
Observe that $\mathrm{Var}(Y\mid E)=\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<b)$ and $\mathrm{Var}(Y\mid E_{1})=\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<a)$. Inserting these two expressions back into (17), we obtain
$\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<b)\geqslant\frac{\mathbb{P}(\mathbf{X}\in A,X_{j}<a)}{\mathbb{P}(\mathbf{X}\in A,X_{j}<b)}\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<a).$
Rearranging gives
$\mathbb{P}(\mathbf{X}\in A,\,X_{j}<a)\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<a)\leqslant\mathbb{P}(\mathbf{X}\in A,\,X_{j}<b)\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<b)$
and therefore (16) (or the first term inside the maximum in (15)) is non-decreasing; and similarly the second term inside the maximum is non-increasing. Moreover, by the marginally atomless property, both quantities are continuous. Therefore, if the $\operatorname*{arg\,min}$ is attained in (15) in the cyclic MinimaxSplit setting, the two terms inside the maximum are equal. Namely,
$\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})=\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}).$ (18)
Since the events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$ form a partition of $\{\mathbf{X}\in A\}$, we have again by the total variance formula,
$\begin{split}&\hskip 14.22636pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\\ &\leqslant\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}].\end{split}$ (19)
Combining (18) and (19), we have
$\begin{split}&\hskip 14.22636pt\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}\}}],\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}\}}]\Big{\}}\\ &\leqslant\frac{1}{2}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}].\end{split}$ (20)
In other words, the remaining risk on each descendant is at most half the risk on the parent node. Inductively, we gain a uniform geometric decay of the maximum risk among the nodes at depth $k$:
$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]\leqslant 2^{-k}\mathrm{Var}(Y).$ (21)
This is the key to controlling the “sizes of the nodes”.
In the case of the VarianceSplit construction, such uniform control is reminiscent—counterexamples exist due to ECP, which explains why assumptions such as the SID condition are favorable to ensure risk decay when using loss functions like (11) (Chi et al., 2022; Mazumder and Wang, 2024). To see the intuition, we consider the event $\{\mathbf{X}\in A\}$ and sub-events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$. Applying the total variance formula to this partition yields
$\text{Var}(Y)=$ $\mathbb{E}[\mathrm{Var}(Y\mid\sigma(A_{L}))]+\mathrm{Var}(\mathbb{E}[Y\mid\sigma(A_{L})])$
$\begin{split}=&\text{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\text{Var}(Y\mid A_{R})\mathbb{P}(A_{R})\\ &+\left(\mathbb{E}[Y\mid A_{L}]-\mathbb{E}[Y\mid A_{R}]\right)^{2}(1-\mathbb{P}(A_{L}))\mathbb{P}(A_{L}).\end{split}$ (22)
From this, we see that when ECP occurs, $\mathbb{P}(A_{L})\approx 1$ or 00, the improvement is dominated by the first two terms of (22): $\text{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\text{Var}(Y\mid A_{R})\mathbb{P}(A_{R})=V_{\text{left}}+V_{\text{right}}$. Therefore, when ECP occurs, the sum of variances does not necessarily decay fast enough for the VarianceSplit. However, for cyclic MinimaxSplit, $\max\{V_{\text{left}},V_{\text{right}}\}$ always displays geometric decay (21), hence “avoiding” the ECP phenomenon. In other words, the fast decay of the sum of variance is not easily guaranteed from the VarianceSplit construction due to ECP; but the geometric decay of the maximum variance can be assured by the MinimaxSplit construction.
Consider the additive function class
$\mathcal{G}:=\{g(\mathbf{x}):=g_{1}(x_{1})+\dots+g_{d}(x_{d})\},$
where $\mathbf{x}=(x_{1},\dots,x_{d})$. For $g\in\mathcal{G}$, define $\left\lVert g\right\rVert_{\mathrm{TV}}$ as the infimum of $\sum_{i=1}^{d}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}$ over all such additive representations of $g$. (A continuous version of) Theorem 4.2 of Klusowski and Tian (2024) states that if $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$ is constructed from the VarianceSplit algorithm, then for each $k\geqslant 1$,
$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\frac{\left\lVert g\right\rVert_{\mathrm{TV}}^{2}}{k+3}\Big{)}.$ (23)
In the next result, we show that the $1/k$ rate in (23) can be significantly improved to an exponential rate for the cyclic MinimaxSplit algorithm. This also partially answers (P-general) from the Introduction. Note that the splitting here is binary because of the marginal non-atomicity condition.
Theorem 6.
Suppose that $\mathbf{X}$ is marginally atomless. Let the coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$ be constructed from the cyclic MinimaxSplit algorithm. Then uniformly for any $\delta>0$ and $k\geqslant 0$,
$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant\inf_{g\in\mathcal{G}}\bigg{(}\Big{(}(1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big{)}\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]\\ &\hskip 56.9055pt+(1+\delta^{-1})\bigg{(}\frac{1}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg{)}.\end{split}$ (24)
Remark 7.
Since the bound (24) is uniform over $\delta,k,d$, one can further optimize over $\delta$ for fixed $k,d$. For example, taking $\delta=2^{-\lfloor k/d\rfloor/10}$, the right-hand side of (24) then has the more explicit (but cruder) upper bound
$\inf_{g\in\mathcal{G}}\Big{(}5\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+2^{1-\lfloor k/d\rfloor/2}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\Big{)}.$
Remark 8.
In the case where $\left\lVert g\right\rVert_{\mathrm{TV}}$ is much larger than the range $\Delta g:=\sup g-\inf g$, (24) can be further improved to
$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant\inf_{g\in\mathcal{G}}\bigg{(}(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+2^{1/3}(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{2/3}\\ &\hskip 113.81102pt+2^{-2/3}(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}(\Delta g)^{4/3}\bigg{)}.\end{split}$ (25)
Similar considerations also apply for Theorems 9 and 10 below.
Let us explain the intuition behind Theorem 6 when $d=1$ and $Y=g(X)$ for some $g\in\mathcal{G}$. Roughly speaking, (24) amounts to
$\mathbb{E}[(Y-M_{k})^{2}]\leqslant L2^{-2k/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}.$ (26)
To verify (26), observe that $\{M_{k}\}_{k\geqslant 0}$ is a martingale converging to $Y$. We employ the tree representation to control $\mathbb{E}[(M_{k}-M_{k+1})^{2}]$. By Hölder’s inequality and (21),
$\mathbb{E}[(M_{k}-M_{k+1})^{2}]$ $=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}$
$\leqslant\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}\bigg{)}^{1/3}$
×$\hskip 28.45274pt\times\bigg{(}\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}$
$\leqslant\mathrm{Var}(Y)^{2/3}2^{-2k/3}\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}.$
Since each split introduces a difference term $|\ell_{v_{k}}-\ell_{v_{k+1}}|$, which is bounded by the values of $Y$ conditioned on a partition in the covariate space, we have
$\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\leqslant\sum_{A\in\pi_{k+1}}\Big{(}\sup_{A}g-\inf_{A}g\Big{)}\leqslant\left\lVert g\right\rVert_{\mathrm{TV}}.$
On the other hand, $\mathrm{Var}(Y)=\mathrm{Var}(g(\mathbf{X}))\leqslant\left\lVert g\right\rVert_{\mathrm{TV}}^{2}$. Altogether (and using the martingale property), we obtain $\mathbb{E}[(Y-M_{k})^{2}]\leqslant L2^{-2k/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}$.
Let us compare the rates between (23) and (24). For $Y=g(\mathbf{X})$ with $g\in\mathcal{G}$ and $d$ fixed, we consider $\varepsilon>0$. For the VarianceSplit algorithm, (23) requires $2^{k}\approx e^{L\varepsilon^{-1}}$ nodes in the decision tree to guarantee that the MSE is smaller than $\varepsilon$ for some $L>0$. On the other hand, for the cyclic MinimaxSplit algorithm, (24) requires only $2^{k}\approx\varepsilon^{-Ld}$ nodes. Therefore, as $\varepsilon\to 0$, the cyclic MinimaxSplit algorithm requires fewer nodes to attain the same risk bound and thus facilitates computation.
Note that Theorem 6 exhibits the curse of dimensionality, which cannot be circumvented in the current setting to the best of our knowledge (Cattaneo et al., 2022; Tan et al., 2022).
3.3 Empirical risk bound for cyclic MinimaxSplit
We now focus on the regression setting and apply Theorem 9 to derive finite-sample performance guarantees. Suppose that under the original law $\mathbb{P}_{*}$, $\mathbf{X}_{*}$ is marginally atomless and
$Y_{*}\mid\mathbf{X}_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon:=\mathbb{E}[Y_{*}\mid\mathbf{X}_{*}]+\varepsilon,$
where the error $\varepsilon=Y_{*}-g_{*}(\mathbf{X}_{*})$ is sub-Gaussian, i.e., for some $\sigma>0$,
$\mathbb{P}_{*}(|\varepsilon|\geqslant u)\leqslant 2\exp\Big{(}-\frac{u^{2}}{2\sigma^{2}}\Big{)},\quad u\geqslant 0.$
Also, we consider the empirical law of $(\mathbf{X},Y)$ of $N$ samples from $\mathbb{P}_{*}$.
A limitation of Theorem 6 is that it applies only if the covariate $\mathbf{X}$ is marginally atomless (for (20) to hold). Therefore, we need the following result that deals with measures whose marginals may contain atoms. This also provides a more complete answer to (P-general).
Theorem 9.
Suppose that $\mathbf{X}$ is purely atomic [^4] The atomicity condition is only required such that the $\operatorname*{arg\,min}$ in (15) is well-defined; see Footnote 3. and for some $N>0$,
$\max_{j\in[d]}\,\max_{u\in\mathbb{R}}\mathbb{P}(X_{j}=u)\leqslant\frac{1}{N}.$ (27)
Let $M:=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$ and the coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$ be constructed from the cyclic MinimaxSplit algorithm. Then uniformly for any $\delta>0$ and $k\geqslant 0$,
$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}+\inf_{g\in\mathcal{G}}\bigg{(}\Big{(}(1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big{)}\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]\\ &\hskip 170.71652pt+(1+\delta^{-1})\bigg{(}\frac{2}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg{)}.\end{split}$ (28)
The compensation term
$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}$
appearing in (28) guarantees that if the number of samples $N$ is much larger than $2^{k}M^{2}$, the asymptotic upper bounds (24) and (28) are comparable.
[…]
Theorem 10.
Assume that $\mathbf{X}_{*}$ is marginally atomless. We have for $\delta\geqslant 2^{-2\lfloor k/d\rfloor/3}$,
$\mathbb{E}\big{[}\left\lVert g_{k}-g_{*}\right\rVert^{2}\big{]}$ $\leqslant\frac{C2^{k}(\log N)^{2}\log(Nd)}{N}+2\inf_{g\in\mathcal{G}}\bigg{(}\Big{(}(1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big{)}\left\lVert g-g_{*}\right\rVert^{2}$
$\hskip 156.49014pt+(1+\delta^{-1})\bigg{(}\frac{2}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg{)},$
where $C>0$ is a constant depending only on $\left\lVert g_{*}\right\rVert_{\infty}$ and $\sigma^{2}$.
For a fixed sample size $N$, if we assume that $g_{*}\in\mathcal{G}$ and $k=3d\log_{2}N/(3d+2)$ is a multiple of $d$, Theorem 10 has the further consequence that (under the same assumptions)
$\mathbb{E}\big{[}\left\lVert g_{k}-g_{*}\right\rVert^{2}\big{]}\leqslant CN^{-\frac{2}{3d+2}}\big{(}\left\lVert g_{*}\right\rVert_{\mathrm{TV}}+(\log N)^{2}\log(Nd)\big{)}.$ (29)
The proof of Theorem 10 follows essentially the same path as Theorem 4.3 for CART of Klusowski and Tian (2024) while replacing Theorem 4.2 therein by our Theorem 9.
3.4 Numerical experiments
In this section, we provide experiments that verify the advantages of MinimaxSplit and cyclic MinimaxSplit methods in one- and two-dimensional domains, as well as integrated ensemble methods applied to approximating a real-valued function, and therefore it can be used to the application of denoising images. Further numerics (including those in high dimensions) are provided in Appendix B.
3.4.1 One-dimensional input domain
Figure 1: The comparison of the optimal tree, Scikit-learn DecisionTreeRegressor, our MinimaxSplit tree, and VarianceSplit tree on a sinusoidal target function $y=f(x)+\varepsilon=\sin(4\pi x)+\varepsilon$, $x\in[0,1],\varepsilon\sim N(0,\sigma_{\varepsilon}^{2})$ with low (first row, noise variance $\sigma_{\varepsilon}^{2}=0.01$) and high (second row, noise variance $\sigma_{\varepsilon}^{2}=0.25$) noise levels. We illustrate the true function $f$ using solid red curves and use the same training set for each of the low-noise and high-noise cases; the training set represented by the solid dots is one of the 10 different training sets of size 100. We display 10 different model fits based on 10 different batches of training sets, and the averaged mean-squared error evaluated on a different training set across 10 batches.
As shown in Figure 1, our first experiment compares decision tree regression methods under varying noise conditions and offers notable insights into their relative performance and characteristics. The optimal tree is obtained using global optimization over all possible split values when conditioning on the tree structure and split coordinates (in our case, split coordinate can only be chosen along one coordinate since $x\in\mathbb{R}$). Specifically, for max_depth=3, it employs a fixed structure with $72^{3}-1=$ 7 split points, optimized simultaneously using sequential least squares programming to minimize the overall MSE on the training data. The topology of the tree is predetermined and enforced through constraints in the optimization process. The optimal tree approach contrasts with traditional variance-based, top-down decision tree algorithms by optimizing all splits at once, potentially capturing global data patterns more effectively. However, its fixed structure may limit adaptability to varying data complexities compared to conventional adaptive tree methods. The optimal tree serves as a baseline, particularly in high-noise scenarios, and suggests that its optimization criteria may be more effective at capturing the underlying function while minimizing overfitting.
On the other hand, our MinimaxSplit’s competitive performance, especially under high noise, aligns with the theoretical strengths of the minimax optimization in decision trees. Its ability to maintain fidelity to the underlying function in noisy conditions underscores the value of robust splitting criteria in challenging environments. This observation may motivate further exploration of minimax-based approaches in domains with uncertain data quality or high noise levels. As another reference, the standard Scikit-learn DecisionTreeRegressor and the decision tree splitting based on variance minimization show similar performance patterns, with notable degradation in high-noise scenarios. This similarity is expected given their likely use of variance reduction as the splitting criterion. However, their performance lags behind that of MinimaxSplit in noisy conditions, highlighting the flexibility of MinimaxSplit in the presence of significant noise. The varying relative performance suggests that ensemble methods leveraging the strengths of different approaches could yield more robust and adaptive models across diverse data conditions, which we detail in Appendix B. In conclusion, this study provides valuable insight into the behavior of decision tree regression methods under varying noise conditions, highlighting the potential for improved performance through custom optimization criteria and robust splitting methods.
3.4.2 Two-dimensional input domain
For this experiment, we use a synthetic dataset generated from a complex, non-linear function of two features, $\mathbf{x}=(x_{1},x_{2})$. The function is designed to simulate realistic challenges encountered in regression tasks, such as non-linearity and interaction between features. Specifically, the true function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ used for generating the target values is defined as:
$\begin{split}f(\mathbf{x})&=\left(2-2.1u_{1}^{2}+\frac{u_{1}^{4}}{3}\right)u_{1}^{2}+u_{1}u_{2}+\left(-4+4u_{2}^{2}\right)u_{2}^{2};\\ &u_{1}=3(x_{1}-0.5),\\ &u_{2}=3(x_{2}-0.5),\end{split}$ (30)
where $x_{1},x_{2}$ are independently and uniformly sampled from $[0,1]$. This function incorporates both quadratic and higher-order polynomial terms, creating a complex surface with multiple peaks and valleys. The visualizations of the prediction surfaces, as shown in Figure 2, further illustrate the improved accuracy and smoother transitions in the predicted values when using the proposed methods.
Figure 2: The predictions and true function values are visualized using heatmaps, allowing for a clear comparison of how well each model approximates the underlying function (30). The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models. The next three subplots depict the predictions of the $L^{1}$ VarianceSplit, $L^{1}$ MinimaxSplit, $L^{2}$ VarianceSplit, and $L^{2}$ MinimaxSplit. Top Row: max depth 2; Middle Row: max depth 6; Bottom Row: max depth 10.
In this example (as well as in Sections 3.4.3 and 3.4.4 below), we also incorporate an $L^{1}$ variation of the VarianceSplit and MinimaxSplit algorithms, where in the splitting rule, we replace the variances in (11), (14), and (15) by the $L^{1}$ distance from the mean. For instance, the $L^{1}$ variation of (11) is
$\begin{split}(j,x_{j})&=\operatorname*{arg\,min}_{(j,\,x_{j})}\Big{(}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathbb{E}\big{[}|Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}]|\mid\mathbf{X}\in A,\,X_{j}<x_{j}\big{]}\\ &\hskip 56.9055pt+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathbb{E}\big{[}|Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}]|\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\big{]}\Big{)}.\end{split}$ (31)
The optimal $L^{1}$ split point may not be unique, in which case we break ties arbitrarily.
3.4.3 Application to denoising images
Another notable experiment involves applying decision trees to denoise noisy images (Luo et al., 2024). We apply the different tree variants to predict pixel values based on their locations, effectively treating this as a regression problem. The experiments include the following configurations:
  1. 1.
    VarianceSplit with $L^{1}$ (see (31)) and $L^{2}$ (see (11)) norms: These trees use either $L^{1}$ or $L^{2}$ norms to measure errors. Variance $L^{1}$ trees tend to capture more outlier noise, while $L^{2}$ norm-based trees provide smoother approximations.
  2. 2.
    MinimaxSplit and cyclic MinimaxSplit: These methods, especially when paired with cyclic padding, are effective in balancing feature splits in high-dimensional image data. The cyclic MinimaxSplit method shows robustness against feature dominance, which often occurred when a few features had much stronger correlations than others.
The results show that mixed methods, where the error calculation alternates between minimax and variance-based splits, lead to a balance between local feature capturing and overall smoothness.
Figure 3: The noiseless image, noisy image and denoised images using different kinds of splitting criteria for a single tree, allowing for a detailed comparison of how well each model approximates the geometric structure. The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models. The next four subplots depict the predictions of the VarianceSplit, Weighted VarianceSplit, MinimaxSplit, cyclic MinimaxSplit for $L^{1}$ (top row) and $L^{2}$ (bottom row) decision trees. All trees are fitted with a max depth of 10.
The experiment shown in Figure 3 compares decision tree regression methods for denoising the classic Astronaut (Van der Walt et al., 2014) grayscale image, which has been preprocessed to 128 by 128. Subsequently, Gaussian noise with a standard deviation of 0.25 was added to create the noisy version. Examining the output plot reveals distinct characteristics of each method. The $L^{1}$ VarianceSplit approach (RMSE=0.140421) produces a piecewise constant approximation with visible rectangular artifacts, indicating a tendency to oversegment the image. The $L^{1}$ MinimaxSplit method (RMSE=0.114668) yields a smoother reconstruction with better preservation of large-scale features and edges, particularly evident in the feather of the subject’s outline. The $L^{2}$ VarianceSplit method (RMSE=0.138452) results in a slightly smoother output compared to its $L^{1}$ counterpart, but still exhibits noticeable blockiness. The $L^{2}$ MinimaxSplit approach (RMSE=0.113193) achieves the lowest RMSE, producing a reconstruction that balances smoothness with feature preservation, especially apparent in the gradual shading of the astronaut’s shoulder and the nuanced details around her head. The superior performance of the MinimaxSplit criteria, especially with the $L^{2}$ norm, can be attributed to its focus on minimizing worst-case errors. This strategy is particularly effective for image denoising, where it helps preserve important structural elements while smoothing out noise.
Unlike traditional methods that select the best feature at each split based solely on immediate error reduction, the cyclic MinimaxSplit method systematically cycles through the available features as the tree grows. At different levels of the tree, different features are chosen for splitting, based on a pre-determined order or cyclic pattern. This ensures that all features are considered fairly and thus reduces the risk of the model becoming overly dependent on a small subset of features. The cyclic nature emphasizes on modeling each dimension of the underlying function while avoiding heterogeneity across dimensions, which can be particularly beneficial in multivariate cases where feature correlations are complex.
3.4.4 From single tree to ensemble
Although a single decision tree can capture patterns in data through hierarchical partitioning, it often suffers from high variance in predictions, which makes it sensitive to small changes in the training dataset. To address these limitations, an ensemble approach can be utilized to combine multiple decision trees, leading to improved stability and predictive accuracy. In this section, we extend the MinimaxSplit and cyclic MinimaxSplit methods to an ensemble context.
The ensemble method we consider here is the random forest (Breiman et al., 1984), which aggregates multiple decision trees to improve the precision and robustness of the prediction. Each tree is grown to a fixed depth $k$, and different combinations of splitting criteria are used to determine how each subset of the data is partitioned. The ensemble predictions are then averaged to provide the final output (see Algorithm 1 in Appendix C).
Specifically, we use the MinimaxSplit, cyclic MinimaxSplit, and VarianceSplit algorithms as the base learners in a random forest setting. Each tree in the ensemble is trained on a bootstrap sample of the original dataset and predictions are made by averaging the output of individual trees, effectively reducing the overall variance of the model. The model aims to maintain a balance between bias and variance by taking advantage of the strengths of each base model while averaging individual weaknesses.
$L^{2}$ norm Refer to caption
$L^{1}$ norm Refer to caption
Figure 4: Predictions and true function values of (30), visualized using heatmaps. Top Row: The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models. The next three subplots depict the predictions of the variance-based decision tree, minimax decision tree, and cyclic minimax decision tree (depth $k=10$ with $L^{1}$ (top) and $L^{2}$ (bottom) norm), respectively. Each plot is annotated with the corresponding MSE to quantitatively assess the accuracy. Bottom Row: The final four subplots visualize the predictions from four different random forest models (using Algorithm 1 of Appendix C), each built with varying combinations of the three error methods. Again, each plot is labeled with the MSE to facilitate direct comparison.
We apply the above ensemble approach in the same setting as in Section 3.4.2. The experiment in Figure 4 involves training multiple regression models (VarianceSplit, MinimaxSplit, and cyclic MinimaxSplit) and ensemble models on the generated dataset. The results demonstrate that the MinimaxSplit+MinimaxSplit ensemble outperforms the traditional variance-based tree in terms of MSE, particularly in scenarios where the data distribution is non-uniform. Further discussions on a weighted aggregation regime will be provided in Appendix B.3.
4 Contribution and Future Work
In this study, we introduced a novel decision tree splitting strategy, the MinimaxSplit algorithm, along with its multivariate variant, the cyclic MinimaxSplit algorithm. Unlike traditional VarianceSplit methods, which aim to reduce overall variance in the children nodes, our MinimaxSplit algorithm seeks to minimize the maximum variance within the split partitions, thereby reducing the risk of overly biased partitions. The cyclic MinimaxSplit algorithm further ensures that each dimension is used in a balanced manner throughout the tree construction process, avoiding dominance by a subset of features.
A key theoretical result of this study is that the cyclic MinimaxSplit algorithm achieves an exponential MSE decay rate of given sufficient data samples, without requiring additional variance decay assumptions (Theorem 9). Furthermore, we derived empirical risk bounds for this method, establishing its robustness in different settings (Theorem 10). Along the same lines of techniques, we prove novel results on the convergence rates of univariate partition-based martingale approximations, which are of their own interest.
In addition to single-tree regression analysis, we explored ensemble learning approaches, where we combined decision trees constructed using MinimaxSplit, cyclic MinimaxSplit, and VarianceSplit methods within a random forest framework. The results demonstrated that a hybrid approach leveraging different splitting techniques yields superior performance, particularly when dealing with non-uniform data distributions (Section 3.4.4). These findings suggest that the MinimaxSplit approach can provide more stable and adaptive decision trees compared to traditional methods, especially in high-dimensional settings.
Future research could explore the MinimaxSplit algorithm in classification tasks, particularly for imbalanced datasets. We expect that an adaptive splitting strategy could replace cyclic selection, dynamically adjusting to feature importance. These directions may refine the theoretical foundations of MinimaxSplit and broaden its practical applications in high-dimensional and complex learning environments.
Acknowledgment
The authors thank Ruodu Wang for his valuable input. We provide our reproducible code at github.com/hrluo. HL was supported by U.S. Department of Energy under Contract DE-AC02-05CH11231 and U.S. National Science Foundation NSF-DMS 2412403.
Appendix A Proofs and further results
A.1 Proof of Theorem 3
Variance martingale
Proof for the variance martingale.
The first observation is that the split location is precisely the middle point of the two means of the two descendants. This is a special case of Theorem 1 of Ishwaran (2015) applied with $f$ monotone and $X\sim U(0,1)$, but for completeness, we provide a full argument here with our notation. To see this, consider (without loss of generality by shifting) the law of $U$ with mean zero, and let $u=u_{*}$ be the optimal split location. Let $p_{1},p_{2}$ and $m_{1},m_{2}$ be the associated one-step probabilities and locations in the binary tree representation. Since the second moment $\mathbb{E}[M_{1}^{2}]$ is maximized, we must have
$0=\frac{\mathrm{d}}{\mathrm{d}u}\mathbb{E}[M_{1}^{2}]\mid_{u_{*}}=\frac{\mathrm{d}}{\mathrm{d}u}(p_{1}m_{1}^{2}+p_{2}m_{2}^{2})\mid_{u_{*}}.$ (32)
Observe that
$\frac{\mathrm{d}p_{1}}{\mathrm{d}u}=-\frac{\mathrm{d}p_{2}}{\mathrm{d}u}=\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}$
and
$\frac{\mathrm{d}m_{1}}{\mathrm{d}u}\mid_{u_{*}}=\frac{(u_{*}-m_{1})}{p_{1}}\,\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}\mid_{u_{*}};\leavevmode\nobreak\ \frac{\mathrm{d}m_{2}}{\mathrm{d}u}\mid_{u_{*}}=\frac{(m_{2}-u_{*})}{p_{2}}\,\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}\mid_{u_{*}}.$
Inserting into (32) leads to
$\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}\mid_{u_{*}}\Big{(}m_{1}^{2}-m_{2}^{2}+2m_{1}(u_{*}-m_{1})-2m_{2}(u_{*}-m_{2})\Big{)}=0,$
which simplifies into $m_{1}+m_{2}=2u_{*}$ (in the case $\mathrm{d}\mathbb{P}_{U}/\mathrm{d}u(u_{*})=0$, we replace $u_{*}$ with the closest point of $u_{*}$ to $\operatorname{supp}U$ and this does not change the values of $p_{1},p_{2},m_{1},m_{2}$). This fact will be frequently used in the following.
At level $k$, $\mathbb{E}[(M_{k-1}-M_{k})^{2}]$ is a sum of $2^{k}$ terms in the binary tree representation (10), which we may rewrite as $\sum_{j}p_{j}d_{j}^{2}$ using a change of variable. This corresponds to the $2^{k}$ edges (indexed by $[2^{k}]$) in the binary tree between levels $k-1$ and $k$, and $p_{j}$ represents the probabilities of the edge $j$ and $d_{j}$ is the location difference between the end vertices of the edge, or the length of the edge. Our goal is to bound from above
$\mathbb{E}[(M_{k-1}-M_{k})^{2}]=\sum_{j=1}^{2^{k}}p_{j}d_{j}^{2}.$
The key is first to bound the quantity
$\max_{1\leqslant j\leqslant 2^{k-1}}(p_{2j-1}d_{2j-1}^{2}+p_{2j}d_{2j}^{2}).$
That is, the variance increases by splitting the $j$-th mass at level $k-1$ (which is the common ancestor of the $(2j-1)$-th and $(2j)$-th masses). By the martingale property,
$p_{2j-1}d_{2j-1}=p_{2j}d_{2j}.$
It follows that
$p_{2j-1}d_{2j-1}^{2}+p_{2j}d_{2j}^{2}=(p_{2j-1}+p_{2j})d_{2j-1}d_{2j}.$ (33)
The general strategy to analyze the quantity (33) is to start from an arbitrary mass/node $\mathbf{M}$ at level $k-1$, trace back its ancestors, and bound from above the probability of the mass (which is $p_{2j-1}+p_{2j}$ in the above expression) and the upper bounds for the length of the edges (which is $d_{2j-1},\,d_{2j}$).
The path leading to a mass $\mathbf{M}$ at level $k-1$ in the binary tree representation (10) consists of $k$ nodes: $\emptyset=\mathbf{N}_{0},\mathbf{N}_{1},\dots,\mathbf{N}_{k-1}=\mathbf{M}$, where $\mathbf{N}_{j}\in V_{j}$. We denote the lengths of the edges connecting the node $\mathbf{N}_{j-1}$ to its two descendants on level $j$ by $a_{j},b_{j},\,j\in[k-1]$, where $b_{j}=|\ell_{\mathbf{N}_{j-1}}-\ell_{\mathbf{N}_{j}}|$. It follows by the martingale property that the mass $\mathbf{M}$ has probability
$\prod_{j=1}^{k-1}q_{j}:=\prod_{j=1}^{k-1}\frac{a_{j}}{a_{j}+b_{j}}.$ (34)
The lengths of the edges of $\mathbf{M}$ are the mean distances of its children, which are bounded by the distances from $\sum_{j=1}^{k-1}b_{j}$ (the location of $\mathbf{M}$) to the endpoints of the interval that $\mathbf{M}$ carries in the splitting process. Therefore, it is natural to update these two distances as we trace the path from the root of the tree down to the node $\mathbf{M}$.
Let $a_{0}=b_{0}=1\geqslant\sup U-\inf U$. We consider the following algorithm, which starts from the tuple $(a_{0},b_{0},a_{1},b_{1},q_{1})$ and updates it $k-1$ times for each $j\in[k-1]$ when we trace the path down to $\mathbf{M}$. At step $j$, the tuple represents the two pairs $(a_{k},b_{k}),\,(a_{j},b_{j})$ that govern the maximum length of the edges of the $j$-th node (that is, the distance from its location to the boundaries of the intervals this node represents) for a certain $j\in[k-1]$, as well as the probability weight $q_{j}$ multiplied at this step. In a generic setting, suppose that we have $(a_{k},b_{k},a_{j},b_{j},q_{j})$ updated at the node $\mathbf{N}_{j}\in V_{j}$. This means that the edges forming $a_{k+1},\dots,a_{j}$ are all on one side of the path since only the last one of these contributes to the bound. Now when updating the node at level $j+1$ there are two cases:
  1. (i)
    $a_{j+1}$ and $a_{j}$ are on the same side. We update $(a_{k},b_{k},a_{j},b_{j},q_{j})$ with $(a_{k},b_{k},a_{j+1},b_{j+1},q_{j+1})$, where $a_{j+1}\leqslant(a_{j}+b_{j})/2$ and $b_{j+1}\leqslant(a_{k}+b_{k})/2$ (they are the two edges) and $q_{j+1}=a_{j+1}/(a_{j+1}+b_{j+1})$;
  2. (ii)
    $a_{j+1}$ and $a_{j}$ are on distinct sides. We update $(a_{k},b_{k},a_{j},b_{j},q_{j})$ with $(a_{j},b_{j},a_{j+1},b_{j+1},q_{j+1})$, where $b_{j+1}\leqslant(a_{j}+b_{j})/2$, $a_{j+1}\leqslant(a_{k}+b_{k})/2$, and $q_{j+1}=a_{j+1}/(a_{j+1}+b_{j+1})$.
Here, we have used (34) and the fact that $m_{1}+m_{2}=2x_{*}$ where $x_{*}$ is the split location.
Let $(a,b,a^{\prime},b^{\prime},q)$ be generic entries of the tuple and $(\hat{a},\hat{b},\hat{a}^{\prime},\hat{b}^{\prime},\hat{q})$ be the updated tuple, where we recall $\hat{q}=\hat{a}^{\prime}/(\hat{a}^{\prime}+\hat{b}^{\prime})$. We claim that
$\frac{(\hat{a}+\hat{b})(\hat{a}^{\prime}+\hat{b}^{\prime})}{(a+b)(a^{\prime}+b^{\prime})}\leqslant\frac{1}{2\hat{q}}.$ (35)
To see this, we separately consider the two cases above.
  • In case (i), $\hat{a}=a$ and $\hat{b}=b$, and
    $\hat{a}^{\prime}+\hat{b}^{\prime}=a_{j+1}+b_{j+1}=\frac{a_{j+1}}{\hat{q}}\leqslant\frac{a_{j}+b_{j}}{2\hat{q}}=\frac{a^{\prime}+b^{\prime}}{2\hat{q}}.$
  • In case (ii), $\hat{a}=a^{\prime}$ and $\hat{b}=b^{\prime}$, and
    $\hat{a}^{\prime}+\hat{b}^{\prime}=a_{j+1}+b_{j+1}=\frac{a_{j+1}}{\hat{q}}\leqslant\frac{a_{k}+b_{k}}{2\hat{q}}=\frac{a+b}{2\hat{q}}.$
This proves the claim.
Recall that our goal is to bound (33). The probability of the mass at each step is multiplied by $q$, and the edge lengths are bounded by $(a+b)/2$ and $(a^{\prime}+b^{\prime})/2$. By the above claim, their product multiplies by at most $1/2$ each time we update the tuple. We have thus shown that
$\max_{1\leqslant j\leqslant 2^{k-1}}(p_{2j-1}d_{2j-1}^{2}+p_{2j}d_{2j}^{2})\leqslant 2^{1-k}.$ (36)
Observe also that $\sum_{j}p_{j}=1$ and $\sum_{j}d_{j}\leqslant\sup U-\inf U\leqslant 1$ (because the edges at a single level cannot intersect each other since they represent the conditional means of $U$ in disjoint intervals). By Hölder’s inequality,
$\begin{split}\mathbb{E}[(M_{k-1}-M_{k})^{2}]=\sum_{j=1}^{2^{k}}p_{j}d_{j}^{2}&\leqslant\Big{(}\sum_{j=1}^{2^{k}}p_{j}d_{j}^{3}\Big{)}^{2/3}\Big{(}\sum_{j=1}^{2^{k}}p_{j}\Big{)}^{1/3}\\ &\leqslant\Big{(}\max_{j\in[2^{k}]}(p_{j}d_{j}^{2})\sum_{j=1}^{2^{k}}d_{j}\Big{)}^{2/3}\leqslant\max_{j\in[2^{k}]}(p_{j}d_{j}^{2})^{2/3}\leqslant 2^{-2(k-1)/3}.\end{split}$ (37)
By the martingale property, we then have
$\mathbb{E}[(U-M_{k})^{2}]=\sum_{j=k}^{\infty}\mathbb{E}[(M_{j}-M_{j+1})^{2}]\leqslant\sum_{j=k}^{\infty}2^{-2j/3}\leqslant\frac{2^{-2k/3}}{1-2^{-2/3}}\leqslant 2.71\cdot 2^{-2k/3},$
as desired. ∎
Simons martingale
Proof for the Simons martingale.
[…]
We start from $(A_{1},B_{1},q_{1})=(1,1,1)$. Again, denote the lengths of the edges connecting the node $\mathbf{N}_{j-1}$ to its two descendants at level $j$ by $a_{j},b_{j},\,j\in[k-1]$, where $b_{j}=|\ell_{\mathbf{N}_{j-1}}-\ell_{\mathbf{N}_{j}}|$. Consider the node $\mathbf{N}_{j}$ at level $j$. Then $(A_{j},B_{j},q_{j})$ must be of the form (while not distinguishing the order of $A_{j}$ and $B_{j}$)
$(A_{j},B_{j},q_{j})=\Big{(}b_{i}-\sum_{\ell=i+1}^{j}b_{\ell},\,b_{j},\,\frac{a_{j}}{a_{j}+b_{j}}\Big{)}$
for some $i\leqslant j-1$. This happens when the edges $a_{i+1},\dots,a_{j}$ lie on the same side of the path $(\mathbf{N}_{1},\dots,\mathbf{N}_{k})$. For the descendant $\mathbf{N}_{j+1}$ of $\mathbf{N}_{j}$, there are two possibilities:
  1. (i)
    if $a_{j+1}$ is on the same side of the path as $a_{j}$, we define
    $(A_{j+1},B_{j+1},q_{j+1})=\Big{(}b_{i}-\sum_{\ell=i+1}^{j+1}b_{\ell},\,b_{j+1},\,\frac{a_{j+1}}{a_{j+1}+b_{j+1}}\Big{)},$
    where $a_{j+1}\leqslant b_{j}$;
  2. (ii)
    otherwise, we define
    $(A_{j+1},B_{j+1},q_{j+1})=\Big{(}b_{j}-b_{j+1},\,b_{j+1},\,\frac{a_{j+1}}{a_{j+1}+b_{j+1}}\Big{)},$
    where $a_{j+1}\leqslant b_{i}-\sum_{\ell=i+1}^{j}b_{\ell}$.
The choice of $(A_{j+1},B_{j+1},q_{j+1})$ is justified by the separated tree property.
Let $q_{j+1}=a_{j+1}/(a_{j+1}+b_{j+1})$. We next express the ratio $A_{j+1}B_{j+1}/(A_{j}B_{j})$ after the update in step $j+1$ using $q_{j+1}$. We have in case (i) that
$\frac{\Big{(}b_{i}-\sum_{\ell=i+1}^{j+1}b_{\ell}\Big{)}b_{j+1}}{\Big{(}b_{i}-\sum_{\ell=i+1}^{j}b_{\ell}\Big{)}b_{j}}\leqslant\frac{b_{j+1}}{b_{j}}\leqslant\frac{b_{j+1}}{a_{j+1}}=\frac{1}{q_{j+1}}-1,$
and in case (ii),
$\frac{(b_{j}-b_{j+1})b_{j+1}}{\Big{(}b_{i}-\sum_{\ell=i+1}^{j}b_{\ell}\Big{)}b_{j}}\leqslant\frac{b_{j+1}}{b_{i}-\sum_{\ell=i+1}^{j}b_{\ell}}\leqslant\frac{b_{j+1}}{a_{j+1}}=\frac{1}{q_{j+1}}-1.$
In both cases, $A_{j+1}B_{j+1}/(A_{j}B_{j})\leqslant 1/q_{j+1}-1$. In other words, the contribution of the two edges of $\mathbf{M}$ is upper bounded by
$A_{k}B_{k}\prod_{\ell=1}^{k-1}q_{\ell}=\prod_{\ell=1}^{k-1}\frac{A_{\ell}B_{\ell}}{A_{\ell-1}B_{\ell-1}}\prod_{\ell=1}^{k-1}q_{\ell}\leqslant\prod_{\ell=1}^{k-1}\Big{(}\frac{1}{q_{\ell}}-1\Big{)}q_{\ell}=\prod_{\ell=1}^{k-1}(1-q_{\ell}),$
where we recall that $\{q_{\ell}\}_{1\leqslant\ell\leqslant k-1}$ are the one-step probabilities leading to $\mathbf{M}$. On the other hand, we also know that the probability that an edge of $\mathbf{M}$ carries must be bounded by $\prod_{\ell=1}^{k-1}q_{\ell}$. Therefore, we conclude that, with the decomposition
$\mathbb{E}[(M_{k-1}-M_{k})^{2}]=\sum_{j=1}^{2^{k}}p_{j}d_{j}^{2},$
it holds that for each $j\in[2^{k}]$, there exist $\{q_{\ell}\}_{1\leqslant\ell\leqslant k-1}$ such that
$p_{j}d_{j}^{2}\leqslant\prod_{\ell=1}^{k-1}(1-q_{\ell})\quad\text{and}\quad p_{j}\leqslant\prod_{\ell=1}^{k-1}q_{\ell}.$
Multiplying the two inequalities leads to
$p_{j}d_{j}\leqslant\sqrt{\prod_{\ell=1}^{k-1}q_{\ell}(1-q_{\ell})}\leqslant 2^{1-k}.$
Since this holds uniformly in $j$, we conclude that
$\mathbb{E}[(M_{k-1}-M_{k})^{2}]=\sum_{j=1}^{2^{k}}p_{j}d_{j}^{2}\leqslant\max_{j\in[2^{k}]}(p_{j}d_{j})\sum_{j=1}^{2^{k}}d_{j}\leqslant 2^{1-k}.$
By the martingale property,
$\mathbb{E}[(M_{k}-U)^{2}]=\sum_{j=k}^{\infty}\mathbb{E}[(M_{j}-M_{j+1})^{2}]\leqslant\sum_{j=k}^{\infty}2^{-j}\leqslant 2^{1-k}.$
This completes the proof. ∎
Minimax martingale
Proof for the minimax martingale.
At level $k$, there are $2^{k}$ vertices (denoted by $j\in[2^{k}]$) in the binary tree representation of the partition-based martingale approximation. Denote the probabilities of the vertices by $p_{j}$ (which correspond to $\mathbb{P}(U\in A_{j}),\leavevmode\nobreak\ A_{j}\in\pi_{k}$) and the locations by $\ell_{j}$. It follows that
$\mathbb{E}[(U-M_{k})^{2}]=\sum_{j=1}^{2^{k}}p_{j}\mathbb{E}[(U-\ell_{j})^{2}\mid U\in A_{j}]=:\sum_{j=1}^{2^{k}}p_{j}d_{j}^{2}.$
Note that $\sum_{j}p_{j}=1$ and $\sum_{j}d_{j}\leqslant\sup U-\inf U\leqslant 1$. It follows analogously as the Hölder’s inequality argument in (37) that
$\mathbb{E}[(U-M_{k})^{2}]$ $\leqslant\max_{j\in[2^{k}]}(p_{j}d_{j}^{2})^{2/3}.$
Let $\phi(u)=\mathbb{P}(U<u)\mathrm{Var}(U\mid U<u)$ and $\psi(u)=\mathbb{P}(U\geqslant u)\mathrm{Var}(U\mid U\geqslant u)$. Note that $\phi$ is increasing and $\psi$ is decreasing in $u$, and both functions are continuous if we assume that $U$ is atomless. Recall by the minimax property that at each step we pick a $u\in\mathbb{R}$ such that $\max\{\phi(u),\psi(u)\}$ is minimized. This means that $\phi(u)=\psi(u)$. Since $\mathrm{Var}(U)\geqslant\phi(u)+\psi(u)$, we have both $\phi(u)\leqslant\mathrm{Var}(U)/2$ and $\psi(u)\leqslant\mathrm{Var}(U)/2$. Inductively, we have for any $k$, $\max(p_{j}d_{j}^{2})\leqslant 2^{-k}\mathrm{Var}(U)\leqslant 2^{-2-k}$. Combining the above, we arrive at
$\mathbb{E}[(U-M_{k})^{2}]\leqslant\max_{j\in[2^{k}]}(p_{j}d_{j}^{2})^{2/3}\leqslant 2^{-4/3}2^{-2k/3}\leqslant 0.4\cdot 2^{-2k/3},$
as desired. ∎
Median martingale
Proof for the median martingale.
In the same setting as in the minimax case, we write
$\mathbb{E}[(U-M_{k})^{2}]=\sum_{j=1}^{2^{k}}p_{j}d_{j}^{2},$
where $\sum_{j}p_{j}=1$ and $\sum_{j}d_{j}\leqslant 1$. We further know that $\max_{j}p_{j}=2^{-k}$ by the median splitting construction. Therefore,
$\mathbb{E}[(U-M_{k})^{2}]=\sum_{j=1}^{2^{k}}p_{j}d_{j}^{2}\leqslant 2^{-k}\sum_{j=1}^{2^{k}}d_{j}^{2}\leqslant 2^{-k},$
as desired. ∎
A.2 Examples of optimal rates for partition-based martingale approximations
In the following two examples, we show that the geometric rates in Theorem 3 are optimal for the Simons and median martingales.
Example 11 (Rate $r=1/2$ is optimal for the Simons martingale).
Fix an arbitrary $s\in(1/2,1)$ and let $U$ satisfy $\mathbb{P}(U=-1)=1-s$ and for $J\geqslant 0$,
$\mathbb{P}\Big{(}U=\sum_{j=1}^{J}(\frac{1-s}{s})^{j}\Big{)}=s^{J+1}(1-s).$
It follows that $U$ is a bounded random variable. Next, we compute that
$\mathbb{E}\Big{[}U\mid U\geqslant\sum_{j=1}^{J}(\frac{1-s}{s})^{j}\Big{]}$ $=\bigg{(}\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\bigg{)}^{-1}\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\sum_{j=1}^{\ell}(\frac{1-s}{s})^{j}$
$=s^{-(J+1)}\sum_{\ell=J}^{\infty}s^{\ell+1}(2s-1)(1+(\frac{1-s}{s})^{\ell})$
$=\frac{s^{-(J+1)}(2s-1)s^{J+1}}{1-s}+\frac{s^{-(J+1)}(2s-1)s}{1-(1-s)}$
$=\sum_{j=1}^{J+1}(\frac{1-s}{s})^{j}.$
It is then straightforward to verify that in the Simons martingale $\{M_{k}\}_{k\geqslant 0}$, each $M_{k+1}-M_{k}$ is supported on the set $\{0,-((1-s)/s)^{k},((1-s)/s)^{k+1}\}$, where $\mathbb{P}(M_{k+1}-M_{k}=-((1-s)/s)^{k})=(1-s)s^{k}$ and $\mathbb{P}(M_{k+1}-M_{k}=((1-s)/s)^{k+1})=s^{k+1}$. It follows that
$\mathbb{E}[(M_{k+1}-M_{k})^{2}]=(1-s)^{2k+1}s^{-(k+1)}\asymp(\frac{(1-s)^{2}}{s})^{k}.$
In other words, the rate $r$ is given by $r=(1-s)^{2}/s$. If $s$ is picked close to $1/2$, $r$ can be made close enough to $1/2$. Therefore, $r=1/2$ is the optimal rate parameter. This example is illustrated with an atomic distribution $U$, but a slight twist also leads to an example for an absolutely continuous law $U$.
Example 12 (Rate $r=1/2$ is optimal for the median martingale).
Let $s\in(0,1)$ and a bounded random variable $U$ satisfy
$\mathbb{P}\Big{(}U=\sum_{j=1}^{k-1}s^{j-1}-s^{k-1}\Big{)}=2^{-k},\leavevmode\nobreak\ k\geqslant 1.$
Next, we compute that
$\mathbb{E}\Big{[}U\mid U\geqslant\sum_{j=1}^{k}s^{j-1}-s^{k}\Big{]}$ $=2^{k}\sum_{\ell=k+1}^{\infty}2^{-\ell}\bigg{(}\sum_{j=1}^{\ell-1}s^{j-1}-s^{\ell-1}\bigg{)}$
$=2^{k}\sum_{\ell=k+1}^{\infty}2^{-\ell}(\frac{1}{1-s}-\frac{s^{\ell-1}(2-s)}{1-s})$
$=\frac{1}{1-s}-\frac{2-s}{s(1-s)}\frac{s^{k+1}/2}{1-s/2}$
$=\sum_{j=1}^{k}s^{j-1}.$
Note also that
$\sum_{j=1}^{k}s^{j-1}-s^{k}<\sum_{j=1}^{k}s^{j-1}<\sum_{j=1}^{k+1}s^{j-1}-s^{k+1}.$
It is then straightforward to verify that the law of $M_{k+1}-M_{k}$ is supported on $\{0,\pm s^{k}\}$ with $\mathbb{P}(M_{k+1}-M_{k}=s^{k})=\mathbb{P}(M_{k+1}-M_{k}=-s^{k})=2^{-(k+1)}$. We then obtain that
$\mathbb{E}[(M_{k+1}-M_{k})^{2}]=2^{-k}s^{2k}.$
In other words, the rate $r=s^{2}/2$ has a limit $1/2$ as $s\to 1$ from below; hence, the optimal rate is $1/2$. Again, a simple twist yields similar examples, where the law of $U$ is absolutely continuous.
A.3 Proofs of Theorems 4 and 5
Proof of Theorem 4.
Let $\operatorname{supp}M_{1}=\{x,y\}$ where $x<y$. Take $M>\sup\{x,y\}$. Let $\tau_{M}^{+}$ be the first hitting time to $\{x:x\geqslant M\}$, $\tau_{M}^{-}$ be the first hitting time to $\{x:x\leqslant-M\}$, and $\tau_{M}=\min\{\tau_{M}^{+},\tau_{M}^{-}\}$. It is straightforward to prove (see Zhang et al. (2024)) that for some $r\in(0,1)$ and $C>0$,
$\mathbb{E}[(U-M_{k})^{2}]\leqslant\mathbb{E}[{\mathbbm{1}}_{\{\tau_{M}<\infty\}}U^{2}]+CM^{2}r^{k}.$
Since the supports of $\{M_{k}\}_{k\geqslant 0}$ correspond to the means of $U$ conditioned on nested partitions of $\mathbb{R}$, we have $\tau_{M}^{+}<\infty$ implies $M_{k}\geqslant x$ for all $k\geqslant 0$ and $M_{1}=y$. Therefore, conditional on the event $\tau_{M}^{+}<\infty$, $X$ is a martingale that is bounded from below by $x$. By Ville’s inequality,
$\mathbb{P}(\tau_{M}^{+}<\infty)$ $\leqslant\mathbb{P}(\tau_{M}^{+}<\infty\mid M_{1}>x)$
$=\mathbb{P}\left(\sup_{k\geqslant 0}M_{k}\geqslant M\mid M_{1}>x\right)\leqslant\frac{\mathbb{E}[U\mid M_{1}>x]-x}{M-x}=O(M^{-1}).$
Similarly, the same analysis holds for $\tau_{M}^{-}$. Hence, we conclude that $\mathbb{P}(\tau_{M}<\infty)=O(M^{-1}).$ The rest follows line by line as in Zhang et al. (2024), which we sketch for completeness. By Hölder’s inequality, $\mathbb{E}[{\mathbbm{1}}_{\{\tau_{M}<\infty\}}U^{2}]=O(M^{-\delta^{\prime}})$ for some $\delta^{\prime}>0$. With the choices $M=r^{-k/(2+\delta^{\prime})}$ and $q=r^{\delta^{\prime}/(2+\delta^{\prime})}\in(0,1)$, we have that for some $C>0$,
$\mathbb{E}[(U-M_{k})^{2}]\leqslant C(M^{-\delta^{\prime}}+M^{2}r^{k})\leqslant Cq^{-k}.$
The final statement follows immediately from the martingale convergence theorem. ∎
Proof of Theorem 5.
First, we claim that the split points for the nested partitions $\{\pi_{k}\}$ are dense. Indeed, for both martingales, we have shown that the variances on each component in the partition $\pi_{k}$ go to zero uniformly as $k\to\infty$. If the split points were not dense, some element in the partition would cover an interval of positive length, on which the variance of $U$ cannot vanish since we assumed $\inf f>0$ and $\sup f<\infty$. Therefore, for any $\varepsilon>0$, there exists $k$ such that the $k$-th partition $\pi_{k}$ is such that on each interval $A\in\pi_{k}$, $\sup_{A}f\leqslant(1+\varepsilon)\inf_{A}f$.
Since one-step variance is always better than one-step minimax, it remains to show that for an interval $A$ with $\sup_{A}f\leqslant(1+\varepsilon)\inf_{A}f$, there exists $C(\varepsilon)\downarrow 1/4$ as $\varepsilon>0$ such that at each step the total variance reduces by a factor of $C(\varepsilon)$ for the minimax martingale.
To see this, without loss of generality we start from $U$ supported on $[0,1]$ (meaning that $\mathbb{P}(U\in[0,1])=1$) whose density takes values in $[1,1+\varepsilon]$ (we may normalize afterwards to a probability measure, which we omit). Then for any interval $[a,b]\subseteq[0,1]$,
$\mathbb{E}[U\mid U\in[a,b]]\in\Big{[}a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)\Big{]}.$
Therefore,
$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])$ $\leqslant\sup_{m\in[a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)]}\int_{a}^{b}(1+\varepsilon)(x-m)^{2}\mathrm{d}x$
$\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(b-a)^{3}.$
Similarly,
$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])\geqslant\inf_{m\in[a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)]}\int_{a}^{b}(x-m)^{2}\mathrm{d}x\geqslant\frac{1}{12}(b-a)^{3}.$
It follows that the split point $x_{*}\in[0,1]$ must satisfy
$\frac{1}{12}x_{*}^{3}\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(1-x_{*})^{3}\quad\text{and}\quad\frac{1}{12}(1-x_{*})^{3}\geqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}x_{*}^{3}.$
Therefore, for some $a_{\varepsilon}\to 0$, $|x_{*}-1/2|\leqslant a_{\varepsilon}$.
A direct computation yields $\mathrm{Var}(U)\geqslant\frac{1}{12}$. The total remaining variance $\mathbb{E}[(U-M_{1})^{2}]$ is bounded by
$\mathbb{E}[(U-M_{1})^{2}]\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(x_{*}^{3}+(1-x_{*})^{3})\leqslant\frac{4(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(\frac{1}{2}+a_{\varepsilon})^{3}.$
Hence, as $\varepsilon\to 0$, the ratio becomes
$\frac{\mathbb{E}[(U-M_{1})^{2}]}{\mathrm{Var}(U)}\leqslant\frac{2(1+\varepsilon)^{4}}{(2+\varepsilon)^{3}}(1+2a_{\varepsilon})^{3}\to\frac{1}{4},$
as desired. ∎
A.4 Proofs of results from Section 3
Lemma A.1.
Let $\delta>0$. Then for any coupling $(X,Y)$ with $X\geqslant\mathbb{E}[Y]$, it holds that
$\mathrm{Var}(Y)\leqslant(1+\delta)\,\mathbb{E}[(X-Y)^{2}]+\frac{(1+\delta)^{3}}{\delta^{3}}(\sup\operatorname{supp}X-\inf\operatorname{supp}X)^{2}.$ (38)
Proof.
Let $C=(1+\delta)^{3}/\delta^{3}$. Without loss of generality, we assume that $\mathbb{E}[Y]=0$. Equivalent to (38) is
$(1+\delta)\,\mathbb{E}[X^{2}]-2(1+\delta)\,\mathbb{E}[XY]+\delta\mathbb{E}[Y^{2}]+C(\sup\operatorname{supp}X-\inf\operatorname{supp}X)^{2}\geqslant 0.$
After completing the square, it remains to show
$\mathbb{E}\Big{[}(\sqrt{\delta}\,Y-\frac{1+\delta}{\sqrt{\delta}}X)^{2}\Big{]}+\Big{(}C(\sup\operatorname{supp}X-\inf\operatorname{supp}X)^{2}-\frac{1+\delta}{\delta}\mathbb{E}[X^{2}]\Big{)}\geqslant 0.$ (39)
If $\sup\operatorname{supp}X/\inf\operatorname{supp}X\geqslant(1-\sqrt{(1+\delta)/(C\delta)})^{-1}$, the second term is always non-negative and hence (39) holds. Otherwise, since $C=(1+\delta)^{3}/\delta^{3}$,
$(1+\delta)(\inf\operatorname{supp}X)^{2}\geqslant(1+\delta)(1-\sqrt{(1+\delta)/(C\delta)})(\sup\operatorname{supp}X)^{2}\geqslant(\sup\operatorname{supp}X)^{2}.$
It follows that
$\mathbb{E}\Big{[}(\sqrt{\delta}\,Y-\frac{1+\delta}{\sqrt{\delta}}X)^{2}\Big{]}$ $\geqslant\mathbb{E}\Big{[}\sqrt{\delta}\,Y-\frac{1+\delta}{\sqrt{\delta}}X\Big{]}^{2}$
$=\frac{(1+\delta)^{2}}{\delta}\mathbb{E}[X]^{2}\geqslant\frac{(1+\delta)^{2}}{\delta}(\inf\operatorname{supp}X)^{2}\geqslant\frac{(1+\delta)}{\delta}(\sup\operatorname{supp}X)^{2}\geqslant\frac{1+\delta}{\delta}\mathbb{E}[X^{2}].$
This shows (39) and thus proves (38). ∎
Proof of Theorem 6.
We divide the proof into three steps.
Step I: reducing to an inequality involving $\mathrm{Var}(Y)$.
We claim that it suffices to prove
$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\bigg{(}(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}(\left\lVert g\right\rVert_{\mathrm{TV}}\mathrm{Var}(Y))^{2/3}\bigg{)}.$ (40)
To this end, we need an upper bound of $\mathrm{Var}(Y)$. Fix $g\in\mathcal{G}$ and let $g_{*}:=(\sup g+\inf g)/2$. By Minkowski’s inequality,
$\begin{split}\mathrm{Var}(Y)^{1/2}=\min_{y\in\mathbb{R}}\mathbb{E}[(Y-y)^{2}]^{1/2}\leqslant\mathbb{E}[(Y-g_{*})^{2}]^{1/2}&\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\mathbb{E}[(g(\mathbf{X})-g_{*})^{2}]^{1/2}\\ &\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\frac{1}{2}\left\lVert g\right\rVert_{\mathrm{TV}}.\end{split}$ (41)
Squaring both sides yields that for any $\delta>0$,
$\mathrm{Var}(Y)\leqslant(1+\delta)\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\frac{1+\delta^{-1}}{4}\left\lVert g\right\rVert_{\mathrm{TV}}^{2},$
where we have used $(a+b)^{2}\leqslant(1+\delta)a^{2}+(1+\delta^{-1})b^{2}$. It follows that
$\begin{split}\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\mathrm{Var}(Y)^{2/3}&\leqslant\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\bigg{(}(1+\delta)^{2/3}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{2/3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{4/3}\bigg{)}\\ &\leqslant\frac{2(1+\delta)}{3}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\bigg{(}\frac{1}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}\left\lVert g\right\rVert_{\mathrm{TV}}^{2},\end{split}$ (42)
where we have used the inequality $x^{1/3}y^{2/3}\leqslant x/3+2y/3$, which follows from the concavity of logarithm and Jensen’s inequality. Inserting (42) into (40) yields (24).
Step II: decomposition of the risk.
It remains to establish (40). Denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ from the cyclic minimax construction. Note that by construction, #$\#J_{k}\leqslant 2^{k}$ and $M_{k}\mid\mathbf{X}\in I_{j}$ is a piecewise constant random variable, which takes a constant value $y_{k,j}$ inside the interval $I_{j}$, so $M_{k}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}=y_{k,j}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}$. By definition, $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$. Define $g_{j}=g|_{I_{j}}$, so that $\left\lVert g\right\rVert_{\mathrm{TV}}=\sum_{j}\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$. Define $\mathrm{Ran}(g_{j})=[\inf g_{j},\sup g_{j}]$ and recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
The key step is to decompose the local mean-squared error $\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ into two parts for each $j$, with the help of Lemma A.1. Next, we claim that
$\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]=\mathbb{E}[(Y-y_{k,j})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\leqslant U_{j}+V_{j},$ (43)
where $U_{j},V_{j}$ are defined in the following. We divide into two separate cases depending on the location of $y_{k,j}$:
  1. (i)
    $y_{k,j}\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and $V_{j}=(1+\delta^{-1})\,\mathbb{E}[(g(\mathbf{X})-y_{k,j})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$.
  2. (ii)
    $y_{k,j}\not\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and $V_{j}=p_{j}(\Delta g_{j})^{2}(1+\delta)^{3}/\delta^{3}$.
In case (i), (43) follows immediately from the polarization identity $(a+b)^{2}\leqslant(1+\delta)a^{2}+(1+\delta^{-1})b^{2}$. On the other hand, for case (ii), (43) follows by applying Lemma A.1 to the coupling $(g(\mathbf{X}),Y)\mid\mathbf{X}\in I_{j}$ and using that $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$. This completes the proof of (43).
Step III: controlling $\sum_{j}U_{j}$ and $\sum_{j}V_{j}$.
We have by construction and (43) that
$\mathbb{E}[(Y-M_{k})^{2}]=\sum_{j\in J_{k}}\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\leqslant\sum_{j\in J_{k}}U_{j}+\sum_{j\in J_{k}}V_{j}.$ (44)
The first term is easy to control by definition:
$\sum_{j\in J_{k}}U_{j}=\sum_{j\in J_{k}}(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}].$ (45)
To control $\sum_{j}V_{j}$, we apply the same Hölder’s inequality argument as in the proof of Theorem 3 in the minimax case. We have
$\begin{split}\sum_{j\in J_{k}}V_{j}&\leqslant\Big{(}\sum_{j\in J_{k}}p_{j}\big{(}\frac{V_{j}}{p_{j}}\big{)}^{3/2}\Big{)}^{2/3}\Big{(}\sum_{j\in J_{k}}p_{j}\Big{)}^{1/3}\\ &\leqslant\Bigg{(}(\max_{j\in J_{k}}V_{j})\sum_{j\in J_{k}}\sqrt{\frac{V_{j}}{p_{j}}}\Bigg{)}^{2/3}\leqslant\Bigg{(}(\max_{j\in J_{k}}\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}])\sum_{j\in J_{k}}\sqrt{\frac{V_{j}}{p_{j}}}\Bigg{)}^{2/3}.\end{split}$ (46)
To further bound the right-hand side of (46), we recall from (21) that as a consequence of the marginally atomless property,
$\max_{j\in J_{k}}\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\leqslant 2^{-k}\mathrm{Var}(Y).$ (47)
Next, we claim that
$\sqrt{\frac{V_{j}}{p_{j}}}\leqslant(1+\delta^{-1})^{3/2}\Delta g_{j}.$ (48)
This is immediate for case (ii) above (when $y_{k,j}\not\in\mathrm{Ran}(g_{j})$). By definition and using $y_{k,j}\in\mathrm{Ran}(g_{j})$, we have for case (i) above (when $y_{k,j}\in\mathrm{Ran}(g_{j})$),
$V_{j}\leqslant(1+\delta^{-1})p_{j}(\Delta g_{j})^{2}<p_{j}(1+\delta^{-1})^{3}(\Delta g_{j})^{2},$
as desired. On the other hand, suppose that $k=dq+r$ for some integer $q$ and $0\leqslant r<d$, and denote by $\pi_{k}$ the resulting partition of $\mathbb{R}^{d}$. Observe (by induction on $q,r$, and elementary geometry) that
$\sum_{A\in\pi_{k}}\Big{(}\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\Big{)}\leqslant 2^{q(d-1)+r}\left\lVert g\right\rVert_{\mathrm{TV}}.$
In other words, we have that for $k\in\mathbb{N}$,
$\sum_{j\in J_{k}}\Delta g_{j}=\sum_{A\in\pi_{k}}\Big{(}\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\Big{)}\leqslant 2^{k-\lfloor k/d\rfloor}\left\lVert g\right\rVert_{\mathrm{TV}},$ (49)
[…]
$\sum_{j\in J_{k}}\sqrt{\frac{V_{j}}{p_{j}}}\leqslant(1+\delta^{-1})^{3/2}2^{k-\lfloor k/d\rfloor}\left\lVert g\right\rVert_{\mathrm{TV}}.$
Inserting into (46) gives
$\sum_{j\in J_{k}}V_{j}\leqslant((1+\delta^{-1})^{3/2}2^{k-\lfloor k/d\rfloor}\left\lVert g\right\rVert_{\mathrm{TV}}2^{-k}\mathrm{Var}(Y))^{2/3}\leqslant(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}(\left\lVert g\right\rVert_{\mathrm{TV}}\mathrm{Var}(Y))^{2/3}.$ (50)
Finally, inserting (45) and (50) into (44) yields
$\mathbb{E}[(Y-M_{k})^{2}]$ $\leqslant(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}(\left\lVert g\right\rVert_{\mathrm{TV}}\mathrm{Var}(Y))^{2/3}.$
This proves (40) and hence concludes the proof. ∎
Proof of Remark 8.
The bound (41) can be improved to
$\mathrm{Var}(Y)^{1/2}\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\frac{1}{2}\Delta g.$
Inserting the following inequality (which follows from $(a+b)^{p}\leqslant 2^{p-1}(a^{p}+b^{p})$ for $a,b\geqslant 0$ and $p\geqslant 1$)
$\mathrm{Var}(Y)^{2/3}\leqslant 2^{1/3}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{2/3}+2^{-2/3}(\Delta g)^{4/3}$
into (40) yields (25). ∎
Proof of Theorem 9.
The only difference from the proof of Theorem 6 is (47), since for non-marginally atomless measures (20) and (21) may not hold. Instead, by our assumptions, (20) is be replaced by
$\begin{split}&\hskip 14.22636pt\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}\}}],\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}\}}]\Big{\}}\\ &\leqslant\frac{1}{2}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]+\frac{M^{2}}{N},\end{split}$
because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $M^{2}/N$. As a consequence, if we denote by
$u_{k}:=\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}],$
then $u_{k}$ satisfies the recursive inequalities
$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{M^{2}}{N},\leavevmode\nobreak\ k\geqslant 0;\quad u_{0}=\mathrm{Var}(Y).$
To solve this, let $v_{k}:=u_{k}-2M^{2}/N$. Suppose that $u_{0}\geqslant 2M^{2}/N$. Then $v_{k}$ satisfies $v_{k+1}\leqslant v_{k}/2$ and $v_{0}\leqslant\mathrm{Var}(Y)$. It follows that $v_{k}\leqslant 2^{-k}\mathrm{Var}(Y)$ for all $k\geqslant 0$, and hence
$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]=u_{k}\leqslant 2^{-k}\mathrm{Var}(Y)+\frac{2M^{2}}{N}.$ (51)
It is also easy to check that (51) is also satisfied for the case $u_{0}<2M^{2}/N$, that is, (51) holds in general. Replacing (47) by (51) leads to
$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\bigg{(}(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\Big{(}\left\lVert g\right\rVert_{\mathrm{TV}}(\mathrm{Var}(Y)+2^{k+1}\frac{M^{2}}{N})\Big{)}^{2/3}\bigg{)}.$
Applying concavity of the function $x\mapsto x^{2/3}$ for $x>0$, (42), and that
$\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\Big{(}2^{k+1}\frac{M^{2}}{N}\Big{)}^{2/3}\leqslant\frac{\left\lVert g\right\rVert_{\mathrm{TV}}^{2}}{3}+\frac{2^{k+2}}{3}\frac{M^{2}}{N},$
we obtain (28). ∎
Proof of Theorem 10.
Note that if $\mathbf{X}_{*}$ is marginally atomless, the assumption (27) for the coupling $(\mathbf{X},Y)$ in Theorem 9 holds $\mathbb{P}_{*}$-almost surely. Consequently, the proof is almost verbatim compared to Theorem 4.3 of Klusowski and Tian (2024), by applying our Theorem 9 instead of Theorem 4.2 therein. The only difference here is the extra term
$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}$
appearing in (28), where we remind the reader that $M=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$. Nevertheless, since the proof of Theorem 4.3 of Klusowski and Tian (2024) deals first with the case with bounded data, we may simply replace $M$ by that bound. Specifically, the first part of the proof of Theorem 4.3 of Klusowski and Tian (2024) assumes $\max_{i}|Y_{i}|\leqslant U$. We may set $M=2U$ and add the term
$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}\leqslant\frac{2^{k+5}U^{2}}{3N}$
in (B.28) therein (where we used $\delta\geqslant 2^{-2\lfloor k/d\rfloor/3}$). This extra term is carried until (B.34) therein, where it can be absorbed by the term $U^{4}2^{k}\log(Nd)/N$ therein. The rest of the proof remains unchanged. ∎
Appendix B Further numerics
B.1 Convergence rates of partition-based martingale approximations
The experiments in Figure 5 are designed to compare the performance of four different partition-based martingale approximation methods defined in Definition 1—minimax, median, Simons, and variance—based on their ability to approximate a target function using nested partitions. The methods are applied to different probability densities over a specified interval, and the results are visualized in a series of plots. We want to evaluate how differently martingale approximation methods reduce the variance (or the second moment) of a random variable $Y$ that is distributed according to various density functions $f$. In Figure 5(a), we consider the density $f$ of $Y$ given by
$f(x)=\begin{cases}1/51&\text{if }x\in[0,0.9];\\ (1+10^{4}(x-0.9))/51&\text{if }x\in[0.9,1];\\ 0&\text{otherwise,}\end{cases}$ (52)
with four types of partition-based martingale approximations. The minimax method generally shows a steeper decline in MSE in the initial stages, suggesting that it is effective at quickly reducing the worst-case variance.
(a) Density given by (52)
Refer to captionRefer to caption
(b) Density given by $f(x)=11x^{10}$ on $[0,1]$
Figure 5: Plots of the log-MSE, $\log_{2}(\mathbb{E}[(Y-M_{k})^{2}])$ versus the approximation depth $k$ for four different methods, where the density of the law $Y$ is given: (a) by (52) and (b) by $f(x)=11x^{10}$ on $[0,1]$ and zero otherwise. The right panel of (b) plots the ratios $\mathbb{E}[(Y-M_{k+1})^{2}]/\mathbb{E}[(Y-M_{k})^{2}]$ for the four methods as a function of $k$.
Furthermore, the second plot in Figure 5(b) shows the ratio of the remaining risk after each step normalized by the variance of the previous step, as an indicator of rates of improvement as the tree grows deeper. The behavior of these curves can be used to infer the asymptotic rates of convergence for each method, which align with the theoretical analysis of Section 2.3.
B.2 Higher dimensional input domain
Consider a dimension $d\geqslant 1$ that is a multiple of $4$. The Powell function on $n$,
$f(x)=\sum_{i=1}^{d/4}[(x_{4i-3}+10x_{4i-2})^{2}+5(x_{4i-1}-x_{4i})^{2}+(x_{4i-2}-2x_{4i-1})^{4}+10(x_{4i-3}-x_{4i})^{4}],$
with uniformly random samples on $[0,1]^{d}$, serves as an exemplary testbed for decision tree methods in high-dimensional regression problems, offering a nuanced landscape of challenges that mirror real-world complexities. In the dense design scenario ($n>d>100$), the function’s inherent non-linearity and intricate variable interactions push the boundaries of decision trees’ capabilities. The quartic terms in the function introduce sharp curvatures in the response surface, necessitating sophisticated splitting strategies to approximate these non-linear relationships accurately. Moreover, the coupling of variables in quartets creates a web of local interactions that tests the trees’ ability to discern and model multivariable dependencies efficiently.
Sample Size Dimension Sklearn VarianceSplit MinimaxSplit
10 4 $4.48\times 10^{2}$ $2.01\times 10^{2}$ $2.57\times 10^{2}$
10 16 $1.60\times 10^{4}$ $1.52\times 10^{4}$ $1.52\times 10^{4}$
10 64 $3.70\times 10^{4}$ $4.24\times 10^{4}$ $3.79\times 10^{4}$
10 256 $1.35\times 10^{5}$ $1.24\times 10^{5}$ $1.14\times 10^{5}$
10 1024 $9.31\times 10^{5}$ $8.04\times 10^{5}$ $4.28\times 10^{5}$
100 4 $5.49\times 10^{1}$ $5.21\times 10^{1}$ $5.31\times 10^{1}$
100 16 $2.70\times 10^{3}$ $2.67\times 10^{3}$ $2.57\times 10^{3}$
100 64 $1.99\times 10^{4}$ $1.95\times 10^{4}$ $2.20\times 10^{4}$
100 256 $1.01\times 10^{5}$ $1.02\times 10^{5}$ $9.39\times 10^{4}$
100 1024 $4.65\times 10^{5}$ $4.29\times 10^{5}$ $4.02\times 10^{5}$
1000 4 $4.68\times 10^{1}$ $4.65\times 10^{1}$ $4.49\times 10^{1}$
1000 16 $1.85\times 10^{3}$ $1.85\times 10^{3}$ $2.03\times 10^{3}$
1000 64 $1.62\times 10^{4}$ $1.62\times 10^{4}$ $1.59\times 10^{4}$
1000 256 $6.46\times 10^{4}$ $6.47\times 10^{4}$ $6.30\times 10^{4}$
1000 1024 $2.94\times 10^{5}$ $2.96\times 10^{5}$ $2.88\times 10^{5}$
10000 4 $4.61\times 10^{1}$ $4.61\times 10^{1}$ $4.63\times 10^{1}$
10000 16 $1.87\times 10^{3}$ $1.87\times 10^{3}$ $1.97\times 10^{3}$
10000 64 $1.36\times 10^{4}$ $1.36\times 10^{4}$ $1.43\times 10^{4}$
10000 256 $6.22\times 10^{4}$ $6.22\times 10^{4}$ $6.11\times 10^{4}$
10000 1024 $2.80\times 10^{5}$ $2.80\times 10^{5}$ $2.78\times 10^{5}$
Table 1: Comparison of Decision Tree Methods (max_depth=3) using the Powell function without noise. Scikit-learn’s DecisionTreeRegressor uses mean-squared Error (MSE) as its default splitting criterion, which is equivalent to minimizing the variance of the target variable within each split and serve as a baseline comparison here.
Table 1 presents a comprehensive comparison of three decision tree methods—Scikit-learn’s standard implementation, VarianceSplit and MinimaxSplit—across various sample sizes $n$ and dimensions $d$ for the Powell function. This data offers valuable insights into the performance characteristics of these methods in both dense and sparse design scenarios. In the dense design regime ($n>d$), particularly evident in the cases where the sample size (100, 1000 or 10000) exceeds the dimensionality (4, 16 or 64), we observe relatively consistent performance across all three methods. This suggests that when data is abundant relative to the input space, the choice of splitting criterion (be it the standard impurity measure used by Scikit-learn, MinimaxSplit, or VarianceSplit) has less impact on the overall performance. However, as we transition into sparse designs ($d>n$), exemplified by cases where dimensions (64, 256 or 1024) exceed the sample size (10 or 100), more pronounced differences emerge. The MinimaxSplit method often demonstrates superior performance in these high-dimensional, data-scarce scenarios. For instance, with 10 samples and 256 dimensions, the MinimaxSplit approach achieves an MSE of 8.91 × 104, outperforming both Scikit-learn (1.07 × 105) and the VarianceSplit method (9.58 × 104). This aligns with the theoretical strengths of the minimax criterion in handling sparse, high-dimensional data, where it can more robustly identify informative splits with limited samples. The VarianceSplit method generally performs comparably to Scikit-learn’s implementation, which is expected given that both likely use variants of variance reduction as their splitting criteria. However, in some sparse cases, the VarianceSplit method shows slight improvements, possibly due to specific implementation details or hyperparameter choices. As dimension increases for a fixed sample size, we observe a consistent increase in MSE across all methods, illustrating the challenges posed by the curse of dimensionality. This effect is particularly pronounced in the transition from 64 to 256 dimensions, where the MSE often increases by an order of magnitude or more. Interestingly, the performance gap between methods narrows as the sample size increases, even in high-dimensional settings. For instance, with 1000 samples and 256 dimensions, all three methods achieve comparable MSE values (around 6.4 × 104). This suggests that with sufficient data, the choice of splitting criterion becomes less critical and all methods can effectively capture the underlying structure of the Powell function.
B.3 A weighted aggregation approach
The variance approach aims to create homogeneous subsets by minimizing the average squared difference between the observed values and the mean value in each node. In contrast, the minimax criterion focuses on minimizing the maximum error within each split, potentially leading to more balanced trees. The variance criterion, similar to Scikit-learn’s MSE, seeks to minimize the overall variance in child nodes but can be implemented with different loss norms ($L^{1}$ or $L^{2}$). Cyclic minimax, a variant of minimax, alternates through features (i.e., the first and second coordinates) in a predetermined order for splitting, which can be beneficial in high-dimensional spaces when the depth is larger than the input dimension, or when feature importance is known a priori (e.g., the first coordinate is more heterogeneous).
These alternative criteria offer different trade-offs: minimax may be more robust to outliers and preserve edges better in image processing tasks, variance (especially with $L^{2}$ norm) often provides good average-case performance, and cyclic minimax can ensure a more diverse use of features. Unlike Scikit-learn’s implementation, which primarily optimizes for average error reduction, these custom criteria allow for more specialized tree structures tailored to specific problem characteristics or performance goals.
Our novel random forest (see Algorithm 1 of Appendix C) implementation diverges from conventional practices by introducing diverse splitting criteria and weighted aggregation. This approach is motivated by our previous discussion that the recognition that different splitting criteria can capture various aspects of data structure, while weighted aggregation can emphasize more accurate models in the final prediction. The implementation incorporates Variance, Minimax, and cyclic Minimax splitting methods for different weak learners within a single ensemble. The variance method optimizes for overall error reduction, minimax focuses on worst-case performance at each split, and cyclic minimax introduces a deterministic feature selection process. This diversity in the splitting criteria aims to mitigate feature importance bias present in standard random forests and enhance ensemble diversity beyond what is typically achieved through bootstrapping and random feature subset selection (Friedman, 2001).
The weighted aggregation scheme, based on the reciprocal of each tree’s training RMSE, addresses the limitation that not all trees in the ensemble are equally reliable or informative. This adaptive ensemble aggregation allows the model to adjust to the varying quality of its constituent trees, potentially leading to more robust predictions. These modifications are primarily motivated by addressing limitations of standard random forests: feature importance bias, limited ensemble diversity, and non-adaptive aggregation. By introducing varied perspectives on feature importance, adding layers of diversification, and allowing for adaptive weighting, our approach aims to create a more flexible and robust ensemble.
B.4 Heterogeneous splitting
We also develop heterogeneous splitting strategies based on our findings. The heterogeneous splitting strategy in decision trees involves using different error metrics (splitting criteria) at various depths, unlike the homogeneous strategy that uses the same metric throughout. This approach allows flexibility and adaptability, essential for handling complex datasets where different regions of the input space may benefit from different strategies. Early splits might use variance reduction for effective data division, while deeper splits could switch to minimax to minimize worst-case errors, enhancing model robustness against outliers and reducing overfitting risks; alternatively we can also alternate between VarianceSplit and MinimaxSplit to control errors (Figure 6). Additionally, strategies like cyclic MinimaxSplit leverage periodic features, potentially improving performance in time series or cyclic data. This tailored approach adapts the splitting criterion based on depth, potentially leading to more accurate and effective splits as different tree levels might require different considerations. In fact, the minimax/variance alternating strategy with $L^{2}$ norm (RMSE=0.112334) further improves the lowest RMSE compared to a homogeneous splitting criteria. Consequently, a heterogeneous strategy can lead to more accurate and robust models by optimally addressing the unique characteristics of the data at each level of the tree, making it particularly beneficial in complex datasets with variable distributions.
[…]
Appendix C Algorithms
Algorithm 1 Weighted Random Forest with Multiple Splitting Criteria
1:Training data $(\mathbf{X},Y)$, number of trees $N$, minimum samples to split, maximum depth, error method, loss norm, use weights flag
2:Trained Random Forest model
3:Initialize $N$ decision trees with specified parameters
4:for $i=1$ to $N$ do
5:     Set tree’s cyclic padding to $i$
6:     Create bootstrap sample $(X_{i},Y_{i})$ from $(X,Y)$
7:     Train tree using VarianceSplit, MinimaxSplit or CyclicMinimaxSplit
8:     if use_weights is True then
9:         Predict on $(X_{i},Y_{i})$ and calculate RMSE
10:         Set tree weight $w_{i}=1/\text{RMSE}$ (when we set equal weights this becomes the regular random forest)
11:     end if
12:end for
13:if use_weights is True then
14:     Normalize weights: $w_{i}=w_{i}/\sum_{j=1}^{N}w_{j}$
15:end if
16:function Predict($X_{\mathrm{test}}$)
17:     for $i=1$ to $N$ do
18:         Get predictions $p_{i}$ from tree $i$ on $X_{\mathrm{test}}$
19:     end for
20:     if use_weights = True then return $\sum_{i=1}^{N}w_{i}p_{i}$
21:     elsereturn $\frac{1}{N}\sum_{i=1}^{N}p_{i}$
22:     end if
23:end function
Algorithm 2 Single Tree fitting with different Criteria
1:function VarianceSplit(node)
2:     for each feature $j$ do
3:         Sort data points in node by feature $j$
4:         for each potential split point do
5:              Calculate $\mathrm{MSE}_{\mathrm{left}}\leftarrow\sum_{x\in\mathrm{left}}(y_{x}-\bar{y}_{\mathrm{left}})^{2}$
6:              Calculate $\mathrm{MSE}_{\mathrm{right}}\leftarrow\sum_{x\in\mathrm{right}}(y_{x}-\bar{y}_{\mathrm{right}})^{2}$
7:              Calculate total error $\leftarrow\mathrm{MSE}_{\mathrm{left}}+\mathrm{MSE}_{\mathrm{right}}$
8:         end for
9:         Find split point with minimum total error (with minimal number of splitting samples)
10:     end for
11:     return feature and split point with overall minimum error
12:end function
13:function MinimaxSplit(node)
14:     for each feature $j$ do
15:         Sort data points in node by feature $j$
16:         for each potential split point do
17:              Calculate $\mathrm{MSE}_{\mathrm{left}}\leftarrow\sum_{x\in\mathrm{left}}(y_{x}-\bar{y}_{\mathrm{left}})^{2}$
18:              Calculate $\mathrm{MSE}_{\mathrm{right}}\leftarrow\sum_{x\in\mathrm{right}}(y_{x}-\bar{y}_{\mathrm{right}})^{2}$
19:              Calculate max error $\leftarrow\max(\mathrm{MSE}_{\mathrm{left}},\mathrm{MSE}_{\mathrm{right}})$
20:         end for
21:         Find split point with minimum max error (with minimal number of splitting samples)
22:     end for
23:     return feature and split point with overall minimum max error
24:end function
25:function CyclicMinimaxSplit(node, depth, cyclic_padding)
[…]
27:     Sort data points in node by feature feature_index
28:     for each potential split point do
29:         Calculate $\mathrm{MSE}_{\mathrm{left}}\leftarrow\sum_{x\in\mathrm{left}}(y_{x}-\bar{y}_{\mathrm{left}})^{2}$
30:         Calculate $\mathrm{MSE}_{\mathrm{right}}\leftarrow\sum_{x\in\mathrm{right}}(y_{x}-\bar{y}_{\mathrm{right}})^{2}$
31:         Calculate max error $\leftarrow\max(\mathrm{MSE}_{\mathrm{left}},\mathrm{MSE}_{\mathrm{right}})$
32:     end for
33:     Find split point with minimum max error (with minimal number of splitting samples)
34:     return (feature_index, best split point)
35:end function

### v2

@@
-Minimax Decision Trees via Martingale Approximations
+Stabilizing the Splits through Minimax Decision Trees
 Abstract
-We develop a martingale-based approach to constructing decision trees that efficiently approximate a target variable through recursive conditioning.
-We introduce MinimaxSplit, a novel splitting criterion that minimizes the worst-case variance at each step, and analyze its cyclic variant, proving an exponential error decay rate under mild conditions.
-Our analysis builds upon partition-based martingale approximations, providing new insights into their convergence behavior.
-Unlike traditional variance-based methods, MinimaxSplit avoids end-cut preference and performs well in noisy settings.
-We derive empirical risk bounds and also explore its integration into random forests.
-Keywords: martingale approximation; convergence rates; tree-based models; non-parametric regression
+By revisiting the end-cut preference (ECP) phenomenon associated with a single CART (Breiman et al.
+(1984)), we introduce MinimaxSplit decision trees, a robust alternative to CART that selects splits by minimizing the worst-case child risk rather than the average risk.
+For regression, we minimize the maximum within-child squared error; for classification, we minimize the maximum child entropy, yielding a C4.5-compatible criterion.
+We also study a cyclic variant that deterministically cycles coordinates, leading to our main method of cyclic MinimaxSplit decision trees.
+We prove oracle inequalities that cover both regression and classification, under mild marginal non-atomicity conditions.
+The bounds control the tree’s global excess risk by local worst-case impurities and yield fast convergence rates compared to CART.
+We extend the analysis to ensembles that subsample coordinates per node.
+Empirically, (cyclic) MinimaxSplit trees and their forests improve on structured heterogeneity data such as EEG amplitude regression over fixed time horizons, seasonal air quality forecasting, and image denoising framed as non-parametric regression on spatial coordinates.
+Keywords: CART; convergence rates; decision tree models; non-parametric regression
 1 Introduction
-Let $d\geqslant 1$ and $(\mathbf{X},Y)$ be a coupling (or a joint distribution) on $\mathbb{R}^{d+1}$, where $\mathbf{X}\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}$.
-Throughout, vectors are denoted by bold symbols.
-We study the general problem of efficiently approximating $Y$ by conditioning on the value of $\mathbf{X}$ using finite partitions.
-If we adopt the $L^{2}$ norm as the minimization criterion, the problem can be formulated as
-$\min_{m\in\mathcal{M}}\mathbb{E}[(Y-m(\mathbf{X}))^{2}],$ (1)
-where $\mathcal{M}$ denotes some class of candidate functions defined on $\mathbb{R}^{d}$.
-In this paper, we focus on the case where $\mathcal{M}$ consists of piecewise constant functions.
-Generally speaking, the motivation is algorithmic: in many applications, such as decision trees (Breiman et al.
-(1984)), the goal is to provide an efficient approximation of $Y$ by a function of $\mathbf{X}$.
-For instance, if the conditioning on $\mathbf{X}$ is based on a partition of cardinality at most $K$, we have
-=$\mathcal{M}=\bigg{\{}m:\mathbb{R}^{d}\to\mathbb{R}\,\Big{|}\,m(\mathbf{x})=\sum_{A\in\pi}m_{A}{\mathbbm{1}}_{\{\mathbf{x}\in A\}},\,\pi\in\mathfrak{P}(\mathbb{R}^{d}),\,|\pi|\leqslant K,\,m_{A}\in\mathbb{R}\bigg{\}},$
-where $\mathfrak{P}(\mathbb{R}^{d})$ denotes the set of all partitions of $\mathbb{R}^{d}$.
-Clearly, the choice $m_{A}=\mathbb{E}[Y\mid\mathbf{X}\in A]$ is always optimal in (1), so it remains to optimize the partition $\pi$.
-However, finding a direct solution to (1) is generally computationally infeasible, so recursive optimizers are favorable.
-The common approach is to construct a sequence of nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ in $\mathfrak{P}(\mathbb{R}^{d})$ that generate the conditioning on $\mathbf{X}$, where the construction is recursive by sequentially optimizing a certain decision criterion, starting from $\pi_{0}=\{\mathbb{R}^{d}\}$.
-It does not hurt to consider binary partitions, as the general case follows analogously.
-We then formulate our main problem as follows.
-(P-general). Construct nested binary partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ such that $\mathbb{E}[(Y-M_{k})^{2}]$ is small for each $k$, where
-$M_{k}(\omega):=\mathbb{E}[Y\mid\mathbf{X}\in A],\quad\text{for }\mathbf{X}(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k}.$ (2)
-Clearly, if $Y$ is not a function of $\mathbf{X}$, the quantity $\mathbb{E}[(Y-M_{k})^{2}]$ has a strictly positive lower bound $\mathbb{E}[\mathrm{Var}(Y\mid\mathbf{X})]$.
-Another obstruction arises from atoms of $\mathbf{X}$, which prevents an effective split of the partitions.
-The relation (2) and its analogues will be essential to this paper.
-The main theme of this paper is providing solutions to (P-general) in various settings. Roughly speaking, our main contribution can be summarized as follows.
-(Solving P-general).
-If $\mathbb{E}[Y\mid\mathbf{X}]$ is well-behaved, then modulo the lower bound $\mathbb{E}[\mathrm{Var}(Y\mid\mathbf{X})]$ and the atoms of $\mathbf{X}$, there exists an explicit construction of $\{\pi_{k}\}_{k\geqslant 0}$, not depending on the law of $(\mathbf{X},Y)$, such that $\mathbb{E}[(Y-M_{k})^{2}]$ decays exponentially in $k$, where $M_{k}$ is defined in (2).
-An instance of (Solving P-general) is Theorem 9 below, where the construction is given by the cyclic MinimaxSplit algorithm detailed in Section 3.1.
-In the rest of the Introduction, we illustrate the motivations of (P-general) and summarize our contribution via two concrete applications: partition-based martingale approximations and regression trees.
-1.1 Partition-based martingale approximations
-Let $U$ be a real-valued atomless random variable. A partition-based martingale approximation of $U$ is a discrete-time martingale $\{M_{k}\}_{k\geqslant 0}$ such that
-$M_{k}(\omega):=\mathbb{E}[U\mid U\in A],\quad\text{for }U(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k},$ (3)
-for some nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}$, where we assume that the partitions are binary and $\pi_{0}=\{\mathbb{R}\}$.
-The martingale property follows from the tower property of conditional expectations.
-The connection to (P-general) is as follows.
-Suppose that $d=1$ (i.e., $\mathbf{X}$ is real-valued, and will be denoted by $X$ in the following), and the coupling $(X,Y)$ satisfies that $X$ is atomless and $Y$ is a strictly increasing function of $X$.
-Since in this case $\mathbb{E}[\mathrm{Var}(Y\mid X)]=0$ and $X$ is atomless, we expect that the approximation error $\mathbb{E}[(U-M_{k})^{2}]$ vanishes as $k\to\infty$.
-Moreover, each partition in $X$ bijectively maps to a partition in $Y$, and hence (2) reduces to (3).
-In other words, in the current setting, (P-general) is equivalent to the following problem:
-(P-martingale).
-Given a real atomless random variable $U$, construct a partition-based martingale approximation $\{M_{k}\}_{k\geqslant 0}$ of $U$ such that $\mathbb{E}[(U-M_{k})^{2}]\to 0$ exponentially in $k$.
+Tree-based predictors are classical non-parametric learners, from early systems such as THAID to the CART formalization (Morgan and Messenger, 1973; Breiman et al., 1984), with surveys charting both breadth and limitations (Gordon and Olshen, 1984; Loh, 2014) and recent work revisiting large-scale aspects (Klusowski and Tian, 2024).
+A persistent and practically important pathology of single-tree greedy splitting is the end-cut preference (ECP): standard impurity/variance criteria favor splits near node boundaries, especially under weak or no signal.
+ECP has long been observed empirically (Morgan and Messenger, 1973; Torgo, 2001; Su et al., 2024; Wang, 2025) and is now rigorously characterized in simple settings, where noisy directions drive splits toward extremes (Breiman et al., 1984; Cattaneo et al., 2022).
+For a single deep tree, ECP can be harmful: aggressive boundary splits create highly unbalanced children, amplify estimation noise, and can stall meaningful refinement along informative coordinates.
+However, Ishwaran (2015) argues that the same tendency can be beneficial in deep random forests: when a candidate split falls on a noise variable, an end-cut keeps one child large, preserving downstream sample size so later splits (possibly on strong variables) can recover the signal; even along strong variables, ECP-like behavior can be adaptive in regions with little local signal.
+Thus, while ECP degrades a single tree’s stability, it may aid an ensemble by lowering tree correlation and enabling recovery via aggregation.
+At the same time, the forest setting introduces its own costs: ECP interacts with candidate-set multiplicity and variable-usage bias, skewing split frequencies and importance metrics, and further reducing model interpretability.
+Existing methods to mitigate ECP include the delta splitting rule (Morgan and Messenger, 1973), regularization of node size (Torgo, 2001), and the Smooth Sigmoid Surrogate (Su et al., 2024).
+We refer to Wang (2025) for a more comprehensive review of the ECP.
+As illustrated in Figure 1, ECP arises from the greedy split search itself.
+At each node, the algorithm scans many thresholds and chooses the one with the largest impurity decay.
+Thresholds near the boundary isolate very few points; the sample mean of such a tiny child has a high variance, so even pure noise can create a large difference between child means and an inflated impurity reduction.
+Taking the maximum over many candidate cut-off points amplifies this selection bias—extremal thresholds are disproportionately likely to ‘win’.
+Our focus is on explicitly mitigating ECP at the algorithmic level for trees, while retaining the transparency that makes single decision trees interpretable.
+Beyond prediction, this design goal is interpretability: balanced, nondegenerate splits yield partitions whose geometry is easier to understand and communicate than the heterogeneous, boundary-driven partitions that arise under ECP in deep trees and ensembles.
+In this paper, we revisit ECP empirically and theoretically, clarifying the cases in which standard CART-style criteria drift toward extreme cut-off points and how this differs between single trees and forests (Breiman et al., 1984; Morgan and Messenger, 1973; Torgo, 2001; Su et al., 2024; Ishwaran, 2015; Wang, 2025; Cattaneo et al., 2022).
+Then, we introduce a Minimax splitting rule with a cyclic coordinate schedule that systematically discourages degenerate child nodes, thereby mitigating ECP while preserving interpretability.
+As a consequence, our splitting rule admits a faster convergence guarantee compared to standard CART — we establish exponential error decay under mild conditions and relate our guarantees to existing theory (Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Mazumder and Wang, 2024; Cattaneo et al., 2024).
+Conceptually, the analysis connects to partition-based martingale approximations (Simons, 1970; Zhang et al., 2024) and complements asymptotic and inferential developments for trees/forests (Klusowski and Tian, 2024; Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Mazumder and Wang, 2024; Cattaneo et al., 2024), as well as recent scalable partitioning frameworks (Liu et al., 2023; Luo and Pratola, 2023; Luo et al., 2024) and complexity-theoretic perspectives on hierarchical splits (Blanc et al., 2020; O’Donnell et al., 2005).
+Figure 1: The comparison of the scikit-learn DecisionTreeRegressor, our MinimaxSplit tree, and VarianceSplit tree on a sinusoidal target function (in red) $Y=f(X)+\varepsilon=\sin(4\pi X)+\varepsilon$, $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1),$ and $\varepsilon\buildrel\mathrm{law}\over{\sim}\mathcal{N}(0,0.1)$, with maximum depth 4 (top row) and 5 (bottom row).
+We display 10 different model fits based on 10 different batches of training sets of size 100, and the average mean squared error (MSE) evaluated on a different training set across 10 batches.
+2 Minimax decision trees for regression
+The goal of this section is to construct MinimaxSplit decision tree algorithms that avoid ECP and attain exponential convergence.
+We first focus on the regression setting and we discuss the classification setting in Section 4.
+In CART-style regression, a decision tree greedily splits in order to minimize the sum of variances over the responses of the child nodes (Liu et al., 2023; Breiman et al., 1984); instead, we propose to greedily minimize the maximum variance among the child nodes.
+In Section 2.1, we rigorously introduce the algorithms (summarized in Figure 2).
+We study exponential rates for marginally atomless joint distributions (that may not arise from empirical measures $\mathbb{P}_{N}$) in Section 2.2.
+Section 2.3 builds upon this case, covering more generally the non-marginally atomless setting and, in particular, the empirical risk bounds.
+All proofs can be found in Appendix B.1.
+Throughout, we use the notation $d]=\{1,2,\dots,d\}$ for $d\in\mathbb{N}$.
+All splits we consider in this paper will be binary.
+2.1 Algorithm formulation
+We begin by recapping the greedy splitting regime in the VarianceSplit algorithm to construct an efficient decision tree.
+We need the following notion of a splittable set.
+A similar notion is also introduced by Definition 3.1 of Klusowski and Tian (2024).
+Definition 2.1.
+Consider a joint distribution $(\mathbf{X},Y)$ on $\mathbb{R}^{d+1}$. We say a set $A\subseteq\mathbb{R}^{d}$ is $(\mathbf{X},Y)$-unsplittable, or unsplittable, if any of the following occurs:
+  • $\mathbb{P}(\mathbf{X}\in A)=0$;
+  • $\mathbb{P}(\mathbf{X}\in A)>0$ and $Y\mid\mathbf{X}\in A$ is a constant;
+  • $\mathbb{P}(\mathbf{X}\in A)>0$ and $\mathbf{X}\mid\mathbf{X}\in A$ is a constant.
+Otherwise, we say that $A$ is $(\mathbf{X},Y)$-splittable, or splittable.
+The greedy VarianceSplit construction (Section 8.4 of Breiman et al.
+(1984)) proceeds by introducing nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ with axis-aligned borders, in a way that sequentially and greedily minimizes the total risk $\mathbb{E}[(Y-M_{k})^{2}]$ at each step $k$ (or depth $k$, while splitting from the fixed partition $\pi_{k-1}$), where $M_{k}$ is taken as the conditional mean of $Y$ given the $\sigma$-algebra generated by the collection ${\mathbbm{1}}_{\{\mathbf{X}\in A\}},\,A\in\pi_{k}$.
+So, the partitions $\{\pi_{k}\}_{k\geqslant 0}$ of the input space $\mathbb{R}^{d}$ are nested and we require that each $\pi_{k}$ consists of sets that split each splittable element $A\in\pi_{k-1}$ into two hyper-rectangles with axis-aligned borders.
+Formally, suppose that a hyper-rectangle $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})$ belongs to the partition $\pi_{k-1}$ and is splittable (if $A$ is unsplittable, we do not split $A$ and $A\in\pi_{\ell}$ for every $\ell\geqslant k$).
+In particular, $M_{k-1}(\omega)=\mathbb{E}[Y\mid\mathbf{X}\in A]$ if $\mathbf{X}(\omega)\in A$.
+To find a split of $A$ at depth $k$, the VarianceSplit algorithm looks for a dimension $j\in[d]$ and covariate $x_{j}\in[a_{j},b_{j})$ such that the remaining risk is the smallest after splitting $A$ at $x_{j}$ in covariate $j$.
+In other words, one seeks for
+$\begin{split}(j,\hat{x}_{j})&=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big{(}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j})\\ &\hskip 56.9055pt+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{)}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big{(}\mathbb{E}\big{[}(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}\big{]}\\ &\hskip 56.9055pt+\mathbb{E}\big{[}(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}\big{]}\Big{)}\\ &=:\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}(V_{\text{left}}(j)+V_{\text{right}}(j)),\end{split}$ (1)
+where we break ties arbitrarily.
+[^1] Strictly speaking, the $\operatorname*{arg\,min}$ in (1) may not always be attained for general joint distributions $(\mathbf{X},Y)$, unless we assume that $\mathbf{X}$ is marginally either atomless or has a finite support, which will always be the case in our analysis (such as the atomicity condition in Theorem 6 below).
+However, in the general setting, if we allow splits in which atoms can be duplicated and assigned to both nodes (instead of only on the right node as in (1)), $\operatorname*{arg\,min}$ can be attained.
+In this paper, we will implicitly assume that $\operatorname*{arg\,min}$ is always attained unless otherwise stated.
+Consequently, the splittable set $A$ splits into the sets $A_{L}:=[a_{1},b_{1})\times\dots\times[a_{j},\hat{x}_{j})\times\dots\times[a_{d},b_{d})$ and $A_{R}:=[a_{1},b_{1})\times\dots\times[\hat{x}_{j},b_{j})\times\dots\times[a_{d},b_{d})$ to form its two descendants in the set $\pi_{k}$, and define
+$M_{k}(\omega)=\begin{cases}\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{L};\\ \mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{R}.\end{cases}$ (2)
+In other words, at depth $k\geqslant 0$ and after constructing the partition $\pi_{k}$, define $M_{k}$ by
+$M_{k}(\omega)=\mathbb{E}[Y\mid\mathbf{X}\in A],\quad\text{if }\mathbf{X}(\omega)\in A,\quad\text{for }A\in\pi_{k}.$ (3)
+This leads to an associated process $\{M_{k}\}_{k\geqslant 0}$ to the joint distribution $(\mathbf{X},Y)$.
+[^2] The process $\{M_{k}\}_{k\geqslant 0}$ is a martingale, but we will not use this fact in the rest of the main text.
+To connect $M_{k}$ with the estimator $\hat{m}_{k}$ of the regression function at step $k$, we note that for any $\mathbf{x}\in\mathbb{R}^{d}$, there exists a unique $A_{\mathbf{x}}\in\pi_{k}$ containing $\mathbf{x}$, then $\hat{m}_{k}(\mathbf{x})=\mathbb{E}[Y\mid\mathbf{X}\in A_{\mathbf{x}}]$.
+Figure 2: The comparison of the VarianceSplit, MinimaxSplit and cyclic MinimaxSplit trees.
+Here we use the notation: $A_{j,s}^{L}=A_{t}\cap\{x_{j}<s\}$$A_{j,s}^{R}=A_{t}\cap\{x_{j}\geqslant s\}$$\mathcal{S}_{j}(A_{t})$ are candidate thresholds (e.g., midpoints of sorted $x_{j}$ in $A_{t}$); $P_{t}(B)=\mathbb{P}(\mathbf{X}\in B\mid\mathbf{X}\in A_{t})$.
+The MinimaxSplit algorithm.
+We now introduce the MinimaxSplit algorithm.
+A common feature of the MinimaxSplit and VarianceSplit algorithms is that they both start from a nested sequence of partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ consisting of (at-most) binary splits into hyper-rectangles with axis-aligned borders.
+The two algorithms differ in the way the partition is constructed: in the MinimaxSplit setting, the split is chosen to minimize the maximum variance among the two descendants.
+Formally, for a splittable set $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})\in\pi_{k-1}$, define the split location as
+$\begin{split}(j,\hat{x}_{j})&=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}],\\ &\hskip 85.35826pt\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}]\Big{\}}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}),\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{\}}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}V_{\text{left}}(j),V_{\text{right}}(j)\Big{\}},\end{split}$ (4)
+where we break ties arbitrarily. We call this the MinimaxSplit rule. After obtaining the partition $\pi_{k}$, define $M_{k}$ through (3).
+The cyclic MinimaxSplit algorithm.
+The cyclic MinimaxSplit algorithm is a variation of the MinimaxSplit algorithm.
+Instead of optimizing over all covariates $j\in[d]$ in (4), we cycle through the $d$ covariates as the tree grows.
+That is, for $k\geqslant 0$, let $j=j_{k}=1+(k\leavevmode\nobreak\ (\text{mod }d))$.
+For all splittable sets $A\in\pi_{k}$, we split $A$ in the $j$-th coordinate to form its two descendants at depth $k+1$.
+The split location is then defined as
+$\begin{split}\hat{x}_{j}&=\operatorname*{arg\,min}_{x_{j}\in[a_{j},b_{j})}\max\Big{\{}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}],\\ &\hskip 85.35826pt\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}]\Big{\}}\\ &=\operatorname*{arg\,min}_{x_{j}\in[a_{j},b_{j})}\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}),\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{\}},\end{split}$ (5)
+where we break ties arbitrarily. Define $M_{k}$ by (3).
+If $d=1$, the cyclic MinimaxSplit algorithm coincides with the MinimaxSplit algorithm.
+If $d>1$, compared to the MinimaxSplit algorithm, the cyclic MinimaxSplit algorithm is technically more tractable.
+Moreover, since all features are considered equally, the cyclic MinimaxSplit algorithm appears more robust in recovering the underlying geometry of the regression function, which we illustrate numerically in Section 5.
+On the other hand, the cyclic MinimaxSplit algorithm may be less effective in higher dimensions because the relevant dimensions may not be easily identified or reached in the first few splits, yet it avoids a pathological example in Section 2.4.
+2.2 Risk bound
+The goal of this section is to justify the exponential convergence under the cyclic MinimaxSplit rule.
+Assuming marginal non-atomicity, we show that up to universal constants and model mis-specification, the approximation error decays exponentially with rate $r^{k}$ for some explicit $r\in(0,1)$ that depends only on dimension $d$.
+We note here a crucial difference from the statistics literature that our analysis of the MinimaxSplit rule first focuses on general joint distributions $(\mathbf{X},Y)$ on $\mathbb{R}^{d+1}$, where $\mathbf{X}\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}$.
+[^3] This can be intuitively regarded as empirical measures with infinitely many samples.
+Empirical measures are a special case.
+The class of joint distributions we consider will satisfy the following marginally atomless property.
+Definition 2.2.
+Let $d\geqslant 1$ and $\mathbf{X}=(X_{1},\dots,X_{d})$ be an $\mathbb{R}^{d}$-valued random vector.
+We say that $\mathbf{X}$ (or its law) is marginally atomless if for any $j\in[d]$ and $x\in\mathbb{R}$, $\mathbb{P}(X_{j}=x)=0$.
+The reason for considering general joint distributions is twofold.
+First, we provide more general probabilistic results that may be of interest in other settings.
+Second, in our derivation of the empirical risk bounds, it would be more natural and technically convenient to first study the marginally atomless setting.
+Although the empirical measure is not marginally atomless, only minor changes will be required to take care of the atomic case (to be discussed in Section 2.3 below).
+Let us start with some simple observations on the MinimaxSplit rule. The following lemma is a simple consequence of the total variance formula and will be proved in Appendix B.1.
+Lemma 2.1.
+Suppose that $\mathbf{X}$ is marginally atomless and $j\in[d]$. Then the function
+$\varphi_{L}:x_{j}\mapsto\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j})$ (6)
+is non-decreasing and continuous. Similarly,
+$\varphi_{R}:x_{j}\mapsto\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})$
+is non-increasing and continuous.
+Choose an arbitrary element $\hat{x}_{j}\in\operatorname*{arg\,min}_{x_{j}}\max\{\varphi_{L}(x_{j}),\varphi_{R}(x_{j})\}.$ Let us first assume that $\varphi_{L}(\hat{x}_{j})\leqslant\varphi_{R}(\hat{x}_{j})$.
+Define $x_{j}^{*}:=\min\{x_{j}:\varphi_{L}(x_{j})=\varphi_{R}(x_{j})\}$.
+By the monotonicity of $\varphi_{L}$ and $\varphi_{R}$, we have $\hat{x}_{j}\leqslant x_{j}^{*}$, and thus
+$\varphi_{L}(\hat{x}_{j})\leqslant\varphi_{L}(x_{j}^{*})=\varphi_{R}(x_{j}^{*})\leqslant\varphi_{R}(\hat{x}_{j}).$ (7)
+Using (7) and the fact that $\hat{x}_{j}\in\operatorname*{arg\,min}_{x_{j}}\max\{\varphi_{L}(x_{j}),\varphi_{R}(x_{j})\}$, we have
+$\varphi_{L}(x_{j}^{*})=\varphi_{R}(x_{j}^{*})=\varphi_{R}(\hat{x}_{j}).$ (8)
+Since the events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<x_{j}^{*}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant x_{j}^{*}\right\}$ form a partition of $\{\mathbf{X}\in A\}$, we have by the total variance formula,
+$\begin{split}\varphi_{L}(\hat{x}_{j})+\varphi_{R}(\hat{x}_{j})\leqslant\varphi_{L}(x_{j}^{*})+\varphi_{R}(x_{j}^{*})\leqslant\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}].\end{split}$ (9)
+Combining (8) and (9), we have
+$\begin{split}\max\Big{\{}\varphi_{L}(\hat{x}_{j}),\varphi_{R}(\hat{x}_{j})\Big{\}}=\varphi_{R}(\hat{x}_{j})\leqslant\frac{1}{2}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}].\end{split}$ (10)
+The other case $\varphi_{L}(\hat{x}_{j})\geqslant\varphi_{R}(\hat{x}_{j})$ is similar.
+In other words, the remaining risk on each descendant is at most half the risk on the parent node.
+Inductively, we gain a uniform exponential decay of the maximum risk among the nodes at depth $k$:
+$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\mathrm{Var}(Y\mid\mathbf{X}\in A)=\max_{A\in\pi_{k}}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]\leqslant 2^{-k}\mathrm{Var}(Y).$ (11)
+This is the key to controlling the “sizes of the nodes”.
+In the case of the VarianceSplit construction, such uniform control is not possible—counterexamples exist due to ECP, which we explain in the next example.
+Example 1.
+We revisit (a special case of) the location regression model analyzed in Cattaneo et al.
+(2022).
+Let $(X,Y)$ be an empirical measure of $\mathrm{Unif}(0,1)\times\mathcal{N}(0,1)$ on $0,1]\times\mathbb{R}$ with $n$ samples.
+[^4] XX is not marginally atomless, but the example can always be modified by smoothing the data without affecting its validity.
+Suppose that we perform VarianceSplit for one step.
+Theorem 4.1 of Cattaneo et al.
+(2022) implies that, for $n$ large enough, the split lies between the first and $\sqrt{n}$-th order statistics with probability exceeding $1/6$.
+With high probability (meaning 1o(1)1-o(1) as $n\to\infty$), the right node has remaining risk at least (1o(1))n(1-o(1))n while the total risk at the parent node is at most $(1+o(1))n$.
+By a union bound, we see that with an arbitrary $\varepsilon>0$, for sufficiently large $n$,
+$\mathbb{P}\Big{(}\max_{A\in\pi_{1}}\mathbb{P}(X\in A)\mathrm{Var}(Y\mid X\in A)\geqslant(1-\varepsilon)\mathrm{Var}(Y)\Big{)}>\frac{1}{7}.$
+This example also explains why assumptions such as the sufficient impurity decay (SID) condition are favorable to ensure risk decay when using the splitting criterion (1) (Chi et al., 2022; Mazumder and Wang, 2024).
+To see further intuition of ECP, we consider the event $\{\mathbf{X}\in A\}$ and sub-events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$.
+Without loss of generality, we may assume $\mathbb{P}(\mathbf{X}\in A)=1$ and apply the total variance formula to this partition, yielding
+$\mathbb{P}(\mathbf{X}\in A)\mathrm{Var}(Y)=\mathrm{Var}(Y)$ $=\mathbb{E}[\mathrm{Var}(Y\mid\sigma(A_{L}))]+\mathrm{Var}(\mathbb{E}[Y\mid\sigma(A_{L})])$
+$\begin{split}&=\mathrm{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\mathrm{Var}(Y\mid A_{R})\mathbb{P}(A_{R})\\ &\hskip 56.9055pt+\left(\mathbb{E}[Y\mid A_{L}]-\mathbb{E}[Y\mid A_{R}]\right)^{2}(1-\mathbb{P}(A_{L}))\mathbb{P}(A_{L}).\end{split}$ (12)
+From this, we see that when ECP occurs, $\mathbb{P}(A_{L})\approx 1$ or 0, the improvement is dominated by the first two terms of (12): $\mathrm{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\mathrm{Var}(Y\mid A_{R})\mathbb{P}(A_{R})=V_{\text{left}}(j)+V_{\text{right}}(j)$.
+Therefore, when ECP occurs, the sum of variances does not necessarily decay fast enough for the VarianceSplit.
+However, for (cyclic) MinimaxSplit, $\max\{V_{\text{left}}(j),V_{\text{right}}(j)\}$ always displays exponential decay (11), hence “avoiding” the ECP phenomenon.
+In other words, the fast decay of the sum of variances is not easily guaranteed from the VarianceSplit construction due to ECP; but the exponential decay of the maximum variance can be assured by the MinimaxSplit construction, as we show in Theorem 2.
+We consider the additive function class
+$\mathcal{G}:=\{g(\mathbf{x}):=g_{1}(x_{1})+\dots+g_{d}(x_{d})\},$
+where $\mathbf{x}=(x_{1},\dots,x_{d})$.
+For $g\in\mathcal{G}$, define $\left\lVert g\right\rVert_{\mathrm{TV}}$ as the infimum of $\sum_{i=1}^{d}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}$ over all such additive representations of $g$, where $\left\lVert\cdot\right\rVert_{\mathrm{TV}}$ denotes the total variation of a univariate function of bounded variation.
+Theorem 2.
+Suppose that $\mathbf{X}$ is marginally atomless.
+Let the associated $\{M_{k}\}_{k\geqslant 0}$ be constructed from the cyclic MinimaxSplit algorithm.
+Then uniformly for any $\delta>0$ and $k\geqslant 0$,
+$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant\inf_{g\in\mathcal{G}}\bigg{(}\Big{(}(1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big{)}\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]\\ &\hskip 56.9055pt+(1+\delta^{-1})\bigg{(}\frac{1}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg{)}.\end{split}$ (13)
+Remark 3 (The MinimaxSplit analogue).
+Theorem 2 does not hold in general for $d\geqslant 2$ if we replace the cyclic MinimaxSplit algorithm by the MinimaxSplit algorithm, a phenomenon we illustrate in Section 2.4 below.
+Nevertheless, the MinimaxSplit algorithm often achieves desirable risk decay in high-dimensional settings in which cyclic MinimaxSplit is inefficient; see Appendix D.2.
+The first general result for the convergence rate of the VarianceSplit algorithm without the variance decay assumption was obtained by Klusowski and Tian (2024).
+In a similar vein, Cattaneo et al.
+(2024) studied the convergence rates of the oblique CART and obtained a $1/k$ upper bound for MSE in the general setting (Theorem 1 therein) as well as an $r^{k}$ upper bound under further assumptions such as node sizes (Theorem 4 therein).
+The unconditional rate guarantee of $1/k$ appears significantly slower than the exponential decay rates under variance decay assumptions.
+(A continuous version of) Theorem 4.2 of Klusowski and Tian (2024) states that if the associated $\{M_{k}\}_{k\geqslant 0}$ is constructed from the VarianceSplit algorithm, then for each $k\geqslant 1$,
+$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\frac{\left\lVert g\right\rVert_{\mathrm{TV}}^{2}}{k+3}\Big{)}.$ (14)
+Let us compare the rates between (14) and (13).
+Consider $Y=g(\mathbf{X})$ with $g\in\mathcal{G}$ and $\varepsilon>0$.
+The following table summarizes the convergence rates of the VarianceSplit and MinimaxSplit algorithms.
+Note that Theorem 2 exhibits the curse of dimensionality, which cannot be circumvented in the current setting to the best of our knowledge (Cattaneo et al., 2022; Tan et al., 2022).
+[…]
+VarianceSplit (1) (14) $O(1/\varepsilon)$
+MinimaxSplit (4) N/A does not exist in general for $d\geqslant 2$
+Cyclic MinimaxSplit (5) Theorem 2 $O(d\log(1/\varepsilon))$
+Since the bound (13) is uniform over $\delta,k,d$, one can further optimize over $\delta$ for fixed $k,d$.
+For example, taking $\delta=2^{-\lfloor k/d\rfloor/10}$, the right-hand side of (13) then has the more explicit (but cruder) upper bound
+$\inf_{g\in\mathcal{G}}\Big{(}5\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+2^{1-\lfloor k/d\rfloor/2}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\Big{)}.$
+Remark 4.
+In the case where $\left\lVert g\right\rVert_{\mathrm{TV}}$ is much larger than the range $\Delta g:=\sup g-\inf g$, (13) can be further improved to
+$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant\inf_{g\in\mathcal{G}}\bigg{(}(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+2^{1/3}(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{2/3}\\ &\hskip 113.81102pt+2^{-2/3}(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}(\Delta g)^{4/3}\bigg{)}.\end{split}$ (15)
+Theorems 6 and 7 below enjoy the same specialized bounds.
+In the remainder of this section, we explain the ideas underlying the proof of Theorem 2, in the special case where $d=1$ and $Y=g(X)$ for some $g\in\mathcal{G}$. In this case, (13) reduces to
+$\mathbb{E}[(Y-M_{k})^{2}]\leqslant L2^{-2k/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}$ (16)
+for some $L>0$, which we verify next.
+[…]
+$\mathbb{E}[(M_{k}-M_{k+1})^{2}]=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2},$ (17)
+where for two vertices $v,w$, we write $v\sim w$ if they are connected.
+Let $\mathcal{F}_{k}$ denote the $\sigma$-algebra generated by the collection ${\mathbbm{1}}_{\{\mathbf{X}\in A\}},\,A\in\pi_{k}$.
+The nodewise predictor $M_{k}=\mathbb{E}[Y\mid\mathcal{F}_{k}]$ is a Doob martingale in $L^{2}$, and its risk satisfies
+$\mathbb{E}\bigl{[}(Y-M_{k})^{2}\bigr{]}=\sum_{i=k}^{\infty}\mathbb{E}\bigl{[}(M_{i+1}-M_{i})^{2}\bigr{]}\quad\text{ and }\quad\mathbb{E}\bigl{[}(M_{k}-M_{k+1})^{2}\bigr{]}=\sum_{A\in\pi_{k}}\sum_{\begin{subarray}{c}A^{\prime}\in\pi_{k+1}\\ A^{\prime}\subset A\end{subarray}}\mathbb{P}(A^{\prime})\bigl{(}\mu_{A^{\prime}}-\mu_{A}\bigr{)}^{2}.$ (18)
+In other words, convergence rates reduce to controlling the one-step increments created by each split.
+The VarianceSplit rule provides no uniform control: under ECP, a child can have arbitrarily small mass, so large within-node fluctuations can be confined to microscopic slivers, and the increment need not contract at a fixed rate.
+Our MinimaxSplit instead minimizes $\max\{\mathbb{P}(A_{L})\mathrm{Var}(Y\mid A_{L}),\,\mathbb{P}(A_{R})\mathrm{Var}(Y\mid A_{R})\}$, ensuring that at every refinement the worst child risk decreases by a constant factor.
+Summed along the tree, this yields an exponential contraction of $\mathbb{E}[(M_{k+1}-M_{k})^{2}]$ and hence of $\mathbb{E}[(Y-M_{k})^{2}]$.
+The cyclic schedule strengthens this guarantee by forcing each coordinate to be refined regularly, preventing all contraction from being spent on a single axis and avoiding symmetry-breaking failures; see Section 2.4.
+Figure 3: Visualizing the nested sequence of partitions $\{\pi_{k}\}_{k\geqslant 0}$ and its associated binary subtree representation.
+To put the above intuition into play, we apply Hölder’s inequality and (11) to obtain
+$\begin{split}\mathbb{E}[(M_{k}-M_{k+1})^{2}]&=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\\ &\leqslant\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}\bigg{)}^{1/3}\\ &\hskip 28.45274pt\times\bigg{(}\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}\\ &\leqslant\mathrm{Var}(Y)^{2/3}2^{-2k/3}\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}.\end{split}$ (19)
+Since each split introduces a difference term $|\ell_{v_{k}}-\ell_{v_{k+1}}|$, which is bounded by the values of $Y$ conditioned on a partition in the covariate space, we have
+$\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\leqslant\sum_{A\in\pi_{k+1}}\Big{(}\sup_{A}g-\inf_{A}g\Big{)}\leqslant\left\lVert g\right\rVert_{\mathrm{TV}}.$ (20)
+On the other hand, $\mathrm{Var}(Y)=\mathrm{Var}(g(X))\leqslant\left\lVert g\right\rVert_{\mathrm{TV}}^{2}$.
+Inserting (20) into (19) and using (18), we obtain $\mathbb{E}[(Y-M_{k})^{2}]\leqslant L2^{-2k/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}$ for some $L>0$.
+Our approach is fundamentally different from Klusowski and Tian (2024) in the following sense.
+On the one hand, the approach of Klusowski and Tian (2024) relies on aggregating one-step lower bounds on the impurity gain (or the decay of the risk).
+On the other hand, recall from (12) that the total risk decreases by the amount
+$\left(\mathbb{E}[Y\mid A_{L}]-\mathbb{E}[Y\mid A_{R}]\right)^{2}(1-\mathbb{P}(A_{L}))\mathbb{P}(A_{L})$ (21)
+after $A$ splits into $A_{L}\cup A_{R}$.
+The intuition is that although the gain (21) can be small at a certain step, it reveals information on the underlying joint distribution that guarantees that either it has already gone through significant risk decay in previous steps, or it is soon happening, or the total variation $\left\lVert g\right\rVert_{\mathrm{TV}}$ must be large.
+Let us illustrate the above intuition by a quick example.
+Example 5.
+Consider the case where $d=1$, $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$, and $Y=\sin(2\pi 2^{p}X)$.
+Figure 4 below plots the remaining risk $\mathbb{E}[(Y-M_{k})^{2}],\leavevmode\nobreak\ 1\leqslant k\leqslant K$ with power parameters $p=1,2,3,4$ and maximum tree depth $K=12$.
+Due to the symmetry of the sine function, the first few splits of the MinimaxSplit do not reduce the remaining risk, since the splitting points are points of symmetry.
+However, as soon as the splitting breaks through the symmetries, the remaining risk decays exponentially to zero.
+The VarianceSplit algorithm makes mild progress at every step, but is quickly surpassed by MinimaxSplit once the latter breaks through the symmetries.
+The reason is that due to ECP, the splitting points are near the boundary of the nodes, leading to unbalanced nodes.
+In other words, while the MinimaxSplit algorithm may not reduce the risk as steadily as VarianceSplit, the steps with little risk reduction often implicitly overcome certain hard constraints inherent to the regression problem (such as the symmetries of the sine function), paving the way for a much faster risk decay in later stages.
+Note that the occurrence of minimal risk deduction from the first few steps of MinimaxSplit does not contradict the risk bound from Theorem 2, because the upper bound already incorporates the total variation of the signal function.
+Figure 4: Comparison of the mean squared errors for the MinimaxSplit and VarianceSplit regimes, in the setting where $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$ and $Y=\sin(2\pi 2^{p}X)$ with $p=1,2,3,4$ and maximum tree depth $12$.
+2.3 Empirical risk bound for cyclic MinimaxSplit
+So far, we have discussed the cases where the joint distribution $(\mathbf{X},Y)$ is marginally atomless, which does not include the class of empirical measures.
+We now focus on the regression setting and derive finite-sample performance guarantees.
+Suppose that under the original law $\mathbb{P}_{*}$, $\mathbf{X}_{*}$ is marginally atomless and $Y_{*}\mid\mathbf{X}_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon:=\mathbb{E}[Y_{*}\mid\mathbf{X}_{*}]+\varepsilon,$ where the error $\varepsilon=Y_{*}-g_{*}(\mathbf{X}_{*})$ is sub-Gaussian, i.e., for some $\sigma>0$, $\mathbb{P}_{*}(|\varepsilon|\geqslant u)\leqslant 2\exp(-{u^{2}}/({2\sigma^{2}})),\leavevmode\nobreak\ u\geqslant 0.$ Also, we consider the empirical law of $(\mathbf{X},Y)$ of $N$ samples from $\mathbb{P}_{*}$.
+A limitation of Theorem 2 is that it applies only if the covariate $\mathbf{X}$ is marginally atomless (for (10) to hold).
+Therefore, we need the following result that deals with measures whose marginals may contain atoms.
+For a random variable $Y$, denote by $\operatorname{supp}Y$ the support of $Y$.
+Theorem 6.
+Suppose that $\mathbf{X}$ is purely atomic and for some $N>0$,
+$\max_{j\in[d]}\,\max_{u\in\mathbb{R}}\mathbb{P}(X_{j}=u)\leqslant\frac{1}{N}.$ (22)
+Let $\Delta Y:=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$, and let the associated $\{M_{k}\}_{k\geqslant 0}$ be constructed from the cyclic MinimaxSplit algorithm.
+Then uniformly for any $\delta>0$ and $k\geqslant 0$,
+$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}(\Delta Y)^{2}}{3N}+\inf_{g\in\mathcal{G}}\bigg{(}\Big{(}(1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big{)}\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]\\ &\hskip 170.71652pt+(1+\delta^{-1})\bigg{(}\frac{2}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg{)}.\end{split}$ (23)
+Our next key result is the following oracle inequality for decision trees under model mis-specification (i.e., when $g_{*}$ does not belong to $\mathcal{G}$).
+Let $g_{k}$ be the output of the cyclic MinimaxSplit construction (that is, $M_{k}=g_{k}(\mathbf{X})$).
+Note that $g_{k}$ is random under the law $\mathbb{P}_{*}$.
+We use $\left\lVert\cdot\right\rVert$ to denote the $L^{2}$ norm in $\mathbb{P}_{*}$, e.g.
+$\left\lVert g_{k}-g_{*}\right\rVert^{2}=\mathbb{E}[(g_{k}(\mathbf{X}_{*})-g_{*}(\mathbf{X}_{*}))^{2}]$, where the expectation is taken in $\mathbb{P}_{*}$.
+Theorem 7.
+Assume that $\mathbf{X}_{*}$ is marginally atomless. We have for $\delta\geqslant 2^{-2\lfloor k/d\rfloor/3}$,
+$\mathbb{E}\big{[}\left\lVert g_{k}-g_{*}\right\rVert^{2}\big{]}$ $\leqslant\frac{C2^{k}(\log N)^{2}\log(Nd)}{N}+2\inf_{g\in\mathcal{G}}\bigg{(}\Big{(}(1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big{)}\left\lVert g-g_{*}\right\rVert^{2}$
+$\hskip 156.49014pt+(1+\delta^{-1})\bigg{(}\frac{2}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg{)},$
+where $C>0$ is a constant depending only on $\left\lVert g_{*}\right\rVert_{\infty}$ and $\sigma^{2}$.
+For a fixed sample size $N$, if we assume that $g_{*}\in\mathcal{G}$ and $k=3d\log_{2}N/(3d+2)$ is a multiple of $d$, Theorem 7 has the further consequence that (under the same assumptions)
+$\mathbb{E}\big{[}\left\lVert g_{k}-g_{*}\right\rVert^{2}\big{]}\leqslant CN^{-\frac{2}{3d+2}}\big{(}\left\lVert g_{*}\right\rVert_{\mathrm{TV}}+(\log N)^{2}\log(Nd)\big{)}.$ (24)
+The proof of Theorem 7 follows essentially the same path as Theorem 4.3 for CART of Klusowski and Tian (2024) while replacing Theorem 4.2 therein by our Theorem 6.
+In addition to the performance and convergence rate guarantees, we briefly comment on the computational efficiency of MinimaxSplit compared to VarianceSplit.
+Typically, the time complexity consists of sorting and identifying the split dimension and location (Louppe, 2014, Section 5).
+The time complexity for sorting does not depend on the splitting criteria.
+In terms of solving for the split dimension and location, our MinimaxSplit algorithm (4) (or its cyclic variant (5)) is advantageous over VarianceSplit (1), because of the monotonicity of the risk as a function of the split location, thus reducing the time complexity from linear in the node size to logarithmic in the node size.
+2.4 Why cyclic MinimaxSplit outperforms MinimaxSplit
+The key reason why the above risk bounds and the oracle inequality are only derived for the cyclic MinimaxSplit algorithm is that there is an anti-symmetry-breaking preference (ASBP) for MinimaxSplit, which we briefly explain in this section.
+The gist is that if the underlying model possesses different symmetry conditions across different dimensions, the MinimaxSplit algorithm is likely to always split over the asymmetric dimensions, thus leaving the symmetric dimensions untouched.
+This leads to an undesirable convergence rate, or even inconsistency in the large-sample regime.
+We start with a simple example in the marginally atomless case.
+Example 8.
+Consider $\mathbf{X}=(X_{1},X_{2})$ uniformly distributed on $1,1]^{2}$ and $Y=X_{1}+|X_{2}|$.
+Clearly, $Y=f(\mathbf{X})$ for some $f\in\mathcal{G}$ with $\left\lVert f\right\rVert_{\mathrm{TV}}=4$.
+Recall (4) and consider a MinimaxSplit step applied to a splittable set $A=[\alpha,\beta)\times[-1,1]$ where $1\leqslant\alpha<\beta\leqslant 1$.
+Along the second dimension $x_{2}$, the split location of MinimaxSplit in (4) is always given by $x_{2}=0$, due to the symmetry of the law of $(\mathbf{X},Y)$ along $x_{2}=0$ on the set $A$.
+However, there is no risk decay given by this split, again due to the above symmetry.
+On the other hand, along the first dimension $x_{1}$, the split location of MinimaxSplit in (4) is given by $x_{1}=(\alpha+\beta)/2$ and the risk decay is strictly positive.
+It follows that the $\operatorname*{arg\,min}$ in (4) always satisfies $j=1$.
+By induction, if we apply the MinimaxSplit algorithm to the joint distribution $(\mathbf{X},Y)$, the split will always be along the first dimension $x_{1}$.
+Therefore, if the associated $\{M_{k}\}_{k\geqslant 0}$ is constructed from the MinimaxSplit algorithm, each $M_{k}$ can be written as a function of $X_{1}$ only, say $M_{k}=h_{k}(X_{1})$.
+But then by the independence of $X_{1}$ and $X_{2}$, for any $k$,
+$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&=\mathbb{E}[(X_{1}-h_{k}(X_{1})+|X_{2}|)^{2}]\\ &=\mathbb{E}[(X_{1}-h_{k}(X_{1}))^{2}]+\mathbb{E}[|X_{2}|^{2}]+2\mathbb{E}[(X_{1}-h_{k}(X_{1}))]\mathbb{E}[|X_{2}|]\\ &\geqslant\mathrm{Var}(X_{1}-h_{k}(X_{1}))+\mathrm{Var}(|X_{2}|)+(\mathbb{E}[X_{1}-h_{k}(X_{1})]+\mathbb{E}[|X_{2}|])^{2}\\ &\geqslant\mathrm{Var}(|X_{2}|)=\frac{1}{12},\end{split}$ (25)
+and thus the MSE does not even decay to zero, meaning that the counterpart of (13) cannot hold.
+Example 8 does not involve samples, but it is intuitively true that ASBP should still hold if one considers empirical measures of $(\mathbf{X},Y)$ with the goal of estimating $f$.
+This can be made rigorous using a Donsker class argument, which we discuss in Appendix C.
+Example 8 also generalizes to higher dimensions, such as using the function $Y=f(\mathbf{X})$ where $f(x_{1},\dots,x_{d})=x_{1}+\dots+x_{d-1}+|x_{d}|$.
+This ASBP phenomenon has an important consequence of the inadmissibility of classic random forests in the context of MinimaxSplit.
+We follow the setting of Section 7 of Klusowski and Tian (2024), which is a slight adaptation of the Breiman’s classic random forest (Breiman, 2001).
+A random forest consists of a collection of independent subsampled trees, where at each node, the split dimensions are optimized among a random pre-selected set $\mathcal{S}\subseteq[d]$.
+The set $\mathcal{S}$ is independently randomly generated with a fixed size $m_{\mathrm{try}}\in[d]$ for each node.
+The notation $m_{\mathrm{try}}$ will be kept throughout this paper.
+Consider a random forest under the MinimaxSplit regime, with $m_{\mathrm{try}}\geqslant 2$ and $g(\mathbf{x})=x_{1}+\dots+x_{d-1}+|x_{d}|$ on $1,1]^{d}$.
+Then each $\mathcal{S}$ must include a dimension in [d1][d-1], along which the split always reduces more risk than the $d$-th dimension.
+This implies that the $d$-th dimension is never reached and the risk is strictly bounded away from zero for any $k$, using an argument similar to (25).
+Therefore, the consistency of the random forest under the MinimaxSplit regime is not always guaranteed in the large-sample regime (even if the signal is noiseless), unless we pick $m_{\mathrm{try}}=1$.
+In the next section, we analyze the case $m_{\mathrm{try}}=1$, which we call a random-dimension random forest, and establish the corresponding consistency for the MinimaxSplit random forest.
+The empirical performance will be presented in Section 5.3.
+3 Ensemble models
+3.1 MinimaxSplit random forests
+Although a single decision tree can capture patterns in data through hierarchical partitioning, it often suffers from high variance in predictions, which makes it sensitive to small changes in the training dataset.
+To address these limitations, an ensemble approach can be utilized to combine multiple decision trees, leading to improved stability and predictive accuracy.
+In this section, we extend the MinimaxSplit and cyclic MinimaxSplit methods to an ensemble context.
+The ensemble method we consider here is the original Breiman’s random forest (Breiman et al., 1984), which aggregates multiple decision trees to improve the precision and robustness of the prediction.
+Each tree is grown to a fixed depth $k$ according to some splitting rule, and the split dimension of each node is optimized among independent selections of random subsets of $[d]$ with a fixed cardinality $m_{\mathrm{try}}$.
+Each tree in the ensemble is trained on a bootstrap sample of the original dataset and predictions are made by averaging the output of individual trees, effectively reducing the overall variance of the model (see Algorithm 1 in Appendix E).
+We call the ensemble the MinimaxSplit (resp.
+VarianceSplit) random forest, if the base learners follow the MinimaxSplit (resp.
+VarianceSplit) algorithm.
+[^6] Note that the cyclic MinimaxSplit case does not apply (unless $m_{\mathrm{try}}=1$) due to the lack of an optimization procedure over the dimensions.
+3.2 Empirical risk bound for MinimaxSplit random-dimension random forests
+Our discussions in Section 2.4 indicate that the random forest approach for the MinimaxSplit algorithms is better performed with $m_{\mathrm{try}}=1$ to guarantee consistency.
+In other words, we randomize the dimension at every node.
+As a consequence, applying the MinimaxSplit and the cyclic MinimaxSplit splitting rules yield the same ensemble in distribution.
+We call the resulting ensemble the MinimaxSplit random-dimension random forest.
+Recall the regression setting from Section 2.3 and consider a random forest consisting of trees in which split dimensions at each node are i.i.d.
+sampled from $[d]$, independently of the data, similar to ExtraTrees (Geurts et al., 2006) but the randomness lies in choosing the dimension instead of the split location.
+Fix a depth $k$ and the number of samples $N$.
+Denote by $n$ the number of trees in the forest.
+Let $\Sigma$ denote the law of the splitting dimensions and $\mathbb{E}_{n}$ denote the expectation with respect to the empirical measure on $n$ samples from $\Sigma$.
+Let $g_{k}$ be the estimator constructed at deapth $k$ from the MinimaxSplit random-dimension random forest.
+Our goal is then to bound the $L^{2}$ error $\mathbb{E}\big{[}\left\lVert\mathbb{E}_{n}[g_{k}]-g^{*}\right\rVert^{2}\big{]},$where the outer expectation is taken in terms of randomness both from the empirical samples from $(\mathbf{X}_{*},Y_{*})$ and from the empirical measure $\mathbb{E}_{n}$.
+We will provide empirical evidence of the effectiveness of such a randomised construction in Section 5.3.
+Theorem 9.
+Assume the same regression setting of Theorem 7, and the above random forest definition. We have
+$\begin{split}\mathbb{E}\big{[}\left\lVert\mathbb{E}_{n}[g_{k}]-g^{*}\right\rVert^{2}\big{]}\leqslant&\frac{C2^{k}(\log N)^{2}\log(Nd)+2^{k}(\log N)dn^{-1/3}}{N}\\ &+2\inf_{g\in\mathcal{G}}\bigg{(}2\left\lVert g-g^{*}\right\rVert^{2}+\frac{8}{3}\Big{(}\left\lVert g-g^{*}\right\rVert^{2}+\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\Big{)}\Big{(}e^{-\frac{k}{6d}}+\frac{d}{n^{1/3}}\Big{)}\bigg{)},\end{split}$ (26)
+where $C>0$ is a constant depending only on $\left\lVert g^{*}\right\rVert_{\infty}$ and $\sigma^{2}$.
+In particular, if $N/(2^{k}(\log N)^{2}\log(Nd))\to\infty$, $k/d\to\infty$, and $n/d^{3}\to\infty$, then the MinimaxSplit random-dimension random forest is consistent.
+Let us compare Theorem 9 with the empirical risk bound for a single tree given by Theorem 7.
+First, the exponent k/(6d)-k/(6d) now does not contain a floor function but has an asymptotically slower rate if $k/d\to\infty$.
+This is due to the non-cyclicity of the (random) selection of dimensions.
+Second, the formula (26) does not involve $\delta$ for simplicity, although the same argument works for more general choices of $\delta>0$.
+4 Minimax decision trees for classification
+Following the C4.5 generalization in Klusowski and Tian (2024), we also provide corresponding results when using MinimaxSplit for entropy loss.
+Throughout this section, we consider binary response $Y\in\{\pm 1\}$ with covariates $\mathbf{X}\in\mathbb{R}^{d}$ (so we consider $(\mathbf{X},Y)$ as a joint distribution), and $\eta(\mathbf{x})\coloneqq\mathbb{P}(Y=1\mid\mathbf{X}=\mathbf{x})$ denotes the regression function.
+For any measurable cell $A\subset\mathbb{R}^{d}$ we write
+$\eta_{A}\;\coloneqq\;\mathbb{P}(Y=1\mid\mathbf{X}\in A),\qquad w(A)\;\coloneqq\;\mathbb{P}(\mathbf{X}\in A),$
+and we measure impurity on $A$ with the Shannon entropy
+$H(Y\mid\mathbf{X}\in A)\;=\;h(\eta_{A})\;\coloneqq\;-\eta_{A}\log\eta_{A}-(1-\eta_{A})\log(1-\eta_{A}).$ (27)
+Axis-aligned splits refine a cell $A$ into left and right children $A_{L}=\{\mathbf{x}\in A:x_{j}<z\}$ and $A_{R}=A\setminus A_{L}$ for some $j\in[d]$ and $z\in\mathbb{R}$.
+Standard C4.5 chooses $(j,z)$ to minimize the weighted sum $w(A_{L})h(\eta_{A_{L}})+w(A_{R})h(\eta_{A_{R}})$.
+In the MinimaxSplit version, we instead choose $(j,z)$ to minimize $\max\{w(A_{L})h(\eta_{A_{L}}),\,w(A_{R})h(\eta_{A_{R}})\}$, which explicitly prevents one child from remaining both large and highly impure.
+As in the regression setting, when splitting at level $k$, the cyclic MinimaxSplit variant fixes the coordinate at level $k$ to $j_{k}=1+(k\ (\mathrm{mod}\ d))$ and optimizes only over the threshold $z$.
+Given a split rule, we obtain a nested sequence of axis-aligned partitions $\{\pi_{k}\}_{k\geqslant 0}$ by applying a single binary split per level.
+For any $\omega$, we let $A_{k}(\omega)\in\pi_{k}$ be the unique cell containing $\mathbf{X}(\omega)$.
+The node statistic we track is the log-odds of the average class-probability in that cell,
+$M_{k}(\omega)\;\coloneqq\;\log\Big{(}\frac{\eta_{A_{k}(\omega)}}{1-\eta_{A_{k}(\omega)}}\Big{)}.$ (28)
+Let $g_{k}(\mathbf{x}):=\log(\eta_{A_{k}}/(1-\eta_{A_{k}}))$ for $\mathbf{x}\in A_{k},\leavevmode\nobreak\ A_{k}\in\pi_{k}$.
+The predictor associated with depth $k$ is the plug-in classifier $2{\mathbbm{1}}_{\{g_{k}(\mathbf{x})\geqslant 0\}}-1$ (or equivalently, $2{\mathbbm{1}}_{\{\eta_{A_{k}}\geqslant{1}/{2}\}}-1$ for the unique $A_{k}\in\pi_{k}$ that contains $\mathbf{x}$) and the canonical convex surrogate is the conditional cross-entropy on the partition.
+With this preparation, the results in this section follow the same logical arc as in the regression setting.
+Theorem 10 treats the marginally atomless case and shows that, under the cyclic MinimaxSplit schedule, the log-odds process $\{M_{k}\}$ contracts in mean square at an exponential rate that depends only on the dimension and the coordinate range.
+Theorem 11 covers the complementary purely atomic setting.
+Here, we need a mild anti-concentration assumption on coordinate masses; the theorem then delivers an explicit upper bound for $\mathbb{E}[(Y-M_{k})^{2}]$ that reflects the discrete resolution $N$ of the atoms.
+Finally, Theorem 12 translates these bounds into an oracle inequality for the excess classification risk of $g_{k}$ under the 0-1 loss via standard calibration between cross-entropy and misclassification.
+Theorem 10.
+Suppose that $\mathbf{X}$ is marginally atomless.
+Let the associated process $\{M_{k}\}_{k\geqslant 0}$ be constructed from (28) using the cyclic MinimaxSplit algorithm.
+Then uniformly for any $k\geqslant 0$,
+$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant\inf_{g\in\mathcal{G}}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}\Big{)}.\end{split}$ (29)
+Theorem 11.
+Suppose that $\mathbf{X}$ is purely atomic and for some $N>0$,
+$\max_{j\in[d]}\,\max_{u\in\mathbb{R}}\mathbb{P}(X_{j}=u)\leqslant\frac{1}{N}.$ (30)
+Let the associated process $\{M_{k}\}_{k\geqslant 0}$ be constructed from (28) using the cyclic MinimaxSplit algorithm. Then uniformly for any $k\geqslant 0$,
+$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant 2^{-\lfloor k/d\rfloor/2}\frac{2^{k}\log N}{N}+\inf_{g\in\mathcal{G}}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}(\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}+\frac{\left\lVert g\right\rVert_{\mathrm{TV}}}{2})\Big{)}.\end{split}$ (31)
+The proofs of Theorems 10 and 11 follow essentially the same principles as in the regression setting, except that now $\{M_{k}\}_{k\geqslant 0}$ no longer forms a martingale.
+The key is to apply the monotonicity of the map $z\mapsto\mathbb{P}(\mathbf{X}\in A,\,X_{j}<z)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A,\,X_{j}<z))$, a fact that mirrors Lemma 2.1 and will be justified in Lemma B.3 below.
+In the next result, we state the corresponding oracle inequality.
+Following Klusowski and Tian (2024), for a map $g$ we consider the misclassification risk $\mathrm{Err}(g):=\mathbb{P}(Y\neq 2{\mathbbm{1}}_{\{g(\mathbf{X})\geqslant 0\}}-1).$ Further, we set $g_{*}(\mathbf{x}):=\log(\mathbb{P}(Y=1\mid\mathbf{X}=\mathbf{x})/\mathbb{P}(Y=-1\mid\mathbf{X}=\mathbf{x}))$.
+Theorem 12.
+Assume that $\mathbf{X}_{*}$ is marginally atomless. We have that under the cyclic MinimaxSplit construction,
+$\begin{split}\mathbb{E}[\mathrm{Err}(g_{k})]-\mathrm{Err}(g_{*})\leqslant L\Big{(}&\frac{2^{k}(\log N)^{2}\log(Nd)}{N}\Big{)}^{1/4}\\ &+\inf_{g\in\mathcal{G}}\Big{(}\left\lVert g-g_{*}\right\rVert^{2}+\Big{(}{\frac{\log 2}{4}}\Big{)}^{1/4}\,2^{1-\lfloor k/d\rfloor/4}(\left\lVert g\right\rVert_{\mathrm{TV}}^{1/4}+\frac{\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}}{\sqrt{2}})\Big{)},\end{split}$ (32)
+where $L>0$ is a positive universal constant.
+Remark 13.
+The bound (32) can be relaxed to
+$\mathbb{E}[\mathrm{Err}(g_{k})]-\mathrm{Err}(g_{*})\leqslant L\Big{(}\frac{2^{k}(\log N)^{2}\log(Nd)}{N}\Big{)}^{1/4}+\inf_{g\in\mathcal{G}}\Big{(}\left\lVert g-g_{*}\right\rVert^{2}+L2^{-\lfloor k/d\rfloor/4}(\left\lVert g\right\rVert_{\mathrm{TV}}+1)\Big{)}$
+for some large constant $L>0$, using the rough inequality $x\leqslant\min\{x^{2}+1,x^{4}+1\}$.
+Compared to the oracle inequality of Klusowski and Tian (2024), we see that we have achieved exponential convergence rates and removed the factor $\left\lVert g\right\rVert_{\infty}$.
+The proof of Theorem 12 follows in exactly the same way as Theorem 4.3 of Klusowski and Tian (2024) in the C4.5 case, as the arguments therein do not depend on how the tree is constructed, as long as the splitting rule is axis-aligned.
+5 Numerical experiments and applications
+In this section, we provide experiments that verify the advantages of MinimaxSplit and cyclic MinimaxSplit methods in low-dimensional domains, as well as integrated ensemble methods applied to approximating a real-valued function and denoising images.
+Performance metrics are defined in Appendix D.1, and further numerical experiments (including those in high dimensions) are provided in Appendix D.
+5.1 Avoiding end-cut preference for stable cuts
+To show that the ECP is remedied in MinimaxSplit, we consider a special case of the location regression model in (Cattaneo et al., 2022), similar to the discussions in Example 1.
+In our experiment, we generate independent replicates with $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$ and $Y$ following the standard normal or Student’s $t(\cdot)$ distributions (pure noise), sort by $X$, and scan all possible splits in the experiments in Figure 5.
+We also provide two baselines: Uniform random cutting picks a split point entirely at random within the current predictor range, ensuring that there is no systematic bias yet, ignoring any data-driven signal.
+Sampling an observed value similarly randomizes split location among actual samples.
+Both serve as baseline benchmarks: they eliminate ECP, but do not offer an impurity-guided improvement mechanism.
+For each replicate, we record ${\min(n_{L},n_{R})}/{n}$, the relative size of the smaller child, and we choose either VarianceSplit (the standard CART impurity reduction) or MinimaxSplit.
+As the tail of the noise becomes heavier, VarianceSplit becomes increasingly prone to end-cuts, isolating tiny child nodes almost every time under Student’s t with one degree of freedom.
+In contrast, the output of MinimaxSplit remains more tightly concentrated near perfectly balanced halves across all three distributions, especially for the light-tailed noise cases, thus preventing small-node failures.
+Random uniform and random observed–point cuts likewise avoid pathological end-cuts but do so at the cost of entirely unguided splits.
+Thus, if one requires a data‐adaptive binary split that resists extreme outliers without surrendering balance, the MinimaxSplit criterion clearly outperforms VarianceSplit (or CART) and random baselines.
+Figure 5: Empirical distributions of sample size $n=500$ over $1000$ replicates of the smaller-child proportion $\min(n_{L},n_{R})/n$ when splitting pure-noise data generated by standard normal (left), $t(3)$ (middle), and $t(1)$ (right) distributions.
+We also use dashed lines to denote the threshold 0.05.
+Figure 6: Average leaf size (mean number of samples per leaf) versus maximum tree depth for four splitting criteria (VarianceSplit (CART), MinimaxSplit, C1, C2), at two noise levels.
+Error bars show one standard deviation across repeated fits.
+The vertical axis is on a log scale to expose imbalance: small and collapsing leaf means with low variability (as for MinimaxSplit and, to a lesser extent, CART) indicate balanced partitioning, whereas the slower decay and large spread for C1/C2 reflect persistent ECP that peels off tiny fragments and leaves dominant branches.
+The signal function is given by (33).
+Increasing the centered Gaussian noise from $\sigma=0.01$ to $\sigma=0.1$ slightly attenuates rates but does not erase the structural contrast.
+End-cut preference propagates imbalance down the tree: a tiny “peeled-off” child at one split creates a cascade of uneven leaves.
+To quantify this, in the next experiment, for each fully grown tree we compute
+$\text{average leaf node size}=\frac{1}{L}\sum_{j=1}^{L}n_{j},$
+[…]
+where $L$ is the number of leaves and $n_{j}$ is the size of leaf $j$.
+Small average leaf node sizes and collapsed error bars signal balanced partitions; large means with huge variance indicate persistent end-cuts leaving one dominant branch and tiny fragments.
+Our main methodological contribution, the MinimaxSplit algorithm can be used with ensembles and is applicable in real data, but also yields a theoretically stronger version of Buja and Lee (2001)’s formula C1 and C2, where they showed that the minimum variance (a.k.a., one-sided purity) approach can identify extreme subsets much better than traditional CART.
+Figure 6 records the leaf distributions as the maximum depth of the tree grows, with $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$ and signal
+$f(x)=\begin{cases}\sin(x),&0\leqslant x<\frac{1}{3};\\ -2x,&\frac{1}{3}\leqslant x<\frac{2}{3};\\ 0,&\text{otherwise}.\end{cases}$ (33)
+MinimaxSplit quickly reduces average leaf size and variability, yielding balanced leaves.
+VarianceSplit (CART), C1, and C2 shrink mean leaf size slowly and retain a large spread: they repeatedly cut tiny end pieces while leaving a large branch, showing ECP.
+5.2 Fixed-horizon EEG amplitude regression
+We now examine whether the theoretical advantages of MinimaxSplit and cyclic MinimaxSplit carry over to a non-parametric time-series regression problem built from the Bonn electroencephalography (EEG) corpus (Andrzejak et al., 2001).777as explained at https://www.mathworks.com/help/wavelet/ug/time-frequency-convolutional-network-for-eeg-data-classification.html Each EEG record is a single-channel segment of length $T\approx 23.6$ seconds sampled at $173.61$ Hz.
+We formulate a scalar regression task: given time $t\in[0,T]$ within a fixed horizon $T$, predict the standardized EEG amplitude $y(t)$.
+This time-only design isolates the splitting behavior along a single coordinate, precisely the regime in which ECP is the most consequential for greedy impurity splits and where our minimax approach is expected to stabilize partitions.
+Our preprocessing follows two principles.
+First, we standardize the amplitudes within each segment to zero mean and unit variance.
+This removes irrelevant scale without obscuring the short-range temporal variability that the tree must fit.
+Second, we optionally downsample by an integer factor to control computational budget while preserving the temporal ordering.
+Both steps are label-free and therefore introduce no leakage.
+For evaluation, we fix a horizon T=20T=20\,s, draw train-test splits inside $[0,T]$ by a random holdout, and keep the split fixed across methods and depths so that any performance difference reflects only the splitting rule.
+These choices echo our theoretical emphasis on how the split criterion shapes the geometry of the partition.
+[…]
+In Figure 7, we compare three trees with identical hyperparameters (maximum split size $(=2)$ and stopping rules kept fixed): standard CART using squared-error reduction (with pruning), its variant VarianceSplit, and our MinimaxSplit which minimizes the worst child variance.
+The experimental protocol is intentionally minimal so that the fitted functions reveal the partition geometry.
+Input features consist only of time $t$ and targets are the standardized amplitudes $y(t)$.
+We train each tree on the training subset within $[0,T]$ and report test RMSE, MAE, and the coefficient of determination $R^{2}$ on the holdout portion, aggregating results across all available EEG segments.
+Because the same split is reused across methods, depth-wise comparisons are matched at the sample level.
+We also visualize the predicted curves against the ground truth on a representative segment to expose the qualitative effect of each split rule.
+The results align with the theoretical expectations.
+As the tree depth increases, CART and the sum-of-variances baseline increasingly display end-cut behavior, producing piecewise constant fits that flatten toward the edges of the interval and degrade extrapolation between local extrema.
+MinimaxSplit, by contrast, maintains near-balanced children throughout the path, which yields visibly tighter fits to the oscillatory structure of the signal and systematically lower errors.
+In addition, MinimaxSplit reduces error and raises $R^{2}$ markedly as depth increases, while CART and the VarianceSplit baseline improve more slowly.
+The trend is most pronounced in the random holdout regime where interpolation dominates; here, the additional leaves available to deeper trees are useful only if the split rule does not collapse sample sizes, and the MinimaxSplit criterion preserves effective sample size by construction.
+This observation matches our finding that MinimaxSplit concentrates the smaller-child proportion away from zero and one across noise laws, thereby eliminating pathological small-node failures observed under impurity-based CART.
+5.3 Image denoising with ensembles
+Figure 8: Comparative analysis of various random forest (depth $K=10$, 50 estimators) approaches for image denoising using different splitting strategies.
+The experiment compares six random forest configurations on a noisy astronaut image: original noiseless image; noisy input; three greedily splitting methods ($m_{\mathrm{try}}=d$) using consistent splitting criteria across all weak learners (trees) — VarianceSplit, MinimaxSplit, and Cyclic MinimaxSplit; three random-dimension methods where feature selection is randomized at each node (i.e., $m_{\mathrm{try}}=1$) while location selection follows the three different criteria, as described in Section 3.2.
+Figure 9: Comparison of decision tree regression methods (depth $K=10$) for image reconstruction using different splitting criteria — VarianceSplit, MinimaxSplit, and cyclic MinimaxSplit.
+Each method was trained on noisy astronaut image data with spatial coordinates as features and pixel intensities as targets.
+Another notable experiment involves applying decision trees to denoise noisy images (Luo et al., 2024).
+We apply the different tree variants to predict pixel values based on their locations, effectively treating this as a regression problem.
+The experiments shown in Figures 8 and 9 compare decision tree and random forest regression methods for denoising the classic Astronaut (Van der Walt et al., 2014) grayscale image, which has been preprocessed to 256 by 256.
+The goal is to recover a clean grayscale image $f:[0,1]^{2}\to[0,1]$ from noisy observations $Y=f(\mathbf{X})+\varepsilon$ at pixel centers $\mathbf{X}=(X_{1},X_{2})$ laid out on a $256\times 256$ grid (equipped with the uniform distribution).
+We normalize intensities to $[0,1]$ and add i.i.d.
+Gaussian noise $\varepsilon\buildrel\mathrm{law}\over{\sim}\mathcal{N}(0,\,0.1^{2})$.
+Each pixel becomes one regression sample, and we fit axis-aligned trees that approximate the unknown $f$ as a piecewise constant function on a dyadic partition, which is precisely the regime covered by our analysis of axis-aligned trees and their ensembles.
+We compare six random forest variants built from our splitting rules: VarianceSplit, MinimaxSplit, and cyclic MinimaxSplit, each with or without random feature subsampling at every node ($m_{\mathrm{try}}=1$ feature per split).
+All forests consist of $50$ trees, each grown to depth $K=10$; trees are bootstrapped and predictions are averaged.
+Figure 8 shows qualitative reconstructions for the six forests.
+The VarianceSplit forest leaves noticeable grain in flat regions and blurs edges at depth $10$.
+In contrast, both Minimax ensembles suppress speckle while preserving high-contrast features such as facial contours and texture edges.
+Further, across all metrics, the MinimaxSplit forests outperforms the VarianceSplit forest, with the largest gains in SSIM, indicating better structural fidelity.
+The improvements are coherent with Theorem 9: the worst-child control given by the MinimaxSplit rule prevents high-variance children from persisting across levels.
+The single-tree performance is shown in Figure 9.
+The experiments support two messages from the theory.
+First, enforcing a worst-child improvement at each split translates into visibly sharper reconstructions and uniformly better pixelwise metrics.
+Second, randomizing the split dimension ($m_{\mathrm{try}}=1$) is not an arbitrary choice and ensures the consistency of Minimax forests and practical stability.
+6 Conclusion
+In this study, we reexamine the well-known phenomenon of ECP, which is associated with traditional CART decision trees Breiman et al.
+(1984).
+To remedy this undesirable behavior for single decision trees Ishwaran and Kogalur (2010), we introduce a novel decision tree splitting strategy, the MinimaxSplit algorithm, along with its multivariate variant, the cyclic MinimaxSplit algorithm.
+Unlike traditional VarianceSplit methods, which aim to reduce overall variance in the child nodes, our MinimaxSplit algorithm seeks to minimize the maximum variance within the split partitions, thereby reducing the risk of overly biased partitions.
+The cyclic MinimaxSplit algorithm further ensures that each dimension is used in a balanced manner throughout the tree construction process, avoiding dominance by a subset of features.
+We do not claim that this new method could outperform the classical CART-like splitting under all circumstances, yet we point out that our construction avoids ECP and attains a better convergence rate guarantee as the tree progresses deeper.
+The (non-cyclic) MinimaxSplit construction features the anti-symmetry-breaking preference—consistency is not guaranteed in certain symmetric cases, which can, however, be alleviated by using ensemble methods.
+A key theoretical methodological development of this study is that the cyclic MinimaxSplit algorithm achieves an exponential MSE decay rate given sufficient data samples, without requiring additional variance decay assumptions (Theorem 6) in regression and classification settings.
+Furthermore, we derive empirical risk bounds for this method, establishing its robustness in different settings (Theorem 7).
+Using the same powerful techniques used for trees, we prove novel results on the convergence rates of univariate partition-based martingale approximations as a by-product (see Appendix A), which are of their own interest.
+In addition to single-tree regression analysis, we explore ensemble learning approaches, where we combine decision trees constructed using MinimaxSplit, cyclic MinimaxSplit, and VarianceSplit methods within a random forest framework.
+The results demonstrate that a hybrid approach leveraging different splitting techniques yields superior performance, particularly when dealing with non-uniform data distributions (Section 3).
+These findings suggest that the MinimaxSplit approach can provide more stable and adaptive decision trees compared to traditional methods, especially in low-dimensional settings.
+Acknowledgment
+The authors thank Ruodu Wang for his valuable input. We provide our reproducible code at github.com/hrluo.
+Appendix A On partition-based martingale approximations
+In this section, we develop the framework of partition-based martingale approximations and show that a number of constructions feature exponential convergence rates.
+These problems are of their own interest (see the end of Appendix A.1) and have implications in the regression tree setting.
+The main takeaway is that:
+  • The techniques developed in this work for analyzing convergence rates of the MinimaxSplit algorithm extend to many other settings, such as convergence rates for martingale approximations;
+  • Precise (asymptotic) convergence rates of martingale approximations depend on the initial law, but a uniform exponential rate is always guaranteed under mild conditions;
+  • Applying VarianceSplit to samples $\{(\mathbf{X}_{i},Y_{i})\}_{1\leqslant i\leqslant n}$ satisfying monotonicity in all $d$ dimensions (meaning that for every $j\in[d]$ and distinct $i,i^{\prime}\in[n]$, $((\mathbf{X}_{i})_{j}-(\mathbf{X}_{i^{\prime}})_{j})(Y_{i}-Y_{i^{\prime}})=\sigma_{j}$ where $\sigma_{j}\in\{\pm 1\}$ depends only on $j$) does not incur ECP.
+A.1 Basic concepts
+Consider nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}$ and a real-valued random variable $U$ with a finite second moment.
+We use the abbreviation $\sigma(\pi_{k})$ to denote the $\sigma$-algebra generated by events of the form $\{U\in A\}_{A\in\pi_{k}}$.
+Define $M_{k}:=\mathbb{E}[U\mid\sigma(\pi_{k})]$.
+If $\Pi:=\{\pi_{k}\}_{k\geqslant 0}$ is a nested sequence of partitions, $\{\sigma(\pi_{k})\}_{k\geqslant 0}$ is a filtration generated by indicator functions of sets in $\pi_{k}$’s (Doob, 1953).
+It follows from the tower property of conditional expectations that the sequence $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ is a martingale (in fact, a Doob martingale (Doob, 1940, 1953)).
+We call this martingale the $\Pi$-based martingale approximation of the random variable $U$, or in general a partition-based martingale approximation that approximates the random variable $U$ in $L^{2}$.
+Unless $U$ is a constant, there are different partition-based martingale approximations depending on the distribution of $U$.
+Our goal in this section is to identify a few partition-based martingale approximations that efficiently approximate $U$, where the construction algorithm is universal.
+The efficiency criterion is given by the decay of the MSE $\mathbb{E}[(U-M_{k})^{2}]$.
+The MSE is non-increasing in $k$ by the nested property of $\{\pi_{k}\}_{k\geqslant 0}$ and the total variance formula.
+Without loss of generality, we give our construction of a partition of a generic interval $a,b)\subset\mathbb{R}$, where $a,b\in\mathbb{R}\cup\{\pm\infty\}$.
+[^8] Here we slightly abuse notation that $a,b)=(-\infty,b)$ if $a=$-$\infty$.
+The sequence of partitions $\Pi$ will be constructed recursively, where $\pi_{0}=\{\mathbb{R}\}$ and for every $k\geqslant 0$, each interval $A\in\pi_{k}$ splits into two intervals, by following the same construction, forming the elements in $\pi_{k+1}$.
+In the following, we introduce four distinct splitting rules that define partition-based martingale approximations.
+For simplicity, we assume that $U$ is atomless, so that the endpoints of the intervals do not matter and that no trivial split occurs.
+The general case with atoms can be analyzed in a similar way but with more technicality.
+Definition 14.
+Suppose we are given an atomless law of $U$ and a non-empty interval $I=[a,b)$, where $a,b\in\mathbb{R}\cup\{\pm\infty\}$.
+  1. (i)
+    Define
+    $u_{\text{var}}=\operatorname*{arg\,min}_{u\in I}\big{(}\mathbb{P}(U\in[a,u))\mathrm{Var}(U\mid U\in[a,u))+\mathbb{P}(U\in[u,b))\mathrm{Var}(U\mid U\in[u,b))\big{)}.$ (34)
+    If the minimizer is not unique, we pick the largest minimizer.
+The variance splitting rule (corresponding to the VarianceSplit algorithm in Section 2.1) splits $I$ into the two sets $a,u_{\text{var}})$ and $u_{\text{var}},b)$.
+  2. (ii)
+    Define $u_{\text{Simons}}=\mathbb{E}[U\mid U\in I]$. The Simons splitting rule splits $I$ into the two sets $a,u_{\text{Simons}})$ and $u_{\text{Simons}},b)$.
+  3. (iii)
+    Define
+    $u_{\text{minimax}}=\operatorname*{arg\,min}_{u\in I}\max\big{\{}\mathbb{P}(U\in[a,u))\mathrm{Var}(U\mid U\in[a,u)),\,\mathbb{P}(U\in[u,b))\mathrm{Var}(U\mid U\in[u,b))\big{\}}.$
+    If the minimizer is not unique, we pick the largest minimizer.
+The minimax splitting rule (corresponding to the MinimaxSplit algorithm in Section 2.1) splits $I$ into the two sets $a,u_{\text{minimax}})$ and $u_{\text{minimax}},b)$.
+  4. (iv)
+    Define
+    $u_{\text{median}}=\sup\{u\in I:\mathbb{P}(a\leqslant U<u)=\mathbb{P}(u\leqslant U<b)\}.$
+    The median splitting rule splits $I$ into two sets $a,u_{\text{median}})$ and $u_{\text{median}},b)$.
+In turn, when applying recursively each splitting rule starting from $\pi_{0}=\{\mathbb{R}\}$, we obtain a nested sequence of partitions $\Pi=\{\pi_{k}\}_{k\geqslant 0}$ depending on the law of $U$.
+We call the resulting $\Pi$-based martingale approximation $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ the variance (resp.
+Simons, minimax, median) martingale (with respect to $U$).
+See Figure 10 for examples of the four martingale approximations and their convergence rates.
+Figure 10: Plots of the log-MSE, $\log_{2}(\mathbb{E}[(U-M_{k})^{2}])$ versus the approximation depth $k$ for four different methods, where the density of the law $U$ is given: (a) by $f(u)\propto{\mathbbm{1}}_{[0,0.9]}(u)+(1+10^{4}(u-0.9)){\mathbbm{1}}_{[0.9,1]}(u)$ and (b) by $f(u)\propto u^{10}{\mathbbm{1}}_{[0,1]}(u)$.
+The right panel of (b) shows the ratios $\mathbb{E}[(U-M_{k+1})^{2}]/\mathbb{E}[(U-M_{k})^{2}]$ for the four methods as a function of $k$.
+Example 15.
+If $U$ is uniformly distributed on a compact interval, all four martingales coincide.
+For example, if $U\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$, each of the four martingales from Definition 14 satisfy $M_{k}=\mathbb{E}[U\mid\sigma(\pi_{k})]$, where $\pi_{k}:=\{[j/2^{k},(j+1)/2^{k}):0\leqslant j<2^{k}\}$ and it follows that $\mathbb{E}[(U-M_{k})^{2}]=4^{-k}/12$.
+The intuition for the minimax and median martingales is that at each splitting step, we balance the “sizes” of $U{\mathbbm{1}}_{\{U\in I\}}$ on the two sets.
+The “size” corresponds to the (unconditional) variance for the minimax martingale and the total probability for the median martingale.
+Intending to minimize the MSE $\mathbb{E}[(U-M_{k})^{2}]$ at each step $k$, the variance martingale naturally arises as an algorithm that greedily reduces the remaining risk within $U$ in each iteration through layers.
 Motivated by a martingale embedding problem, Simons (1970) first introduced the Simons martingale and established the a.s.
 convergence $M_{k}\to U$ (and hence also in $L^{2}$), but did not analyze the convergence rate.
-Another recent motivation for studying the convergence rate of the Simons martingale arises from the construction of powerful e-values in hypothesis testing (Ramdas and Wang, 2024).
+A recent motivation for studying the convergence rate of the Simons martingale arises from the construction of powerful e-values in hypothesis testing (Ramdas and Wang, 2025).
 Zhang et al.
@@
 Similar techniques are also used in the study of stationary ergodic sequences (Wu and Woodroofe, 2004; Zhao and Woodroofe, 2008) and Markovian walks (Grama et al., 2018).
-Note that in our setting, $Y$ is a sum of martingale differences, but whose variance is bounded, and hence one cannot apply the martingale CLT and its convergence rates.
-1.2 Regression trees
-Consider a regression problem where we have a data set $(\mathbf{X}_{i},Y_{i})_{1\leqslant i\leqslant N}$, where $\mathbf{X}_{i}\in\mathbb{R}^{d}$ represents the covariates (or features) and $Y_{i}\in\mathbb{R}$ represents the responses.
-We follow van der Vaart and Wellner (2013) and assume the regression model:
-$Y_{i}=g_{*}(\mathbf{X}_{i})+\varepsilon_{i},$ (4)
-where $g_{*}:\mathbb{R}^{d}\to\mathbb{R}$ is the true signal function, $\{\mathbf{X}_{i}\}_{1\leqslant i\leqslant N}$ are i.i.d.
-sampled from a certain law, and $\varepsilon_{i}$’s are i.i.d.
-random (Gaussian) errors with zero mean.
-In other words, the data set is sampled from some coupling $(\mathbf{X}_{*},Y_{*})$ where $Y_{*}\mid\mathbf{X}_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon$ with $\mathbf{X}_{*}$ and $\varepsilon$ independent.
-The function $g_{*}$ can be estimated by minimizing the empirical $L^{2}$ risk (i.e., sum of squares; also called $L^{2}$ loss or mean-squared error (MSE) in this paper):
-$\hat{m}_{N}=\operatorname*{arg\,min}_{m\in\mathcal{M}}\frac{1}{N}\sum_{i=1}^{N}\left(Y_{i}-m(\mathbf{X}_{i})\right)^{2},$ (5)
-where $\mathcal{M}$ denotes the class of candidate functions for the mean $m$.
-The quantity (5) serves as an approximation to $\mathbb{E}[\left(Y_{*}-m(\mathbf{X}_{*})\right)^{2}]$.
-We usually establish the consistency of the estimator $\hat{m}_{N}$ by showing the convergence in probability:
-[…]
-Here, the underlying probability measure is the true data generating mechanism $\mathbb{P}$, the joint law of $(\mathbf{X}_{*},Y_{*})$.
-A decision tree regression model for (4) is a non-parametric model for constructing $\hat{m}_{N}$ by partitioning the input space into distinct regions $A_{1},A_{2},\dots$ and fitting a simple model $\hat{m}_{N,i}$ to each region $A_{i}$ (Breiman et al., 1984).
-In this case, $\hat{m}_{N}$ takes the form of
-[…]
-and the $\mathcal{M}$ in (5) is the class of all piecewise constant functions.
-The construction procedures of $A_{i}$’ s and $\hat{m}_{N,i}$’ s are data dependent, where the regions are associated to a tree grown sequentially by maximizing the chosen decision criteria, and hence the parameters associated with each tree node are chosen recursively.
-The recursive nature of the decision tree also indicates the connection to (P-general).
-Let $\mathbb{P}_{N}$ denote the empirical measure of the training data and $(\mathbf{X},Y)\buildrel\mathrm{law}\over{\sim}\mathbb{P}_{N}$.
-A (simplified variant of the) decision tree algorithm produces nested binary partitions $\{\pi_{k}\}_{k\geqslant 0}$ with axis-aligned borders of the input space $\mathbb{R}^{d}$ based on the law of $(\mathbf{X},Y)$, where $\pi_{0}=\{\mathbb{R}^{d}\}$ and the index $k$ is the depth of the tree (Breiman et al., 1984).
-The prediction $M_{k}$ given by a tree of depth $k$ is defined by (2):
-$M_{k}(\omega)=M_{k(N)}(\omega):=\mathbb{E}[Y\mid\mathbf{X}\in A],\quad\text{for }\mathbf{X}(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k}.$
-Note that this notation $M_{k}$ can be interpreted as substituting a random input $\mathbf{X}$ into (7).
-The estimator of $g_{*}$ is then given by $\hat{m}_{k}(\mathbf{x})=\mathbb{E}[Y\mid\mathbf{X}\in A]$ for $\mathbf{x}\in A,\leavevmode\nobreak\ A\in\pi_{k}$.
-To facilitate our discussion, we assume that the number of samples in any $A_{i}\in\pi_{k}$ is nonzero and tends to $\infty$ as $k\rightarrow\infty$ uniformly, and moreover, $N/2^{k}$ is exponentially growing in $k$ [^1] Gordon and Olshen (1984) (Theorem 3.6) imposed a similar set of stronger assumptions that require specific growth rates.
-Such an assumption will not be required in any of the main results in this paper, but is convenient for the explanation here.
-In this case, we are interested in obtaining the fast convergence rates of (6), a problem formulated as follows.
-(P-regression).
-Under the model (7) and the above assumption, construct a decision tree algorithm such that if the true signal function $g_{*}$ belongs to a certain wide function class, $\mathbb{E}[\left\lVert\hat{m}_{k}-g_{*}\right\rVert^{2}]$ decays exponentially in $k$.
-The classic CART algorithm (Breiman et al.
-(1984)) for decision trees involves the following widely adopted greedy algorithm that recursively constructs partitions $\{\pi_{k}\}_{k\geqslant 0}$.
-To partition a set $A\in\pi_{k}$, the greedy algorithm selects a dimension and a threshold where the set $A$ is split into two subsets such that the total remaining risk is minimized.
-See (11) below for a formal definition.
-A simplified variant of CART without pruning, called VarianceSplit, is described in Section 3.1.
-In the theoretical analysis of the convergence rate of CART in the regression setting, researchers often assume a lower bound on the variance decay when the algorithm splits each node in the decision tree (Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Mazumder and Wang, 2024; Cattaneo et al., 2024).
-For example, a prevalent variance decay assumption is known as the sufficient impurity decrease (SID) condition introduced by Chi et al.
-(2022), which requires that each split in the CART algorithm decreases the variance of the parent node by at least a constant factor, leading to exponential decay rates of the MSE of the form $r^{k}$ for some $r\in(0,1)$.
-The first general result without the variance decay assumption was obtained by Klusowski and Tian (2024).
-Their main result asserts that modulo model mis-specification, the MSE of CART is upper bounded by $1/k$ at depth $k$; see (23) below.
-In a similar vein, Cattaneo et al.
-(2024) studied the convergence rates of the oblique CART and obtained a $1/k$ upper bound for MSE in the general setting (Theorem 1 therein) as well as an $r^{k}$ upper bound under further assumptions such as node sizes (Theorem 4 therein).
-The unconditional rate guarantee of $1/k$ appears significantly slower than the exponential decay rates under variance decay assumptions.
-On the other hand, Cattaneo et al.
-(2022) discussed why CART may have slower-than-polynomial convergence.
-This phenomenon was known as the end-cut preference (ECP) in the literature (see Section 11.8 of Breiman et al.
-(1984)).
-We have chosen to work on the model (7) in the simple regression setting (4), where sequential construction is prevalent (Luo and Pratola, 2023; Luo et al., 2024; Liu et al., 2023).
-Although there are multiple approaches to construct $\hat{m}_{N}$ in (7) sequentially (Loh, 2014), we focus on the CART-like decision tree model (Breiman et al., 1984) in the regression setting above.
-The classification setting has been well studied in the computational complexity and encoding literature (Blanc et al., 2020; O’Donnell et al., 2005) in terms of using uniform bounds on Boolean functions between partial trees and optimal trees.
-1.3 Our contribution
-Partition-based martingale approximations.
-We show that various martingale constructions are feasible and enjoy exponential convergence rates.
-For instance, a classic construction by Simons (1970) is the following: start from $\pi_{0}=\{\mathbb{R}\}$ and for each $k\geqslant 0$ and $A=[a,b)\in\pi_{k}$ where $-\infty\leqslant a<b\leqslant\infty$, split $A$ into $[a,\mathbb{E}[U\mid U\in A])$ and $[\mathbb{E}[U\mid U\in A],b)$ to form the sets in $\pi_{k+1}$.
-We show that if $U$ is bounded, the convergence rate of the Simons martingale is $C2^{-k}$, where $C$ does not depend on the law of $U$.
-Moreover, the rate $2^{-k}$ can be shown to be asymptotically optimal.
-Other constructions of partition-based martingale approximations involve replacing the conditional mean in the Simons martingale by other criteria such as the median.
-The explicit constructions are given formally by Definition 1, whose convergence rates are summarized in Theorem 3.
-All constructions yield explicit exponential rates of convergence in both the bounded and the unbounded cases under moment constraints (Theorem 4).
-Under further assumptions on the law of $U$, we also establish exponential rates of $Cr^{k}$ where $r$ can be made arbitrarily close to $1/4$ (a threshold that cannot be surpassed with binary partitions) while allowing $C$ to depend on the law of $U$ (Theorem 5).
-Regression trees.
-We introduce new splitting criteria, namely the MinimaxSplit algorithm and its multivariate variation, the cyclic MinimaxSplit algorithm (see (14) and (15) below for formal definitions).
-On one hand, the VarianceSplit algorithm is based on variance reduction; that is, for each split, the remaining total variance within the children nodes is minimized (among the splits over all dimensions).
-On the other hand, by definition, the MinimaxSplit algorithm selects the split that minimizes the maximum variance within the children nodes (among splits over all dimensions), hence the term minimax.
-The cyclic MinimaxSplit algorithm selects the same active dimension on each level of the tree and the active dimension cycles over all dimensions.
-We show that for the cyclic MinimaxSplit algorithm, the MSE at depth $k$ decays exponentially with rate $2^{-2k/(3d)}$ (given sufficient samples), without any further assumption (Theorem 9).
-We further develop empirical risk bounds for the cyclic MinimaxSplit algorithm (Theorem 10).
-Paper outline.
-The respective solutions to (P-general) in the partition-based martingale approximation and the regression tree settings will be detailed in Sections 2 and 3.
-All technical proofs are given in Appendix A.
-2 On partition-based martingale approximations
-In this section, we develop the framework of partition-based martingale approximations and show that a number of constructions feature exponential convergence rates, thus answering (P-martingale).
-2.1 Basic concepts
-We say that a sequence of finite partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ is nested if for each $k\geqslant 0$ and $A\in\pi_{k+1}$, there exists (a unique) $B\in\pi_{k}$ such that $A\subseteq B$.
-Consider nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}$ and a real-valued random variable $U$ with a finite second moment.
-Recall (3):
-$M_{k}(\omega):=\mathbb{E}[U\mid U\in A],\quad\text{for }U(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k},$ (8)
-where the MSE $\mathbb{E}[(U-M_{k})^{2}]$ is non-increasing in $k$ by the nested property of $\{\pi_{k}\}_{k\geqslant 0}$ and the total variance formula.
-A discrete-time real-valued stochastic process $\{M_{k}\}_{k\geqslant 0}$ is a martingale if for any $0\leqslant\ell<k$, $\mathbb{E}[M_{k}\mid\sigma(M_{0},\dots,M_{\ell})]=M_{\ell}$.
-We use the abbreviation $\sigma(\pi_{k})$ to denote the $\sigma$-algebra generated by events of the form $\{U\in A\}_{A\in\pi_{k}}$.
-If $\Pi:=\{\pi_{k}\}_{k\geqslant 0}$ is a nested sequence of partitions, $\{\sigma(\pi_{k})\}_{k\geqslant 0}$ is a filtration generated by indicator functions of sets in $\pi_{k}$’s (Doob, 1953).
-It follows from the tower property of conditional expectations that the sequence $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ is a martingale (in fact, a Doob martingale (Doob, 1940, 1953)).
-We call this martingale the $\Pi$-based martingale approximation of the random variable $U$, or in general a partition-based martingale approximation that approximates the random variable in terms of MSE.
-Unless $U$ is a constant, there are different kinds of partition, so there are different partition-based martingale approximations depending on the distribution of $U$.
-Our goal in this section is to identify a few partition-based martingale approximations that efficiently approximate $U$, where the construction algorithm is universal; see (P-martingale).
-The efficiency criterion is given by the decay of the MSE $\mathbb{E}[(U-M_{k})^{2}]$.
-Without loss of generality, we give our construction of a partition of a generic interval $[a,b)\subset\mathbb{R}$, where $a,b\in\mathbb{R}\cup\{\pm\infty\}$.
-[^2] Here we slightly abuse notation that $[a,b)=(-\infty,b)$ if $a=-\infty$.
-The sequence of partitions $\Pi$ in the response space will be constructed recursively, where $\pi_{0}=\{\mathbb{R}\}$ and for every $k\geqslant 0$, each interval $A\in\pi_{k}$ splits into two intervals, by following the same construction, forming the elements in $\pi_{k+1}$.
-In the following, we introduce four distinct splitting rules that define partition-based martingale approximations.
-For simplicity, we assume that $U$ is atomless, so that the endpoints of the intervals do not matter and that no trivial split occurs.
-Definition 1.
-Suppose we are given an atomless law of $U$ and a non-empty interval $I=[a,b)$, where $a,b\in\mathbb{R}\cup\{\pm\infty\}$.
-  1. (i)
-    Define
-    $u_{\text{var}}=\operatorname*{arg\,min}_{u\in I}\big{(}\mathbb{P}(U\in[a,u))\mathrm{Var}(U\mid U\in[a,u))+\mathbb{P}(U\in[u,b))\mathrm{Var}(U\mid U\in[u,b))\big{)}.$ (9)
-    If the minimizer is not unique, we pick the largest minimizer.
-The variance splitting rule (corresponding to the VarianceSplit algorithm in Section 3.1) splits $I$ into the two sets $[a,u_{\text{var}})$ and $[u_{\text{var}},b)$.
-  2. (ii)
-    Define $u_{\text{Simons}}=\mathbb{E}[U\mid U\in I]$. The Simons splitting rule splits $I$ into the two sets $[a,u_{\text{Simons}})$ and $[u_{\text{Simons}},b)$.
-  3. (iii)
-    Define
-    $u_{\text{minimax}}=\operatorname*{arg\,min}_{u\in I}\max\big{\{}\mathbb{P}(U\in[a,u))\mathrm{Var}(U\mid U\in[a,u)),\,\mathbb{P}(U\in[u,b))\mathrm{Var}(U\mid U\in[u,b))\big{\}}.$
-    If the minimizer is not unique, we pick the largest minimizer.
-The minimax splitting rule (corresponding to the MinimaxSplit algorithm in Section 3.1) splits $I$ into the two sets $[a,u_{\text{minimax}})$ and $[u_{\text{minimax}},b)$.
-  4. (iv)
-    Define
-    $u_{\text{median}}=\sup\{u\in I:\mathbb{P}(a\leqslant U<u)=\mathbb{P}(u\leqslant U<b)\}.$
-    The median splitting rule splits $I$ into two sets $[a,u_{\text{median}})$ and $[u_{\text{median}},b)$.
-In turn, when applying recursively the variance (resp.
-Simons, minimax, median) splitting rule starting from $\pi_{0}=\{\mathbb{R}\}$, we obtain a nested sequence of partitions $\Pi=\{\pi_{k}\}_{k\geqslant 0}$ depending on the law of $U$.
-We call the resulting $\Pi$-based martingale approximation $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ the variance (resp.
-Simons, minimax, median) martingale (with respect to $U$).
-Example 2.
-If $U$ is uniformly distributed on a compact interval, all four martingales coincide.
-For example, if $U\sim\mathrm{Uni}[0,1]$, each of the four martingales from Definition 1 satisfy $M_{k}=\mathbb{E}[U\mid\sigma(\pi_{k})]$, where $\pi_{k}:=\{[j/2^{k},(j+1)/2^{k}):0\leqslant j<2^{k}\}$ and it follows that $\mathbb{E}[(U-M_{k})^{2}]\asymp 4^{-k}$ (where $\asymp$ means up to universal constants).
-The intuition for the minimax and median martingales is that at each splitting step, we balance the “sizes” of $U{\mathbbm{1}}_{\{U\in I\}}$ on the two sets.
-The “size” corresponds to the (unconditional) variance for the minimax martingale and the total probability for the median martingale.
-Intending to minimize the MSE $\mathbb{E}[(U-M_{k})^{2}]$ at each step $k$, the variance martingale naturally arises as an algorithm that greedily reduces the remaining risk within $U$ in each iteration through layers.
-In a partition-based martingale, the nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ can be naturally identified as the vertices of a binary tree.
-A binary tree is the unique infinite tree $(V,E)$ such that all vertices have degree $3$, except for a unique vertex $\emptyset$ called the root, which has degree $2$.
-Denote by $V_{k}$ the set of all vertices in $V$ with the graph distance from the root equal to $k$.
-Then, each vertex $v\in V_{k}$ can be identified as a set $A\in\pi_{k}$.
-With each edge $e=(A_{k},A_{k+1})\in E$ where $A_{k}\in\pi_{k}$ and $A_{k+1}\in\pi_{k+1}$, we may associate a coefficient $p_{e}:=\mathbb{P}(U\in A_{k+1})\in[0,1]$.
-With each vertex $v=A_{k}$, we may associate a location $\ell_{v}:=\mathbb{E}[U\mid U\in A_{k}]$.
-It follows that $M_{k}$ is supported on the discrete points $\{\ell_{v}\}_{v\in V_{k}}$ and furthermore,
-$\mathbb{E}[(M_{k}-M_{k+1})^{2}]=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}.$ (10)
-This binary tree representation (10) of the risk will be frequently used in this paper.
-In the next two sections, we discuss two types of results: uniform rates for bounded $U$ (Section 2.2) and non-uniform rates (Section 2.3) where the asymptotic constant may depend on the possibly unbounded law of $U$.
-2.2 Uniform convergence rates
+Note that in our setting, $Y$ is a sum of martingale differences, but whose variance is bounded, and hence one cannot apply the martingale CLT or its convergence rates.
+In the next two sections, we discuss two types of results: uniform rates for bounded $U$ (Appendix A.2) and non-uniform rates (Appendix A.3) where the asymptotic constant may depend on the possibly unbounded law of $U$.
+A.2 Uniform convergence rates
 Theorem 16.
-Let $U$ be a $[0,1]$-valued atomless random variable and $\{M_{k}\}_{k\geqslant 0}$ be one of the four martingale approximations with respect to $U$ given by Definition 1.
-The following statements hold.
+Let $U$ be a $[0,1]$-valued atomless random variable. The following statements hold.
   1. (i)
@@
     If $\{M_{k}\}_{k\geqslant 0}$ is the median martingale, $\mathbb{E}[(U-M_{k})^{2}]\leqslant 2^{-k}$.
+The proof of Theorem 16 is lengthy and technical, which we defer to Appendix A.4.
 The general case of a bounded $U$ can be derived by a scaling argument, since the constructions in Definition 14 are scale-invariant.
-Let us sketch the arguments for the minimax martingale, as the same idea becomes crucial when applied to minimax decision trees we develop in Section 3.
-Assuming that $U$ is atomless and supported in $[0,1]$, we apply the representation (10).
-By construction, $\sum p_{(v_{k},v_{k+1})}=1$ and $\sum|\ell_{v_{k}}-\ell_{v_{k+1}}|\leqslant\sup\operatorname{supp}U-\inf\operatorname{supp}U=1$.
-Moreover, by the minimax property and the law of total variance, it is straightforward to verify (with details in Section 3.2) that the variances decay at least geometrically by half at each split at $u_{\text{minimax}}$, i.e., for any $v\in V_{k}$ with $v^{\prime}\sim w$ for $v^{\prime}\in V_{k-1}$,
-$\max_{w\in V_{k+1},\,v\sim w}p_{(v,w)}(\ell_{v}-\ell_{w})^{2}\leqslant\frac{1}{2}p_{(v^{\prime},v)}(\ell_{v^{\prime}}-\ell_{v})^{2}.$
-Therefore, by induction,
-$\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\leqslant L2^{-k}\mathrm{Var}(U)\leqslant L2^{-k},$
-where $L>0$ is a universal constant that may not be the same on each occurrence. By Hölder’s inequality,
-$\mathbb{E}[(M_{k}-M_{k+1})^{2}]$ $=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}$
-$\leqslant\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}\bigg{)}^{1/3}$
-×$\hskip 28.45274pt\times\bigg{(}\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}$
-$\leqslant\bigg{(}\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\bigg{)}^{2/3}\leqslant L2^{-2k/3}.$
-The rest then follows from the martingale property: $\mathbb{E}[(U-M_{k})^{2}]\leqslant\sum_{\ell\geqslant k}\mathbb{E}[(M_{\ell}-M_{\ell+1})^{2}]\leqslant L2^{-2k/3}$.
 Finding the optimal rate $r\in(0,1)$ (where $\mathbb{E}[(U-M_{k})^{2}]\leqslant Lr^{k}$ for some universal constant $L$) is also an intriguing question.
 Observe that one cannot have $r<1/4$ by Example 15 above.
-In Examples 11 and 12 below, we show that the rate $r$ in Theorem 3 is indeed optimal for the Simons and median martingales.
-2.3 Non-uniform convergence rates
-The above Theorem 3 requires that the support of $U$ is bounded and the asymptotic constant depends only on the range of $U$, $\sup U-\inf U$.
+In Examples 17 and 18 below, we show that the rate $r$ in Theorem 16 is indeed optimal for the Simons and median martingales.
+Example 17 (Rate $r=1/2$ is optimal for the Simons martingale).
+Fix an arbitrary $s\in(1/2,1)$ and let $U$ satisfy $\mathbb{P}(U=-1)=1-s$ and for $J\geqslant 0$,
+$\mathbb{P}\bigg{(}U=\sum_{j=1}^{J}\Big{(}\frac{1-s}{s}\Big{)}^{j}\bigg{)}=s^{J+1}(1-s).$
+It follows that $U$ is a bounded random variable. Next, we compute that
+$\mathbb{E}\bigg{[}U\mid U\geqslant\sum_{j=1}^{J}\Big{(}\frac{1-s}{s}\Big{)}^{j}\bigg{]}$ $=\bigg{(}\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\bigg{)}^{-1}\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\sum_{j=1}^{\ell}\Big{(}\frac{1-s}{s}\Big{)}^{j}$
+$=s^{-(J+1)}\sum_{\ell=J}^{\infty}s^{\ell+1}(2s-1)\bigg{(}1+\Big{(}\frac{1-s}{s}\Big{)}^{\ell}\bigg{)}$
+$=\frac{s^{-(J+1)}(2s-1)s^{J+1}}{1-s}+\frac{s^{-(J+1)}(2s-1)s}{1-(1-s)}$
+$=\sum_{j=1}^{J+1}\Big{(}\frac{1-s}{s}\Big{)}^{j}.$
+It is then straightforward to verify that in the Simons martingale $\{M_{k}\}_{k\geqslant 0}$, each $M_{k+1}-M_{k}$ is supported on the set $\{0,-((1-s)/s)^{k},((1-s)/s)^{k+1}\}$, where $\mathbb{P}(M_{k+1}-M_{k}=-((1-s)/s)^{k})=(1-s)s^{k}$ and $\mathbb{P}(M_{k+1}-M_{k}=((1-s)/s)^{k+1})=s^{k+1}$.
+It follows that
+$\mathbb{E}[(M_{k+1}-M_{k})^{2}]=(1-s)^{2k+1}s^{-(k+1)}\asymp\Big{(}\frac{(1-s)^{2}}{s}\Big{)}^{k}.$
+In other words, the rate $r$ is given by $r=(1-s)^{2}/s$.
+If $s$ is picked close to $1/2$, $r$ can be made close enough to $1/2$.
+Therefore, $r=1/2$ is the optimal rate parameter.
+This example is illustrated with an atomic distribution $U$, but a slight twist also leads to an example for an absolutely continuous law $U$.
+Example 18 (Rate $r=1/2$ is optimal for the median martingale).
+Let $s\in(0,1)$ and a bounded random variable $U$ satisfy
+$\mathbb{P}\Big{(}U=\sum_{j=1}^{k-1}s^{j-1}-s^{k-1}\Big{)}=2^{-k},\leavevmode\nobreak\ k\geqslant 1.$
+Next, we compute that
+$\mathbb{E}\Big{[}U\mid U\geqslant\sum_{j=1}^{k}s^{j-1}-s^{k}\Big{]}$ $=2^{k}\sum_{\ell=k+1}^{\infty}2^{-\ell}\bigg{(}\sum_{j=1}^{\ell-1}s^{j-1}-s^{\ell-1}\bigg{)}$
+$=2^{k}\sum_{\ell=k+1}^{\infty}2^{-\ell}\Big{(}\frac{1}{1-s}-\frac{s^{\ell-1}(2-s)}{1-s}\Big{)}$
+$=\frac{1}{1-s}-\frac{2-s}{s(1-s)}\frac{s^{k+1}/2}{1-s/2}$
+$=\sum_{j=1}^{k}s^{j-1}.$
+Note also that
+$\sum_{j=1}^{k}s^{j-1}-s^{k}<\sum_{j=1}^{k}s^{j-1}<\sum_{j=1}^{k+1}s^{j-1}-s^{k+1}.$
+It is then straightforward to verify that the law of $M_{k+1}-M_{k}$ is supported on $\{0,\pm s^{k}\}$ with $\mathbb{P}(M_{k+1}-M_{k}=s^{k})=\mathbb{P}(M_{k+1}-M_{k}=-s^{k})=2^{-(k+1)}$.
+We then obtain that
+$\mathbb{E}[(M_{k+1}-M_{k})^{2}]=2^{-k}s^{2k}.$
+In other words, the rate $r=s^{2}/2$ has a limit $1/2$ as $s\to 1$ from below; hence, the optimal rate is $1/2$.
+Again, a simple twist yields similar examples, where the law of $U$ is absolutely continuous.
+A.3 Non-uniform convergence rates
+The above Theorem 16 requires that the support of $U$ is bounded and the asymptotic constant depends only on $\sup\operatorname{supp}U-\inf\operatorname{supp}U$.
 We show in the next result that, without the bounded range assumption, non-uniform convergence rates can still be guaranteed.
@@
 Let $U$ be an atomless random variable with a finite $(2+\varepsilon)$-th moment for some $\varepsilon>0$, and $\{M_{k}\}_{k\geqslant 0}$ be one of the four partition-based martingale approximations given by Definition 14.
-Then there exist constants $r\in(0,1)$ and $C>0$, both possibly depending on the law of $U$, such that
+Then there exists a constant $r\in(0,1)$ depending only on $\varepsilon$ and another constant $C>0$ depending on the law of $U$, such that
 $\mathbb{E}[(U-M_{k})^{2}]\leqslant Cr^{k}.$
 In particular, $M_{k}\to U$ both a.s. and in $L^{2}$.
+Proof.
+Let $\operatorname{supp}M_{1}=\{x,y\}$ where $x<y$.
+Take $M>\sup\{x,y\}$.
+Let $\tau_{M}^{+}$ be the first hitting time to $\{x:x\geqslant M\}$, $\tau_{M}^{-}$ be the first hitting time to $\{x:x\leqslant-M\}$, and $\tau_{M}=\min\{\tau_{M}^{+},\tau_{M}^{-}\}$.
+It is straightforward to prove (see the proof of Lemma 5.6 of Zhang et al.
+(2024)) that for some $r\in(0,1)$ and $C>0$,
+$\mathbb{E}[(U-M_{k})^{2}]\leqslant\mathbb{E}[{\mathbbm{1}}_{\{\tau_{M}<\infty\}}U^{2}]+CM^{2}r^{k}.$
+Since the supports of $\{M_{k}\}_{k\geqslant 0}$ correspond to the means of $U$ conditioned on nested partitions of $\mathbb{R}$, we have $\tau_{M}^{+}<\infty$ implies $M_{k}\geqslant x$ for all $k\geqslant 0$ and $M_{1}=y$.
+Therefore, conditional on the event $\tau_{M}^{+}<\infty$, $X$ is a martingale that is bounded from below by $x$.
+By Ville’s inequality,
+$\mathbb{P}(\tau_{M}^{+}<\infty)$ $\leqslant\mathbb{P}(\tau_{M}^{+}<\infty\mid M_{1}>x)$
+$=\mathbb{P}\left(\sup_{k\geqslant 0}M_{k}\geqslant M\mid M_{1}>x\right)\leqslant\frac{\mathbb{E}[U\mid M_{1}>x]-x}{M-x}=O(M^{-1}).$
+Similarly, the same analysis holds for $\tau_{M}^{-}$.
+Hence, we conclude that $\mathbb{P}(\tau_{M}<\infty)=O(M^{-1}).$ The rest follows line by line as in Zhang et al.
+(2024), which we sketch for completeness.
+By Hölder’s inequality, $\mathbb{E}[{\mathbbm{1}}_{\{\tau_{M}<\infty\}}U^{2}]=O(M^{-\delta^{\prime}})$ for some $\delta^{\prime}>0$ depending only on $\varepsilon$.
+With the choices $M=r^{-k/(2+\delta^{\prime})}$ and $q=r^{\delta^{\prime}/(2+\delta^{\prime})}\in(0,1)$, we have that for some $C>0$,
+$\mathbb{E}[(U-M_{k})^{2}]\leqslant C(M^{-\delta^{\prime}}+M^{2}r^{k})\leqslant Cq^{-k}.$
+The final statement follows immediately from the martingale convergence theorem. ∎
 If $U$ is bounded, Theorem 19 is a special case of Theorem 16.
 The case of the Simons martingale has been established by Zhang et al.
-(2024), and our proof of Theorem 4 will follow a similar route.
+(2024), and our proof of Theorem 19 follows a similar route.
 Since the asymptotic constant is allowed to depend on the law of $U$, the optimal non-uniform rate $r$ might be smaller than the uniform ones (in Theorem 16).
-Recall from Section 2.2 that the rates obtained in Theorem 3 are asymptotically optimal for the Simons and median martingales (see Examples 11 and 12).
+Recall from Appendix A.2 that the rates obtained in Theorem 16 are asymptotically optimal for the Simons and median martingales (see Examples 17 and 18).
 For the variance and minimax martingales, we show in the next result that the rate $r=1/4$ is optimal under certain regularity conditions on the law of $U$.
@@
 $\mathbb{E}[|U-M_{k}|^{2}]\leqslant Cr^{k}.$
-In other words, under the assumptions in Theorem 5, the MSE has an asymptotic convergence rate of $1/4+\varepsilon$.
+Proof.
+First, we claim that the split points for the nested partitions $\{\pi_{k}\}$ are dense.
+Indeed, for both martingales, we have shown that the variances on each component in the partition $\pi_{k}$ go to zero uniformly as $k\to\infty$.
+If the split points were not dense, some element in the partition would cover an interval of positive length, on which the variance of $U$ cannot vanish since we assumed $\inf f>0$ and $\sup f<\infty$.
+Therefore, for any $\varepsilon>0$, there exists $k$ such that the $k$-th partition $\pi_{k}$ is such that on each interval $A\in\pi_{k}$, $\sup_{A}f\leqslant(1+\varepsilon)\inf_{A}f$.
+Since one-step variance is always better than one-step minimax, it remains to show that for an interval $A$ with $\sup_{A}f\leqslant(1+\varepsilon)\inf_{A}f$, there exists $C(\varepsilon)\downarrow 1/4$ as $\varepsilon>0$ such that at each step the total variance reduces by a factor of $C(\varepsilon)$ for the minimax martingale.
+To see this, without loss of generality we start from $U$ supported on $[0,1]$ (meaning that $\mathbb{P}(U\in[0,1])=1$) whose density takes values in $1,1+\varepsilon]$ (we may normalize afterwards to a probability measure, which we omit).
+Then for any interval $a,b]\subseteq[0,1]$,
+$\mathbb{E}[U\mid U\in[a,b]]\in\Big{[}a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)\Big{]}.$
+Therefore,
+$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])$ $\leqslant\sup_{m\in[a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)]}\int_{a}^{b}(1+\varepsilon)(x-m)^{2}\mathrm{d}x$
+$\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(b-a)^{3}.$
+Similarly,
+$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])\geqslant\inf_{m\in[a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)]}\int_{a}^{b}(x-m)^{2}\mathrm{d}x\geqslant\frac{1}{12}(b-a)^{3}.$
+It follows that the split point $x_{*}\in[0,1]$ must satisfy
+$\frac{1}{12}x_{*}^{3}\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(1-x_{*})^{3}\quad\text{and}\quad\frac{1}{12}(1-x_{*})^{3}\geqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}x_{*}^{3}.$
+Therefore, for some $a_{\varepsilon}\to 0$, $|x_{*}-1/2|\leqslant a_{\varepsilon}$.
+A direct computation yields $\mathrm{Var}(U)\geqslant\frac{1}{12}$. The total remaining variance $\mathbb{E}[(U-M_{1})^{2}]$ is bounded by
+$\mathbb{E}[(U-M_{1})^{2}]\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(x_{*}^{3}+(1-x_{*})^{3})\leqslant\frac{4(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(\frac{1}{2}+a_{\varepsilon})^{3}.$
+Hence, as $\varepsilon\to 0$, the ratio becomes
+$\frac{\mathbb{E}[(U-M_{1})^{2}]}{\mathrm{Var}(U)}\leqslant\frac{2(1+\varepsilon)^{4}}{(2+\varepsilon)^{3}}(1+2a_{\varepsilon})^{3}\to\frac{1}{4},$
+as desired. ∎
+In other words, under the assumptions in Theorem 20, the MSE has an asymptotic convergence rate of $1/4+\varepsilon$”.
 Unfortunately, we are unable to remove the assumptions for the density of $U$ in Theorem 20 nor give counterexamples.
 We conjecture that the same conclusion of Theorem 20 holds without those assumptions.
-3 The MinimaxSplit algorithm
-The goal of this section is to provide a solution to (P-regression) by constructing decision tree algorithms with exponential convergence.
-Section 3.1 gives the constructions inspired by the minimax martingale construction in Section 2.
-In traditional regression decision tree settings that minimize (1), models greedily split in order to minimize the sum of variances over the responses of the children nodes (Liu et al., 2023; Breiman et al., 1984), yet we propose to greedily minimize the maximum variance among the children nodes.
-In the following, we say that a law on $\mathbb{R}^{d}$ is marginally atomless if its projection onto any of the $d$ dimensions is atomless.
-We will start from a marginally atomless coupling (that may not arise from empirical measures $\mathbb{P}_{N}$) and prove exponential rates in Section 3.2.
-Section 3.3 builds upon this case, covering more generally the non-marginally atomless setting and, in particular, the empirical risk bounds.
-Further numerical analysis will be presented in Section 3.4.
-All proofs can be found in Appendix A.4.
-3.1 Formulation of the algorithms
-We begin by recapping the greedy splitting regime in the VarianceSplit algorithm to construct an efficient decision tree.
-Subsequently, we propose alternative CART algorithms, namely the MinimaxSplit algorithm, along with its variation, the cyclic MinimaxSplit algorithm, and make comparisons against the VarianceSplit algorithm.
-Roughly speaking, the (cyclic) MinimaxSplit algorithms are analogues of the minimax martingale from Definition 1.
-Hereafter, we denote by $[d]=\{1,\dots,d\}$ for $d\in\mathbb{N}$.
-We also need the following notion of a splittable set.
-If the coupling is defined on a continuous space, the split can always be performed; yet, for atomic measures like empirical measures, we may halt decision splits when there is no probability mass remained.
-Intuitively, in the following decision tree algorithms, the decision trees only split a node if it corresponds to a splittable set.
-A similar notion is also introduced by Definition 3.1 of Klusowski and Tian (2024).
-Definition 3.1.
-Consider a coupling $(\mathbf{X},Y)$ on $\mathbb{R}^{d+1}$. We say a set $A\subseteq\mathbb{R}^{d}$ is $(\mathbf{X},Y)$-non-splittable, or non-splittable, if any of the following occurs:
-  • $\mathbb{P}(\mathbf{X}\in A)=0$;
-  • $\mathbb{P}(\mathbf{X}\in A)>0$ and $Y\mid\mathbf{X}\in A$ is a constant;
-  • $\mathbb{P}(\mathbf{X}\in A)>0$ and $\mathbf{X}\mid\mathbf{X}\in A$ is a constant.
-Otherwise, we say that $A$ is $(\mathbf{X},Y)$-splittable, or splittable.
-Revisiting the VarianceSplit algorithm.
-The greedy VarianceSplit construction (Section 8.4 of Breiman et al.
-(1984) and Luo and Li (2024)) proceeds by introducing nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ with axis-aligned borders, in a way that sequentially and greedily minimizes the total risk $\mathbb{E}[(Y-M_{k})^{2}]$ at each step $k$ (while splitting from the fixed partition $\pi_{k-1}$), where $M_{k}$ is taken as the conditional mean of $Y$ given the $\sigma$-algebra generated by the collection ${\mathbbm{1}}_{\{\mathbf{X}\in A\}},\,A\in\pi_{k}$.
-So, the partitions $\{\pi_{k}\}_{k\geqslant 0}$ of the input space are nested and we require that each $\pi_{k}$ consists of sets that split each element $A\in\pi_{k-1}$ into two hyper-rectangles with axis-aligned borders, unless all covariates and/or all response values on the event $\{\mathbf{X}\in A\}$ are the same, in which case we do not perform split and $A\in\pi_{k}$.
-Formally, suppose that a hyper-rectangle $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})$ belongs to the partition $\pi_{k-1}$ and is splittable.
-In particular, $M_{k-1}(\omega)=\mathbb{E}[Y\mid\mathbf{X}\in A]$ if $\mathbf{X}(\omega)\in A$.
-To find a split of $A$ at depth $k$, the VarianceSplit algorithm looks for a covariate $j\in[d]$ and $x_{j}\in[a_{j},b_{j})$ such that the remaining risk $\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]$ is the smallest after splitting the $A$ at $x_{j}$ in covariate $j$.
-In other words, one seeks for
-$\begin{split}(j,x_{j})&=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big{(}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j})\\ &\hskip 56.9055pt+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{)}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big{(}\mathbb{E}\big{[}(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}\big{]}\\ &\hskip 56.9055pt+\mathbb{E}\big{[}(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}\big{]}\Big{)}\\ &=:\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}(V_{\text{left}}+V_{\text{right}}),\end{split}$ (11)
-where we break ties arbitrarily.
-[^3] Strictly speaking, the $\operatorname*{arg\,min}$ in (11) may not always be attained for general couplings $(\mathbf{X},Y)$, unless we assume that $\mathbf{X}$ is marginally either atomless or has a finite support, which will always be the case in our analysis.
-However, in the general setting, if we allow splits in which atoms can be duplicated and assigned to both nodes (instead of only on the right node as in (11)), $\operatorname*{arg\,min}$ can be attained.
-In this paper, we will implicitly assume that $\operatorname*{arg\,min}$ is always attained.
-Consequently, the splittable set $A$ splits into the sets $A_{L}:=[a_{1},b_{1})\times\dots\times[a_{j},x_{j})\times\dots\times[a_{d},b_{d})$ and $A_{R}:=[a_{1},b_{1})\times\dots\times[x_{j},b_{j})\times\dots\times[a_{d},b_{d})$ to form its two descendants in the set $\pi_{k}$, and define
-$M_{k}(\omega)=\begin{cases}\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{L};\\ \mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{R}.\end{cases}$ (12)
-In other words, at depth $k\geqslant 0$ and after constructing the partition $\pi_{k}$, define $M_{k}$ by
-$M_{k}(\omega)=\mathbb{E}[Y\mid\mathbf{X}\in A],\quad\text{if }\mathbf{X}(\omega)\in A,\quad\text{for }A\in\pi_{k}.$ (13)
-This leads to the desired coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$, where $\{M_{k}\}_{k\geqslant 0}$ is a martingale.
-It is also common practice to weigh the terms $V_{\text{left}}$ and $V_{\text{right}}$ in (11) by the sample sizes of the left and right children nodes (resulting in a minimization of $n_{\text{left}}V_{\text{left}}+n_{\text{right}}V_{\text{right}}$), for which we call the Weighted Variance algorithm.
-The MinimaxSplit algorithm.
-A common feature of the MinimaxSplit and VarianceSplit algorithms is that they both start from a nested sequence of partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ consisting of (at-most) binary splits into hyper-rectangles with axis-aligned borders.
-The $k$-th approximation $M_{k}$ is defined by (13).
-The two algorithms differ in the way the partition is constructed: in the MinimaxSplit setting, the split is no longer the best split that reduces total variance but instead is the best split that minimizes the maximum variance within the two descendants.
-Formally, for a splittable set $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})\in\pi_{k-1}$, define the split location as
-$\begin{split}(j,\hat{x}_{j})&=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}],\\ &\hskip 85.35826pt\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}]\Big{\}}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}),\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{\}}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}V_{\text{left}},V_{\text{right}}\Big{\}},\end{split}$ (14)
-where we break ties arbitrarily.
-Consequently, the splittable set $A$ splits into the two hyper-rectangles $[a_{1},b_{1})\times\dots\times[a_{j},\hat{x}_{j})\times\dots\times[a_{d},b_{d})$ and $[a_{1},b_{1})\times\dots\times[\hat{x}_{j},b_{j})\times\dots\times[a_{d},b_{d})$ to form its two descendants in the set $\pi_{k}$.
-We call this the MinimaxSplit rule.
-After obtaining the partition $\pi_{k}$, define $M_{k}$ through (13).
-This leads to a coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$, where $\{M_{k}\}_{k\geqslant 0}$ is a martingale.
-The cyclic MinimaxSplit algorithm.
-The cyclic MinimaxSplit algorithm is a variation of the MinimaxSplit algorithm.
-Instead of optimizing over all covariates $j\in[d]$ in (14), we cycle through the $d$ covariates as the tree grows.
-That is, for $k\geqslant 1$, let $j=j_{k}=(k-1\text{ (mod }d))+1\in[d]$.
-For all splittable sets $A\in\pi_{k-1}$, we split $A$ in the $j$-th coordinate to form its two descendants at depth $k$.
-The split location is then defined as
-$\begin{split}\hat{x}_{j}&=\operatorname*{arg\,min}_{x_{j}\in[a_{j},b_{j})}\max\Big{\{}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}],\\ &\hskip 85.35826pt\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}]\Big{\}}\\ &=\operatorname*{arg\,min}_{x_{j}\in[a_{j},b_{j})}\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}),\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{\}},\end{split}$ (15)
-where we break ties arbitrarily.
-Note that if $\mathbf{X}$ is marginally atomless, the minimizer $\hat{x}_{j}$ is essentially unique, in the sense that $\mathbb{P}(\mathbf{X}\in[a_{1},b_{1})\times\dots\times(\hat{x}_{j},\hat{x}_{j}^{\prime})\times\dots\times[a_{d},b_{d}))=0$ between two minimizers $\hat{x}_{j}<\hat{x}_{j}^{\prime}$.
-Define $M_{k}$ by (13).
-This leads to the desired coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$, where again, $\{M_{k}\}_{k\geqslant 0}$ is a martingale aiming to approximate $Y$.
-If $d=1$, the cyclic MinimaxSplit algorithm coincides with the MinimaxSplit algorithm.
-If $d>1$, compared to the MinimaxSplit algorithm, the cyclic MinimaxSplit algorithm is technically more tractable.
-Moreover, since all features are considered equally, the cyclic MinimaxSplit algorithm appears more robust in recovering the underlying geometry of the regression function in the setting of (4) in low dimensions, which we illustrate numerically in Section 3.4.
-On the other hand, the cyclic MinimaxSplit algorithm may be less effective in higher dimensions because the major dimensions may not be easily identified or reached in the first few splits.
-(At-most-binary) tree representation.
-In the context of regression trees, the same development of the tree representation (10) holds, except that the tree may not be a binary tree due to possible non-splittable sets.
-Instead, there exists a tree $(V,E)$ that is a sub-tree of a binary tree whose vertices are identified as the sets in the nested partitions.
-With the same notation leading to (10), we have
-$\mathbb{E}[(M_{k}-M_{k+1})^{2}]=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}.$
-The goal of the next section is to provide a partial answer to (P-general) under the cyclic MinimaxSplit rule.
-Assuming marginal non-atomicity, we show that up to universal constants and model mis-specification, the approximation error decays exponentially with rate $r^{k}$ for some explicit $r\in(0,1)$ that depends only on dimension $d$.
-3.2 Cyclic MinimaxSplit risk bound
-We first claim that if $\mathbf{X}$ is marginally atomless and $j\in[d]$, the $L^{2}$ loss
-$\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j})$ (16)
-is non-decreasing and continuous in $x_{j}$.
-To see this, we take $a<b$ and define events $E=E_{1}\cup E_{2}$, where $E_{1}=\{\mathbf{X}\in A,\,X_{j}<a\}$ and $E_{2}=E_{1}^{c}=\{\mathbf{X}\in A,\,a\leqslant X_{j}<b\}.$ By the total variance formula,
-$\begin{split}\mathrm{Var}(Y\mid E)&=\mathbb{E}[\mathrm{Var}(Y\mid\sigma(E_{1},E_{2}))\mid E]+\mathrm{Var}(\mathbb{E}[Y\mid\sigma(E_{1},E_{2})]\mid E)\\ &\geqslant\mathbb{E}[\mathrm{Var}(Y\mid\sigma(E_{1},E_{2}))\mid E]\geqslant\mathbb{P}(E_{1}\mid E)\mathrm{Var}(Y\mid E_{1}).\end{split}$ (17)
-Observe that $\mathrm{Var}(Y\mid E)=\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<b)$ and $\mathrm{Var}(Y\mid E_{1})=\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<a)$.
-Inserting these two expressions back into (17), we obtain
-$\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<b)\geqslant\frac{\mathbb{P}(\mathbf{X}\in A,X_{j}<a)}{\mathbb{P}(\mathbf{X}\in A,X_{j}<b)}\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<a).$
-Rearranging gives
-$\mathbb{P}(\mathbf{X}\in A,\,X_{j}<a)\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<a)\leqslant\mathbb{P}(\mathbf{X}\in A,\,X_{j}<b)\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<b)$
-and therefore (16) (or the first term inside the maximum in (15)) is non-decreasing; and similarly the second term inside the maximum is non-increasing.
-Moreover, by the marginally atomless property, both quantities are continuous.
-Therefore, if the $\operatorname*{arg\,min}$ is attained in (15) in the cyclic MinimaxSplit setting, the two terms inside the maximum are equal.
-Namely,
-$\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})=\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}).$ (18)
-Since the events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$ form a partition of $\{\mathbf{X}\in A\}$, we have again by the total variance formula,
-$\begin{split}&\hskip 14.22636pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\\ &\leqslant\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}].\end{split}$ (19)
-Combining (18) and (19), we have
-$\begin{split}&\hskip 14.22636pt\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}\}}],\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}\}}]\Big{\}}\\ &\leqslant\frac{1}{2}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}].\end{split}$ (20)
-In other words, the remaining risk on each descendant is at most half the risk on the parent node. Inductively, we gain a uniform geometric decay of the maximum risk among the nodes at depth $k$:
-$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]\leqslant 2^{-k}\mathrm{Var}(Y).$ (21)
-This is the key to controlling the “sizes of the nodes”.
-In the case of the VarianceSplit construction, such uniform control is reminiscent—counterexamples exist due to ECP, which explains why assumptions such as the SID condition are favorable to ensure risk decay when using loss functions like (11) (Chi et al., 2022; Mazumder and Wang, 2024).
-To see the intuition, we consider the event $\{\mathbf{X}\in A\}$ and sub-events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$.
-Applying the total variance formula to this partition yields
-$\text{Var}(Y)=$ $\mathbb{E}[\mathrm{Var}(Y\mid\sigma(A_{L}))]+\mathrm{Var}(\mathbb{E}[Y\mid\sigma(A_{L})])$
-$\begin{split}=&\text{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\text{Var}(Y\mid A_{R})\mathbb{P}(A_{R})\\ &+\left(\mathbb{E}[Y\mid A_{L}]-\mathbb{E}[Y\mid A_{R}]\right)^{2}(1-\mathbb{P}(A_{L}))\mathbb{P}(A_{L}).\end{split}$ (22)
-From this, we see that when ECP occurs, $\mathbb{P}(A_{L})\approx 1$ or 00, the improvement is dominated by the first two terms of (22): $\text{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\text{Var}(Y\mid A_{R})\mathbb{P}(A_{R})=V_{\text{left}}+V_{\text{right}}$.
-Therefore, when ECP occurs, the sum of variances does not necessarily decay fast enough for the VarianceSplit.
-However, for cyclic MinimaxSplit, $\max\{V_{\text{left}},V_{\text{right}}\}$ always displays geometric decay (21), hence “avoiding” the ECP phenomenon.
-In other words, the fast decay of the sum of variance is not easily guaranteed from the VarianceSplit construction due to ECP; but the geometric decay of the maximum variance can be assured by the MinimaxSplit construction.
-Consider the additive function class
-$\mathcal{G}:=\{g(\mathbf{x}):=g_{1}(x_{1})+\dots+g_{d}(x_{d})\},$
-where $\mathbf{x}=(x_{1},\dots,x_{d})$.
-For $g\in\mathcal{G}$, define $\left\lVert g\right\rVert_{\mathrm{TV}}$ as the infimum of $\sum_{i=1}^{d}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}$ over all such additive representations of $g$.
-(A continuous version of) Theorem 4.2 of Klusowski and Tian (2024) states that if $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$ is constructed from the VarianceSplit algorithm, then for each $k\geqslant 1$,
-$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\frac{\left\lVert g\right\rVert_{\mathrm{TV}}^{2}}{k+3}\Big{)}.$ (23)
-In the next result, we show that the $1/k$ rate in (23) can be significantly improved to an exponential rate for the cyclic MinimaxSplit algorithm.
-This also partially answers (P-general) from the Introduction.
-Note that the splitting here is binary because of the marginal non-atomicity condition.
-Theorem 6.
-Suppose that $\mathbf{X}$ is marginally atomless.
-Let the coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$ be constructed from the cyclic MinimaxSplit algorithm.
-Then uniformly for any $\delta>0$ and $k\geqslant 0$,
-$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant\inf_{g\in\mathcal{G}}\bigg{(}\Big{(}(1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big{)}\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]\\ &\hskip 56.9055pt+(1+\delta^{-1})\bigg{(}\frac{1}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg{)}.\end{split}$ (24)
+A.4 Proof of Theorem 16
+Variance
+Lemma A.1.
+Let $U$ be an atomless random variable supported on $I=[a,b]$. Then there exists
+$u_{*}\in\operatorname*{arg\,min}_{u\in I}\big{(}\mathbb{P}(U\in[a,u))\mathrm{Var}(U\mid U\in[a,u))+\mathbb{P}(U\in[u,b))\mathrm{Var}(U\mid U\in[u,b))\big{)}$
+such that
+$2u_{*}=\mathbb{E}[U\mid U\leqslant u_{*}]+\mathbb{E}[U\mid U>u_{*}]$ (35)
+and moreover, splitting at $u_{*}$ and at $u_{\mathrm{var}}$ defined through (34) have the same effect, [^9] Recall that we break ties in (34) by choosing the largest minimizer.
+i.e., $\mathbb{P}(U\leqslant u_{*})=\mathbb{P}(U\leqslant u_{\mathrm{var}})$.
+Proof.
+Let $Q$ be the quantile function of $U$. Since $U$ is atomless, we have if $u=Q(p)$,
+$\mathbb{P}(U\in[a,u))\mathrm{Var}(U\mid U\in[a,u))+\mathbb{P}(U\in[u,b))\mathrm{Var}(U\mid U\in[u,b))$
+$=p\mathrm{Var}(U\mid U\in[a,Q(p)))+(1-p)\mathrm{Var}(U\mid U\in[Q(p),b)).$
+Let $u_{\mathrm{var}}$ be defined through (34) and $p_{\mathrm{var}}=\mathbb{P}(U\leqslant u_{\mathrm{var}})$.
+Then with the short-hand notation $m_{a}(p)=\mathbb{E}[U\mid U\in[a,Q(p))]$ and $m_{b}(p)=\mathbb{E}[U\mid U\in[Q(p),b)]$, we have
+$p_{\mathrm{var}}$ $\in\operatorname*{arg\,min}_{p\in[0,1]}\big{(}p\mathrm{Var}(U\mid U\in[a,Q(p)))+(1-p)\mathrm{Var}(U\mid U\in[Q(p),b))\big{)}$
+$=\operatorname*{arg\,max}_{p\in[0,1]}\big{(}p\,m_{a}(p)^{2}+(1-p)m_{b}(p)^{2}\big{)}=:\operatorname*{arg\,max}_{p\in[0,1]}\Psi_{U}(p).$
+Observe that for $p\in(0,1)$,
+$\frac{\mathrm{d}}{\mathrm{d}p}m_{a}(p)=\frac{Q(p)-m_{a}(p)}{p}\qquad\text{ and }\qquad\frac{\mathrm{d}}{\mathrm{d}p}m_{b}(p)=\frac{m_{b}(p)-Q(p)}{1-p}.$
+We then have
+$\frac{\mathrm{d}}{\mathrm{d}p}\Psi_{U}(p)$ $=m_{a}(p)^{2}+2m_{a}(p)(Q(p)-m_{a}(p))-m_{b}(p)^{2}+2m_{b}(p)(m_{b}(p)-Q(p))$
+$=(m_{b}(p)-m_{a}(p))(m_{a}(p)+m_{b}(p)-2Q(p)).$
+Since $U$ is atomless, it is non-degenerate, and hence $m_{b}(p_{\mathrm{var}})-m_{a}(p_{\mathrm{var}})>0$ and $p_{\mathrm{var}}\in(0,1)$.
+Note that the functions $m_{a},m_{b}$ are differentiable and hence continuous; $Q$ may have jumps but have limits on both sides at every point.
+Since $p_{\mathrm{var}}\in\operatorname*{arg\,max}_{p\in[0,1]}\Psi_{U}(p)$, there are two cases:
+  • $Q$ is continuous at $p_{\mathrm{var}}$.
+Then $\Psi_{U}^{\prime}(p_{\mathrm{var}})=0$, so $m_{a}(p_{\mathrm{var}})+m_{b}(p_{\mathrm{var}})=2Q(p_{\mathrm{var}})$.
+In this case, we also have $Q(p_{\mathrm{var}})=u_{\mathrm{var}}$.
+Therefore, taking $u_{*}=u_{\mathrm{var}}$ gives (35).
+  • $Q$ is discontinuous at $p_{\mathrm{var}}$.
+Then we must have $\Psi_{U}^{\prime}(p_{\mathrm{var}}^{-})\geqslant 0$ and $\Psi_{U}^{\prime}(p_{\mathrm{var}}^{+})\leqslant 0$, where for a generic function $f$ defined on $\mathbb{R}$ we use the notation $f(x^{-})=\lim_{y\to x^{-}}f(y)$ and $f(x^{+})=\lim_{y\to x^{+}}f(y)$.
+Equivalently,
+    $2Q(p_{\mathrm{var}}^{-})\leqslant m_{a}(p_{\mathrm{var}})+m_{b}(p_{\mathrm{var}})\leqslant 2Q(p_{\mathrm{var}}^{+}).$ (36)
+    Let $u_{1}=Q(p_{\mathrm{var}}^{-})$ and $u_{2}=Q(p_{\mathrm{var}}^{+})$.
+Since $Q$ is discontinuous at $p_{\mathrm{var}}$, we have $\mathbb{P}(u_{1}\leqslant U\leqslant u_{2})=0$.
+Therefore, the map $u\mapsto\mathbb{P}(U\leqslant u)$ is constant on $u_{1},u_{2}]$ and for each $u\in[u_{1},u_{2}]$, $(\mathbb{E}[U\mid U\leqslant u],\mathbb{E}[U\mid U>u])=(m_{a}(p_{\mathrm{var}}),m_{b}(p_{\mathrm{var}}))$.
+It follows from (36) that (35) holds for some $u_{*}\in[u_{1},u_{2}]$.
+Using the definition (34) and $p_{\mathrm{var}}=\mathbb{P}(U\leqslant u_{\mathrm{var}})$, we have $u_{\mathrm{var}}=\max_{u\in[u_{1},u_{2}]}u=u_{2}$, which implies that $\mathbb{P}(U\leqslant u_{*})=\mathbb{P}(U\leqslant u_{\mathrm{var}})$.
+Therefore, in both cases, the desired claim is verified. ∎
 Remark 21.
-Since the bound (24) is uniform over $\delta,k,d$, one can further optimize over $\delta$ for fixed $k,d$.
-For example, taking $\delta=2^{-\lfloor k/d\rfloor/10}$, the right-hand side of (24) then has the more explicit (but cruder) upper bound
-$\inf_{g\in\mathcal{G}}\Big{(}5\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+2^{1-\lfloor k/d\rfloor/2}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\Big{)}.$
-Remark 8.
-In the case where $\left\lVert g\right\rVert_{\mathrm{TV}}$ is much larger than the range $\Delta g:=\sup g-\inf g$, (24) can be further improved to
-$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant\inf_{g\in\mathcal{G}}\bigg{(}(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+2^{1/3}(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{2/3}\\ &\hskip 113.81102pt+2^{-2/3}(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}(\Delta g)^{4/3}\bigg{)}.\end{split}$ (25)
-Similar considerations also apply for Theorems 9 and 10 below.
-Let us explain the intuition behind Theorem 6 when $d=1$ and $Y=g(X)$ for some $g\in\mathcal{G}$. Roughly speaking, (24) amounts to
-$\mathbb{E}[(Y-M_{k})^{2}]\leqslant L2^{-2k/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}.$ (26)
-To verify (26), observe that $\{M_{k}\}_{k\geqslant 0}$ is a martingale converging to $Y$.
-We employ the tree representation to control $\mathbb{E}[(M_{k}-M_{k+1})^{2}]$.
-By Hölder’s inequality and (21),
-$\mathbb{E}[(M_{k}-M_{k+1})^{2}]$ $=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}$
-$\leqslant\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}\bigg{)}^{1/3}$
-×$\hskip 28.45274pt\times\bigg{(}\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}$
-$\leqslant\mathrm{Var}(Y)^{2/3}2^{-2k/3}\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}.$
-Since each split introduces a difference term $|\ell_{v_{k}}-\ell_{v_{k+1}}|$, which is bounded by the values of $Y$ conditioned on a partition in the covariate space, we have
-$\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\leqslant\sum_{A\in\pi_{k+1}}\Big{(}\sup_{A}g-\inf_{A}g\Big{)}\leqslant\left\lVert g\right\rVert_{\mathrm{TV}}.$
-On the other hand, $\mathrm{Var}(Y)=\mathrm{Var}(g(\mathbf{X}))\leqslant\left\lVert g\right\rVert_{\mathrm{TV}}^{2}$.
-Altogether (and using the martingale property), we obtain $\mathbb{E}[(Y-M_{k})^{2}]\leqslant L2^{-2k/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}$.
-Let us compare the rates between (23) and (24).
-For $Y=g(\mathbf{X})$ with $g\in\mathcal{G}$ and $d$ fixed, we consider $\varepsilon>0$.
-For the VarianceSplit algorithm, (23) requires $2^{k}\approx e^{L\varepsilon^{-1}}$ nodes in the decision tree to guarantee that the MSE is smaller than $\varepsilon$ for some $L>0$.
-On the other hand, for the cyclic MinimaxSplit algorithm, (24) requires only $2^{k}\approx\varepsilon^{-Ld}$ nodes.
-Therefore, as $\varepsilon\to 0$, the cyclic MinimaxSplit algorithm requires fewer nodes to attain the same risk bound and thus facilitates computation.
-Note that Theorem 6 exhibits the curse of dimensionality, which cannot be circumvented in the current setting to the best of our knowledge (Cattaneo et al., 2022; Tan et al., 2022).
-3.3 Empirical risk bound for cyclic MinimaxSplit
-We now focus on the regression setting and apply Theorem 9 to derive finite-sample performance guarantees.
-Suppose that under the original law $\mathbb{P}_{*}$, $\mathbf{X}_{*}$ is marginally atomless and
-$Y_{*}\mid\mathbf{X}_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon:=\mathbb{E}[Y_{*}\mid\mathbf{X}_{*}]+\varepsilon,$
-where the error $\varepsilon=Y_{*}-g_{*}(\mathbf{X}_{*})$ is sub-Gaussian, i.e., for some $\sigma>0$,
-$\mathbb{P}_{*}(|\varepsilon|\geqslant u)\leqslant 2\exp\Big{(}-\frac{u^{2}}{2\sigma^{2}}\Big{)},\quad u\geqslant 0.$
-Also, we consider the empirical law of $(\mathbf{X},Y)$ of $N$ samples from $\mathbb{P}_{*}$.
-A limitation of Theorem 6 is that it applies only if the covariate $\mathbf{X}$ is marginally atomless (for (20) to hold).
-Therefore, we need the following result that deals with measures whose marginals may contain atoms.
-This also provides a more complete answer to (P-general).
-Theorem 9.
-Suppose that $\mathbf{X}$ is purely atomic [^4] The atomicity condition is only required such that the $\operatorname*{arg\,min}$ in (15) is well-defined; see Footnote 3. and for some $N>0$,
-$\max_{j\in[d]}\,\max_{u\in\mathbb{R}}\mathbb{P}(X_{j}=u)\leqslant\frac{1}{N}.$ (27)
-Let $M:=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$ and the coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$ be constructed from the cyclic MinimaxSplit algorithm.
-Then uniformly for any $\delta>0$ and $k\geqslant 0$,
-$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}+\inf_{g\in\mathcal{G}}\bigg{(}\Big{(}(1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big{)}\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]\\ &\hskip 170.71652pt+(1+\delta^{-1})\bigg{(}\frac{2}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg{)}.\end{split}$ (28)
-The compensation term
-$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}$
-appearing in (28) guarantees that if the number of samples $N$ is much larger than $2^{k}M^{2}$, the asymptotic upper bounds (24) and (28) are comparable.
-[…]
-Theorem 10.
-Assume that $\mathbf{X}_{*}$ is marginally atomless. We have for $\delta\geqslant 2^{-2\lfloor k/d\rfloor/3}$,
-$\mathbb{E}\big{[}\left\lVert g_{k}-g_{*}\right\rVert^{2}\big{]}$ $\leqslant\frac{C2^{k}(\log N)^{2}\log(Nd)}{N}+2\inf_{g\in\mathcal{G}}\bigg{(}\Big{(}(1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big{)}\left\lVert g-g_{*}\right\rVert^{2}$
-$\hskip 156.49014pt+(1+\delta^{-1})\bigg{(}\frac{2}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg{)},$
-where $C>0$ is a constant depending only on $\left\lVert g_{*}\right\rVert_{\infty}$ and $\sigma^{2}$.
-For a fixed sample size $N$, if we assume that $g_{*}\in\mathcal{G}$ and $k=3d\log_{2}N/(3d+2)$ is a multiple of $d$, Theorem 10 has the further consequence that (under the same assumptions)
-$\mathbb{E}\big{[}\left\lVert g_{k}-g_{*}\right\rVert^{2}\big{]}\leqslant CN^{-\frac{2}{3d+2}}\big{(}\left\lVert g_{*}\right\rVert_{\mathrm{TV}}+(\log N)^{2}\log(Nd)\big{)}.$ (29)
-The proof of Theorem 10 follows essentially the same path as Theorem 4.3 for CART of Klusowski and Tian (2024) while replacing Theorem 4.2 therein by our Theorem 9.
-3.4 Numerical experiments
-In this section, we provide experiments that verify the advantages of MinimaxSplit and cyclic MinimaxSplit methods in one- and two-dimensional domains, as well as integrated ensemble methods applied to approximating a real-valued function, and therefore it can be used to the application of denoising images.
-Further numerics (including those in high dimensions) are provided in Appendix B.
-3.4.1 One-dimensional input domain
-Figure 1: The comparison of the optimal tree, Scikit-learn DecisionTreeRegressor, our MinimaxSplit tree, and VarianceSplit tree on a sinusoidal target function $y=f(x)+\varepsilon=\sin(4\pi x)+\varepsilon$, $x\in[0,1],\varepsilon\sim N(0,\sigma_{\varepsilon}^{2})$ with low (first row, noise variance $\sigma_{\varepsilon}^{2}=0.01$) and high (second row, noise variance $\sigma_{\varepsilon}^{2}=0.25$) noise levels.
-We illustrate the true function $f$ using solid red curves and use the same training set for each of the low-noise and high-noise cases; the training set represented by the solid dots is one of the 10 different training sets of size 100.
-We display 10 different model fits based on 10 different batches of training sets, and the averaged mean-squared error evaluated on a different training set across 10 batches.
-As shown in Figure 1, our first experiment compares decision tree regression methods under varying noise conditions and offers notable insights into their relative performance and characteristics.
-The optimal tree is obtained using global optimization over all possible split values when conditioning on the tree structure and split coordinates (in our case, split coordinate can only be chosen along one coordinate since $x\in\mathbb{R}$).
-Specifically, for max_depth=3, it employs a fixed structure with $72^{3}-1=$ 7 split points, optimized simultaneously using sequential least squares programming to minimize the overall MSE on the training data.
-The topology of the tree is predetermined and enforced through constraints in the optimization process.
-The optimal tree approach contrasts with traditional variance-based, top-down decision tree algorithms by optimizing all splits at once, potentially capturing global data patterns more effectively.
-However, its fixed structure may limit adaptability to varying data complexities compared to conventional adaptive tree methods.
-The optimal tree serves as a baseline, particularly in high-noise scenarios, and suggests that its optimization criteria may be more effective at capturing the underlying function while minimizing overfitting.
-On the other hand, our MinimaxSplit’s competitive performance, especially under high noise, aligns with the theoretical strengths of the minimax optimization in decision trees.
-Its ability to maintain fidelity to the underlying function in noisy conditions underscores the value of robust splitting criteria in challenging environments.
-This observation may motivate further exploration of minimax-based approaches in domains with uncertain data quality or high noise levels.
-As another reference, the standard Scikit-learn DecisionTreeRegressor and the decision tree splitting based on variance minimization show similar performance patterns, with notable degradation in high-noise scenarios.
-This similarity is expected given their likely use of variance reduction as the splitting criterion.
-However, their performance lags behind that of MinimaxSplit in noisy conditions, highlighting the flexibility of MinimaxSplit in the presence of significant noise.
-The varying relative performance suggests that ensemble methods leveraging the strengths of different approaches could yield more robust and adaptive models across diverse data conditions, which we detail in Appendix B.
-In conclusion, this study provides valuable insight into the behavior of decision tree regression methods under varying noise conditions, highlighting the potential for improved performance through custom optimization criteria and robust splitting methods.
-3.4.2 Two-dimensional input domain
-For this experiment, we use a synthetic dataset generated from a complex, non-linear function of two features, $\mathbf{x}=(x_{1},x_{2})$.
-The function is designed to simulate realistic challenges encountered in regression tasks, such as non-linearity and interaction between features.
-Specifically, the true function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ used for generating the target values is defined as:
-$\begin{split}f(\mathbf{x})&=\left(2-2.1u_{1}^{2}+\frac{u_{1}^{4}}{3}\right)u_{1}^{2}+u_{1}u_{2}+\left(-4+4u_{2}^{2}\right)u_{2}^{2};\\ &u_{1}=3(x_{1}-0.5),\\ &u_{2}=3(x_{2}-0.5),\end{split}$ (30)
-where $x_{1},x_{2}$ are independently and uniformly sampled from $[0,1]$.
-This function incorporates both quadratic and higher-order polynomial terms, creating a complex surface with multiple peaks and valleys.
-The visualizations of the prediction surfaces, as shown in Figure 2, further illustrate the improved accuracy and smoother transitions in the predicted values when using the proposed methods.
-Figure 2: The predictions and true function values are visualized using heatmaps, allowing for a clear comparison of how well each model approximates the underlying function (30).
-The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models.
-The next three subplots depict the predictions of the $L^{1}$ VarianceSplit, $L^{1}$ MinimaxSplit, $L^{2}$ VarianceSplit, and $L^{2}$ MinimaxSplit.
-Top Row: max depth 2; Middle Row: max depth 6; Bottom Row: max depth 10.
-In this example (as well as in Sections 3.4.3 and 3.4.4 below), we also incorporate an $L^{1}$ variation of the VarianceSplit and MinimaxSplit algorithms, where in the splitting rule, we replace the variances in (11), (14), and (15) by the $L^{1}$ distance from the mean.
-For instance, the $L^{1}$ variation of (11) is
-$\begin{split}(j,x_{j})&=\operatorname*{arg\,min}_{(j,\,x_{j})}\Big{(}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathbb{E}\big{[}|Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}]|\mid\mathbf{X}\in A,\,X_{j}<x_{j}\big{]}\\ &\hskip 56.9055pt+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathbb{E}\big{[}|Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}]|\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\big{]}\Big{)}.\end{split}$ (31)
-The optimal $L^{1}$ split point may not be unique, in which case we break ties arbitrarily.
-3.4.3 Application to denoising images
-Another notable experiment involves applying decision trees to denoise noisy images (Luo et al., 2024).
-We apply the different tree variants to predict pixel values based on their locations, effectively treating this as a regression problem.
-The experiments include the following configurations:
-  1. 1.
-    VarianceSplit with $L^{1}$ (see (31)) and $L^{2}$ (see (11)) norms: These trees use either $L^{1}$ or $L^{2}$ norms to measure errors.
-Variance $L^{1}$ trees tend to capture more outlier noise, while $L^{2}$ norm-based trees provide smoother approximations.
-  2. 2.
-    MinimaxSplit and cyclic MinimaxSplit: These methods, especially when paired with cyclic padding, are effective in balancing feature splits in high-dimensional image data.
-The cyclic MinimaxSplit method shows robustness against feature dominance, which often occurred when a few features had much stronger correlations than others.
-The results show that mixed methods, where the error calculation alternates between minimax and variance-based splits, lead to a balance between local feature capturing and overall smoothness.
-Figure 3: The noiseless image, noisy image and denoised images using different kinds of splitting criteria for a single tree, allowing for a detailed comparison of how well each model approximates the geometric structure.
-The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models.
-The next four subplots depict the predictions of the VarianceSplit, Weighted VarianceSplit, MinimaxSplit, cyclic MinimaxSplit for $L^{1}$ (top row) and $L^{2}$ (bottom row) decision trees.
-All trees are fitted with a max depth of 10.
-The experiment shown in Figure 3 compares decision tree regression methods for denoising the classic Astronaut (Van der Walt et al., 2014) grayscale image, which has been preprocessed to 128 by 128.
-Subsequently, Gaussian noise with a standard deviation of 0.25 was added to create the noisy version.
-Examining the output plot reveals distinct characteristics of each method.
-The $L^{1}$ VarianceSplit approach (RMSE=0.140421) produces a piecewise constant approximation with visible rectangular artifacts, indicating a tendency to oversegment the image.
-The $L^{1}$ MinimaxSplit method (RMSE=0.114668) yields a smoother reconstruction with better preservation of large-scale features and edges, particularly evident in the feather of the subject’s outline.
-The $L^{2}$ VarianceSplit method (RMSE=0.138452) results in a slightly smoother output compared to its $L^{1}$ counterpart, but still exhibits noticeable blockiness.
-The $L^{2}$ MinimaxSplit approach (RMSE=0.113193) achieves the lowest RMSE, producing a reconstruction that balances smoothness with feature preservation, especially apparent in the gradual shading of the astronaut’s shoulder and the nuanced details around her head.
-The superior performance of the MinimaxSplit criteria, especially with the $L^{2}$ norm, can be attributed to its focus on minimizing worst-case errors.
-This strategy is particularly effective for image denoising, where it helps preserve important structural elements while smoothing out noise.
-Unlike traditional methods that select the best feature at each split based solely on immediate error reduction, the cyclic MinimaxSplit method systematically cycles through the available features as the tree grows.
-At different levels of the tree, different features are chosen for splitting, based on a pre-determined order or cyclic pattern.
-This ensures that all features are considered fairly and thus reduces the risk of the model becoming overly dependent on a small subset of features.
-The cyclic nature emphasizes on modeling each dimension of the underlying function while avoiding heterogeneity across dimensions, which can be particularly beneficial in multivariate cases where feature correlations are complex.
-3.4.4 From single tree to ensemble
-Although a single decision tree can capture patterns in data through hierarchical partitioning, it often suffers from high variance in predictions, which makes it sensitive to small changes in the training dataset.
-To address these limitations, an ensemble approach can be utilized to combine multiple decision trees, leading to improved stability and predictive accuracy.
-In this section, we extend the MinimaxSplit and cyclic MinimaxSplit methods to an ensemble context.
-The ensemble method we consider here is the random forest (Breiman et al., 1984), which aggregates multiple decision trees to improve the precision and robustness of the prediction.
-Each tree is grown to a fixed depth $k$, and different combinations of splitting criteria are used to determine how each subset of the data is partitioned.
-The ensemble predictions are then averaged to provide the final output (see Algorithm 1 in Appendix C).
-Specifically, we use the MinimaxSplit, cyclic MinimaxSplit, and VarianceSplit algorithms as the base learners in a random forest setting.
-Each tree in the ensemble is trained on a bootstrap sample of the original dataset and predictions are made by averaging the output of individual trees, effectively reducing the overall variance of the model.
-The model aims to maintain a balance between bias and variance by taking advantage of the strengths of each base model while averaging individual weaknesses.
-$L^{2}$ norm Refer to caption
-$L^{1}$ norm Refer to caption
-Figure 4: Predictions and true function values of (30), visualized using heatmaps.
-Top Row: The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models.
-The next three subplots depict the predictions of the variance-based decision tree, minimax decision tree, and cyclic minimax decision tree (depth $k=10$ with $L^{1}$ (top) and $L^{2}$ (bottom) norm), respectively.
-Each plot is annotated with the corresponding MSE to quantitatively assess the accuracy.
-Bottom Row: The final four subplots visualize the predictions from four different random forest models (using Algorithm 1 of Appendix C), each built with varying combinations of the three error methods.
-Again, each plot is labeled with the MSE to facilitate direct comparison.
-We apply the above ensemble approach in the same setting as in Section 3.4.2.
-The experiment in Figure 4 involves training multiple regression models (VarianceSplit, MinimaxSplit, and cyclic MinimaxSplit) and ensemble models on the generated dataset.
-The results demonstrate that the MinimaxSplit+MinimaxSplit ensemble outperforms the traditional variance-based tree in terms of MSE, particularly in scenarios where the data distribution is non-uniform.
-Further discussions on a weighted aggregation regime will be provided in Appendix B.3.
-4 Contribution and Future Work
-In this study, we introduced a novel decision tree splitting strategy, the MinimaxSplit algorithm, along with its multivariate variant, the cyclic MinimaxSplit algorithm.
-Unlike traditional VarianceSplit methods, which aim to reduce overall variance in the children nodes, our MinimaxSplit algorithm seeks to minimize the maximum variance within the split partitions, thereby reducing the risk of overly biased partitions.
-The cyclic MinimaxSplit algorithm further ensures that each dimension is used in a balanced manner throughout the tree construction process, avoiding dominance by a subset of features.
-A key theoretical result of this study is that the cyclic MinimaxSplit algorithm achieves an exponential MSE decay rate of given sufficient data samples, without requiring additional variance decay assumptions (Theorem 9).
-Furthermore, we derived empirical risk bounds for this method, establishing its robustness in different settings (Theorem 10).
-Along the same lines of techniques, we prove novel results on the convergence rates of univariate partition-based martingale approximations, which are of their own interest.
-In addition to single-tree regression analysis, we explored ensemble learning approaches, where we combined decision trees constructed using MinimaxSplit, cyclic MinimaxSplit, and VarianceSplit methods within a random forest framework.
-The results demonstrated that a hybrid approach leveraging different splitting techniques yields superior performance, particularly when dealing with non-uniform data distributions (Section 3.4.4).
-These findings suggest that the MinimaxSplit approach can provide more stable and adaptive decision trees compared to traditional methods, especially in high-dimensional settings.
-Future research could explore the MinimaxSplit algorithm in classification tasks, particularly for imbalanced datasets.
-We expect that an adaptive splitting strategy could replace cyclic selection, dynamically adjusting to feature importance.
-These directions may refine the theoretical foundations of MinimaxSplit and broaden its practical applications in high-dimensional and complex learning environments.
-Acknowledgment
-The authors thank Ruodu Wang for his valuable input.
-We provide our reproducible code at github.com/hrluo.
-HL was supported by U.S.
-Department of Energy under Contract DE-AC02-05CH11231 and U.S.
-National Science Foundation NSF-DMS 2412403.
-Appendix A Proofs and further results
-A.1 Proof of Theorem 3
-Variance martingale
+If the support of $U$ is a connected interval, Lemma A.1 is a special case of Theorem 1 of Ishwaran (2015) applied with $f$ monotone and $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$.
 Proof for the variance martingale.
-The first observation is that the split location is precisely the middle point of the two means of the two descendants.
-This is a special case of Theorem 1 of Ishwaran (2015) applied with $f$ monotone and $X\sim U(0,1)$, but for completeness, we provide a full argument here with our notation.
-To see this, consider (without loss of generality by shifting) the law of $U$ with mean zero, and let $u=u_{*}$ be the optimal split location.
-Let $p_{1},p_{2}$ and $m_{1},m_{2}$ be the associated one-step probabilities and locations in the binary tree representation.
-Since the second moment $\mathbb{E}[M_{1}^{2}]$ is maximized, we must have
-$0=\frac{\mathrm{d}}{\mathrm{d}u}\mathbb{E}[M_{1}^{2}]\mid_{u_{*}}=\frac{\mathrm{d}}{\mathrm{d}u}(p_{1}m_{1}^{2}+p_{2}m_{2}^{2})\mid_{u_{*}}.$ (32)
-Observe that
-$\frac{\mathrm{d}p_{1}}{\mathrm{d}u}=-\frac{\mathrm{d}p_{2}}{\mathrm{d}u}=\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}$
-and
-$\frac{\mathrm{d}m_{1}}{\mathrm{d}u}\mid_{u_{*}}=\frac{(u_{*}-m_{1})}{p_{1}}\,\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}\mid_{u_{*}};\leavevmode\nobreak\ \frac{\mathrm{d}m_{2}}{\mathrm{d}u}\mid_{u_{*}}=\frac{(m_{2}-u_{*})}{p_{2}}\,\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}\mid_{u_{*}}.$
-Inserting into (32) leads to
-$\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}\mid_{u_{*}}\Big{(}m_{1}^{2}-m_{2}^{2}+2m_{1}(u_{*}-m_{1})-2m_{2}(u_{*}-m_{2})\Big{)}=0,$
-which simplifies into $m_{1}+m_{2}=2u_{*}$ (in the case $\mathrm{d}\mathbb{P}_{U}/\mathrm{d}u(u_{*})=0$, we replace $u_{*}$ with the closest point of $u_{*}$ to $\operatorname{supp}U$ and this does not change the values of $p_{1},p_{2},m_{1},m_{2}$).
-This fact will be frequently used in the following.
-At level $k$, $\mathbb{E}[(M_{k-1}-M_{k})^{2}]$ is a sum of $2^{k}$ terms in the binary tree representation (10), which we may rewrite as $\sum_{j}p_{j}d_{j}^{2}$ using a change of variable.
-This corresponds to the $2^{k}$ edges (indexed by $[2^{k}]$) in the binary tree between levels $k-1$ and $k$, and $p_{j}$ represents the probabilities of the edge $j$ and $d_{j}$ is the location difference between the end vertices of the edge, or the length of the edge.
+Recall the binary subtree representation (17).
+Since $U$ is atomless, the binary subtree is in fact always a binary tree.
+At level $k$, $\mathbb{E}[(M_{k-1}-M_{k})^{2}]$ is a sum of $2^{k}$ terms, which we may rewrite as $\sum_{j}p_{j}d_{j}^{2}$ using a change of variable.
+This corresponds to the $2^{k}$ edges (indexed by $j\in[2^{k}]$) in the binary tree between levels k1k-1 and $k$, and $p_{j}$ represents the probabilities of the edge $j$ and $d_{j}$ is the location difference between the end vertices of the edge, or the length of the edge.
 Our goal is to bound from above
@@
 $\max_{1\leqslant j\leqslant 2^{k-1}}(p_{2j-1}d_{2j-1}^{2}+p_{2j}d_{2j}^{2}).$
-That is, the variance increases by splitting the $j$-th mass at level $k-1$ (which is the common ancestor of the $(2j-1)$-th and $(2j)$-th masses). By the martingale property,
+That is, the variance increases by splitting the $j$-th mass at level k1k-1 (which is the common ancestor of the (2j1)(2j-1)-th and $(2j)$-th masses). By the martingale property,
 $p_{2j-1}d_{2j-1}=p_{2j}d_{2j}.$
@@
 $p_{2j-1}d_{2j-1}^{2}+p_{2j}d_{2j}^{2}=(p_{2j-1}+p_{2j})d_{2j-1}d_{2j}.$ (37)
-The general strategy to analyze the quantity (33) is to start from an arbitrary mass/node $\mathbf{M}$ at level $k-1$, trace back its ancestors, and bound from above the probability of the mass (which is $p_{2j-1}+p_{2j}$ in the above expression) and the upper bounds for the length of the edges (which is $d_{2j-1},\,d_{2j}$).
-The path leading to a mass $\mathbf{M}$ at level $k-1$ in the binary tree representation (10) consists of $k$ nodes: $\emptyset=\mathbf{N}_{0},\mathbf{N}_{1},\dots,\mathbf{N}_{k-1}=\mathbf{M}$, where $\mathbf{N}_{j}\in V_{j}$.
+The general strategy to analyze the quantity (37) is to start from an arbitrary mass/node $\mathbf{M}$ at level k1k-1, trace back its ancestors, bound from above the probability of the mass (which is $p_{2j-1}+p_{2j}$ in the above expression), and establish upper bounds for the length of the edges (which is $d_{2j-1},\,d_{2j}$).
+The path leading to a mass $\mathbf{M}$ at level k1k-1 in the binary tree representation (17) consists of $k$ nodes: $\emptyset=\mathbf{N}_{0},\mathbf{N}_{1},\dots,\mathbf{N}_{k-1}=\mathbf{M}$, where $\mathbf{N}_{j}\in V_{j}$.
 We denote the lengths of the edges connecting the node $\mathbf{N}_{j-1}$ to its two descendants on level $j$ by $a_{j},b_{j},\,j\in[k-1]$, where $b_{j}=|\ell_{\mathbf{N}_{j-1}}-\ell_{\mathbf{N}_{j}}|$.
@@
 Let $a_{0}=b_{0}=1\geqslant\sup U-\inf U$.
-We consider the following algorithm, which starts from the tuple $(a_{0},b_{0},a_{1},b_{1},q_{1})$ and updates it $k-1$ times for each $j\in[k-1]$ when we trace the path down to $\mathbf{M}$.
+We consider the following algorithm, which starts from the tuple $(a_{0},b_{0},a_{1},b_{1},q_{1})$ and updates it k1k-1 times for each $j\in[k-1]$ when we trace the path down to $\mathbf{M}$.
 At step $j$, the tuple represents the two pairs $(a_{k},b_{k}),\,(a_{j},b_{j})$ that govern the maximum length of the edges of the $j$-th node (that is, the distance from its location to the boundaries of the intervals this node represents) for a certain $j\in[k-1]$, as well as the probability weight $q_{j}$ multiplied at this step.
@@
 We update $(a_{k},b_{k},a_{j},b_{j},q_{j})$ with $(a_{j},b_{j},a_{j+1},b_{j+1},q_{j+1})$, where $b_{j+1}\leqslant(a_{j}+b_{j})/2$, $a_{j+1}\leqslant(a_{k}+b_{k})/2$, and $q_{j+1}=a_{j+1}/(a_{j+1}+b_{j+1})$.
-Here, we have used (34) and the fact that $m_{1}+m_{2}=2x_{*}$ where $x_{*}$ is the split location.
-Let $(a,b,a^{\prime},b^{\prime},q)$ be generic entries of the tuple and $(\hat{a},\hat{b},\hat{a}^{\prime},\hat{b}^{\prime},\hat{q})$ be the updated tuple, where we recall $\hat{q}=\hat{a}^{\prime}/(\hat{a}^{\prime}+\hat{b}^{\prime})$.
-We claim that
+Here, we have used (38) and Lemma A.1.
+[…]
 $\frac{(\hat{a}+\hat{b})(\hat{a}^{\prime}+\hat{b}^{\prime})}{(a+b)(a^{\prime}+b^{\prime})}\leqslant\frac{1}{2\hat{q}}$. (39)
@@
     $\hat{a}^{\prime}+\hat{b}^{\prime}=a_{j+1}+b_{j+1}=\frac{a_{j+1}}{\hat{q}}\leqslant\frac{a_{k}+b_{k}}{2\hat{q}}=\frac{a+b}{2\hat{q}}$.
-This proves the claim.
+This proves (39).
 Recall that our goal is to bound (37).
 The probability of the mass at each step is multiplied by $q$, and the edge lengths are bounded by $(a+b)/2$ and $(a^{\prime}+b^{\prime})/2$.
-By the above claim, their product multiplies by at most $1/2$ each time we update the tuple.
+By (39), their product multiplies by at most $1/2$ each time we update the tuple.
 We have thus shown that
@@
 as desired. ∎
-Simons martingale
+Simons
 Proof for the Simons martingale.
-[…]
+It is noted in Zhang et al.
+(2024) that the Simons martingale satisfies the separated tree condition (i.e., the branches of the tree representation from all different levels, when projected onto the real line as intervals, are either disjoint or have a containment relationship), which is a consequence of the construction.
+We follow a similar argument to the variance martingale.
+Consider a node $\mathbf{M}$ at level k1k-1 and a path $(\emptyset=\mathbf{N}_{0},\dots,\mathbf{N}_{k-1}=\mathbf{M})$ leading to $\mathbf{M}$, where $\emptyset$ denotes the root of a binary tree.
+We provide an algorithm that recursively defines the tuples $(A_{j},B_{j},q_{j}),\leavevmode\nobreak\ 0\leqslant j\leqslant k-1$ for each node $\mathbf{N}_{j},\leavevmode\nobreak\ 0\leqslant j\leqslant k-1$ on that path, from the root to $\mathbf{M}$.
+The quantities $A_{j}$ and $B_{j}$ represent the maximum possible lengths of the two edges emanating from a node $\mathbf{N}_{j}$, and $q$ represents the one-step splitting probability leading to $\mathbf{N}_{j}$, based on the constraints from the separated tree property.
 We start from $(A_{1},B_{1},q_{1})=(1,1,1)$.
@@
 This completes the proof. ∎
-Minimax martingale
+Minimax
 Proof for the minimax martingale.
@@
 Note that $\phi$ is increasing and $\psi$ is decreasing in $u$, and both functions are continuous if we assume that $U$ is atomless.
-Recall by the minimax property that at each step we pick a $u\in\mathbb{R}$ such that $\max\{\phi(u),\psi(u)\}$ is minimized.
+Recall by the minimax property that at each step we pick $u\in\mathbb{R}$ such that $\max\{\phi(u),\psi(u)\}$ is minimized.
 This means that $\phi(u)=\psi(u)$.
@@
 as desired. ∎
-Median martingale
+Median
 Proof for the median martingale.
@@
 as desired. ∎
-A.2 Examples of optimal rates for partition-based martingale approximations
-In the following two examples, we show that the geometric rates in Theorem 3 are optimal for the Simons and median martingales.
-Example 11 (Rate $r=1/2$ is optimal for the Simons martingale).
-Fix an arbitrary $s\in(1/2,1)$ and let $U$ satisfy $\mathbb{P}(U=-1)=1-s$ and for $J\geqslant 0$,
-$\mathbb{P}\Big{(}U=\sum_{j=1}^{J}(\frac{1-s}{s})^{j}\Big{)}=s^{J+1}(1-s).$
-It follows that $U$ is a bounded random variable. Next, we compute that
-$\mathbb{E}\Big{[}U\mid U\geqslant\sum_{j=1}^{J}(\frac{1-s}{s})^{j}\Big{]}$ $=\bigg{(}\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\bigg{)}^{-1}\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\sum_{j=1}^{\ell}(\frac{1-s}{s})^{j}$
-$=s^{-(J+1)}\sum_{\ell=J}^{\infty}s^{\ell+1}(2s-1)(1+(\frac{1-s}{s})^{\ell})$
-$=\frac{s^{-(J+1)}(2s-1)s^{J+1}}{1-s}+\frac{s^{-(J+1)}(2s-1)s}{1-(1-s)}$
-$=\sum_{j=1}^{J+1}(\frac{1-s}{s})^{j}.$
-It is then straightforward to verify that in the Simons martingale $\{M_{k}\}_{k\geqslant 0}$, each $M_{k+1}-M_{k}$ is supported on the set $\{0,-((1-s)/s)^{k},((1-s)/s)^{k+1}\}$, where $\mathbb{P}(M_{k+1}-M_{k}=-((1-s)/s)^{k})=(1-s)s^{k}$ and $\mathbb{P}(M_{k+1}-M_{k}=((1-s)/s)^{k+1})=s^{k+1}$.
-It follows that
-$\mathbb{E}[(M_{k+1}-M_{k})^{2}]=(1-s)^{2k+1}s^{-(k+1)}\asymp(\frac{(1-s)^{2}}{s})^{k}.$
-In other words, the rate $r$ is given by $r=(1-s)^{2}/s$.
-If $s$ is picked close to $1/2$, $r$ can be made close enough to $1/2$.
-Therefore, $r=1/2$ is the optimal rate parameter.
-This example is illustrated with an atomic distribution $U$, but a slight twist also leads to an example for an absolutely continuous law $U$.
-Example 12 (Rate $r=1/2$ is optimal for the median martingale).
-Let $s\in(0,1)$ and a bounded random variable $U$ satisfy
-$\mathbb{P}\Big{(}U=\sum_{j=1}^{k-1}s^{j-1}-s^{k-1}\Big{)}=2^{-k},\leavevmode\nobreak\ k\geqslant 1.$
-Next, we compute that
-$\mathbb{E}\Big{[}U\mid U\geqslant\sum_{j=1}^{k}s^{j-1}-s^{k}\Big{]}$ $=2^{k}\sum_{\ell=k+1}^{\infty}2^{-\ell}\bigg{(}\sum_{j=1}^{\ell-1}s^{j-1}-s^{\ell-1}\bigg{)}$
-$=2^{k}\sum_{\ell=k+1}^{\infty}2^{-\ell}(\frac{1}{1-s}-\frac{s^{\ell-1}(2-s)}{1-s})$
-$=\frac{1}{1-s}-\frac{2-s}{s(1-s)}\frac{s^{k+1}/2}{1-s/2}$
-$=\sum_{j=1}^{k}s^{j-1}.$
-Note also that
-$\sum_{j=1}^{k}s^{j-1}-s^{k}<\sum_{j=1}^{k}s^{j-1}<\sum_{j=1}^{k+1}s^{j-1}-s^{k+1}.$
-It is then straightforward to verify that the law of $M_{k+1}-M_{k}$ is supported on $\{0,\pm s^{k}\}$ with $\mathbb{P}(M_{k+1}-M_{k}=s^{k})=\mathbb{P}(M_{k+1}-M_{k}=-s^{k})=2^{-(k+1)}$.
-We then obtain that
-$\mathbb{E}[(M_{k+1}-M_{k})^{2}]=2^{-k}s^{2k}.$
-In other words, the rate $r=s^{2}/2$ has a limit $1/2$ as $s\to 1$ from below; hence, the optimal rate is $1/2$.
-Again, a simple twist yields similar examples, where the law of $U$ is absolutely continuous.
-A.3 Proofs of Theorems 4 and 5
-Proof of Theorem 4.
-Let $\operatorname{supp}M_{1}=\{x,y\}$ where $x<y$.
-Take $M>\sup\{x,y\}$.
-Let $\tau_{M}^{+}$ be the first hitting time to $\{x:x\geqslant M\}$, $\tau_{M}^{-}$ be the first hitting time to $\{x:x\leqslant-M\}$, and $\tau_{M}=\min\{\tau_{M}^{+},\tau_{M}^{-}\}$.
-It is straightforward to prove (see Zhang et al.
-(2024)) that for some $r\in(0,1)$ and $C>0$,
-$\mathbb{E}[(U-M_{k})^{2}]\leqslant\mathbb{E}[{\mathbbm{1}}_{\{\tau_{M}<\infty\}}U^{2}]+CM^{2}r^{k}.$
-Since the supports of $\{M_{k}\}_{k\geqslant 0}$ correspond to the means of $U$ conditioned on nested partitions of $\mathbb{R}$, we have $\tau_{M}^{+}<\infty$ implies $M_{k}\geqslant x$ for all $k\geqslant 0$ and $M_{1}=y$.
-Therefore, conditional on the event $\tau_{M}^{+}<\infty$, $X$ is a martingale that is bounded from below by $x$.
-By Ville’s inequality,
-$\mathbb{P}(\tau_{M}^{+}<\infty)$ $\leqslant\mathbb{P}(\tau_{M}^{+}<\infty\mid M_{1}>x)$
-$=\mathbb{P}\left(\sup_{k\geqslant 0}M_{k}\geqslant M\mid M_{1}>x\right)\leqslant\frac{\mathbb{E}[U\mid M_{1}>x]-x}{M-x}=O(M^{-1}).$
-Similarly, the same analysis holds for $\tau_{M}^{-}$.
-Hence, we conclude that $\mathbb{P}(\tau_{M}<\infty)=O(M^{-1}).$ The rest follows line by line as in Zhang et al.
-(2024), which we sketch for completeness.
-By Hölder’s inequality, $\mathbb{E}[{\mathbbm{1}}_{\{\tau_{M}<\infty\}}U^{2}]=O(M^{-\delta^{\prime}})$ for some $\delta^{\prime}>0$.
-With the choices $M=r^{-k/(2+\delta^{\prime})}$ and $q=r^{\delta^{\prime}/(2+\delta^{\prime})}\in(0,1)$, we have that for some $C>0$,
-$\mathbb{E}[(U-M_{k})^{2}]\leqslant C(M^{-\delta^{\prime}}+M^{2}r^{k})\leqslant Cq^{-k}.$
-The final statement follows immediately from the martingale convergence theorem. ∎
-Proof of Theorem 5.
-First, we claim that the split points for the nested partitions $\{\pi_{k}\}$ are dense.
-Indeed, for both martingales, we have shown that the variances on each component in the partition $\pi_{k}$ go to zero uniformly as $k\to\infty$.
-If the split points were not dense, some element in the partition would cover an interval of positive length, on which the variance of $U$ cannot vanish since we assumed $\inf f>0$ and $\sup f<\infty$.
-Therefore, for any $\varepsilon>0$, there exists $k$ such that the $k$-th partition $\pi_{k}$ is such that on each interval $A\in\pi_{k}$, $\sup_{A}f\leqslant(1+\varepsilon)\inf_{A}f$.
-Since one-step variance is always better than one-step minimax, it remains to show that for an interval $A$ with $\sup_{A}f\leqslant(1+\varepsilon)\inf_{A}f$, there exists $C(\varepsilon)\downarrow 1/4$ as $\varepsilon>0$ such that at each step the total variance reduces by a factor of $C(\varepsilon)$ for the minimax martingale.
-To see this, without loss of generality we start from $U$ supported on $[0,1]$ (meaning that $\mathbb{P}(U\in[0,1])=1$) whose density takes values in $[1,1+\varepsilon]$ (we may normalize afterwards to a probability measure, which we omit).
-Then for any interval $[a,b]\subseteq[0,1]$,
-$\mathbb{E}[U\mid U\in[a,b]]\in\Big{[}a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)\Big{]}.$
-Therefore,
-$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])$ $\leqslant\sup_{m\in[a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)]}\int_{a}^{b}(1+\varepsilon)(x-m)^{2}\mathrm{d}x$
-$\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(b-a)^{3}.$
-Similarly,
-$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])\geqslant\inf_{m\in[a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)]}\int_{a}^{b}(x-m)^{2}\mathrm{d}x\geqslant\frac{1}{12}(b-a)^{3}.$
-It follows that the split point $x_{*}\in[0,1]$ must satisfy
-$\frac{1}{12}x_{*}^{3}\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(1-x_{*})^{3}\quad\text{and}\quad\frac{1}{12}(1-x_{*})^{3}\geqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}x_{*}^{3}.$
-Therefore, for some $a_{\varepsilon}\to 0$, $|x_{*}-1/2|\leqslant a_{\varepsilon}$.
-A direct computation yields $\mathrm{Var}(U)\geqslant\frac{1}{12}$. The total remaining variance $\mathbb{E}[(U-M_{1})^{2}]$ is bounded by
-$\mathbb{E}[(U-M_{1})^{2}]\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(x_{*}^{3}+(1-x_{*})^{3})\leqslant\frac{4(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(\frac{1}{2}+a_{\varepsilon})^{3}.$
-Hence, as $\varepsilon\to 0$, the ratio becomes
-$\frac{\mathbb{E}[(U-M_{1})^{2}]}{\mathrm{Var}(U)}\leqslant\frac{2(1+\varepsilon)^{4}}{(2+\varepsilon)^{3}}(1+2a_{\varepsilon})^{3}\to\frac{1}{4},$
-as desired. ∎
-A.4 Proofs of results from Section 3
-Lemma A.1.
-Let $\delta>0$. Then for any coupling $(X,Y)$ with $X\geqslant\mathbb{E}[Y]$, it holds that
+Appendix B Proofs of main results
+B.1 Proofs of results in Section 2
+Proof of Lemma 2.1.
+We take arbitrary $a<b$ and define events $E=E_{1}\cup E_{2}$, where $E_{1}=\{\mathbf{X}\in A,\,X_{j}<a\}$ and $E_{2}=E\setminus E_{1}=\{\mathbf{X}\in A,\,a\leqslant X_{j}<b\}.$ By the total variance formula,
+$\begin{split}\mathrm{Var}(Y\mid E)&=\mathbb{E}[\mathrm{Var}(Y\mid\sigma(E_{1},E_{2}))\mid E]+\mathrm{Var}(\mathbb{E}[Y\mid\sigma(E_{1},E_{2})]\mid E)\\ &\geqslant\mathbb{E}[\mathrm{Var}(Y\mid\sigma(E_{1},E_{2}))\mid E]\geqslant\mathbb{P}(E_{1}\mid E)\mathrm{Var}(Y\mid E_{1}).\end{split}$ (42)
+Observe that $\mathrm{Var}(Y\mid E)=\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<b)$ and $\mathrm{Var}(Y\mid E_{1})=\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<a)$.
+Inserting these two expressions back into (42), we obtain
+$\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<b)\geqslant\frac{\mathbb{P}(\mathbf{X}\in A,X_{j}<a)}{\mathbb{P}(\mathbf{X}\in A,X_{j}<b)}\mathrm{Var}(Y\mid\mathbf{X}\in A,X_{j}<a).$
+Rearranging gives
+$\mathbb{P}(\mathbf{X}\in A,\,X_{j}<a)\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<a)\leqslant\mathbb{P}(\mathbf{X}\in A,\,X_{j}<b)\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<b)$
+and therefore $\varphi_{L}$ is non-decreasing; and similarly $\varphi_{R}$ is non-increasing. The continuity of $\varphi_{L}$ and $\varphi_{R}$ follows from the marginally atomless condition. ∎
+We also need the following risk decomposition lemma.
+Lemma B.1.
+Let $\delta>0$. Then for any joint distribution $(X,Y)$ with $X\geqslant\mathbb{E}[Y]$, it holds that
 $\mathrm{Var}(Y)\leqslant(1+\delta)\,\mathbb{E}[(X-Y)^{2}]+\frac{(1+\delta)^{3}}{\delta^{3}}(\sup\operatorname{supp}X-\inf\operatorname{supp}X)^{2}.$ (43)
@@
 $\mathbb{E}\Big{[}(\sqrt{\delta}\,Y-\frac{1+\delta}{\sqrt{\delta}}X)^{2}\Big{]}+\Big{(}C(\sup\operatorname{supp}X-\inf\operatorname{supp}X)^{2}-\frac{1+\delta}{\delta}\mathbb{E}[X^{2}]\Big{)}\geqslant 0.$ (44)
-If $\sup\operatorname{supp}X/\inf\operatorname{supp}X\geqslant(1-\sqrt{(1+\delta)/(C\delta)})^{-1}$, the second term is always non-negative and hence (39) holds.
+If $\inf\operatorname{supp}X=0$ or $\sup\operatorname{supp}X/\inf\operatorname{supp}X\geqslant(1-\sqrt{(1+\delta)/(C\delta)})^{-1}$, the second term is always non-negative and hence (44) holds.
 Otherwise, since $C=(1+\delta)^{3}/\delta^{3}$,
-$(1+\delta)(\inf\operatorname{supp}X)^{2}\geqslant(1+\delta)(1-\sqrt{(1+\delta)/(C\delta)})(\sup\operatorname{supp}X)^{2}\geqslant(\sup\operatorname{supp}X)^{2}.$
+$(1+\delta)(\inf\operatorname{supp}X)^{2}\geqslant(1+\delta)(1-\sqrt{(1+\delta)/(C\delta)})(\sup\operatorname{supp}X)^{2}=(\sup\operatorname{supp}X)^{2}.$
 It follows that
@@
 It remains to establish (45).
-Denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ from the cyclic minimax construction.
-Note that by construction, #$\#J_{k}\leqslant 2^{k}$ and $M_{k}\mid\mathbf{X}\in I_{j}$ is a piecewise constant random variable, which takes a constant value $y_{k,j}$ inside the interval $I_{j}$, so $M_{k}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}=y_{k,j}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}$.
+Denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ by the cyclic minimax construction.
+Note that by construction, $J_{k}\leqslant 2^{k}$ and $M_{k}\mid\mathbf{X}\in I_{j}$ is a piecewise constant random variable, which takes a constant value $y_{k,j}$ inside the box $I_{j}$, so $M_{k}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}=y_{k,j}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}$.
 By definition, $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$.
@@
 Define $\mathrm{Ran}(g_{j})=[\inf g_{j},\sup g_{j}]$ and recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
-The key step is to decompose the local mean-squared error $\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ into two parts for each $j$, with the help of Lemma A.1.
+The key step is to decompose the local mean squared error $\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ into two parts for each $j$, with the help of Lemma B.1.
 Next, we claim that
@@
   1. (i)
-    $y_{k,j}\in\mathrm{Ran}(g_{j})$.
-Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and $V_{j}=(1+\delta^{-1})\,\mathbb{E}[(g(\mathbf{X})-y_{k,j})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$.
+    $y_{k,j}\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and
+    $V_{j}=\min\big{\{}(1+\delta^{-1})\,\mathbb{E}[(g(\mathbf{X})-y_{k,j})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}],\,\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\big{\}}.$
   2. (ii)
-    $y_{k,j}\not\in\mathrm{Ran}(g_{j})$.
-Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and $V_{j}=p_{j}(\Delta g_{j})^{2}(1+\delta)^{3}/\delta^{3}$.
+    $y_{k,j}\not\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and
+    $V_{j}=\min\Big{\{}\frac{(1+\delta)^{3}}{\delta^{3}}\,p_{j}(\Delta g_{j})^{2},\,\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\Big{\}}.$
 In case (i), (48) follows immediately from the polarization identity $(a+b)^{2}\leqslant(1+\delta)a^{2}+(1+\delta^{-1})b^{2}$.
-On the other hand, for case (ii), (43) follows by applying Lemma A.1 to the coupling $(g(\mathbf{X}),Y)\mid\mathbf{X}\in I_{j}$ and using that $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$.
+On the other hand, for case (ii), (48) follows by applying Lemma B.1 to the joint distribution $(g(\mathbf{X}),Y)\mid\mathbf{X}\in I_{j}$ and using that $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$.
 This completes the proof of (48).
@@
 $\sum_{j\in J_{k}}U_{j}=\sum_{j\in J_{k}}(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}].$ (50)
-To control $\sum_{j}V_{j}$, we apply the same Hölder’s inequality argument as in the proof of Theorem 3 in the minimax case. We have
+To control $\sum_{j}V_{j}$, we apply Hölder’s inequality. We have
 $\begin{split}\sum_{j\in J_{k}}V_{j}&\leqslant\Big{(}\sum_{j\in J_{k}}p_{j}\big{(}\frac{V_{j}}{p_{j}}\big{)}^{3/2}\Big{)}^{2/3}\Big{(}\sum_{j\in J_{k}}p_{j}\Big{)}^{1/3}\\ &\leqslant\Bigg{(}(\max_{j\in J_{k}}V_{j})\sum_{j\in J_{k}}\sqrt{\frac{V_{j}}{p_{j}}}\Bigg{)}^{2/3}\leqslant\Bigg{(}(\max_{j\in J_{k}}\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}])\sum_{j\in J_{k}}\sqrt{\frac{V_{j}}{p_{j}}}\Bigg{)}^{2/3}.\end{split}$ (51)
-To further bound the right-hand side of (46), we recall from (21) that as a consequence of the marginally atomless property,
+To further bound the right-hand side of (51), we first recall from (11) that as a consequence of the marginally atomless property,
 $\max_{j\in J_{k}}\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\leqslant 2^{-k}\mathrm{Var}(Y).$ (52)
@@
 $\sqrt{\frac{V_{j}}{p_{j}}}\leqslant(1+\delta^{-1})^{3/2}\Delta g_{j}.$ (53)
-This is immediate for case (ii) above (when $y_{k,j}\not\in\mathrm{Ran}(g_{j})$).
-By definition and using $y_{k,j}\in\mathrm{Ran}(g_{j})$, we have for case (i) above (when $y_{k,j}\in\mathrm{Ran}(g_{j})$),
+This is immediate for case (ii) above (when $y_{k,j}\not\in\mathrm{Ran}(g_{j})$). By definition, we have for case (i) above (when $y_{k,j}\in\mathrm{Ran}(g_{j})$),
 $V_{j}\leqslant(1+\delta^{-1})p_{j}(\Delta g_{j})^{2}<p_{j}(1+\delta^{-1})^{3}(\Delta g_{j})^{2},$
@@
 $\sum_{j\in J_{k}}\Delta g_{j}=\sum_{A\in\pi_{k}}\Big{(}\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\Big{)}\leqslant 2^{k-\lfloor k/d\rfloor}\left\lVert g\right\rVert_{\mathrm{TV}},$ (54)
-[…]
+where $\lfloor\cdot\rfloor$ is the floor function. Combined with (53), we arrive at
 $\sum_{j\in J_{k}}\sqrt{\frac{V_{j}}{p_{j}}}\leqslant(1+\delta^{-1})^{3/2}2^{k-\lfloor k/d\rfloor}\left\lVert g\right\rVert_{\mathrm{TV}}.$
@@
 Proof of Remark 4.
-The bound (41) can be improved to
+In the proof of Theorem 2, the bound (46) can be improved to
 $\mathrm{Var}(Y)^{1/2}\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\frac{1}{2}\Delta g.$
@@
 Proof of Theorem 6.
-The only difference from the proof of Theorem 6 is (47), since for non-marginally atomless measures (20) and (21) may not hold. Instead, by our assumptions, (20) is be replaced by
-$\begin{split}&\hskip 14.22636pt\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}\}}],\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}\}}]\Big{\}}\\ &\leqslant\frac{1}{2}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]+\frac{M^{2}}{N},\end{split}$
-because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $M^{2}/N$. As a consequence, if we denote by
-$u_{k}:=\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}],$
+Recall that $\Delta Y:=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$.
+The only difference from the proof of Theorem 2 is (52), since for non-marginally atomless measures (10) and (11) may not hold.
+Instead, by our assumptions, (10) is be replaced by
+$\max\Big{\{}\varphi_{L}(\hat{x}_{j}),\varphi_{R}(\hat{x}_{j})\Big{\}}\leqslant\frac{1}{2}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]+\frac{(\Delta Y)^{2}}{N},$
+because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $(\Delta Y)^{2}/N$. As a consequence, if we denote by
+$u_{k}:=\max_{A\in\pi_{k}}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}],$
 then $u_{k}$ satisfies the recursive inequalities
-$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{M^{2}}{N},\leavevmode\nobreak\ k\geqslant 0;\quad u_{0}=\mathrm{Var}(Y).$
-To solve this, let $v_{k}:=u_{k}-2M^{2}/N$.
-Suppose that $u_{0}\geqslant 2M^{2}/N$.
+$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{(\Delta Y)^{2}}{N},\leavevmode\nobreak\ k\geqslant 0;\quad u_{0}=\mathrm{Var}(Y).$
+To solve this, let $v_{k}:=u_{k}-2(\Delta Y)^{2}/N$.
+Suppose that $u_{0}\geqslant 2(\Delta Y)^{2}/N$.
 Then $v_{k}$ satisfies $v_{k+1}\leqslant v_{k}/2$ and $v_{0}\leqslant\mathrm{Var}(Y)$.
-It follows that $v_{k}\leqslant 2^{-k}\mathrm{Var}(Y)$ for all $k\geqslant 0$, and hence
-$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]=u_{k}\leqslant 2^{-k}\mathrm{Var}(Y)+\frac{2M^{2}}{N}.$ (51)
-It is also easy to check that (51) is also satisfied for the case $u_{0}<2M^{2}/N$, that is, (51) holds in general. Replacing (47) by (51) leads to
-$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\bigg{(}(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\Big{(}\left\lVert g\right\rVert_{\mathrm{TV}}(\mathrm{Var}(Y)+2^{k+1}\frac{M^{2}}{N})\Big{)}^{2/3}\bigg{)}.$
+It follows that $v_{k}\leqslant 2^{-k}\mathrm{Var}(Y)$ for all $k\geqslant 0$, and hence the maximum risk among the nodes at level $k$ is
+$\max_{A\in\pi_{k}}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]=u_{k}\leqslant 2^{-k}\mathrm{Var}(Y)+\frac{2(\Delta Y)^{2}}{N}.$ (56)
+It is also easy to check that (56) is also satisfied for the case $u_{0}<2(\Delta Y)^{2}/N$, that is, (56) holds in general. Replacing (52) by (56) leads to
+$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\bigg{(}(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\Big{(}\left\lVert g\right\rVert_{\mathrm{TV}}(\mathrm{Var}(Y)+2^{k+1}\frac{(\Delta Y)^{2}}{N})\Big{)}^{2/3}\bigg{)}.$
 Applying concavity of the function $x\mapsto x^{2/3}$ for $x>0$, (47), and that
-$\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\Big{(}2^{k+1}\frac{M^{2}}{N}\Big{)}^{2/3}\leqslant\frac{\left\lVert g\right\rVert_{\mathrm{TV}}^{2}}{3}+\frac{2^{k+2}}{3}\frac{M^{2}}{N},$
+$\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\Big{(}2^{k+1}\frac{(\Delta Y)^{2}}{N}\Big{)}^{2/3}\leqslant\frac{\left\lVert g\right\rVert_{\mathrm{TV}}^{2}}{3}+\frac{2^{k+2}}{3}\frac{(\Delta Y)^{2}}{N},$
 we obtain (23). ∎
 Proof of Theorem 7.
-Note that if $\mathbf{X}_{*}$ is marginally atomless, the assumption (27) for the coupling $(\mathbf{X},Y)$ in Theorem 9 holds $\mathbb{P}_{*}$-almost surely.
+Note that if $\mathbf{X}_{*}$ is marginally atomless, the assumption (22) for the joint distribution $(\mathbf{X},Y)$ in Theorem 6 holds $\mathbb{P}_{*}$-almost surely.
 Consequently, the proof is almost verbatim compared to Theorem 4.3 of Klusowski and Tian (2024), by applying our Theorem 6 instead of Theorem 4.2 therein.
 The only difference here is the extra term
-$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}$
-appearing in (28), where we remind the reader that $M=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$.
-Nevertheless, since the proof of Theorem 4.3 of Klusowski and Tian (2024) deals first with the case with bounded data, we may simply replace $M$ by that bound.
+$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}(\Delta Y)^{2}}{3N}$
+appearing in (23), where we remind the reader that $\Delta Y=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$.
+Nevertheless, since the proof of Theorem 4.3 of Klusowski and Tian (2024) deals first with the case with bounded data, we may simply replace $\Delta Y$ by that bound.
 Specifically, the first part of the proof of Theorem 4.3 of Klusowski and Tian (2024) assumes $\max_{i}|Y_{i}|\leqslant U$.
-We may set $M=2U$ and add the term
-$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}\leqslant\frac{2^{k+5}U^{2}}{3N}$
+We may set $\Delta Y=2U$ and add the term
+$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}(\Delta Y)^{2}}{3N}\leqslant\frac{2^{k+5}U^{2}}{3N}$
 in (B.28) therein (where we used $\delta\geqslant 2^{-2\lfloor k/d\rfloor/3}$).
@@
 ∎
-Appendix B Further numerics
-B.1 Convergence rates of partition-based martingale approximations
-The experiments in Figure 5 are designed to compare the performance of four different partition-based martingale approximation methods defined in Definition 1—minimax, median, Simons, and variance—based on their ability to approximate a target function using nested partitions.
-The methods are applied to different probability densities over a specified interval, and the results are visualized in a series of plots.
-We want to evaluate how differently martingale approximation methods reduce the variance (or the second moment) of a random variable $Y$ that is distributed according to various density functions $f$.
-In Figure 5(a), we consider the density $f$ of $Y$ given by
-$f(x)=\begin{cases}1/51&\text{if }x\in[0,0.9];\\ (1+10^{4}(x-0.9))/51&\text{if }x\in[0.9,1];\\ 0&\text{otherwise,}\end{cases}$ (52)
-with four types of partition-based martingale approximations.
-The minimax method generally shows a steeper decline in MSE in the initial stages, suggesting that it is effective at quickly reducing the worst-case variance.
-(a) Density given by (52)
-Refer to captionRefer to caption
-(b) Density given by $f(x)=11x^{10}$ on $[0,1]$
-Figure 5: Plots of the log-MSE, $\log_{2}(\mathbb{E}[(Y-M_{k})^{2}])$ versus the approximation depth $k$ for four different methods, where the density of the law $Y$ is given: (a) by (52) and (b) by $f(x)=11x^{10}$ on $[0,1]$ and zero otherwise.
-The right panel of (b) plots the ratios $\mathbb{E}[(Y-M_{k+1})^{2}]/\mathbb{E}[(Y-M_{k})^{2}]$ for the four methods as a function of $k$.
-Furthermore, the second plot in Figure 5(b) shows the ratio of the remaining risk after each step normalized by the variance of the previous step, as an indicator of rates of improvement as the tree grows deeper.
-The behavior of these curves can be used to infer the asymptotic rates of convergence for each method, which align with the theoretical analysis of Section 2.3.
-B.2 Higher dimensional input domain
-Consider a dimension $d\geqslant 1$ that is a multiple of $4$. The Powell function on $n$,
-$f(x)=\sum_{i=1}^{d/4}[(x_{4i-3}+10x_{4i-2})^{2}+5(x_{4i-1}-x_{4i})^{2}+(x_{4i-2}-2x_{4i-1})^{4}+10(x_{4i-3}-x_{4i})^{4}],$
-with uniformly random samples on $[0,1]^{d}$, serves as an exemplary testbed for decision tree methods in high-dimensional regression problems, offering a nuanced landscape of challenges that mirror real-world complexities.
-In the dense design scenario ($n>d>100$), the function’s inherent non-linearity and intricate variable interactions push the boundaries of decision trees’ capabilities.
+B.2 Proofs of results in Section 3
+Before we formally prove Theorem 9, we point out a key ingredient that distinguishes it from previous proofs where only a single decision tree is involved.
+Recall that $n$ is the number of trees in the forest, $\Sigma$ is the law of the splitting dimensions, and $\mathbb{E}_{n}$ denotes the expectation with respect to the empirical measure on $n$ samples from $\Sigma$.
+In Theorem 2, the cyclicity of the cyclic MinimaxSplit algorithm is essential to proving (54).
+However, a stronger upper bound for
+$\sum_{j\in J_{k}}\Delta g_{j}=\sum_{A\in\pi_{k}}\Big{(}\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\Big{)}$
+exists by counting the (random) number of splits in each dimension, which is generally of the form
+$\sum_{A\in\pi_{k}}\Big{(}\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\Big{)}\leqslant\sum_{i=1}^{d}\xi_{k,i}\left\lVert g_{i}\right\rVert_{\mathrm{TV}},$ (57)
+for some (random) positive integers $\xi_{k,i},\leavevmode\nobreak\ i\in[d]$ depending on the samples from $\Sigma$.
+See also (61) below.
+The intuition in proving Theorem 9 is that we keep (57) as an upper bound, instead of (54).
+Then, with enough randomness injected into the splitting dimensions (and hence the random variables $\{\xi_{k,i}\}$), the right-hand side of (57) is small if averaged over all trees in the forest, as shown by the next lemma.
+Lemma B.2.
+Fix $d\geqslant 2$.
+Consider a stochastic recursion defined as follows.
+Let $(\xi_{0,1},\dots,\xi_{0,d})=(1,\dots,1)$ and for each $k\geqslant 1$, given $(\xi_{k-1,1},\dots,\xi_{k-1,d})$, we define $(\xi_{k,1},\dots,\xi_{k,d})$ through the following procedure:
+  • sample independent copies $({\xi}^{\prime}_{k-1,1},\dots,{\xi}^{\prime}_{k-1,d})$ and $({\xi}^{\prime\prime}_{k-1,1},\dots,{\xi}^{\prime\prime}_{k-1,d})$ of $(\xi_{k-1,1},\dots,\xi_{k-1,d})$;
+  • uniformly sample $\zeta_{k}\in[d]$;
+  • define $\xi_{k,\zeta_{k}}=\max\{\xi^{\prime}_{k-1,\zeta_{k}},{\xi}^{\prime\prime}_{k-1,\zeta_{k}}\}$;
+  • for $j\in[d]$ and $j\neq\zeta_{k}$, define $\xi_{k,j}=\xi^{\prime}_{k-1,j}+{\xi}^{\prime\prime}_{k-1,j}$.
+Then $\mathbb{E}[\xi_{k,j}]\leqslant 2^{k}e^{-k/(4d)}$ for all $j\in[d]$.
+Remark 22.
+The rate $e^{-1/(4d)}$ is not tight. Finding the tight exponent appears to be a non-trivial task, even for $d=2$.
+Proof of Lemma B.2.
+By construction, for each fixed $k\geqslant 0$, the law of each vector $(\xi_{k,1},\dots,\xi_{k,d})$ is exchangeable.
+As a consequence, $\xi_{k,j}\overset{\mathrm{law}}{=}W_{k}$ for all $j\in[d]$, where $\{W_{k}\}_{k\geqslant 0}$ is defined as follows: $W_{0}=1$, and for each $k\geqslant 1$, we take independent copies ${W}^{\prime}_{k-1},W^{\prime\prime}_{k-1}$ of $W_{k-1}$ and define
+$W_{k}=\begin{cases}{W}^{\prime}_{k-1}+{W}^{\prime\prime}_{k-1}&\text{ with probability }1-1/d;\\ \max\{{W}^{\prime}_{k-1},{W}^{\prime\prime}_{k-1}\}&\text{ with probability }1/d.\end{cases}$
+It remains to find $\mathbb{E}[W_{k}]$. It follows from construction that $\mathbb{E}[(W_{0})^{2}]=1$ and for each $k\geqslant 1$,
+$\mathbb{E}[(W_{k})^{2}]$ $=(1-\frac{1}{d})\mathbb{E}\Big{[}\Big{(}{W}^{\prime}_{k-1}+{W}^{\prime\prime}_{k-1}\Big{)}^{2}\Big{]}+\frac{1}{d}\mathbb{E}\Big{[}\Big{(}\max\{{W}^{\prime}_{k-1},{W}^{\prime\prime}_{k-1}\}\Big{)}^{2}\Big{]}$
+$\leqslant 4(1-\frac{1}{d})\mathbb{E}[(W_{k-1})^{2}]+\frac{2}{d}\mathbb{E}[(W_{k-1})^{2}]$
+$=(4-\frac{2}{d})\mathbb{E}[(W_{k-1})^{2}].$
+where we have used the inequality $(a+b)^{2}\leqslant 2(a^{2}+b^{2})$. It follows that
+$\mathbb{E}[W_{k}]\leqslant\sqrt{\mathbb{E}[(W_{k})^{2}]}\leqslant(4-\frac{2}{d})^{k/2}\leqslant 2^{k}e^{-k/(4d)},$
+as desired. ∎
+Proof of Theorem 9.
+In the following, we use $S$ to denote a random variable with law $\Sigma$. We first establish an upper bound of
+$\mathbb{E}[(Y-\mathbb{E}_{n}[M_{k}^{S}])^{2}].$
+Since the randomness from $\mathbb{E}_{n}$ is independent from that of $(\mathbf{X},Y)$, we have by Jensen’s inequality,
+$\mathbb{E}[(Y-\mathbb{E}_{n}[M_{k}^{S}])^{2}]\leqslant\mathbb{E}[\mathbb{E}_{n}[\mathbb{E}[(Y-M_{k}^{S})^{2}]]].$ (58)
+Here, the outer expectation is with respect to the randomness from the empirical measure $\mathbb{E}_{n}$ (the law of $(\mathbf{X},Y)$ is fixed in the settings of Theorems 2 and 6), and the inner expectation is with respect to the joint law of $(\mathbf{X},Y,\{M_{k}^{S}\})$.
+In other words, we trivially bound the risk of the average by the average of the risks among the $n$ trees.
+Denote the empirical samples from $\Sigma$ by $S_{1},\dots,S_{n}$. Our next goal is to control
+$\mathbb{E}_{n}[\mathbb{E}[(Y-M_{k}^{S})^{2}]]=\frac{1}{n}\sum_{\ell=1}^{n}\mathbb{E}[(Y-M_{k}^{S_{\ell}})^{2}].$
+We apply the same decomposition as in step II of the proof of Theorem 2, with the specific choice of $\delta=1$.
+For $j\in J_{k}$, we let $g_{j}^{\ell}=g|_{I_{j}^{\ell}}$, where $\{I_{j}^{\ell}\}_{j\in J_{k}}$ is the partition at level $k$ of a tree whose splitting rule is given by $S_{\ell}$.
+Then (53) now has the form
+$\sqrt{\frac{V_{j}^{\ell}}{p_{j}^{\ell}}}\leqslant 2^{3/2}\Delta g_{j}^{\ell}.$
+Proceeding as in step III of the proof of Theorem 2, (55) becomes
+$\sum_{j\in J_{k}^{\ell}}V_{j}^{\ell}\leqslant\Big{(}2^{-k}\mathrm{Var}(Y)2^{3/2}\sum_{j\in J_{k}^{\ell}}\Delta g_{j}^{\ell}\Big{)}^{2/3},$ (59)
+and hence for all $g\in\mathcal{G}$ with the representation $g=g_{1}+\dots+g_{d}$,
+$\mathbb{E}_{n}[\mathbb{E}[(Y-M_{k}^{S})^{2}]]\leqslant 2\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+2\mathrm{Var}(Y)^{2/3}2^{-2k/3}\frac{1}{n}\sum_{\ell=1}^{n}\Big{(}\sum_{j\in J_{k}^{\ell}}\Delta g_{j}^{\ell}\Big{)}^{2/3}.$ (60)
+Next, we construct the (random) values $\{\xi_{k,i}^{\ell}\}_{\ell\in[n],\,i\in[d]}$ such that for each $k\in\mathbb{N}$ and $\ell\in[n]$,
+$\sum_{j\in J_{k}^{\ell}}\Delta g_{j}^{\ell}\preceq_{\mathrm{st}}\sum_{i=1}^{d}\xi_{k,i}^{\ell}\left\lVert g_{i}\right\rVert_{\mathrm{TV}},$ (61)
+where $\preceq_{\mathrm{st}}$ denotes stochastic dominance. Define $\{\xi_{k,i}^{\ell}\}_{\ell\in[n],\,i\in[d]}$ as follows:
+  • $\{\xi_{k,i}^{\ell}\}_{i\in[d]},\,\ell\in[n]$ are i.i.d. copies of $\{\xi_{k,i}\}_{i\in[d]}$.
+  • $\xi_{0,i}=1$ for all $i\in[d]$.
+  • For $k\geqslant 1$, given the law of $\{\xi_{k-1,i}\}_{i\in[d]}$, generate two i.i.d.
+copies $\{\xi^{\prime}_{k-1,i}\}_{i\in[d]}$ and $\{\xi^{\prime\prime}_{k-1,i}\}_{i\in[d]}$, independently sample a uniform $\zeta_{k}\in[d]$, and define
+    $\xi_{k,\zeta_{k}}=\max\{\xi^{\prime}_{k-1,\zeta_{k}},\xi^{\prime\prime}_{k-1,\zeta_{k}}\};\quad\xi_{k,i}=\xi^{\prime}_{k-1,i}+\xi^{\prime\prime}_{k-1,i},\leavevmode\nobreak\ i\neq\zeta_{k}.$
+See Figure 11 for an illustration of the first few values of $\{\xi_{k,i}\}$.
+Figure 11: Showcasing the possibilities of the splits for levels $k=0,1,2$, with $d=2$.
+To see that this definition yields (61), we apply induction.
+The base case $k=0$ is easy as $\sup g-\inf g\leqslant\sum_{i=1}^{d}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}$.
+Suppose that (61) holds with k1k-1.
+Then given a tree of depth $k$ (where we recall that the split dimensions of all nodes are i.i.d.
+uniform in $[d]$), we condition on the split dimension at level zero (node $\emptyset$, say into child nodes $u,v$), which we denote by $\zeta_{k}$.
+Next, we decompose each variation $\Delta g_{j}^{\ell}$ in the $d$ dimensions by $\Delta g_{j}^{\ell}\leqslant\sum_{i=1}^{d}\Delta g_{j,i}^{\ell}$, where $\Delta g_{j,i}^{\ell}$ is the variation of the $i$-th coordinate of $g$ in $I_{j}$ for sample $\ell$.
+  • For a dimension $i\neq\zeta_{k}$, we apply the trivial bound that the variations over dimension $i$ for $j\in J_{k}^{\ell}$ is at most the sum of the variations among the descendants of $u$ and of $v$.
+By induction hypothesis, we have
+    $\sum_{j\in J_{k}}\Delta g_{j,i}^{\ell}$ $\leqslant\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ u\end{subarray}}\Delta g_{j,i}^{\ell}+\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ v\end{subarray}}\Delta g_{j,i}^{\ell}$
+    $\preceq_{\mathrm{st}}\xi^{\prime}_{k-1,i}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}+\xi^{\prime\prime}_{k-1,i}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}=\xi_{k,i}\left\lVert g_{i}\right\rVert_{\mathrm{TV}},$
+    where $j\succ u$ means that the set $I_{j}$ is a descendant of $u$ in the binary subtree representation.
+  • For dimension $\zeta_{k}$, we again use the fact that the variation $\sum_{j\in J_{k}}\Delta g_{j,\zeta_{k}}^{\ell}$ is at most the sum of variations among the descendants of $u,v$, but the total variation $\left\lVert g_{\zeta_{k}}\right\rVert_{\mathrm{TV}}$ is already decomposed into two, say $\left\lVert g_{\zeta_{k}}^{(i)}\right\rVert_{\mathrm{TV}},\,i=1,2$, corresponding to $u,v$.
+Therefore, by inductive hypothesis,
+    $\sum_{j\in J_{k}}\Delta g_{j,\zeta_{k}}^{\ell}$ $\leqslant\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ u\end{subarray}}\Delta g_{j,\zeta_{k}}^{\ell}+\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ v\end{subarray}}\Delta g_{j,\zeta_{k}}^{\ell}$
+    $\preceq_{\mathrm{st}}\xi^{\prime}_{k-1,{\zeta_{k}}}\left\lVert g_{\zeta_{k}}^{(1)}\right\rVert_{\mathrm{TV}}+\xi^{\prime\prime}_{k-1,{\zeta_{k}}}\left\lVert g_{\zeta_{k}}^{(2)}\right\rVert_{\mathrm{TV}}$
+    $\leqslant\max\{\xi^{\prime}_{k-1,{\zeta_{k}}},\xi^{\prime\prime}_{k-1,{\zeta_{k}}}\}\left\lVert g_{\zeta_{k}}\right\rVert_{\mathrm{TV}}$
+    $=\xi_{k,\zeta_{k}}\left\lVert g_{\zeta_{k}}\right\rVert_{\mathrm{TV}}.$
+This proves (61).
+Applying (61) and Hölder’s inequality, we obtain
+$\frac{1}{n}\sum_{\ell=1}^{n}\Big{(}\sum_{j\in J_{k}^{\ell}}\Delta g_{j}^{\ell}\Big{)}^{2/3}$ $\preceq_{\mathrm{st}}\frac{1}{n}\sum_{\ell=1}^{n}\Big{(}\sum_{i=1}^{d}\xi_{k,i}^{\ell}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}\Big{)}^{2/3}$
+$\leqslant 2^{2k/3}\Big{(}\frac{1}{n}\sum_{\ell=1}^{n}\sum_{i=1}^{d}2^{-k}\xi_{k,i}^{\ell}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}\Big{)}^{2/3}$
+$\leqslant 2^{2k/3}\Big{(}\frac{1}{n}\sum_{i=1}^{d}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}\max_{i\in[d]}\sum_{\ell=1}^{n}2^{-k}\xi_{k,i}^{\ell}\Big{)}^{2/3}$
+$\leqslant 2^{2k/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\Big{(}\frac{1}{n}\max_{i\in[d]}\sum_{\ell=1}^{n}2^{-k}\xi_{k,i}^{\ell}\Big{)}^{2/3},$
+where in the last step we recall that for $g\in\mathcal{G}$, $\left\lVert g\right\rVert_{\mathrm{TV}}=\sum_{i=1}^{d}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}$.
+The next step is to bound from above the inner maximum, where we recall that under the outer expectation (as $S_{1},\dots,S_{n}$ are i.i.d.
+sampled from $\Sigma$), the random variables $\{\xi_{k,i}^{\ell}\}_{\ell\in[n],i\in[d]}$ are i.i.d.
+in $\ell$.
+Employing Lemma B.2, we have for each $i\in[d]$ that
+$\mathbb{E}\bigg{[}\sum_{\ell=1}^{n}2^{-k}\xi_{k,i}^{\ell}\bigg{]}$ $=n2^{-k}\mathbb{E}[\xi_{k,1}^{\ell}]\leqslant ne^{-\frac{k}{4d}}.$
+By the union bound and the Hoeffding inequality (note that $2^{-k}\xi_{k,i}^{\ell}\in[0,1]$), we have for $u\geqslant 0$ that
+$\mathbb{P}\Big{(}\max_{i\in[d]}\frac{1}{n}\sum_{\ell=1}^{n}2^{-k}\xi_{k,i}^{\ell}\geqslant e^{-\frac{k}{4d}}+u\Big{)}$ $\leqslant d\,\mathbb{P}\Big{(}\frac{1}{n}\sum_{\ell=1}^{n}2^{-k}\xi_{k,i}^{\ell}\geqslant e^{-\frac{k}{4d}}+u\Big{)}$
+$\leqslant d\,\mathbb{P}\Big{(}\sum_{\ell=1}^{n}2^{-k}\xi_{k,i}^{\ell}\geqslant\mathbb{E}\Big{[}\sum_{\ell=1}^{n}2^{-k}\xi_{k,i}^{\ell}\Big{]}+nu\Big{)}\leqslant de^{-2nu^{2}}.$
+Therefore, a standard computation yields
+$\mathbb{E}\Big{[}\Big{(}\max_{i\in[d]}\frac{1}{n}\sum_{\ell=1}^{n}2^{-k}\xi_{k,i}^{\ell}\Big{)}^{2/3}\Big{]}\leqslant e^{-\frac{k}{6d}}+\frac{d}{n^{1/3}}.$
+Inserting into (60) and taking expectation, we see that for all $g\in\mathcal{G}$,
+$\mathbb{E}[\mathbb{E}_{n}[\mathbb{E}[(Y-M_{k}^{S})^{2}]]]\leqslant 2\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+2\mathrm{Var}(Y)^{2/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\Big{(}e^{-\frac{k}{6d}}+\frac{d}{n^{1/3}}\Big{)}.$
+Using (58) and (47), we obtain that for all $g\in\mathcal{G}$,
+$\mathbb{E}[(Y-\mathbb{E}_{n}[M_{k}^{S}])^{2}]\leqslant 2\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+2\Big{(}\frac{4}{3}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\Big{)}\Big{(}e^{-\frac{k}{6d}}+\frac{d}{n^{1/3}}\Big{)}.$
+Following the same lines as the proof of Theorem 6, we obtain that in the presence of marginal atoms of weights at most $1/N$, it holds for all $g\in\mathcal{G}$,
+$\mathbb{E}[(Y-\mathbb{E}_{n}[M_{k}^{S}])^{2}]\leqslant 2\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\frac{8}{3}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\left\lVert g\right\rVert_{\mathrm{TV}}^{2}+\frac{2^{k}M^{2}}{N}\Big{)}\Big{(}e^{-\frac{k}{6d}}+\frac{d}{n^{1/3}}\Big{)}.$
+Finally, we proceed with the arguments in proving Theorem 7.
+Again we follow the arguments in the proof of Theorem 4.3 for CART of Klusowski and Tian (2024).
+By setting $M=2U$ where $U=\max|Y_{i}|$, the extra term of
+$\frac{8}{3}\frac{2^{k}M^{2}}{N}\Big{(}e^{-\frac{k}{6d}}+\frac{d}{n^{1/3}}\Big{)}\leqslant C\Big{(}\frac{2^{k}U^{2}}{N}+\frac{2^{k}U^{2}d}{Nn^{1/3}}\Big{)}$
+is carried until (B.34) of Klusowski and Tian (2024).
+The term $2^{k}U^{2}/N$ is absorbed by the term $U^{4}2^{k}\log(Nd)/N$ therein.
+The other term $2^{k}U^{2}d/(Nn^{1/3})$ is carried into (B.35), and is bounded by $C2^{k}(\log N)d/(Nn^{1/3})$ under the choice of $U=\left\lVert g^{*}\right\rVert_{\infty}+2\sigma\sqrt{2\log N}$ at the end of that proof, where $C>0$ may depend on $\left\lVert g^{*}\right\rVert_{\infty}$ and $\sigma$.
+This concludes (26).
+∎
+B.3 Proofs of results in Section 4
+The analysis in the classification setting also relies on two monotonicity envelopes for the conditional probability.
+Recall (27).
+Lemma B.3 provides left and right bounds for the probability of an order event inside a cell in terms of the functions
+$\varphi_{L}(x)\coloneqq\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x)\,h\left(\mathbb{P}(Y=1\mid\mathbf{X}\in A,\,X_{j}<x)\right);$
+$\varphi_{R}(x)\coloneqq\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x)\,h\left(\mathbb{P}(Y=1\mid\mathbf{X}\in A,\,X_{j}\geqslant x)\right),$
+which are nondecreasing and nonincreasing, respectively.
+Lemma B.4 further provides an information inequality that upper bounds the entropy.
+These two lemmas are the entropic analogues of the unconditional variance contraction used in the regression section; they are the technical bridges that allow us to carry over the exponential decay arguments from squared loss to cross-entropy.
+Lemma B.3.
+Suppose that $\mathbf{X}$ is marginally atomless and $j\in[d]$.
+Then the function $\varphi_{L}$ is non-decreasing and continuous, and $\varphi_{R}$ is non-increasing and continuous.
+Moreover, under the (cyclic) MinimaxSplit construction,
+$\max_{j\in J_{k}}\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\leqslant 2^{-k}\mathbb{E}[\log(1+e^{-YM_{0}})]=2^{-k}H(Y).$
+Proof.
+To see the monotonicity of $\varphi_{L}$, we take arbitrary $a<b$. Since the conditional entropy is bounded from above by the entropy, we have
+$h\left(\mathbb{P}(Y=1\mid\mathbf{X}\in A,\,X_{j}<b)\right)\geqslant\mathbb{P}(X_{j}<a\mid\mathbf{X}\in A,\,X_{j}<b)\,h\left(\mathbb{P}(Y=1\mid\mathbf{X}\in A,\,X_{j}<a)\right).$
+Rearranging yields that $\varphi_{L}(a)\leqslant\varphi_{L}(b)$.
+The continuity of $\varphi_{L}$ and $\varphi_{R}$ follows from the marginally atomless condition.
+The final claim follows in the same way as the derivation of (11).
+∎
+Lemma B.4.
+Let $Y$ be a $\{\pm 1\}$-valued random variable. For any joint distribution $(X,Y)$, it holds that
+$H(Y)\leqslant\mathbb{E}[\log(1+e^{-XY})]+\frac{1}{4}|\sup\operatorname{supp}X-\inf\operatorname{supp}X|.$ (62)
+Proof.
+We may assume that $\sup\operatorname{supp}X-\inf\operatorname{supp}X$, otherwise the claim is trivial.
+Given the marginal law of $X$, the right-hand side of (62) is minimized if $X$ and $Y$ are comonotonic.
+This is achieved by setting, with $a=\inf\operatorname{supp}X\leqslant\sup\operatorname{supp}X=b$, that $p=\mathbb{P}(X=b,Y=1)=1-\mathbb{P}(X=a,Y=-1)$, for some $p\in[0,1]$.
+It remains to show the deterministic inequality
+$\kappa_{a,b}(p):=h(p)+(p\log(1+e^{-b})+(1-p)\log(1+e^{a}))+\frac{1}{4}|b-a|\geqslant 0.$
+Note that $\kappa_{a,b}$ is convex in $p$ and $\kappa_{a,b}^{\prime}(p)=0$ if and only if
+$\log\Big{(}\frac{1-p}{p}\Big{)}=\log\Big{(}\frac{1+e^{-b}}{1+e^{a}}\Big{)},$
+or $p=(1+e^{a})/((1+e^{a})+(1+e^{-b}))$. It remains to prove
+$\log\Big{(}\frac{(1+e^{a})(1+e^{-b})}{(1+e^{a})+(1+e^{-b})}\Big{)}\geqslant-\frac{1}{4}|b-a|,$
+or equivalently,
+$\log((1+e^{a})^{-1}+(1+e^{-b})^{-1})\leqslant\frac{1}{4}|b-a|.$ (63)
+With D:=ba>0D:=b-a>0 fixed, the left-hand side of (63) is maximized at a=D/2a=-D/2 and $b=D/2$ by elementary calculus.
+On the other hand, it is easy to verify by elementary calculus that $\log(2/(1+e^{-D/2}))\leqslant D/4=|b-a|/4$.
+This proves (63) and hence completes the proof.
+∎
+Proof of Theorem 10.
+We first recall the notations from the proof of Theorem 2.
+Denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ from the cyclic MinimaxSplit construction.
+Note that $M_{k}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}=y_{k,j}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}$, where $y_{k,j}=\log(\mathbb{P}(Y=1\mid\mathbf{X}\in I_{j})/\mathbb{P}(Y=-1\mid\mathbf{X}\in I_{j}))$.
+Define $g_{j}=g|_{I_{j}}$, so that $\left\lVert g\right\rVert_{\mathrm{TV}}=\sum_{j}\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
+Recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
+By applying Lemma B.4 to the joint distribution $(g(\mathbf{X}),Y)\mid\mathbf{X}\in I_{j}$, decompose
+$\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]=\mathbb{E}[\log(1+e^{-Yy_{k,j}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\leqslant U_{j}+V_{j},$ (64)
+where $U_{j}=\mathbb{E}[\log(1+e^{-Yg(\mathbf{X})}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and
+$V_{j}=\min\Big{\{}\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}],\frac{p_{j}\Delta g_{j}}{4}\Big{\}}.$
+We have by construction and (64) that
+$\mathbb{E}[\log(1+e^{-YM_{k}})]=\sum_{j\in J_{k}}\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\leqslant\sum_{j\in J_{k}}U_{j}+\sum_{j\in J_{k}}V_{j}.$ (65)
+The first term is easy to control by definition:
+$\sum_{j\in J_{k}}U_{j}\leqslant\sum_{j\in J_{k}}\mathbb{E}[\log(1+e^{-Yg(\mathbf{X})}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]=\mathbb{E}[\log(1+e^{-Yg(\mathbf{X})})].$ (66)
+To control $\sum_{j}V_{j}$, we apply the same Hölder’s inequality argument as in the proof of Theorem 16 in the regression setting. More precisely, we have
+$\begin{split}\sum_{j\in J_{k}}V_{j}&\leqslant\Big{(}\sum_{j\in J_{k}}p_{j}\big{(}\frac{V_{j}}{p_{j}}\big{)}^{2}\Big{)}^{1/2}\Big{(}\sum_{j\in J_{k}}p_{j}\Big{)}^{1/2}\\ &\leqslant\Bigg{(}(\max_{j\in J_{k}}V_{j})\sum_{j\in J_{k}}\frac{V_{j}}{p_{j}}\Bigg{)}^{1/2}\leqslant\Bigg{(}(\max_{j\in J_{k}}\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}])\sum_{j\in J_{k}}\frac{V_{j}}{p_{j}}\Bigg{)}^{1/2}.\end{split}$ (67)
+To further bound the right-hand side of (67), we recall from Lemma B.3 that as a consequence of the marginally atomless property,
+$\max_{j\in J_{k}}\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\leqslant 2^{-k}\mathbb{E}[\log(1+e^{-YM_{0}})]=2^{-k}H(Y).$ (68)
+Note that ${V_{j}}/{p_{j}}\leqslant\Delta g_{j}/4.$ Following the proof of Theorem 16, we arrive at
+$\sum_{j\in J_{k}}\frac{V_{j}}{p_{j}}\leqslant 2^{-2+k-\lfloor k/d\rfloor}\left\lVert g\right\rVert_{\mathrm{TV}}.$
+Inserting into (67) gives (using that $H(Y)\leqslant\log 2$)
+$\sum_{j\in J_{k}}V_{j}\leqslant(2^{k-2-\lfloor k/d\rfloor}\left\lVert g\right\rVert_{\mathrm{TV}}2^{-k}H(Y))^{1/2}\leqslant\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}.$ (69)
+Finally, inserting (66) and (69) into (65) yields (29). ∎
+Proof of Theorem 11.
+The only difference from the proof of Theorem 10 is (68). Denoting by $\hat{x}_{j}$ the optimal MinimaxSplit location on $A$ in dimension $j$, we have
+$\max\Big{\{}\varphi_{L}(\hat{x}_{j}),\varphi_{R}(\hat{x}_{j})\Big{\}}\leqslant\frac{1}{2}\mathbb{P}(\mathbf{X}\in A)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A))+\frac{\log N}{N},$
+because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $\log N/N$ (which can be seen from the extreme case with a sample $1$ added to samples of 1-1).
+As a consequence, if we denote by
+$u_{k}:=\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A)),$
+then $u_{k}$ satisfies the recursive inequalities
+$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{\log N}{N},\leavevmode\nobreak\ k\geqslant 0;\quad u_{0}=H(Y).$
+Solving this yields
+$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A))=u_{k}\leqslant 2^{-k}H(Y)+\frac{2\log N}{N}.$ (70)
+It is also easy to check that (70) is also satisfied for the case $u_{0}<2\log N/N$, that is, (70) holds in general. Replacing (68) by (70) leads to
+$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\bigg{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\Big{(}\left\lVert g\right\rVert_{\mathrm{TV}}(1+2^{k+1}\frac{\log N}{N})\Big{)}^{1/2}\bigg{)}.$
+Applying concavity of the function $x\mapsto x^{1/2}$ for $x>0$ and that
+$\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}\Big{(}2^{k+1}\frac{\log N}{N}\Big{)}^{1/2}\leqslant\frac{\left\lVert g\right\rVert_{\mathrm{TV}}}{2}+\frac{2^{k}\log N}{N},$
+we obtain (31). ∎
+Appendix C Additional example for anti-symmetry-breaking preference
+In this appendix, we provide an empirical version of Example 8, showing that the ASBP persists even when samples are involved and the symmetry is not perfect.
+Example 23.
+Define $f(\mathbf{x})=x_{1}+|x_{2}|,\leavevmode\nobreak\ \mathbf{x}=(x_{1},x_{2})\in[-1,1]^{2}$, which is Lipschitz and bounded.
+Let $\mathbb{P}_{*}$ denote the law of $\mathbf{X}$ in Example 8 and $\mathbb{P}_{n}$ denote the empirical law of $\{(\mathbf{X}_{i},Y_{i})\}_{1\leqslant i\leqslant n}$, where each $\mathbf{X}_{i}$ is sampled from $\mathbb{P}_{*}$ and $Y_{i}=f(\mathbf{X}_{i})$.
+Denote by $\mathcal{R}$ the set of all axis-aligned rectangles in $1,1]^{2}$.
+It follows from standard arguments in van der Vaart and Wellner (2013) that the class $\{f{\mathbbm{1}}_{R}\}_{R\in\mathcal{R}}$ is a Donsker class, so with $m_{R}=\mathbb{E}^{\mathbb{P}_{*}}[f(\mathbf{X})\mid\mathbf{X}\in R]$, we have $\sup_{R\in\mathcal{R}}|\mathbb{E}^{\mathbb{P}_{n}}[f(\mathbf{X})\mid\mathbf{X}\in R]-m_{R}|=O_{\mathbb{P}}(1/\sqrt{n})$.
+Here, we use the notation $\xi_{n}=O_{\mathbb{P}}(\alpha_{n})$ if the sequence $\{\xi_{n}/\alpha_{n}\}$ is tight.
+Let $g_{R}(\mathbf{x})=(f(\mathbf{x})-m_{R})^{2}$ for $R\in\mathcal{R}$.
+The class $\{g_{R}{\mathbbm{1}}_{R}\}_{R\in\mathcal{R}}$ is also Donsker.
+For $R\in\mathcal{R}$, let $V_{R}$ denote the empirical risk on $R$, i.e.,
+$V_{R}=\frac{1}{n}\sum_{i=1}^{n}((f(\mathbf{X}_{i})-\mathbb{E}^{\mathbb{P}_{n}}[f(\mathbf{X})\mid\mathbf{X}\in R])^{2}{\mathbbm{1}}_{\mathbf{X}_{i}\in R}).$
+It follows from the Donsker property and triangle inequality that
+$\sup_{R\in\mathcal{R}}|V_{R}-\mathbb{E}[V_{R}]|=O_{\mathbb{P}}\Big{(}\frac{1}{\sqrt{n}}\Big{)}.$ (71)
+Next, we apply (71) to analyze the split dimensions and locations when applying MinimaxSplit to $\mathbb{P}_{n}$.
+For any $\varepsilon>0$, there exists $C=C(\varepsilon)>0$ such that $\mathbb{P}(\forall R\in\mathcal{R},\,|V_{R}-\mathbb{E}[V_{R}]|<C/\sqrt{n})>1-\varepsilon$.
+Our goal is to show the deterministic fact that for some $C_{0}=C_{0}(\varepsilon)>0$, if $n\geqslant C_{0}64^{K}$, on the event
+$\Big{\{}\forall R\in\mathcal{R},\,|V_{R}-\mathbb{E}[V_{R}]|<\frac{C}{\sqrt{n}}\Big{\}},$ (72)
+the splits of the first $K$ levels are all along the first coordinate $x_{1}$. Consider a rectangle $R_{a,b}=[a,b]\times[-1,1]\in\mathcal{R}$ where $1\leqslant a<b\leqslant 1$. Define
+$R_{a,b;u}^{\uparrow}$ $=[a,b]\times[u,1],\quad R_{a,b;u}^{\downarrow}=[a,b]\times[-1,u],\quad u\in[-1,1];$
+$R_{a,b;u}^{\leftarrow}$ $=[a,u]\times[-1,1],\quad R_{a,b;u}^{\rightarrow}=[u,b]\times[-1,1],\quad u\in[a,b].$
+Also denote by $m(p)=\mathbb{E}[(M-1){\mathbbm{1}}_{\{M\geqslant 1\}}/M]$ where $M\buildrel\mathrm{law}\over{\sim}\mathrm{Bin}(n,p)$.
+A standard computation (by first conditioning on the number of samples in the rectangle) yields that
+$\begin{split}\mathbb{E}\left[V_{R_{a,b}}\right]&=m\Big{(}\frac{b-a}{2}\Big{)}\frac{b-a}{2}\frac{(b-a)^{2}+1}{12};\\ \mathbb{E}\left[V_{R_{a,b;u}^{\uparrow}}\right]&=m\Big{(}\frac{(1-u)(b-a)}{4}\Big{)}\frac{(1-u)(b-a)}{4}\Big{(}\frac{(b-a)^{2}}{12}+\phi(u)\Big{)};\\ \mathbb{E}\left[V_{R_{a,b;u}^{\downarrow}}\right]&=m\Big{(}\frac{(u+1)(b-a)}{4}\Big{)}\frac{(u+1)(b-a)}{4}\Big{(}\frac{(b-a)^{2}}{12}+\phi(-u)\Big{)};\\ \mathbb{E}\left[V_{R_{a,b;u}^{\leftarrow}}\right]&=m\Big{(}\frac{u-a}{2}\Big{)}\frac{u-a}{2}\frac{(u-a)^{2}+1}{12};\\ \mathbb{E}\left[V_{R_{a,b;u}^{\rightarrow}}\right]&=m\Big{(}\frac{b-u}{2}\Big{)}\frac{b-u}{2}\frac{(b-u)^{2}+1}{12},\end{split}$ (73)
+where
+$\phi(u)=\begin{cases}\frac{(1-u)^{2}}{12}&\text{ if }u\geqslant 0;\\ \frac{u^{4}-4u^{3}-6u^{2}-4u+1}{12(1-u)^{2}}&\text{ if }u<0.\end{cases}$
+In the following, the constants $C$ (depending on $\varepsilon$ only) may vary from line to line.
+Note that m(p)=1O(1/(np))m(p)=1-O(1/(np)) jointly as $np\to\infty$ (see e.g., Corollary 1 of Wang et al.
+(2008)).
+The following results hold if $b-a\geqslant C/n^{1/6}$ for $n$ large enough.
+  • Consider a split of $R_{a,b}$ into $R_{a,b;u}^{\leftarrow}$ and $R_{a,b;u}^{\rightarrow}$.
+The MinimaxSplit location satisfies $|u-(a+b)/2|\leqslant C/\sqrt{n}$ by (72) and (73).
+The risk decay is at least
+    $\mathbb{E}\left[V_{R_{a,b}}\right]-\mathbb{E}\left[V_{R_{a,b;u}^{\leftarrow}}\right]-\mathbb{E}\left[V_{R_{a,b;u}^{\rightarrow}}\right]-\frac{3C}{\sqrt{n}}\geqslant\frac{(b-a)^{3}}{32}-\frac{C}{\sqrt{n}}.$
+  • Consider a split of $R_{a,b}$ into $R_{a,b;u}^{\uparrow}$ and $R_{a,b;u}^{\downarrow}$.
+The function $u\mapsto(1-u)\phi(u)$ is decreasing and differentiable on [1,1][-1,1] with a strictly negative derivative at $u=0$.
+Therefore, the MinimaxSplit location satisfies $|u|\leqslant C/((b-a)\sqrt{n})$ by (72) and (73).
+The risk decay is at most
+    $\mathbb{E}\left[V_{R_{a,b}}\right]-\mathbb{E}\left[V_{R_{a,b;u}^{\downarrow}}\right]-\mathbb{E}\left[V_{R_{a,b;u}^{\uparrow}}\right]+\frac{3C}{\sqrt{n}}\leqslant\frac{C}{\sqrt{n}}.$
+Therefore, if $b-a\geqslant C/n^{1/6}$, the MinimaxSplit dimension is always $x_{1}$.
+Next, we show that under our assumption $n\geqslant C_{0}64^{K}$, we always have $b-a\geqslant C/n^{1/6}$ for the first $K$ levels of splits.
+The split at level $k=0$ is obvious.
+Denote by $a_{k}$ the smallest size (measured in dimension $x_{1}$) of the rectangles at level $k$.
+It follows that $a_{0}=2$ and for each $k$, $a_{k+1}\geqslant a_{k}/2-C/\sqrt{n}$.
+Solving this recursion yields $a_{k}\geqslant 2^{1-k}-2C/\sqrt{n}$ for all $k$.
+In particular, $a_{K}\geqslant 2^{1-K}-2C/\sqrt{n}\geqslant C/n^{1/6}$.
+In summary, we have shown that if $n\geqslant C_{0}64^{K}$, with high probability, the MinimaxSplit dimension is always along the first dimension for the first $K$ levels of the tree.
+As a consequence, the consistency of the MinimaxSplit algorithm is not always guaranteed in the large-sample regime, unlike VarianceSplit (Theorem 4.3 of Klusowski and Tian (2024)) and cyclic MinimaxSplit (Theorem 7).
+Appendix D Further numerics
+D.1 Performance metrics
+Mean Squared Error (MSE).
+Given a matrix (or image) $y\in\mathbb{R}^{H\times W}$ and an estimate $\hat{y}$, we define
+$\mathrm{MSE}(y,\hat{y})=\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}\bigl{(}y_{ij}-\hat{y}_{ij}\bigr{)}^{2}.$
+MSE emphasizes large errors (squaring) and is simple to optimize. We also define the Root Mean Squared Error (RMSE) as
+$\operatorname{RMSE}(y,\hat{y})\;=\;\sqrt{\operatorname{MSE}(y,\hat{y})}\;=\;\sqrt{\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}\bigl{(}y_{ij}-\hat{y}_{ij}\bigr{)}^{2}}\,.$
+Mean Absolute Error (MAE).
+We define
+$\mathrm{MAE}(y,\hat{y})=\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}\lvert y_{ij}-\hat{y}_{ij}\rvert$.
+MAE (an $\ell_{1}$ criterion) is more robust to outliers than MSE and often produces sharper restorations with less over-smoothing in image restoration tasks; in practice, it can outperform $\ell_{2}$ metrics on denoising and related problems (Zhao et al., 2017; Jadon et al., 2024).
+Coefficient of Determination ($R^{2}$).
+Let the spatial mean be
+$\bar{y}\;=\;\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}y_{ij}.$
+Define the residual and total sums of squares
+$\mathrm{SS}_{\mathrm{res}}\;=\;\sum_{i=1}^{H}\sum_{j=1}^{W}\bigl{(}y_{ij}-\hat{y}_{ij}\bigr{)}^{2},\qquad\mathrm{SS}_{\mathrm{tot}}\;=\;\sum_{i=1}^{H}\sum_{j=1}^{W}\bigl{(}y_{ij}-\bar{y}\bigr{)}^{2}.$
+Then
+$R^{2}(y,\hat{y})\;=\;1-\frac{\mathrm{SS}_{\mathrm{res}}}{\mathrm{SS}_{\mathrm{tot}}}\;=\;1-\frac{\operatorname{MSE}(y,\hat{y})}{\operatorname{Var}(y)}\,,\quad\text{where }\;\operatorname{Var}(y)\;=\;\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}\bigl{(}y_{ij}-\bar{y}\bigr{)}^{2}.$
+Structural Similarity (SSIM) (Wang et al., 2004).
+For a window $(\Omega\subset\{1,\ldots,H\}\times\{1,\ldots,W\})$ with non-negative weights $(\{w_{uv}\}_{(u,v)\in\Omega})$ satisfying $\sum_{(u,v)\in\Omega}w_{uv}=1$, define the local means, variances and covariance
+$\mu_{y}(\Omega)=\sum_{(u,v)\in\Omega}w_{uv}\,y_{uv},\qquad\mu_{\hat{y}}(\Omega)=\sum_{(u,v)\in\Omega}w_{uv}\,\hat{y}_{uv},$
+$\sigma_{y}^{2}(\Omega)=\sum_{(u,v)\in\Omega}w_{uv}\,\bigl{(}y_{uv}-\mu_{y}(\Omega)\bigr{)}^{2},\qquad\sigma_{\hat{y}}^{2}(\Omega)=\sum_{(u,v)\in\Omega}w_{uv}\,\bigl{(}\hat{y}_{uv}-\mu_{\hat{y}}(\Omega)\bigr{)}^{2},$
+$\sigma_{y\hat{y}}(\Omega)=\sum_{(u,v)\in\Omega}w_{uv}\,\bigl{(}y_{uv}-\mu_{y}(\Omega)\bigr{)}\bigl{(}\hat{y}_{uv}-\mu_{\hat{y}}(\Omega)\bigr{)}.$
+With stability constants $C_{1}=(K_{1}L)^{2}$ and $C_{2}=(K_{2}L)^{2}$ (dynamic range $L$; typically $K_{1}=0.01$, $K_{2}=0.03$), the window-level structural similarity is
+$\mathrm{SSIM}_{\Omega}(y,\hat{y})\;=\;\frac{\bigl{(}2\,\mu_{y}(\Omega)\,\mu_{\hat{y}}(\Omega)+C_{1}\bigr{)}\bigl{(}2\,\sigma_{y\hat{y}}(\Omega)+C_{2}\bigr{)}}{\bigl{(}\mu_{y}(\Omega)^{2}+\mu_{\hat{y}}(\Omega)^{2}+C_{1}\bigr{)}\bigl{(}\sigma_{y}^{2}(\Omega)+\sigma_{\hat{y}}^{2}(\Omega)+C_{2}\bigr{)}}.$
+The image-level SSIM is the average over all window centers $\{\Omega_{m}\}_{m=1}^{M}$,
+[…]
+In our experiments the images are normalized to $[0,1]$, hence $L=1$, and we use Gaussian weights $w_{uv}$ and mean aggregation as in skimage.metrics.structural_similarity.
+D.2 Higher-dimensional input domain
+Consider a dimension $d\geqslant 1$ that is a multiple of $4$. The Powell function on $0,1]^{d}$,
+$f(\mathbf{x})=\sum_{i=1}^{d/4}[(x_{4i-3}+10x_{4i-2})^{2}+5(x_{4i-1}-x_{4i})^{2}+(x_{4i-2}-2x_{4i-1})^{4}+10(x_{4i-3}-x_{4i})^{4}],$
+with uniformly random samples on $0,1]^{d}$, serves as an exemplary testbed for decision tree methods in high-dimensional regression problems due to its inherent non-linearity and intricate variable interactions.
 The quartic terms in the function introduce sharp curvatures in the response surface, necessitating sophisticated splitting strategies to approximate these non-linear relationships accurately.
-Moreover, the coupling of variables in quartets creates a web of local interactions that tests the trees’ ability to discern and model multivariable dependencies efficiently.
 Sample Size Dimension Sklearn VarianceSplit MinimaxSplit
-10 4 $4.48\times 10^{2}$ $2.01\times 10^{2}$ $2.57\times 10^{2}$
-10 16 $1.60\times 10^{4}$ $1.52\times 10^{4}$ $1.52\times 10^{4}$
-10 64 $3.70\times 10^{4}$ $4.24\times 10^{4}$ $3.79\times 10^{4}$
-10 256 $1.35\times 10^{5}$ $1.24\times 10^{5}$ $1.14\times 10^{5}$
-10 1024 $9.31\times 10^{5}$ $8.04\times 10^{5}$ $4.28\times 10^{5}$
-100 4 $5.49\times 10^{1}$ $5.21\times 10^{1}$ $5.31\times 10^{1}$
-100 16 $2.70\times 10^{3}$ $2.67\times 10^{3}$ $2.57\times 10^{3}$
-100 64 $1.99\times 10^{4}$ $1.95\times 10^{4}$ $2.20\times 10^{4}$
-100 256 $1.01\times 10^{5}$ $1.02\times 10^{5}$ $9.39\times 10^{4}$
-100 1024 $4.65\times 10^{5}$ $4.29\times 10^{5}$ $4.02\times 10^{5}$
-1000 4 $4.68\times 10^{1}$ $4.65\times 10^{1}$ $4.49\times 10^{1}$
-1000 16 $1.85\times 10^{3}$ $1.85\times 10^{3}$ $2.03\times 10^{3}$
-1000 64 $1.62\times 10^{4}$ $1.62\times 10^{4}$ $1.59\times 10^{4}$
-1000 256 $6.46\times 10^{4}$ $6.47\times 10^{4}$ $6.30\times 10^{4}$
-1000 1024 $2.94\times 10^{5}$ $2.96\times 10^{5}$ $2.88\times 10^{5}$
-10000 4 $4.61\times 10^{1}$ $4.61\times 10^{1}$ $4.63\times 10^{1}$
-10000 16 $1.87\times 10^{3}$ $1.87\times 10^{3}$ $1.97\times 10^{3}$
-10000 64 $1.36\times 10^{4}$ $1.36\times 10^{4}$ $1.43\times 10^{4}$
-10000 256 $6.22\times 10^{4}$ $6.22\times 10^{4}$ $6.11\times 10^{4}$
-10000 1024 $2.80\times 10^{5}$ $2.80\times 10^{5}$ $2.78\times 10^{5}$
-Table 1: Comparison of Decision Tree Methods (max_depth=3) using the Powell function without noise.
-Scikit-learn’s DecisionTreeRegressor uses mean-squared Error (MSE) as its default splitting criterion, which is equivalent to minimizing the variance of the target variable within each split and serve as a baseline comparison here.
-Table 1 presents a comprehensive comparison of three decision tree methods—Scikit-learn’s standard implementation, VarianceSplit and MinimaxSplit—across various sample sizes $n$ and dimensions $d$ for the Powell function.
-This data offers valuable insights into the performance characteristics of these methods in both dense and sparse design scenarios.
-In the dense design regime ($n>d$), particularly evident in the cases where the sample size (100, 1000 or 10000) exceeds the dimensionality (4, 16 or 64), we observe relatively consistent performance across all three methods.
-This suggests that when data is abundant relative to the input space, the choice of splitting criterion (be it the standard impurity measure used by Scikit-learn, MinimaxSplit, or VarianceSplit) has less impact on the overall performance.
-However, as we transition into sparse designs ($d>n$), exemplified by cases where dimensions (64, 256 or 1024) exceed the sample size (10 or 100), more pronounced differences emerge.
+10 4 $4.48\times 10^{2}$ $\boxed{2.01\times 10^{2}}$ $2.57\times 10^{2}$
+10 16 $1.60\times 10^{4}$ $1.52\times 10^{4}$} $\boxed{1.52\times 10^{4}}$
+10 64 $3.70\times 10^{4}$ $4.24\times 10^{4}$ $\boxed{3.79\times 10^{4}}$
+[…]
+10 1024 $9.31\times 10^{5}$ $8.04\times 10^{5}$ $\boxed{4.28\times 10^{5}}$
+100 4 $5.49\times 10^{1}$ $\boxed{5.21\times 10^{1}}$ $5.31\times 10^{1}$
+100 16 $2.70\times 10^{3}$ $2.67\times 10^{3}$ $\boxed{2.57\times 10^{3}}$
+100 64 $1.99\times 10^{4}$ $\boxed{1.95\times 10^{4}}$ $2.20\times 10^{4}$
+100 256 $1.01\times 10^{5}$ $1.02\times 10^{5}$ $\boxed{9.39\times 10^{4}}$
+100 1024 $4.65\times 10^{5}$ $4.29\times 10^{5}$ $\boxed{4.02\times 10^{5}}$
+[…]
+1000 16 $\boxed{1.85\times 10^{3}}$ $1.85\times 10^{3}$ $2.03\times 10^{3}$
+1000 64 $1.62\times 10^{4}$ $1.62\times 10^{4}$ $\boxed{1.59\times 10^{4}}$
+1000 256 $6.46\times 10^{4}$ $6.47\times 10^{4}$ $\boxed{6.30\times 10^{4}}$
+1000 1024 $2.94\times 10^{5}$ $2.96\times 10^{5}$ $\boxed{2.88\times 10^{5}}$
+10000 4 $\boxed{4.61\times 10^{1}}$ $4.61\times 10^{1}$ $4.63\times 10^{1}$
+10000 16 $\boxed{1.87\times 10^{3}}$ $1.87\times 10^{3}$ $1.97\times 10^{3}$
+10000 64 $\boxed{1.36\times 10^{4}}$ $1.36\times 10^{4}$ $1.43\times 10^{4}$
+10000 256 $6.22\times 10^{4}$ $6.22\times 10^{4}$ $\boxed{6.11\times 10^{4}}$
+10000 1024 $2.80\times 10^{5}$ $2.80\times 10^{5}$ $\boxed{2.78\times 10^{5}}$
+Table 1: Comparison of Decision Tree Methods ($\text{maximum depth}=3$) using the Powell function without noise.
+Scikit-learn’s DecisionTreeRegressor uses MSE as its default splitting criterion, which is equivalent to minimizing the variance of the target variable within each split and serves as a baseline comparison here.
+The winner in each setting is boxed.
+Table 1 compares three decision tree methods—scikit-learn’s standard implementation, VarianceSplit and MinimaxSplit—across various sample sizes $n$ and dimensions $d$ for the Powell function.
+In the dense design regime ($n>d$), we observe relatively consistent performance across all three methods.
+However, as we transition into sparse designs ($d>n$), more pronounced differences emerge.
 The MinimaxSplit method often demonstrates superior performance in these high-dimensional, data-scarce scenarios.
 For instance, with 10 samples and 256 dimensions, the MinimaxSplit approach achieves an MSE of 8.91 × 104, outperforming both scikit-learn (1.07 × 105) and the VarianceSplit method (9.58 × 104).
-This aligns with the theoretical strengths of the minimax criterion in handling sparse, high-dimensional data, where it can more robustly identify informative splits with limited samples.
-The VarianceSplit method generally performs comparably to Scikit-learn’s implementation, which is expected given that both likely use variants of variance reduction as their splitting criteria.
-However, in some sparse cases, the VarianceSplit method shows slight improvements, possibly due to specific implementation details or hyperparameter choices.
-As dimension increases for a fixed sample size, we observe a consistent increase in MSE across all methods, illustrating the challenges posed by the curse of dimensionality.
-This effect is particularly pronounced in the transition from 64 to 256 dimensions, where the MSE often increases by an order of magnitude or more.
-Interestingly, the performance gap between methods narrows as the sample size increases, even in high-dimensional settings.
-For instance, with 1000 samples and 256 dimensions, all three methods achieve comparable MSE values (around 6.4 × 104).
-This suggests that with sufficient data, the choice of splitting criterion becomes less critical and all methods can effectively capture the underlying structure of the Powell function.
-B.3 A weighted aggregation approach
-The variance approach aims to create homogeneous subsets by minimizing the average squared difference between the observed values and the mean value in each node.
-In contrast, the minimax criterion focuses on minimizing the maximum error within each split, potentially leading to more balanced trees.
-The variance criterion, similar to Scikit-learn’s MSE, seeks to minimize the overall variance in child nodes but can be implemented with different loss norms ($L^{1}$ or $L^{2}$).
-Cyclic minimax, a variant of minimax, alternates through features (i.e., the first and second coordinates) in a predetermined order for splitting, which can be beneficial in high-dimensional spaces when the depth is larger than the input dimension, or when feature importance is known a priori (e.g., the first coordinate is more heterogeneous).
-These alternative criteria offer different trade-offs: minimax may be more robust to outliers and preserve edges better in image processing tasks, variance (especially with $L^{2}$ norm) often provides good average-case performance, and cyclic minimax can ensure a more diverse use of features.
-Unlike Scikit-learn’s implementation, which primarily optimizes for average error reduction, these custom criteria allow for more specialized tree structures tailored to specific problem characteristics or performance goals.
-Our novel random forest (see Algorithm 1 of Appendix C) implementation diverges from conventional practices by introducing diverse splitting criteria and weighted aggregation.
-This approach is motivated by our previous discussion that the recognition that different splitting criteria can capture various aspects of data structure, while weighted aggregation can emphasize more accurate models in the final prediction.
-The implementation incorporates Variance, Minimax, and cyclic Minimax splitting methods for different weak learners within a single ensemble.
-The variance method optimizes for overall error reduction, minimax focuses on worst-case performance at each split, and cyclic minimax introduces a deterministic feature selection process.
-This diversity in the splitting criteria aims to mitigate feature importance bias present in standard random forests and enhance ensemble diversity beyond what is typically achieved through bootstrapping and random feature subset selection (Friedman, 2001).
-The weighted aggregation scheme, based on the reciprocal of each tree’s training RMSE, addresses the limitation that not all trees in the ensemble are equally reliable or informative.
-This adaptive ensemble aggregation allows the model to adjust to the varying quality of its constituent trees, potentially leading to more robust predictions.
-These modifications are primarily motivated by addressing limitations of standard random forests: feature importance bias, limited ensemble diversity, and non-adaptive aggregation.
-By introducing varied perspectives on feature importance, adding layers of diversification, and allowing for adaptive weighting, our approach aims to create a more flexible and robust ensemble.
-B.4 Heterogeneous splitting
-We also develop heterogeneous splitting strategies based on our findings.
-The heterogeneous splitting strategy in decision trees involves using different error metrics (splitting criteria) at various depths, unlike the homogeneous strategy that uses the same metric throughout.
-This approach allows flexibility and adaptability, essential for handling complex datasets where different regions of the input space may benefit from different strategies.
-Early splits might use variance reduction for effective data division, while deeper splits could switch to minimax to minimize worst-case errors, enhancing model robustness against outliers and reducing overfitting risks; alternatively we can also alternate between VarianceSplit and MinimaxSplit to control errors (Figure 6).
-Additionally, strategies like cyclic MinimaxSplit leverage periodic features, potentially improving performance in time series or cyclic data.
-This tailored approach adapts the splitting criterion based on depth, potentially leading to more accurate and effective splits as different tree levels might require different considerations.
-In fact, the minimax/variance alternating strategy with $L^{2}$ norm (RMSE=0.112334) further improves the lowest RMSE compared to a homogeneous splitting criteria.
-Consequently, a heterogeneous strategy can lead to more accurate and robust models by optimally addressing the unique characteristics of the data at each level of the tree, making it particularly beneficial in complex datasets with variable distributions.
+Appendix E Pseudocode for algorithms
+Input: Node $A_{t}$ at depth $k$ with sample $D_{t}=\{(x_{i},y_{i}):x_{i}\in A_{t}\}$; feature set $d]=\{1,\dots,d\}$ and $m_{try}=d$ by default
+Output: $S_{j}(A_{t})$: candidate thresholds for feature $j$ at $A_{t}$; $A^{L}_{j,s}=A_{t}\cap\{x_{j}\leqslant s\}$, $A^{R}_{j,s}=A_{t}\cap\{x_{j}>s\}$; $P_{t}(B)=\Pr(X\in B\mid X\in A_{t})\approx n(B)/n(A_{t})$
+1
+VarianceSplit (CART)
+  // cf. Fig. 2, left
+2 for $(j,s)\in\{(j,s):j\in[d],\ s\in S_{j}(A_{t})\}$ do
+3 $\mathrm{Crit}(j,s)\leftarrow P_{t}(A^{L}_{j,s})\,\mathrm{Var}(Y\mid A^{L}_{j,s})+P_{t}(A^{R}_{j,s})\,\mathrm{Var}(Y\mid A^{R}_{j,s})$
+4
 […]
-Appendix C Algorithms
-Algorithm 1 Weighted Random Forest with Multiple Splitting Criteria
-1:Training data $(\mathbf{X},Y)$, number of trees $N$, minimum samples to split, maximum depth, error method, loss norm, use weights flag
-2:Trained Random Forest model
-3:Initialize $N$ decision trees with specified parameters
-4:for $i=1$ to $N$ do
-5:     Set tree’s cyclic padding to $i$
-6:     Create bootstrap sample $(X_{i},Y_{i})$ from $(X,Y)$
-7:     Train tree using VarianceSplit, MinimaxSplit or CyclicMinimaxSplit
-8:     if use_weights is True then
-9:         Predict on $(X_{i},Y_{i})$ and calculate RMSE
-10:         Set tree weight $w_{i}=1/\text{RMSE}$ (when we set equal weights this becomes the regular random forest)
-11:     end if
-12:end for
-13:if use_weights is True then
-14:     Normalize weights: $w_{i}=w_{i}/\sum_{j=1}^{N}w_{j}$
-15:end if
-16:function Predict($X_{\mathrm{test}}$)
-17:     for $i=1$ to $N$ do
-18:         Get predictions $p_{i}$ from tree $i$ on $X_{\mathrm{test}}$
-19:     end for
-20:     if use_weights = True then return $\sum_{i=1}^{N}w_{i}p_{i}$
-21:     elsereturn $\frac{1}{N}\sum_{i=1}^{N}p_{i}$
-22:     end if
-23:end function
-Algorithm 2 Single Tree fitting with different Criteria
-1:function VarianceSplit(node)
-2:     for each feature $j$ do
-3:         Sort data points in node by feature $j$
-4:         for each potential split point do
-5:              Calculate $\mathrm{MSE}_{\mathrm{left}}\leftarrow\sum_{x\in\mathrm{left}}(y_{x}-\bar{y}_{\mathrm{left}})^{2}$
-6:              Calculate $\mathrm{MSE}_{\mathrm{right}}\leftarrow\sum_{x\in\mathrm{right}}(y_{x}-\bar{y}_{\mathrm{right}})^{2}$
-7:              Calculate total error $\leftarrow\mathrm{MSE}_{\mathrm{left}}+\mathrm{MSE}_{\mathrm{right}}$
-8:         end for
-9:         Find split point with minimum total error (with minimal number of splitting samples)
-10:     end for
-11:     return feature and split point with overall minimum error
-12:end function
-13:function MinimaxSplit(node)
-14:     for each feature $j$ do
-15:         Sort data points in node by feature $j$
-16:         for each potential split point do
-17:              Calculate $\mathrm{MSE}_{\mathrm{left}}\leftarrow\sum_{x\in\mathrm{left}}(y_{x}-\bar{y}_{\mathrm{left}})^{2}$
-18:              Calculate $\mathrm{MSE}_{\mathrm{right}}\leftarrow\sum_{x\in\mathrm{right}}(y_{x}-\bar{y}_{\mathrm{right}})^{2}$
-19:              Calculate max error $\leftarrow\max(\mathrm{MSE}_{\mathrm{left}},\mathrm{MSE}_{\mathrm{right}})$
-20:         end for
-21:         Find split point with minimum max error (with minimal number of splitting samples)
-22:     end for
-23:     return feature and split point with overall minimum max error
-24:end function
-25:function CyclicMinimaxSplit(node, depth, cyclic_padding)
+MinimaxSplit
+  // cf. Fig. 2, middle
+6 for $(j,s)\in\{(j,s):j\in[d],\ s\in S_{j}(A_{t})\}$ do
+7 $\mathrm{Crit}(j,s)\leftarrow\max\big{\{}P_{t}(A^{L}_{j,s})\,\mathrm{Var}(Y\mid A^{L}_{j,s}),\ P_{t}(A^{R}_{j,s})\,\mathrm{Var}(Y\mid A^{R}_{j,s})\big{\}}$
+8
 […]
-27:     Sort data points in node by feature feature_index
-28:     for each potential split point do
-29:         Calculate $\mathrm{MSE}_{\mathrm{left}}\leftarrow\sum_{x\in\mathrm{left}}(y_{x}-\bar{y}_{\mathrm{left}})^{2}$
-30:         Calculate $\mathrm{MSE}_{\mathrm{right}}\leftarrow\sum_{x\in\mathrm{right}}(y_{x}-\bar{y}_{\mathrm{right}})^{2}$
-31:         Calculate max error $\leftarrow\max(\mathrm{MSE}_{\mathrm{left}},\mathrm{MSE}_{\mathrm{right}})$
-32:     end for
-33:     Find split point with minimum max error (with minimal number of splitting samples)
-34:     return (feature_index, best split point)
-35:end function
+Cyclic MinimaxSplit
+  // cf. Fig. 2, right
+$j_{k}\leftarrow 1+(k\bmod d)$
+  // cyclic feature at depth $k$
+10 for $s\in S_{j_{k}}(A_{t})$ do
+11 $\mathrm{Crit}(j_{k},s)\leftarrow\max\big{\{}P_{t}(A^{L}_{j_{k},s})\,\mathrm{Var}(Y\mid A^{L}_{j_{k},s}),\ P_{t}(A^{R}_{j_{k},s})\,\mathrm{Var}(Y\mid A^{R}_{j_{k},s})\big{\}}$
+12
+$s^{\star}\leftarrow\arg\min_{s\in S_{j_{k}}(A_{t})}\mathrm{Crit}(j_{k},s)$;  split $A_{t}\to A^{L}_{j_{k},s^{\star}},\,A^{R}_{j_{k},s^{\star}}$.
+Algorithm 1 Node-level splitting rules in the notation of Section 2
+Figure 12: Comparison of node-level split selection in two random forest variants.
+Left: a random-dimension random forest selects a random feature uniformly for each node and choose the best split over the chosen feature.
+Right: a regular RF node samples $m_{\mathrm{try}}$ coordinates uniformly from the $d$ features and chooses the best split over all features according to the chosen impurity/criterion (the highlighted feature illustrates the winner).
+Input: Sample $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{N}$, number of trees $B\in\mathbb{N}$, maximum depth $K\in\mathbb{N}$, minimum leaf size $n_{\min}\in\mathbb{N}$, per-node feature budget $m_{\mathrm{try}}$, split rule $\mathrm{SplitRule}$.
+Output: Ensemble predictor $\hat{f}_{B}(x)=\frac{1}{B}\sum_{b=1}^{B}\hat{f}^{(b)}(x)$.
+1
+2for $b\leftarrow 1$ to $B$ do
+ // Bootstrap sample
+3   Draw NN indices $I^{(b)}=(i_{1},\dots,i_{N})$ uniformly with replacement from $[N]$ and set $\mathcal{D}^{(b)}=\{(X_{i_{\ell}},Y_{i_{\ell}})\}_{\ell=1}^{N}$.
+ // Grow a full tree on $\mathcal{D}^{(b)}$
+4   Initialize root node with index set $S_{\text{root}}=[N]$ and depth 0.
+5 while there exists a node $(S,\text{depth})$ with $|S|>n_{\min}$ and $\text{depth}<K$ do
+    // Sample a random subset of coordinates at this node
+6      Draw without replacement a subset $\mathcal{J}\subseteq[d]$ with $|\mathcal{J}|=m_{\mathrm{try}}\in[d]$.
+    // Choose best coordinate/threshold according to the chosen impurity criterion
+7    $(j^{\star},t^{\star})\leftarrow\mathrm{SplitRule}(S,\mathcal{J},\mathcal{D}^{(b)})$.
+    // Partition the node
+8    $S_{L}\leftarrow\{i\in S:\,X_{ij^{\star}}\leqslant t^{\star}\}$, $S_{R}\leftarrow\{i\in S:\,X_{ij^{\star}}>t^{\star}\}$.
+    // Create children; stop if a child is too small
+9    if $|S_{L}|\leqslant n_{\min}$ or $\text{depth}+1=K$ then
+10         mark left child as a leaf with prediction $\hat{\mu}_{S_{L}}=\frac{1}{|S_{L}|}\sum_{i\in S_{L}}Y_{i}$;
+11    else
+12         push $(S_{L},\text{depth}+1)$ to the queue;
+13    if $|S_{R}|\leqslant n_{\min}$ or $\text{depth}+1=K$ then
+14         mark right child as a leaf with prediction $\hat{\mu}_{S_{R}}=\frac{1}{|S_{R}|}\sum_{i\in S_{R}}Y_{i}$;
+15    else
+16         push $(S_{R},\text{depth}+1)$ to the queue;
+17
+18
+ // Define the tree predictor by leaf means
+19   For any $x\in\mathbb{R}^{d}$, let $\hat{f}^{(b)}(x)$ be the mean $\hat{\mu}_{S}$ of the leaf whose cell contains $x$.
+Algorithm 2 Random Forest with generic split rule
+Input: Sample $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{N}$, number of trees $B$, maximum depth $K$, minimum leaf size $n_{\min}$, per-node feature budget $m_{\mathrm{try}}=1$, split rule $\mathrm{SplitRule}$.
+Output: Ensemble predictor $\hat{f}_{B}(x)=\frac{1}{B}\sum_{b=1}^{B}\hat{f}^{(b)}(x)$.
+1
+2for $b\leftarrow 1$ to $B$ do
+3   Draw a bootstrap sample $\mathcal{D}^{(b)}$ of size NN from $\mathcal{D}$.
+4  Initialize the root node with index set $S_{\text{root}}=[N]$ and depth 0.
+5 while there exists a node $(S,\text{depth})$ with $|S|>n_{\min}$ and $\text{depth}<K$ do
+    // Sample a random subset of coordinates at this node
+6      Draw without replacement a subset $\mathcal{J}\subseteq[d]$ with $|\mathcal{J}|=m_{\mathrm{try}}$, where we set $m_{\mathrm{try}}=1$.
+    // Choose best coordinate/threshold using only $\mathcal{J}$
+7    $(j^{\star},t^{\star})\leftarrow\mathrm{SplitRule}(S,\mathcal{J},\mathcal{D}^{(b)})$.
+    // Partition and create children
+8    $S_{L}\leftarrow\{i\in S:\,X_{ij^{\star}}\leqslant t^{\star}\}$, $S_{R}\leftarrow\{i\in S:\,X_{ij^{\star}}>t^{\star}\}$.
+9    if $|S_{L}|\leqslant n_{\min}$ or $\text{depth}+1=K$ then
+10         make left child a leaf with prediction $\hat{\mu}_{S_{L}}$;
+11    else
+12         push $(S_{L},\text{depth}+1)$;
+13    if $|S_{R}|\leqslant n_{\min}$ or $\text{depth}+1=K$ then
+14         make right child a leaf with prediction $\hat{\mu}_{S_{R}}$;
+15    else
+16         push $(S_{R},\text{depth}+1)$;
+17
+18
+19  Define $\hat{f}^{(b)}(x)$ by the mean of the leaf containing $x$.
+Algorithm 3 Random-Dimension Random Forest (feature subsampling at each node)

### v3

@@
 The bounds control the tree’s global excess risk by local worst-case impurities and yield fast convergence rates compared to CART.
-We extend the analysis to ensembles that subsample coordinates per node.
-Empirically, (cyclic) MinimaxSplit trees and their forests improve on structured heterogeneity data such as EEG amplitude regression over fixed time horizons, seasonal air quality forecasting, and image denoising framed as non-parametric regression on spatial coordinates.
-Keywords: CART; convergence rates; decision tree models; non-parametric regression
+We extend the analysis to a random-dimension forest variant that subsamples coordinates per node.
+Empirically, (cyclic) MinimaxSplit trees and their forests improve over baselines on structured heterogeneous data such as EEG amplitude regression over fixed time horizons and image denoising, framed as non-parametric regression on spatial coordinates.
+Keywords: CART; convergence rates; decision tree models; non-parametric regression; classification
 1 Introduction
@@
 ECP has long been observed empirically (Morgan and Messenger, 1973; Torgo, 2001; Su et al., 2024; Wang, 2025) and is now rigorously characterized in simple settings, where noisy directions drive splits toward extremes (Breiman et al., 1984; Cattaneo et al., 2022).
+A simple example of ECP is given by Figure 1 below; see also Figure 4.
 For a single deep tree, ECP can be harmful: aggressive boundary splits create highly unbalanced children, amplify estimation noise, and can stall meaningful refinement along informative coordinates.
@@
 Thus, while ECP degrades a single tree’s stability, it may aid an ensemble by lowering tree correlation and enabling recovery via aggregation.
-At the same time, the forest setting introduces its own costs: ECP interacts with candidate-set multiplicity and variable-usage bias, skewing split frequencies and importance metrics, and further reducing model interpretability.
+At the same time, the forest setting introduces its own costs: ECP interacts with candidate-set multiplicity and variable-usage bias, skewing split frequencies and importance metrics, and further reducing model stability.
 Existing methods to mitigate ECP include the delta splitting rule (Morgan and Messenger, 1973), regularization of node size (Torgo, 2001), and the Smooth Sigmoid Surrogate (Su et al., 2024).
@@
 Taking the maximum over many candidate cut-off points amplifies this selection bias—extremal thresholds are disproportionately likely to ‘win’.
-Our focus is on explicitly mitigating ECP at the algorithmic level for trees, while retaining the transparency that makes single decision trees interpretable.
-Beyond prediction, this design goal is interpretability: balanced, nondegenerate splits yield partitions whose geometry is easier to understand and communicate than the heterogeneous, boundary-driven partitions that arise under ECP in deep trees and ensembles.
+Our focus is on explicitly mitigating ECP at the algorithmic level for trees, while retaining the transparency that makes single decision trees stable.
+Beyond prediction, this design goal is to generate balanced, nondegenerate splits that yield partitions whose geometry is easier to understand and communicate than the heterogeneous, boundary-driven partitions that arise under ECP in deep trees and ensembles.
 In this paper, we revisit ECP empirically and theoretically, clarifying the cases in which standard CART-style criteria drift toward extreme cut-off points and how this differs between single trees and forests (Breiman et al., 1984; Morgan and Messenger, 1973; Torgo, 2001; Su et al., 2024; Ishwaran, 2015; Wang, 2025; Cattaneo et al., 2022).
-Then, we introduce a Minimax splitting rule with a cyclic coordinate schedule that systematically discourages degenerate child nodes, thereby mitigating ECP while preserving interpretability.
-As a consequence, our splitting rule admits a faster convergence guarantee compared to standard CART — we establish exponential error decay under mild conditions and relate our guarantees to existing theory (Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Mazumder and Wang, 2024; Cattaneo et al., 2024).
-Conceptually, the analysis connects to partition-based martingale approximations (Simons, 1970; Zhang et al., 2024) and complements asymptotic and inferential developments for trees/forests (Klusowski and Tian, 2024; Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Mazumder and Wang, 2024; Cattaneo et al., 2024), as well as recent scalable partitioning frameworks (Liu et al., 2023; Luo and Pratola, 2023; Luo et al., 2024) and complexity-theoretic perspectives on hierarchical splits (Blanc et al., 2020; O’Donnell et al., 2005).
-Figure 1: The comparison of the scikit-learn DecisionTreeRegressor, our MinimaxSplit tree, and VarianceSplit tree on a sinusoidal target function (in red) $Y=f(X)+\varepsilon=\sin(4\pi X)+\varepsilon$, $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1),$ and $\varepsilon\buildrel\mathrm{law}\over{\sim}\mathcal{N}(0,0.1)$, with maximum depth 4 (top row) and 5 (bottom row).
+Then, we introduce a Minimax splitting rule with a cyclic coordinate schedule that systematically discourages degenerate child nodes, thereby mitigating ECP while preserving stability.
+As a consequence, our splitting rule admits a faster convergence guarantee compared to standard CART — we establish exponential error decay under mild conditions and relate our guarantees to existing theory (Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Wager and Walther, 2015; Mazumder and Wang, 2024; Cattaneo et al., 2024).
+Conceptually, the analysis connects to partition-based martingale approximations (Simons, 1970; Zhang et al., 2024) and complements asymptotic and inferential developments for trees/forests (Klusowski and Tian, 2024; Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Wager and Walther, 2015; Mazumder and Wang, 2024; Cattaneo et al., 2024), as well as recent scalable partitioning frameworks (Liu et al., 2023; Luo and Pratola, 2023; Luo et al., 2024; Liu and Ma, 2024) and complexity-theoretic perspectives on hierarchical splits (Blanc et al., 2020; O’Donnell et al., 2005).
+Figure 1: The comparison of the scikit-learn DecisionTreeRegressor, our MinimaxSplit tree, and VarianceSplit tree on a sinusoidal target function (in red) $Y=f(X)+\varepsilon=\sin(4\pi X)+\varepsilon$, $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1),$ and $\varepsilon\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathcal{N}(0,0.1)$, with maximum depth 4 (top row) and 5 (bottom row).
 We display 10 different model fits based on 10 different batches of training sets of size 100, and the average mean squared error (MSE) evaluated on a different training set across 10 batches.
 2 Minimax decision trees for regression
-The goal of this section is to construct MinimaxSplit decision tree algorithms that avoid ECP and attain exponential convergence.
+The goal of this section is to construct (cyclic) MinimaxSplit decision tree algorithms that avoid ECP and attain exponential convergence.
 We first focus on the regression setting and we discuss the classification setting in Section 4.
-In CART-style regression, a decision tree greedily splits in order to minimize the sum of variances over the responses of the child nodes (Liu et al., 2023; Breiman et al., 1984); instead, we propose to greedily minimize the maximum variance among the child nodes.
+In short, in CART-style regression, a decision tree greedily splits in order to minimize the sum of variances over the responses of the child nodes (Liu et al., 2023; Breiman et al., 1984); instead, we propose to greedily minimize the maximum variance among the child nodes.
+The roadmap of this section is as follows.
   • In Section 2.1, we rigorously introduce the algorithms (summarized in Figure 2).
-We study exponential rates for marginally atomless joint distributions (that may not arise from empirical measures $\mathbb{P}_{N}$) in Section 2.2.
-Section 2.3 builds upon this case, covering more generally the non-marginally atomless setting and, in particular, the empirical risk bounds.
-All proofs can be found in Appendix B.1.
-Throughout, we use the notation $d]=\{1,2,\dots,d\}$ for $d\in\mathbb{N}$.
-All splits we consider in this paper will be binary.
+Crucially, we distinguish two distinct algorithms, where we may optimize over all dimensions at each split (the MinimaxSplit algorithm) or we may pre-select a cyclic schedule of dimensions for each depth (the cyclic MinimaxSplit algorithm).
+  • In Section 2.2, we study exponential rates for marginally atomless joint distributions (that may not arise from empirical measures $\mathbb{P}_{N}$) for the cyclic MinimaxSplit algorithm.
+  • Section 2.3 builds upon the marginally atomless case, covering more generally the non-marginally atomless setting and, in particular, the empirical risk bounds for the cyclic MinimaxSplit algorithm.
+  • In Section 2.4, we discuss why there are no general analogous results for the MinimaxSplit algorithm, even though it is shown to perform well empirically even in high-dimensional settings (Appendix E.3).
+Intuitively, the MinimaxSplit algorithm may fail to break symmetry if strong symmetry is imposed in the underlying samples.
+All proofs can be found in Appendix B.1. Throughout, we use the notation $d]=\{1,2,\dots,d\}$ for $d\in\mathbb{N}$. All splits we consider in this paper will be binary.
 2.1 Algorithm formulation
@@
 To find a split of $A$ at depth $k$, the VarianceSplit algorithm looks for a dimension $j\in[d]$ and covariate $x_{j}\in[a_{j},b_{j})$ such that the remaining risk is the smallest after splitting $A$ at $x_{j}$ in covariate $j$.
-In other words, one seeks for
+In other words, one looks for
 $\begin{split}(j,\hat{x}_{j})&=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big(\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j})\\ &\hskip 56.9055pt+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big)\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big(\mathbb{E}\big[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}\big]\\ &\hskip 56.9055pt+\mathbb{E}\big[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}\big]\Big)\\ &=:\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}(V_{\text{left}}(j)+V_{\text{right}}(j)),\end{split}$ (1)
@@
 Consequently, the splittable set $A$ splits into the sets $A_{L}:=[a_{1},b_{1})\times\dots\times[a_{j},\hat{x}_{j})\times\dots\times[a_{d},b_{d})$ and $A_{R}:=[a_{1},b_{1})\times\dots\times[\hat{x}_{j},b_{j})\times\dots\times[a_{d},b_{d})$ to form its two descendants in the set $\pi_{k}$, and define
-$M_{k}(\omega)=\begin{cases}\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{L};\\ \mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{R}.\end{cases}$ (2)
+$M_{k}(\omega)=\begin{cases}\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{L};\\ \mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{R}.\end{cases}$ (2)
 In other words, at depth $k\geqslant 0$ and after constructing the partition $\pi_{k}$, define $M_{k}$ by
 $M_{k}(\omega)=\mathbb{E}[Y\mid\mathbf{X}\in A],\quad\text{if }\mathbf{X}(\omega)\in A,\quad\text{for }A\in\pi_{k}.$ (3)
-This leads to an associated process $\{M_{k}\}_{k\geqslant 0}$ to the joint distribution $(\mathbf{X},Y)$.
+This leads to an associated process $\{M_{k}\}_{k\geqslant 0}$ with the joint distribution $(\mathbf{X},Y)$.
 [^2] The process $\{M_{k}\}_{k\geqslant 0}$ is a martingale, but we will not use this fact in the rest of the main text.
-To connect $M_{k}$ with the estimator $\hat{m}_{k}$ of the regression function at step $k$, we note that for any $\mathbf{x}\in\mathbb{R}^{d}$, there exists a unique $A_{\mathbf{x}}\in\pi_{k}$ containing $\mathbf{x}$, then $\hat{m}_{k}(\mathbf{x})=\mathbb{E}[Y\mid\mathbf{X}\in A_{\mathbf{x}}]$.
+To connect $M_{k}$ with the estimator $\hat{m}_{k}$ of the regression function at step $k$, we note that for any $\mathbf{x}\in\mathbb{R}^{d}$, there exists a unique $A_{\mathbf{x}}\in\pi_{k}$ containing $\mathbf{x}$, and then $\hat{m}_{k}(\mathbf{x})=\mathbb{E}[Y\mid\mathbf{X}\in A_{\mathbf{x}}]$.
 Figure 2: The comparison of the VarianceSplit, MinimaxSplit and cyclic MinimaxSplit trees.
-Here we use the notation: $A_{j,s}^{L}=A_{t}\cap\{x_{j}<s\}$$A_{j,s}^{R}=A_{t}\cap\{x_{j}\geqslant s\}$$\mathcal{S}_{j}(A_{t})$ are candidate thresholds (e.g., midpoints of sorted $x_{j}$ in $A_{t}$); $P_{t}(B)=\mathbb{P}(\mathbf{X}\in B\mid\mathbf{X}\in A_{t})$.
+Here we use the notation: $A_{j,s}^{L}=A_{t}\cap\{x_{j}<s\}$, $A_{j,s}^{R}=A_{t}\cap\{x_{j}\geqslant s\}$; $\mathcal{S}_{j}(A_{t})$ are candidate thresholds (e.g., midpoints of sorted $x_{j}$ in $A_{t}$); $P_{t}(B)=\mathbb{P}(\mathbf{X}\in B\mid\mathbf{X}\in A_{t})$.
 The MinimaxSplit algorithm.
@@
 Instead of optimizing over all covariates $j\in[d]$ in (4), we cycle through the $d$ covariates as the tree grows.
-That is, for $k\geqslant 0$, let $j=j_{k}=1+(k\leavevmode\nobreak\ (\text{mod }d))$.
+That is, for $k\geqslant 0$, let $j=j_{k}=1+(k\penalty 10000\ (\text{mod }d))$.
 For all splittable sets $A\in\pi_{k}$, we split $A$ in the $j$-th coordinate to form its two descendants at depth $k+1$.
@@
 Assuming marginal non-atomicity, we show that up to universal constants and model mis-specification, the approximation error decays exponentially with rate $r^{k}$ for some explicit $r\in(0,1)$ that depends only on dimension $d$.
-We note here a crucial difference from the statistics literature that our analysis of the MinimaxSplit rule first focuses on general joint distributions $(\mathbf{X},Y)$ on $\mathbb{R}^{d+1}$, where $\mathbf{X}\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}$.
+We note here that a crucial difference from the statistics literature is that our analysis of the MinimaxSplit rule first focuses on general joint distributions $(\mathbf{X},Y)$ on $\mathbb{R}^{d+1}$, where $\mathbf{X}\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}$.
 [^3] This can be intuitively regarded as empirical measures with infinitely many samples.
 Empirical measures are a special case.
+Note that this does not mean that our analysis is restricted to the noiseless case; see the setting in Section 2.3.
 The class of joint distributions we consider will satisfy the following marginally atomless property.
@@
 is non-increasing and continuous.
+Remark 1.
+The marginally atomless property of $\mathbf{X}$ is not required for the monotonicity properties of the maps $\varphi_{L}$ and $\varphi_{R}$, but it is necessary for the continuity.
+The continuity is essential for an exact $1/2$ risk decay bound (see below), but this argument will be relaxed later in the derivation of Theorem 7.
 Choose an arbitrary element $\hat{x}_{j}\in\operatorname*{arg\,min}_{x_{j}}\max\{\varphi_{L}(x_{j}),\varphi_{R}(x_{j})\}.$ Let us first assume that $\varphi_{L}(\hat{x}_{j})\leqslant\varphi_{R}(\hat{x}_{j})$.
@@
 (2022).
-Let $(X,Y)$ be an empirical measure of $\mathrm{Unif}(0,1)\times\mathcal{N}(0,1)$ on $0,1]\times\mathbb{R}$ with $n$ samples.
+Let $(X,Y)$ follow the law of an empirical measure of $\mathrm{Unif}(0,1)\times\mathcal{N}(0,1)$ on $0,1]\times\mathbb{R}$ with $n$ samples.
 [^4] XX is not marginally atomless, but the example can always be modified by smoothing the data without affecting its validity.
@@
 (2022) implies that, for $n$ large enough, the split lies between the first and $\sqrt{n}$-th order statistics with probability exceeding $1/6$.
-With high probability (meaning 1o(1)1-o(1) as $n\to\infty$), the right node has remaining risk at least (1o(1))n(1-o(1))n while the total risk at the parent node is at most $(1+o(1))n$.
+With high probability (meaning 1o(1)1-o(1) as $n\to\infty$), the right node has remaining risk at least (1o(1))(1-o(1)) while the total risk at the parent node is at most $(1+o(1))$.
 By a union bound, we see that with an arbitrary $\varepsilon>0$, for sufficiently large $n$,
@@
 This example also explains why assumptions such as the sufficient impurity decay (SID) condition are favorable to ensure risk decay when using the splitting criterion (1) (Chi et al., 2022; Mazumder and Wang, 2024).
-To see further intuition of ECP, we consider the event $\{\mathbf{X}\in A\}$ and sub-events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$.
+Roughly speaking, the SID condition assumes that the best population split removes at least a positive fraction of the within-node variance, leading to a lower bound on local CART progress and eventually an exponential convergence rate in depth for CART and random forests.
+To gain further intuition about ECP, we consider the event $\{\mathbf{X}\in A\}$ and sub-events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$.
 Without loss of generality, we may assume $\mathbb{P}(\mathbf{X}\in A)=1$ and apply the total variance formula to this partition, yielding
@@
 $\begin{split}&=\mathrm{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\mathrm{Var}(Y\mid A_{R})\mathbb{P}(A_{R})\\ &\hskip 56.9055pt+\left(\mathbb{E}[Y\mid A_{L}]-\mathbb{E}[Y\mid A_{R}]\right)^{2}(1-\mathbb{P}(A_{L}))\mathbb{P}(A_{L}).\end{split}$ (12)
-From this, we see that when ECP occurs, $\mathbb{P}(A_{L})\approx 1$ or 0, the improvement is dominated by the first two terms of (12): $\mathrm{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\mathrm{Var}(Y\mid A_{R})\mathbb{P}(A_{R})=V_{\text{left}}(j)+V_{\text{right}}(j)$.
-Therefore, when ECP occurs, the sum of variances does not necessarily decay fast enough for the VarianceSplit.
-However, for (cyclic) MinimaxSplit, $\max\{V_{\text{left}}(j),V_{\text{right}}(j)\}$ always displays exponential decay (11), hence “avoiding” the ECP phenomenon.
-In other words, the fast decay of the sum of variances is not easily guaranteed from the VarianceSplit construction due to ECP; but the exponential decay of the maximum variance can be assured by the MinimaxSplit construction, as we show in Theorem 2.
+From this, we see that when ECP occurs, $\mathbb{P}(A_{L})\approx 1$ or 0, the total remaining risk is dominated by the first two terms of (12): $\mathrm{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\mathrm{Var}(Y\mid A_{R})\mathbb{P}(A_{R})=V_{\text{left}}(j)+V_{\text{right}}(j)$.
+Therefore, when ECP occurs, the sum of variances does not necessarily decay fast enough under VarianceSplit.
 We consider the additive function class
@@
 For $g\in\mathcal{G}$, define $\left\lVert g\right\rVert_{\mathrm{TV}}$ as the infimum of $\sum_{i=1}^{d}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}$ over all such additive representations of $g$, where $\left\lVert\cdot\right\rVert_{\mathrm{TV}}$ denotes the total variation of a univariate function of bounded variation.
+Note that this is the same type of approximation target used by Klusowski and Tian (2024); Mazumder and Wang (2024) for CART/C4.5: their oracle inequality is also stated as an infimum over the additive function class $\mathcal{G}$, with a penalty depending on the additive total-variation norm.
+However, for (cyclic) MinimaxSplit, $\max\{V_{\text{left}}(j),V_{\text{right}}(j)\}$ always displays exponential decay (11), hence “avoiding” the ECP phenomenon.
+In other words, the fast decay of the sum of variances is not easily guaranteed from the VarianceSplit construction due to ECP; but the exponential decay of the maximum variance can be assured by the MinimaxSplit construction, as we show in Theorem 3.
+The MinimaxSplit construction approximately equalizes the two child risks, which implies that the larger child risk is at most one half of the parent risk.
+This local worst-child contraction is the basic mechanism behind the global error bound below.
 Theorem 3.
@@
 Theorem 3 does not hold in general for $d\geqslant 2$ if we replace the cyclic MinimaxSplit algorithm by the MinimaxSplit algorithm, a phenomenon we illustrate in Section 2.4 below.
-Nevertheless, the MinimaxSplit algorithm often achieves desirable risk decay in high-dimensional settings in which cyclic MinimaxSplit is inefficient; see Appendix D.2.
+Nevertheless, the MinimaxSplit algorithm often achieves desirable risk decay in high-dimensional settings in which cyclic MinimaxSplit is inefficient; see Appendix E.3.
 The first general result for the convergence rate of the VarianceSplit algorithm without the variance decay assumption was obtained by Klusowski and Tian (2024).
@@
 $\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant\inf_{g\in\mathcal{G}}\bigg((1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+2^{1/3}(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{2/3}\\ &\hskip 113.81102pt+2^{-2/3}(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}(\Delta g)^{4/3}\bigg).\end{split}$ (15)
-Theorems 6 and 7 below enjoy the same specialized bounds.
+Theorems 7 and 8 below enjoy the same specialized bounds.
 In the remainder of this section, we explain the ideas underlying the proof of Theorem 3, in the special case where $d=1$ and $Y=g(X)$ for some $g\in\mathcal{G}$. In this case, (13) reduces to
@@
 The nodewise predictor $M_{k}=\mathbb{E}[Y\mid\mathcal{F}_{k}]$ is a Doob martingale in $L^{2}$, and its risk satisfies
-$\mathbb{E}\bigl{[}(Y-M_{k})^{2}\bigr{]}=\sum_{i=k}^{\infty}\mathbb{E}\bigl{[}(M_{i+1}-M_{i})^{2}\bigr{]}\quad\text{ and }\quad\mathbb{E}\bigl{[}(M_{k}-M_{k+1})^{2}\bigr{]}=\sum_{A\in\pi_{k}}\sum_{\begin{subarray}{c}A^{\prime}\in\pi_{k+1}\\ A^{\prime}\subset A\end{subarray}}\mathbb{P}(A^{\prime})\bigl{(}\mu_{A^{\prime}}-\mu_{A}\bigr{)}^{2}.$ (18)
+$\mathbb{E}\bigl[(Y-M_{k})^{2}\bigr]=\sum_{i=k}^{\infty}\mathbb{E}\bigl[(M_{i+1}-M_{i})^{2}\bigr]\quad\text{ and }\quad\mathbb{E}\bigl[(M_{k}-M_{k+1})^{2}\bigr]=\sum_{A\in\pi_{k}}\sum_{\begin{subarray}{c}A^{\prime}\in\pi_{k+1}\\ A^{\prime}\subset A\end{subarray}}\mathbb{P}(A^{\prime})\bigl(\mu_{A^{\prime}}-\mu_{A}\bigr)^{2},$ (18)
+where we have used that $Y$ is $\sigma(X)$-measurable.
 In other words, convergence rates reduce to controlling the one-step increments created by each split.
@@
 Example 6.
-Consider the case where $d=1$, $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$, and $Y=\sin(2\pi 2^{p}X)$.
-Figure 4 below plots the remaining risk $\mathbb{E}[(Y-M_{k})^{2}],\leavevmode\nobreak\ 1\leqslant k\leqslant K$ with power parameters $p=1,2,3,4$ and maximum tree depth $K=12$.
+Consider the case where $d=1$, $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$, and $Y=\sin(2\pi 2^{p}X)$.
+Figure 4 below plots the remaining risk $\mathbb{E}[(Y-M_{k})^{2}],\penalty 10000\ 1\leqslant k\leqslant K$ with power parameters $p=1,2,3,4$ and maximum tree depth $K=12$.
 Due to the symmetry of the sine function, the first few splits of the MinimaxSplit do not reduce the remaining risk, since the splitting points are points of symmetry.
@@
 In other words, while the MinimaxSplit algorithm may not reduce the risk as steadily as VarianceSplit, the steps with little risk reduction often implicitly overcome certain hard constraints inherent to the regression problem (such as the symmetries of the sine function), paving the way for a much faster risk decay in later stages.
-Note that the occurrence of minimal risk deduction from the first few steps of MinimaxSplit does not contradict the risk bound from Theorem 2, because the upper bound already incorporates the total variation of the signal function.
-Figure 4: Comparison of the mean squared errors for the MinimaxSplit and VarianceSplit regimes, in the setting where $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$ and $Y=\sin(2\pi 2^{p}X)$ with $p=1,2,3,4$ and maximum tree depth $12$.
+Note that the occurrence of minimal risk reduction from the first few steps of MinimaxSplit does not contradict the risk bound from Theorem 3, because the upper bound already incorporates the total variation of the signal function.
+Figure 4: Comparison of the mean squared errors for the MinimaxSplit and VarianceSplit regimes, in the setting where $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$ and $Y=\sin(2\pi 2^{p}X)$ with $p=1,2,3,4$ and maximum tree depth $12$.
 2.3 Empirical risk bound for cyclic MinimaxSplit
@@
 We now focus on the regression setting and derive finite-sample performance guarantees.
-Suppose that under the original law $\mathbb{P}_{*}$, $\mathbf{X}_{*}$ is marginally atomless and $Y_{*}\mid\mathbf{X}_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon:=\mathbb{E}[Y_{*}\mid\mathbf{X}_{*}]+\varepsilon,$ where the error $\varepsilon=Y_{*}-g_{*}(\mathbf{X}_{*})$ is sub-Gaussian, i.e., for some $\sigma>0$, $\mathbb{P}_{*}(|\varepsilon|\geqslant u)\leqslant 2\exp(-{u^{2}}/({2\sigma^{2}})),\leavevmode\nobreak\ u\geqslant 0.$ Also, we consider the empirical law of $(\mathbf{X},Y)$ of $N$ samples from $\mathbb{P}_{*}$.
+Suppose that under the original law $\mathbb{P}_{*}$, $\mathbf{X}_{*}$ is marginally atomless and $Y_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon,$ where $g_{*}(\mathbf{X}_{*}):=\mathbb{E}[Y_{*}\mid\mathbf{X}_{*}]$ and the error $\varepsilon=Y_{*}-g_{*}(\mathbf{X}_{*})$ is sub-Gaussian, i.e., for some $\sigma>0$, $\mathbb{P}_{*}(|\varepsilon|\geqslant u)\leqslant 2\exp(-{u^{2}}/({2\sigma^{2}})),\penalty 10000\ u\geqslant 0.$ Also, we consider the empirical law of $(\mathbf{X},Y)$ based on $N$ samples from $\mathbb{P}_{*}$.
 A limitation of Theorem 3 is that it applies only if the covariate $\mathbf{X}$ is marginally atomless (for (10) to hold).
@@
 For a random variable $Y$, denote by $\operatorname{supp}Y$ the support of $Y$.
+When the marginal law of $X$ has atoms, the continuity argument used in the proof of the exact worst-child halving property is no longer available.
+Nevertheless, if every coordinate atom has mass at most $1/N$, then the failure of exact balancing can be quantified by an additive perturbation of order $N^{-1}$.
+The next theorem shows that the population bound from the atomless case survives, up to a discrete-resolution correction term.
 Theorem 7.
-Suppose that $\mathbf{X}$ is purely atomic and for some $N>0$,
+Suppose that $\mathbf{X}$ is purely atomic and has a finite support, and for some $N>0$,
 $\max_{j\in[d]}\,\max_{u\in\mathbb{R}}\mathbb{P}(X_{j}=u)\leqslant\frac{1}{N}.$ (22)
@@
 $\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}(\Delta Y)^{2}}{3N}+\inf_{g\in\mathcal{G}}\bigg(\Big((1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]\\ &\hskip 170.71652pt+(1+\delta^{-1})\bigg(\frac{2}{3}+\Big(\frac{1+\delta^{-1}}{4}\Big)^{2/3}\bigg)2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg).\end{split}$ (23)
-Our next key result is the following oracle inequality for decision trees under model mis-specification (i.e., when $g_{*}$ does not belong to $\mathcal{G}$).
-Let $g_{k}$ be the output of the cyclic MinimaxSplit construction (that is, $M_{k}=g_{k}(\mathbf{X})$).
-Note that $g_{k}$ is random under the law $\mathbb{P}_{*}$.
-We use $\left\lVert\cdot\right\rVert$ to denote the $L^{2}$ norm in $\mathbb{P}_{*}$, e.g.
-$\left\lVert g_{k}-g_{*}\right\rVert^{2}=\mathbb{E}[(g_{k}(\mathbf{X}_{*})-g_{*}(\mathbf{X}_{*}))^{2}]$, where the expectation is taken in $\mathbb{P}_{*}$.
+Comparing (23) and (13), we see that there is an additional compensation term of $(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}(\Delta Y)^{2}}{3N}$ for the atomic case.
+This arises because the functions $\varphi_{L}$ and $\varphi_{R}$ are no longer continuous in Lemma 2.1, meaning that the maximum risk decay may not always achieve the $1/2$ rate; see Remark 1.
+The previous results control the approximation error of the population partition generated by cyclic MinimaxSplit.
+To obtain a finite-sample oracle inequality, one combines that deterministic approximation bound with the same truncation and empirical-process steps used in the CART analysis of Klusowski and Tian (2024).
+The effect of the MinimaxSplit rule enters only through the improved approximation term.
 Theorem 8.
@@
 where $C>0$ is a constant depending only on $\left\lVert g_{*}\right\rVert_{\infty}$ and $\sigma^{2}$.
-For a fixed sample size $N$, if we assume that $g_{*}\in\mathcal{G}$ and $k=3d\log_{2}N/(3d+2)$ is a multiple of $d$, Theorem 7 has the further consequence that (under the same assumptions)
-$\mathbb{E}\big{[}\left\lVert g_{k}-g_{*}\right\rVert^{2}\big{]}\leqslant CN^{-\frac{2}{3d+2}}\big{(}\left\lVert g_{*}\right\rVert_{\mathrm{TV}}+(\log N)^{2}\log(Nd)\big{)}.$ (24)
+For a fixed sample size $N$, if we assume that $g_{*}\in\mathcal{G}$ and $k=\lceil(3d\log_{2}N)/(3d+2)\rceil$ is a multiple of $d$, Theorem 8 has the further consequence that (under the same assumptions)
+$\mathbb{E}\big[\left\lVert g_{k}-g_{*}\right\rVert^{2}\big]\leqslant CN^{-\frac{2}{3d+2}}\big(\left\lVert g_{*}\right\rVert^{2}_{\mathrm{TV}}+(\log N)^{2}\log(Nd)\big).$ (24)
 The proof of Theorem 8 follows essentially the same path as Theorem 4.3 for CART of Klusowski and Tian (2024) while replacing Theorem 4.2 therein by our Theorem 7.
+As explained in Remark 4, Theorem 8 does not hold without extra assumptions if the cyclic MinimaxSplit algorithm is replaced by the MinimaxSplit algorithm.
+On another note, in high-dimensional settings, the MSE decay rate $2^{-2\lfloor k/d\rfloor/3}$ deteriorates under the cyclic schedule.
+Nevertheless, we provide numerical support in Appendix E.3 that the MinimaxSplit algorithm in high dimensions still often outperforms other baseline models such as scikit-learn and VarianceSplit.
 In addition to the performance and convergence rate guarantees, we briefly comment on the computational efficiency of MinimaxSplit compared to VarianceSplit.
@@
 Example 9 also generalizes to higher dimensions, such as using the function $Y=f(\mathbf{X})$ where $f(x_{1},\dots,x_{d})=x_{1}+\dots+x_{d-1}+|x_{d}|$.
-This ASBP phenomenon has an important consequence of the inadmissibility of classic random forests in the context of MinimaxSplit.
-We follow the setting of Section 7 of Klusowski and Tian (2024), which is a slight adaptation of the Breiman’s classic random forest (Breiman, 2001).
+This ASBP phenomenon has an important consequence for the inadmissibility of classic random forests in the context of MinimaxSplit.
+We follow the setting of Section 7 of Klusowski and Tian (2024), which is a slight adaptation of Breiman’s classic random forest (Breiman, 2001).
 A random forest consists of a collection of independent subsampled trees, where at each node, the split dimensions are optimized among a random pre-selected set $\mathcal{S}\subseteq[d]$.
@@
 In this section, we extend the MinimaxSplit and cyclic MinimaxSplit methods to an ensemble context.
-The ensemble method we consider here is the original Breiman’s random forest (Breiman et al., 1984), which aggregates multiple decision trees to improve the precision and robustness of the prediction.
+The ensemble method we consider here is Breiman’s classic random forest (Breiman, 2001), which aggregates multiple decision trees to improve the precision and robustness of the prediction.
 Each tree is grown to a fixed depth $k$ according to some splitting rule, and the split dimension of each node is optimized among independent selections of random subsets of $[d]$ with a fixed cardinality $m_{\mathrm{try}}$.
-Each tree in the ensemble is trained on a bootstrap sample of the original dataset and predictions are made by averaging the output of individual trees, effectively reducing the overall variance of the model (see Algorithm 1 in Appendix E).
+Each tree in the ensemble is trained on a bootstrap sample, yet our theoretical results do not consider bootstrap for simplicity.
+Predictions are made by averaging the output of individual trees, effectively reducing the overall variance of the model (see Algorithm 1 in Appendix F).
 We call the ensemble the MinimaxSplit (resp.
@@
 3.2 Empirical risk bound for MinimaxSplit random-dimension random forests
-Our discussions in Section 2.4 indicate that the random forest approach for the MinimaxSplit algorithms is better performed with $m_{\mathrm{try}}=1$ to guarantee consistency.
+Our discussion in Section 2.4 indicates that the random forest approach for the MinimaxSplit algorithms is better implemented with $m_{\mathrm{try}}=1$ to guarantee consistency.
 In other words, we randomize the dimension at every node.
-As a consequence, applying the MinimaxSplit and the cyclic MinimaxSplit splitting rules yield the same ensemble in distribution.
+As a consequence, applying the MinimaxSplit and the cyclic MinimaxSplit rules yield the same ensemble in distribution.
 We call the resulting ensemble the MinimaxSplit random-dimension random forest.
@@
 Let $\Sigma$ denote the law of the splitting dimensions and $\mathbb{E}_{n}$ denote the expectation with respect to the empirical measure on $n$ samples from $\Sigma$.
-Let $g_{k}$ be the estimator constructed at deapth $k$ from the MinimaxSplit random-dimension random forest.
-Our goal is then to bound the $L^{2}$ error $\mathbb{E}\big{[}\left\lVert\mathbb{E}_{n}[g_{k}]-g^{*}\right\rVert^{2}\big{]},$where the outer expectation is taken in terms of randomness both from the empirical samples from $(\mathbf{X}_{*},Y_{*})$ and from the empirical measure $\mathbb{E}_{n}$.
-We will provide empirical evidence of the effectiveness of such a randomised construction in Section 5.3.
+Let $g_{k}$ be the estimator constructed at depth $k$ from the MinimaxSplit random-dimension random forest.
+Our goal is then to bound the $L^{2}$ error $\mathbb{E}\big[\left\lVert\mathbb{E}_{n}[g_{k}]-g_{*}\right\rVert^{2}\big],$ where the outer expectation is taken in terms of randomness both from the empirical samples from $(\mathbf{X}_{*},Y_{*})$ and from the empirical measure $\mathbb{E}_{n}$.
+We will provide empirical evidence of the effectiveness of such a randomized construction in Section 5.3.
 Theorem 10.
-Assume the same regression setting of Theorem 7, and the above random forest definition. We have
-$\begin{split}\mathbb{E}\big{[}\left\lVert\mathbb{E}_{n}[g_{k}]-g^{*}\right\rVert^{2}\big{]}\leqslant&\frac{C2^{k}(\log N)^{2}\log(Nd)+2^{k}(\log N)dn^{-1/3}}{N}\\ &+2\inf_{g\in\mathcal{G}}\bigg{(}2\left\lVert g-g^{*}\right\rVert^{2}+\frac{8}{3}\Big{(}\left\lVert g-g^{*}\right\rVert^{2}+\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\Big{)}\Big{(}e^{-\frac{k}{6d}}+\frac{d}{n^{1/3}}\Big{)}\bigg{)},\end{split}$ (26)
-where $C>0$ is a constant depending only on $\left\lVert g^{*}\right\rVert_{\infty}$ and $\sigma^{2}$.
+Assume the same regression setting as in Theorem 8, and the above random forest definition. We have
+$\begin{split}\mathbb{E}\big[\left\lVert\mathbb{E}_{n}[g_{k}]-g_{*}\right\rVert^{2}\big]\leqslant&\frac{C2^{k}(\log N)^{2}\log(Nd)+2^{k}(\log N)dn^{-1/3}}{N}\\ &+2\inf_{g\in\mathcal{G}}\bigg(2\left\lVert g-g_{*}\right\rVert^{2}+\frac{8}{3}\Big(\left\lVert g-g_{*}\right\rVert^{2}+\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\Big)\Big(e^{-\frac{k}{6d}}+\frac{d}{n^{1/3}}\Big)\bigg),\end{split}$ (26)
+where $C>0$ is a constant depending only on $\left\lVert g_{*}\right\rVert_{\infty}$ and $\sigma^{2}$.
 In particular, if $N/(2^{k}(\log N)^{2}\log(Nd))\to\infty$, $k/d\to\infty$, and $n/d^{3}\to\infty$, then the MinimaxSplit random-dimension random forest is consistent.
@@
 Second, the formula (26) does not involve $\delta$ for simplicity, although the same argument works for more general choices of $\delta>0$.
+The restriction $m_{\mathrm{try}}=1$ is theoretically motivated: for $m_{\mathrm{try}}\geqslant 2$, Example 9 shows that the obstruction to extending Theorem 3 from cyclic MinimaxSplit to the unrestricted MinimaxSplit rule is not the local minimax contraction itself, but rather a coordinate-starvation phenomenon: a signal-bearing coordinate may never be selected along some root-to-leaf paths.
+The next condition rules this out at the level of the selected split coordinate.
+Let $\Xi$ denote all randomization used by the tree construction (candidate-set sampling and tie-breaking), and let $\{\pi_{k}^{\Xi}\}_{k\geqslant 0}$ be the associated random partition sequence.
+For a splittable cell $A\in\pi_{k}^{\Xi}$, write $J^{\Xi}(A)\in[d]$ for the coordinate along which $A$ is split.
+Definition 3.1 (Positive visibility on $S$).
+Fix $S\subseteq[d]$ and $\rho\in(0,1]$.
+We say that the randomized MinimaxSplit tree has positive visibility $\rho$ on $S$ if for every $k\geqslant 0$, every splittable cell $A\in\pi_{k}^{\Xi}$, and every $j\in S$,
+$\mathbb{P}\!\left(J^{\Xi}(A)=j\,\middle|\,\pi_{k}^{\Xi}\right)\geqslant\rho\qquad\text{a.s.}$
+In the following, for $S\subseteq[d]$, we consider the additive function class
+$\mathcal{G}_{S}:=\Big\{g(\mathbf{x}):=\sum_{j\in S}g_{j}(x_{j})\Big\},$
+where $\mathbf{x}=(x_{1},\dots,x_{d})$.
+Theorem 11.
+Suppose that $\mathbf{X}$ is marginally atomless.
+Let $\{M_{k}^{\Xi}\}_{k\geqslant 0}$ be the predictor sequence associated with a randomized MinimaxSplit tree, and assume that the tree has positive visibility $\rho\in(0,1]$ on $S\subseteq[d]$ in the sense of Definition 3.1.
+Then uniformly for any $\delta>0$ and $k\geqslant 0$,
+$\begin{split}\mathbb{E}[(Y-M_{k}^{\Xi})^{2}]\leqslant\inf_{g\in\mathcal{G}_{S}}\Bigg(\left((1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3}e^{-\rho k/3}\right)\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\\ (1+\delta^{-1})\left(\frac{1}{3}+\left(\frac{1+\delta^{-1}}{4}\right)^{2/3}\right)e^{-\rho k/3}\|g\|_{\mathrm{TV}}^{2}\Bigg),\end{split}$ (27)
+where the expectation is taken over both $(\mathbf{X},Y)$ and the tree randomization $\Xi$.
+In particular, if $\rho$ is bounded below by a positive multiple of $1/s$, where $s:=|S|$, then the approximation error decays at rate $\exp(-ck/s)$ for a universal constant $c>0$.
+This reveals that the stochastic visibility scale $1/\rho$ for randomized trees with $m_{\mathrm{try}}\geqslant 2$ is essential to ensure risk bounds.
+This kind of assumption shares the spirit of mild design assumptions used by (Ročková and Saha, 2019).
+The same observation also shows that randomized MinimaxSplit adapts to exact sparsity under the positive visibility condition.
+Indeed, as the approximation error decays at rate $\exp(-ck/s)$, the effective dimension is $s$ rather than the ambient dimension $d$.
 4 Minimax decision trees for classification
@@
 As in the regression setting, when splitting at level $k$, the cyclic MinimaxSplit variant fixes the coordinate at level $k$ to $j_{k}=1+(k\ (\mathrm{mod}\ d))$ and optimizes only over the threshold $z$.
-Given a split rule, we obtain a nested sequence of axis-aligned partitions $\{\pi_{k}\}_{k\geqslant 0}$ by applying a single binary split per level.
+Given a split rule, we obtain a nested sequence of axis-aligned partitions $\{\pi_{k}\}_{k\geqslant 0}$ by applying a single binary split per node per level.
 For any $\omega$, we let $A_{k}(\omega)\in\pi_{k}$ be the unique cell containing $\mathbf{X}(\omega)$.
@@
 $M_{k}(\omega)\;\coloneqq\;\log\Big(\frac{\eta_{A_{k}(\omega)}}{1-\eta_{A_{k}(\omega)}}\Big).$ (29)
-Let $g_{k}(\mathbf{x}):=\log(\eta_{A_{k}}/(1-\eta_{A_{k}}))$ for $\mathbf{x}\in A_{k},\leavevmode\nobreak\ A_{k}\in\pi_{k}$.
+If a cell is pure, we allow $M_{k}=\pm\infty$ in the natural way; then $\log(1+e^{-YM_{k}})=0$ on that cell.
+Let $g_{k}(\mathbf{x}):=\log(\eta_{A_{k}}/(1-\eta_{A_{k}}))$ for $\mathbf{x}\in A_{k},\penalty 10000\ A_{k}\in\pi_{k}$.
 The predictor associated with depth $k$ is the plug-in classifier $2{\mathbbm{1}}_{\{g_{k}(\mathbf{x})\geqslant 0\}}-1$ (or equivalently, $2{\mathbbm{1}}_{\{\eta_{A_{k}}\geqslant{1}/{2}\}}-1$ for the unique $A_{k}\in\pi_{k}$ that contains $\mathbf{x}$) and the canonical convex surrogate is the conditional cross-entropy on the partition.
 With this preparation, the results in this section follow the same logical arc as in the regression setting.
-Theorem 10 treats the marginally atomless case and shows that, under the cyclic MinimaxSplit schedule, the log-odds process $\{M_{k}\}$ contracts in mean square at an exponential rate that depends only on the dimension and the coordinate range.
+Theorem 12 treats the marginally atomless case and shows that, under the cyclic MinimaxSplit schedule, the cross entropy contracts exponentially in $\lfloor k/d\rfloor$.
 Theorem 13 covers the complementary purely atomic setting.
-Here, we need a mild anti-concentration assumption on coordinate masses; the theorem then delivers an explicit upper bound for $\mathbb{E}[(Y-M_{k})^{2}]$ that reflects the discrete resolution $N$ of the atoms.
+Here, we need a mild anti-concentration assumption on coordinate masses; the theorem then delivers an explicit upper bound for $\mathbb{E}[\log(1+e^{-YM_{k}})]$ that reflects the discrete resolution $N$ of the atoms.
 Finally, Theorem 14 translates these bounds into an oracle inequality for the excess classification risk of $g_{k}$ under the 0-1 loss via standard calibration between cross-entropy and misclassification.
@@
 Then uniformly for any $k\geqslant 0$,
-$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant\inf_{g\in\mathcal{G}}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}\Big{)}.\end{split}$ (29)
+$\begin{split}\mathbb{E}[\log(1+e^{-YM_{k}})]&\leqslant\inf_{g\in\mathcal{G}}\Big(\mathbb{E}[\log(1+e^{-Yg(\mathbf{X})})]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}\Big).\end{split}$ (30)
 Theorem 13.
-Suppose that $\mathbf{X}$ is purely atomic and for some $N>0$,
+Suppose that $\mathbf{X}$ is purely atomic and has a finite support, and for some $N>0$,
 $\max_{j\in[d]}\,\max_{u\in\mathbb{R}}\mathbb{P}(X_{j}=u)\leqslant\frac{1}{N}.$ (31)
 Let the associated process $\{M_{k}\}_{k\geqslant 0}$ be constructed from (29) using the cyclic MinimaxSplit algorithm. Then uniformly for any $k\geqslant 0$,
-$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant 2^{-\lfloor k/d\rfloor/2}\frac{2^{k}\log N}{N}+\inf_{g\in\mathcal{G}}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}(\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}+\frac{\left\lVert g\right\rVert_{\mathrm{TV}}}{2})\Big{)}.\end{split}$ (31)
-The proofs of Theorems 10 and 11 follow essentially the same principles as in the regression setting, except that now $\{M_{k}\}_{k\geqslant 0}$ no longer forms a martingale.
+$\begin{split}\mathbb{E}[\log(1+e^{-YM_{k}})]&\leqslant 2^{-\lfloor k/d\rfloor/2}\frac{2^{k}\log(eN)}{N}+\inf_{g\in\mathcal{G}}\Big(\mathbb{E}[\log(1+e^{-Yg(\mathbf{X})})]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}(\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}+\frac{\left\lVert g\right\rVert_{\mathrm{TV}}}{2})\Big).\end{split}$ (32)
+The proofs of Theorems 12 and 13 follow essentially the same principles as in the regression setting, except that now $\{M_{k}\}_{k\geqslant 0}$ no longer forms a martingale.
 The key is to apply the monotonicity of the map $z\mapsto\mathbb{P}(\mathbf{X}\in A,\,X_{j}<z)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A,\,X_{j}<z))$, a fact that mirrors Lemma 2.1 and will be justified in Lemma B.5 below.
@@
 Following Klusowski and Tian (2024), for a map $g$ we consider the misclassification risk $\mathrm{Err}(g):=\mathbb{P}(Y\neq 2{\mathbbm{1}}_{\{g(\mathbf{X})\geqslant 0\}}-1).$ Further, we set $g_{*}(\mathbf{x}):=\log(\mathbb{P}(Y=1\mid\mathbf{X}=\mathbf{x})/\mathbb{P}(Y=-1\mid\mathbf{X}=\mathbf{x}))$.
+Recall that $N$ is the number of samples, $d$ is the underlying dimension, and $g_{k}$ is the output of the cyclic MinimaxSplit construction with depth $k$.
 Theorem 14.
@@
 Compared to the oracle inequality of Klusowski and Tian (2024), we see that we have achieved exponential convergence rates and removed the factor $\left\lVert g\right\rVert_{\infty}$.
-The proof of Theorem 12 follows in exactly the same way as Theorem 4.3 of Klusowski and Tian (2024) in the C4.5 case, as the arguments therein do not depend on how the tree is constructed, as long as the splitting rule is axis-aligned.
+We provide additional classification experiments using MinimaxSplit in Appendix D.
 5 Numerical experiments and applications
 In this section, we provide experiments that verify the advantages of MinimaxSplit and cyclic MinimaxSplit methods in low-dimensional domains, as well as integrated ensemble methods applied to approximating a real-valued function and denoising images.
-Performance metrics are defined in Appendix D.1, and further numerical experiments (including those in high dimensions) are provided in Appendix D.
+Performance metrics are defined in Appendix E.1, and further numerical experiments (including those in high dimensions) are provided in Appendix E.
 5.1 Avoiding end-cut preference for stable cuts
-To show that the ECP is remedied in MinimaxSplit, we consider a special case of the location regression model in (Cattaneo et al., 2022), similar to the discussions in Example 1.
-In our experiment, we generate independent replicates with $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$ and $Y$ following the standard normal or Student’s $t(\cdot)$ distributions (pure noise), sort by $X$, and scan all possible splits in the experiments in Figure 5.
-We also provide two baselines: Uniform random cutting picks a split point entirely at random within the current predictor range, ensuring that there is no systematic bias yet, ignoring any data-driven signal.
+To show that the ECP is remedied in MinimaxSplit, we consider a special case of the location regression model in (Cattaneo et al., 2022), similar to the discussion in Example 2.
+In our experiment, we generate independent replicates with $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$ and $Y$ following the standard normal or Student’s $t$ distributions (pure noise), sort by $X$, and scan all possible splits in the experiments in Figure 5.
+We also provide two baselines: Uniform random cutting picks a split point entirely at random within the current predictor range, ensuring that there is no systematic bias, yet ignoring any data-driven signal.
 Sampling an observed value similarly randomizes split location among actual samples.
-Both serve as baseline benchmarks: they eliminate ECP, but do not offer an impurity-guided improvement mechanism.
+Both serve as benchmarks: they eliminate ECP, but do not offer an impurity-guided improvement mechanism.
 For each replicate, we record ${\min(n_{L},n_{R})}/{n}$, the relative size of the smaller child, and we choose either VarianceSplit (the standard CART impurity reduction) or MinimaxSplit.
@@
 Thus, if one requires a data‐adaptive binary split that resists extreme outliers without surrendering balance, the MinimaxSplit criterion clearly outperforms VarianceSplit (or CART) and random baselines.
-Figure 5: Empirical distributions of sample size $n=500$ over $1000$ replicates of the smaller-child proportion $\min(n_{L},n_{R})/n$ when splitting pure-noise data generated by standard normal (left), $t(3)$ (middle), and $t(1)$ (right) distributions.
+Figure 5: Empirical CDFs of the smaller child proportion $\min(n_{L},n_{R})/n$ for sample size $n=500$, based on $1000$ replicates when splitting pure-noise data generated by standard normal (left), $t(3)$ (middle), and $t(1)$ (right) distributions.
 We also use dashed lines to denote the threshold 0.05.
-Figure 6: Average leaf size (mean number of samples per leaf) versus maximum tree depth for four splitting criteria (VarianceSplit (CART), MinimaxSplit, C1, C2), at two noise levels.
+Figure 6: Average leaf size (mean number of samples per leaf) versus maximum tree depth for six splitting criteria (VarianceSplit (CART), MinimaxSplit, C1, C2 (Buja and Lee, 2001), CovRT (Zhang and Ma, 2026), and alpha-CART Wager and Walther (2015)) at two noise levels.
 Error bars show one standard deviation across repeated fits.
@@
 Small average leaf node sizes and collapsed error bars signal balanced partitions; large means with huge variance indicate persistent end-cuts leaving one dominant branch and tiny fragments.
-Our main methodological contribution, the MinimaxSplit algorithm can be used with ensembles and is applicable in real data, but also yields a theoretically stronger version of Buja and Lee (2001)’s formula C1 and C2, where they showed that the minimum variance (a.k.a., one-sided purity) approach can identify extreme subsets much better than traditional CART.
-Figure 6 records the leaf distributions as the maximum depth of the tree grows, with $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$ and signal
+Our main methodological contribution, the MinimaxSplit algorithm, can be used with ensembles and is applicable to real data.
+It is also related in spirit to the C1 and C2 criteria of Buja and Lee (2001), which emphasize one-sided purity and can better identify extreme subsets than traditional CART.
+Figure 6 records the leaf distributions as the maximum depth of the tree grows, with $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$ and signal
 $f(x)=\begin{cases}\sin(x),&0\leqslant x<\frac{1}{3};\\ -2x,&\frac{1}{3}\leqslant x<\frac{2}{3};\\ 0,&\text{otherwise}.\end{cases}$ (34)
@@
 The results align with the theoretical expectations.
-As the tree depth increases, CART and the sum-of-variances baseline increasingly display end-cut behavior, producing piecewise constant fits that flatten toward the edges of the interval and degrade extrapolation between local extrema.
+As the tree depth increases, CART and the sum-of-variances baselines increasingly display end-cut behavior, producing piecewise constant fits that flatten toward the edges of the interval and degrade extrapolation between local extrema.
 MinimaxSplit, by contrast, maintains near-balanced children throughout the path, which yields visibly tighter fits to the oscillatory structure of the signal and systematically lower errors.
@@
 Figure 8: Comparative analysis of various random forest (depth $K=10$, 50 estimators) approaches for image denoising using different splitting strategies.
-The experiment compares six random forest configurations on a noisy astronaut image: original noiseless image; noisy input; three greedily splitting methods ($m_{\mathrm{try}}=d$) using consistent splitting criteria across all weak learners (trees) — VarianceSplit, MinimaxSplit, and Cyclic MinimaxSplit; three random-dimension methods where feature selection is randomized at each node (i.e., $m_{\mathrm{try}}=1$) while location selection follows the three different criteria, as described in Section 3.2.
+The experiment compares six random forest configurations on a noisy astronaut image: original noiseless image; noisy input; three greedy splitting methods ($m_{\mathrm{try}}=d$) using consistent splitting criteria across all weak learners (trees) — VarianceSplit, MinimaxSplit, and Cyclic MinimaxSplit; three random-dimension methods where feature selection is randomized at each node (i.e., $m_{\mathrm{try}}=1$) while location selection follows the three different criteria, as described in Section 3.2.
 Figure 9: Comparison of decision tree regression methods (depth $K=10$) for image reconstruction using different splitting criteria — VarianceSplit, MinimaxSplit, and cyclic MinimaxSplit.
 Each method was trained on noisy astronaut image data with spatial coordinates as features and pixel intensities as targets.
-Another notable experiment involves applying decision trees to denoise noisy images (Luo et al., 2024).
+Another notable experiment involves applying decision trees to denoise images (Luo et al., 2024).
 We apply the different tree variants to predict pixel values based on their locations, effectively treating this as a regression problem.
-The experiments shown in Figures 8 and 9 compare decision tree and random forest regression methods for denoising the classic Astronaut (Van der Walt et al., 2014) grayscale image, which has been preprocessed to 256 by 256.
+The experiments shown in Figures 8 and 9 compare decision tree and random forest regression methods for denoising the classic Astronaut (Van der Walt et al., 2014) grayscale image, which has been preprocessed to a $256\times 256$ image.
 The goal is to recover a clean grayscale image $f:[0,1]^{2}\to[0,1]$ from noisy observations $Y=f(\mathbf{X})+\varepsilon$ at pixel centers $\mathbf{X}=(X_{1},X_{2})$ laid out on a $256\times 256$ grid (equipped with the uniform distribution).
 We normalize intensities to $[0,1]$ and add i.i.d.
-Gaussian noise $\varepsilon\buildrel\mathrm{law}\over{\sim}\mathcal{N}(0,\,0.1^{2})$.
+Gaussian noise $\varepsilon\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathcal{N}(0,\,0.1^{2})$.
 Each pixel becomes one regression sample, and we fit axis-aligned trees that approximate the unknown $f$ as a piecewise constant function on a dyadic partition, which is precisely the regime covered by our analysis of axis-aligned trees and their ensembles.
@@
 In contrast, both Minimax ensembles suppress speckle while preserving high-contrast features such as facial contours and texture edges.
-Further, across all metrics, the MinimaxSplit forests outperforms the VarianceSplit forest, with the largest gains in SSIM, indicating better structural fidelity.
+Further, across all metrics, the MinimaxSplit forests outperform the VarianceSplit forest, with the largest gains in SSIM, indicating better structural fidelity.
 The improvements are coherent with Theorem 10: the worst-child control given by the MinimaxSplit rule prevents high-variance children from persisting across levels.
@@
 The cyclic MinimaxSplit algorithm further ensures that each dimension is used in a balanced manner throughout the tree construction process, avoiding dominance by a subset of features.
-We do not claim that this new method could outperform the classical CART-like splitting under all circumstances, yet we point out that our construction avoids ECP and attains a better convergence rate guarantee as the tree progresses deeper.
+We do not claim that this new method could outperform the classical CART-like splitting under all circumstances, yet we point out that our construction avoids ECP and attains a better convergence rate guarantee as the tree grows deeper.
 The (non-cyclic) MinimaxSplit construction features the anti-symmetry-breaking preference—consistency is not guaranteed in certain symmetric cases, which can, however, be alleviated by using ensemble methods.
-A key theoretical methodological development of this study is that the cyclic MinimaxSplit algorithm achieves an exponential MSE decay rate given sufficient data samples, without requiring additional variance decay assumptions (Theorem 6) in regression and classification settings.
-Furthermore, we derive empirical risk bounds for this method, establishing its robustness in different settings (Theorem 7).
-Using the same powerful techniques used for trees, we prove novel results on the convergence rates of univariate partition-based martingale approximations as a by-product (see Appendix A), which are of their own interest.
+A key theoretical methodological development of this study is that the cyclic MinimaxSplit algorithm achieves an exponential MSE (or logistic loss) decay rate given sufficient data samples, without requiring additional variance decay assumptions in regression and classification settings (Theorems 7 and 13).
+Furthermore, we derive empirical risk bounds for this method, establishing its robustness in different settings (Theorems 8 and 14).
+Using the same powerful techniques used for trees, we prove novel results on the convergence rates of univariate partition-based martingale approximations as a by-product (see Appendix A), which are of independent interest.
 In addition to single-tree regression analysis, we explore ensemble learning approaches, where we combine decision trees constructed using MinimaxSplit, cyclic MinimaxSplit, and VarianceSplit methods within a random forest framework.
-The results demonstrate that a hybrid approach leveraging different splitting techniques yields superior performance, particularly when dealing with non-uniform data distributions (Section 3).
+We also demonstrate a hybrid approach with different splitting techniques at different layers when dealing with non-uniform data distributions (Section E.2).
 These findings suggest that the MinimaxSplit approach can provide more stable and adaptive decision trees compared to traditional methods, especially in low-dimensional settings.
+In future work, we hope to introduce changes in the cut family and search scope (Ishwaran, 2026) to enhance its high-dimensional behavior.
 Acknowledgment
@@
 In this section, we develop the framework of partition-based martingale approximations and show that a number of constructions feature exponential convergence rates.
-These problems are of their own interest (see the end of Appendix A.1) and have implications in the regression tree setting.
+These problems are of independent interest (see the end of Appendix A.1) and have implications in the regression tree setting.
 The main takeaway is that:
@@
 Define $M_{k}:=\mathbb{E}[U\mid\sigma(\pi_{k})]$.
-If $\Pi:=\{\pi_{k}\}_{k\geqslant 0}$ is a nested sequence of partitions, $\{\sigma(\pi_{k})\}_{k\geqslant 0}$ is a filtration generated by indicator functions of sets in $\pi_{k}$’s (Doob, 1953).
+If $\Pi:=\{\pi_{k}\}_{k\geqslant 0}$ is a nested sequence of partitions, $\{\sigma(\pi_{k})\}_{k\geqslant 0}$ is a filtration generated by indicator functions of sets in the partitions $\{\pi_{k}\}_{k\geqslant 0}$ (Doob, 1953).
 It follows from the tower property of conditional expectations that the sequence $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ is a martingale (in fact, a Doob martingale (Doob, 1940, 1953)).
@@
     The median splitting rule splits $I$ into two sets $a,u_{\text{median}})$ and $u_{\text{median}},b)$.
-In turn, when applying recursively each splitting rule starting from $\pi_{0}=\{\mathbb{R}\}$, we obtain a nested sequence of partitions $\Pi=\{\pi_{k}\}_{k\geqslant 0}$ depending on the law of $U$.
+In turn, when applying each splitting rule recursively starting from $\pi_{0}=\{\mathbb{R}\}$, we obtain a nested sequence of partitions $\Pi=\{\pi_{k}\}_{k\geqslant 0}$ depending on the law of $U$.
 We call the resulting $\Pi$-based martingale approximation $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ the variance (resp.
@@
 If $U$ is uniformly distributed on a compact interval, all four martingales coincide.
-For example, if $U\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$, each of the four martingales from Definition 14 satisfy $M_{k}=\mathbb{E}[U\mid\sigma(\pi_{k})]$, where $\pi_{k}:=\{[j/2^{k},(j+1)/2^{k}):0\leqslant j<2^{k}\}$ and it follows that $\mathbb{E}[(U-M_{k})^{2}]=4^{-k}/12$.
+For example, if $U\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$, each of the four martingales from Definition 16 satisfies $M_{k}=\mathbb{E}[U\mid\sigma(\pi_{k})]$, where $\pi_{k}:=\{[j/2^{k},(j+1)/2^{k}):0\leqslant j<2^{k}\}$ and it follows that $\mathbb{E}[(U-M_{k})^{2}]=4^{-k}/12$.
 The intuition for the minimax and median martingales is that at each splitting step, we balance the “sizes” of $U{\mathbbm{1}}_{\{U\in I\}}$ on the two sets.
@@
 Similar techniques are also used in the study of stationary ergodic sequences (Wu and Woodroofe, 2004; Zhao and Woodroofe, 2008) and Markovian walks (Grama et al., 2018).
-Note that in our setting, $Y$ is a sum of martingale differences, but whose variance is bounded, and hence one cannot apply the martingale CLT or its convergence rates.
+Note that in our setting, $U$ is a sum of martingale differences, but whose variance is bounded, and hence one cannot apply the martingale CLT or its convergence rates.
 In the next two sections, we discuss two types of results: uniform rates for bounded $U$ (Appendix A.2) and non-uniform rates (Appendix A.3) where the asymptotic constant may depend on the possibly unbounded law of $U$.
@@
   1. (i)
-    If $\{M_{k}\}_{k\geqslant 0}$ is the variance martingale, $\mathbb{E}[(U-M_{k})^{2}]\leqslant 2.71\cdot 2^{-2k/3}$.
+    If $\{M_{k}\}_{k\geqslant 0}$ is the variance martingale, $\mathbb{E}[(U-M_{k})^{2}]\leqslant 2^{-2-2k/3}$.
   2. (ii)
@@
 Observe that one cannot have $r<1/4$ by Example 17 above.
-In Examples 17 and 18 below, we show that the rate $r$ in Theorem 16 is indeed optimal for the Simons and median martingales.
+In Examples 19 and 20 below, we show that the rate $r$ in Theorem 18 is indeed optimal for the Simons and median martingales.
 Example 19 (Rate $r=1/2$ is optimal for the Simons martingale).
@@
 $\mathbb{E}\bigg[U\mid U\geqslant\sum_{j=1}^{J}\Big(\frac{1-s}{s}\Big)^{j}\bigg]$ $=\bigg(\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\bigg)^{-1}\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\sum_{j=1}^{\ell}\Big(\frac{1-s}{s}\Big)^{j}$
-$=s^{-(J+1)}\sum_{\ell=J}^{\infty}s^{\ell+1}(2s-1)\bigg{(}1+\Big{(}\frac{1-s}{s}\Big{)}^{\ell}\bigg{)}$
-$=\frac{s^{-(J+1)}(2s-1)s^{J+1}}{1-s}+\frac{s^{-(J+1)}(2s-1)s}{1-(1-s)}$
+$=s^{-(J+1)}\sum_{\ell=J}^{\infty}s^{\ell+1}\frac{(1-s)^{2}}{2s-1}\bigg(1-\Big(\frac{1-s}{s}\Big)^{\ell}\bigg)$
+$=\frac{1-s}{2s-1}\bigg(1-\Big(\frac{1-s}{s}\Big)^{J+1}\bigg)$
 $=\sum_{j=1}^{J+1}\Big(\frac{1-s}{s}\Big)^{j}.$
@@
 Let $s\in(0,1)$ and a bounded random variable $U$ satisfy
-$\mathbb{P}\Big{(}U=\sum_{j=1}^{k-1}s^{j-1}-s^{k-1}\Big{)}=2^{-k},\leavevmode\nobreak\ k\geqslant 1.$
+$\mathbb{P}\Big(U=\sum_{j=1}^{k-1}s^{j-1}-s^{k-1}\Big)=2^{-k},\penalty 10000\ k\geqslant 1.$
 Next, we compute that
@@
 Since the supports of $\{M_{k}\}_{k\geqslant 0}$ correspond to the means of $U$ conditioned on nested partitions of $\mathbb{R}$, we have $\tau_{M}^{+}<\infty$ implies $M_{k}\geqslant x$ for all $k\geqslant 0$ and $M_{1}=y$.
-Therefore, conditional on the event $\tau_{M}^{+}<\infty$, $X$ is a martingale that is bounded from below by $x$.
+Therefore, conditional on the event $\tau_{M}^{+}<\infty$, $\{M_{k}\}_{k\geqslant 0}$ is a martingale that is bounded from below by $x$.
 By Ville’s inequality,
@@
 With the choices $M=r^{-k/(2+\delta^{\prime})}$ and $q=r^{\delta^{\prime}/(2+\delta^{\prime})}\in(0,1)$, we have that for some $C>0$,
-$\mathbb{E}[(U-M_{k})^{2}]\leqslant C(M^{-\delta^{\prime}}+M^{2}r^{k})\leqslant Cq^{-k}.$
+$\mathbb{E}[(U-M_{k})^{2}]\leqslant C(M^{-\delta^{\prime}}+M^{2}r^{k})\leqslant Cq^{k}.$
 The final statement follows immediately from the martingale convergence theorem. ∎
@@
 Since the asymptotic constant is allowed to depend on the law of $U$, the optimal non-uniform rate $r$ might be smaller than the uniform ones (in Theorem 18).
-Recall from Appendix A.2 that the rates obtained in Theorem 16 are asymptotically optimal for the Simons and median martingales (see Examples 17 and 18).
+Recall from Appendix A.2 that the rates obtained in Theorem 18 are asymptotically optimal for the Simons and median martingales (see Examples 19 and 20).
 For the variance and minimax martingales, we show in the next result that the rate $r=1/4$ is optimal under certain regularity conditions on the law of $U$.
@@
 Therefore, for any $\varepsilon>0$, there exists $k$ such that the $k$-th partition $\pi_{k}$ is such that on each interval $A\in\pi_{k}$, $\sup_{A}f\leqslant(1+\varepsilon)\inf_{A}f$.
-Since one-step variance is always better than one-step minimax, it remains to show that for an interval $A$ with $\sup_{A}f\leqslant(1+\varepsilon)\inf_{A}f$, there exists $C(\varepsilon)\downarrow 1/4$ as $\varepsilon>0$ such that at each step the total variance reduces by a factor of $C(\varepsilon)$ for the minimax martingale.
-To see this, without loss of generality we start from $U$ supported on $[0,1]$ (meaning that $\mathbb{P}(U\in[0,1])=1$) whose density takes values in $1,1+\varepsilon]$ (we may normalize afterwards to a probability measure, which we omit).
+Since one-step variance is always better than one-step minimax, it remains to show that for an interval $A$ with $(1-\varepsilon^{2})\sup_{A}f\leqslant\inf_{A}f$, there exists $C(\varepsilon)\downarrow 1/4$ as $\varepsilon>0$ such that at each step the total variance reduces by a factor of $C(\varepsilon)$ for the minimax martingale.
+To see this, without loss of generality we start from $U$ supported on $[0,1]$ (meaning that $\mathbb{P}(U\in[0,1])=1$) whose density takes values in $1-\varepsilon,1+\varepsilon]$, where $\varepsilon\in(0,1)$.
 Then for any interval $a,b]\subseteq[0,1]$,
-$\mathbb{E}[U\mid U\in[a,b]]\in\Big{[}a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)\Big{]}.$
+$\mathbb{E}[U\mid U\in[a,b]]\in\Big[a+\frac{1-\varepsilon}{2}(b-a),\,a+\frac{1+\varepsilon}{2}(b-a)\Big].$
 Therefore,
-$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])$ $\leqslant\sup_{m\in[a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)]}\int_{a}^{b}(1+\varepsilon)(x-m)^{2}\mathrm{d}x$
-$\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(b-a)^{3}.$
+$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])$ $\leqslant\sup_{m\in[a+\frac{1-\varepsilon}{2}(b-a),\,a+\frac{1+\varepsilon}{2}(b-a)]}\int_{a}^{b}(1+\varepsilon)(x-m)^{2}\,\mathrm{d}x$
+$\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{12}(b-a)^{3}.$
 Similarly,
-$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])\geqslant\inf_{m\in[a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)]}\int_{a}^{b}(x-m)^{2}\mathrm{d}x\geqslant\frac{1}{12}(b-a)^{3}.$
+$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])$ $\geqslant\inf_{m\in[a+\frac{1-\varepsilon}{2}(b-a),\,a+\frac{1+\varepsilon}{2}(b-a)]}\int_{a}^{b}(1-\varepsilon)(x-m)^{2}\,\mathrm{d}x$
+$\geqslant\frac{1-\varepsilon}{12}(b-a)^{3}.$
 It follows that the split point $x_{*}\in[0,1]$ must satisfy
-$\frac{1}{12}x_{*}^{3}\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(1-x_{*})^{3}\quad\text{and}\quad\frac{1}{12}(1-x_{*})^{3}\geqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}x_{*}^{3}.$
+$\frac{1-\varepsilon}{12}x_{*}^{3}\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{12}(1-x_{*})^{3}\quad\text{and}\quad\frac{1-\varepsilon}{12}(1-x_{*})^{3}\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{12}x_{*}^{3}.$
 Therefore, for some $a_{\varepsilon}\to 0$, $|x_{*}-1/2|\leqslant a_{\varepsilon}$.
-A direct computation yields $\mathrm{Var}(U)\geqslant\frac{1}{12}$. The total remaining variance $\mathbb{E}[(U-M_{1})^{2}]$ is bounded by
-$\mathbb{E}[(U-M_{1})^{2}]\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(x_{*}^{3}+(1-x_{*})^{3})\leqslant\frac{4(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(\frac{1}{2}+a_{\varepsilon})^{3}.$
+A direct computation yields $\mathrm{Var}(U)\geqslant\frac{1-\varepsilon}{12}$. The total remaining variance $\mathbb{E}[(U-M_{1})^{2}]$ is bounded by
+$\mathbb{E}[(U-M_{1})^{2}]$ $\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{12}\bigl(x_{*}^{3}+(1-x_{*})^{3}\bigr)$
+$\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{6}\Big(\frac{1}{2}+a_{\varepsilon}\Big)^{3}.$
 Hence, as $\varepsilon\to 0$, the ratio becomes
-$\frac{\mathbb{E}[(U-M_{1})^{2}]}{\mathrm{Var}(U)}\leqslant\frac{2(1+\varepsilon)^{4}}{(2+\varepsilon)^{3}}(1+2a_{\varepsilon})^{3}\to\frac{1}{4},$
+$\frac{\mathbb{E}[(U-M_{1})^{2}]}{\mathrm{Var}(U)}\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{4(1-\varepsilon)}(1+2a_{\varepsilon})^{3}\to\frac{1}{4},$
 as desired. ∎
@@
 Proof.
-Let $Q$ be the quantile function of $U$. Since $U$ is atomless, we have if $u=Q(p)$,
+Let $Q$ be the quantile function of $U$. Since $U$ is atomless, if $u=Q(p)$, then
 $\mathbb{P}(U\in[a,u))\mathrm{Var}(U\mid U\in[a,u))+\mathbb{P}(U\in[u,b))\mathrm{Var}(U\mid U\in[u,b))$
@@
 Since $U$ is atomless, it is non-degenerate, and hence $m_{b}(p_{\mathrm{var}})-m_{a}(p_{\mathrm{var}})>0$ and $p_{\mathrm{var}}\in(0,1)$.
-Note that the functions $m_{a},m_{b}$ are differentiable and hence continuous; $Q$ may have jumps but have limits on both sides at every point.
+Note that the functions $m_{a},m_{b}$ are differentiable and hence continuous; $Q$ may have jumps but has limits on both sides at every point.
 Since $p_{\mathrm{var}}\in\operatorname*{arg\,max}_{p\in[0,1]}\Psi_{U}(p)$, there are two cases:
@@
 Remark 23.
-If the support of $U$ is a connected interval, Lemma A.1 is a special case of Theorem 1 of Ishwaran (2015) applied with $f$ monotone and $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$.
+If the support of $U$ is a connected interval, Lemma A.1 is a special case of Theorem 1 of Ishwaran (2015) applied with $f$ monotone and $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$.
 Proof for the variance martingale.
@@
 This proves (40).
-Recall that our goal is to bound (37).
-The probability of the mass at each step is multiplied by $q$, and the edge lengths are bounded by $(a+b)/2$ and $(a^{\prime}+b^{\prime})/2$.
-By (39), their product multiplies by at most $1/2$ each time we update the tuple.
-We have thus shown that
-$\max_{1\leqslant j\leqslant 2^{k-1}}(p_{2j-1}d_{2j-1}^{2}+p_{2j}d_{2j}^{2})\leqslant 2^{1-k}.$ (40)
-Observe also that $\sum_{j}p_{j}=1$ and $\sum_{j}d_{j}\leqslant\sup U-\inf U\leqslant 1$ (because the edges at a single level cannot intersect each other since they represent the conditional means of $U$ in disjoint intervals).
-By Hölder’s inequality,
-$\begin{split}\mathbb{E}[(M_{k-1}-M_{k})^{2}]=\sum_{j=1}^{2^{k}}p_{j}d_{j}^{2}&\leqslant\Big{(}\sum_{j=1}^{2^{k}}p_{j}d_{j}^{3}\Big{)}^{2/3}\Big{(}\sum_{j=1}^{2^{k}}p_{j}\Big{)}^{1/3}\\ &\leqslant\Big{(}\max_{j\in[2^{k}]}(p_{j}d_{j}^{2})\sum_{j=1}^{2^{k}}d_{j}\Big{)}^{2/3}\leqslant\max_{j\in[2^{k}]}(p_{j}d_{j}^{2})^{2/3}\leqslant 2^{-2(k-1)/3}.\end{split}$ (41)
-By the martingale property, we then have
-$\mathbb{E}[(U-M_{k})^{2}]=\sum_{j=k}^{\infty}\mathbb{E}[(M_{j}-M_{j+1})^{2}]\leqslant\sum_{j=k}^{\infty}2^{-2j/3}\leqslant\frac{2^{-2k/3}}{1-2^{-2/3}}\leqslant 2.71\cdot 2^{-2k/3},$
-as desired. ∎
+For each interval $I\in\pi_{k}$, write
+$p_{I}:=\mathbb{P}(U\in I),\qquad m_{I}:=\mathbb{E}[U\mid U\in I],\qquad L_{I}:=m_{I}-\inf I,\qquad R_{I}:=\sup I-m_{I}.$
+The same recursive argument leading to (40) yields
+$\max_{I\in\pi_{k}}p_{I}L_{I}R_{I}\leqslant 2^{-2-k}.$ (41)
+Indeed, along each update, the quantity $p_{I}L_{I}R_{I}$ is multiplied by at most $1/2$, while initially $L_{I}+R_{I}\leqslant\sup U-\inf U\leqslant 1$, so $L_{I}R_{I}\leqslant 1/4$.
+Since $U\mid(U\in I)$ is supported on the interval $I$, we have
+$\mathrm{Var}(U\mid U\in I)\leqslant(m_{I}-\inf I)(\sup I-m_{I})=L_{I}R_{I}.$
+Therefore,
+$\mathbb{E}[(U-M_{k})^{2}]$ $=\sum_{I\in\pi_{k}}p_{I}\mathrm{Var}(U\mid U\in I)\leqslant\sum_{I\in\pi_{k}}p_{I}L_{I}R_{I}.$
+Set $s_{I}:=\sqrt{L_{I}R_{I}}$. By Hölder’s inequality and since $\sum_{I\in\pi_{k}}p_{I}=1$,
+$\sum_{I\in\pi_{k}}p_{I}L_{I}R_{I}=\sum_{I\in\pi_{k}}p_{I}s_{I}^{2}$ $\leqslant\Big(\sum_{I\in\pi_{k}}p_{I}s_{I}^{3}\Big)^{2/3}\Big(\sum_{I\in\pi_{k}}p_{I}\Big)^{1/3}$
+$\leqslant\Big(\max_{I\in\pi_{k}}p_{I}s_{I}^{2}\sum_{I\in\pi_{k}}s_{I}\Big)^{2/3}.$
+Moreover,
+$2s_{I}\leqslant L_{I}+R_{I}=\sup I-\inf I,$
+so, since the intervals in $\pi_{k}$ are disjoint and contained in $\inf U,\sup U]$,
+$\sum_{I\in\pi_{k}}s_{I}\leqslant\frac{1}{2}\sum_{I\in\pi_{k}}(\sup I-\inf I)\leqslant\frac{1}{2}(\sup U-\inf U)\leqslant\frac{1}{2}.$
+Combining this with (41), we obtain
+$\mathbb{E}[(U-M_{k})^{2}]\leqslant\Big(2^{-2-k}\cdot\frac{1}{2}\Big)^{2/3}\leqslant 2^{-2-2k/3},$
+and the proof is complete. ∎
 Simons
@@
 Consider a node $\mathbf{M}$ at level k1k-1 and a path $(\emptyset=\mathbf{N}_{0},\dots,\mathbf{N}_{k-1}=\mathbf{M})$ leading to $\mathbf{M}$, where $\emptyset$ denotes the root of a binary tree.
-We provide an algorithm that recursively defines the tuples $(A_{j},B_{j},q_{j}),\leavevmode\nobreak\ 0\leqslant j\leqslant k-1$ for each node $\mathbf{N}_{j},\leavevmode\nobreak\ 0\leqslant j\leqslant k-1$ on that path, from the root to $\mathbf{M}$.
+We provide an algorithm that recursively defines the tuples $(A_{j},B_{j},q_{j}),\penalty 10000\ 0\leqslant j\leqslant k-1$ for each node $\mathbf{N}_{j},\penalty 10000\ 0\leqslant j\leqslant k-1$ on that path, from the root to $\mathbf{M}$.
 The quantities $A_{j}$ and $B_{j}$ represent the maximum possible lengths of the two edges emanating from a node $\mathbf{N}_{j}$, and $q$ represents the one-step splitting probability leading to $\mathbf{N}_{j}$, based on the constraints from the separated tree property.
-We start from $(A_{1},B_{1},q_{1})=(1,1,1)$.
+We start from $(A_{0},B_{0},q_{0})=(1,1,1)$.
 Again, denote the lengths of the edges connecting the node $\mathbf{N}_{j-1}$ to its two descendants at level $j$ by $a_{j},b_{j},\,j\in[k-1]$, where $b_{j}=|\ell_{\mathbf{N}_{j-1}}-\ell_{\mathbf{N}_{j}}|$.
@@
 In both cases, $A_{j+1}B_{j+1}/(A_{j}B_{j})\leqslant 1/q_{j+1}-1$. In other words, the contribution of the two edges of $\mathbf{M}$ is upper bounded by
-$A_{k}B_{k}\prod_{\ell=1}^{k-1}q_{\ell}=\prod_{\ell=1}^{k-1}\frac{A_{\ell}B_{\ell}}{A_{\ell-1}B_{\ell-1}}\prod_{\ell=1}^{k-1}q_{\ell}\leqslant\prod_{\ell=1}^{k-1}\Big{(}\frac{1}{q_{\ell}}-1\Big{)}q_{\ell}=\prod_{\ell=1}^{k-1}(1-q_{\ell}),$
+$A_{k-1}B_{k-1}\prod_{\ell=0}^{k-1}q_{\ell}=\prod_{\ell=1}^{k-1}\frac{A_{\ell}B_{\ell}}{A_{\ell-1}B_{\ell-1}}\prod_{\ell=1}^{k-1}q_{\ell}\leqslant\prod_{\ell=1}^{k-1}\Big(\frac{1}{q_{\ell}}-1\Big)q_{\ell}=\prod_{\ell=1}^{k-1}(1-q_{\ell}),$
 where we recall that $\{q_{\ell}\}_{1\leqslant\ell\leqslant k-1}$ are the one-step probabilities leading to $\mathbf{M}$.
@@
 $\mathbb{E}[(M_{k-1}-M_{k})^{2}]=\sum_{j=1}^{2^{k}}p_{j}d_{j}^{2}\leqslant\max_{j\in[2^{k}]}(p_{j}d_{j})\sum_{j=1}^{2^{k}}d_{j}\leqslant 2^{1-k}.$
-By the martingale property,
+By the martingale property and the main result of Simons (1970),
 $\mathbb{E}[(M_{k}-U)^{2}]=\sum_{j=k}^{\infty}\mathbb{E}[(M_{j}-M_{j+1})^{2}]\leqslant\sum_{j=k}^{\infty}2^{-j}\leqslant 2^{1-k}.$
@@
 At level $k$, there are $2^{k}$ vertices (denoted by $j\in[2^{k}]$) in the binary tree representation of the partition-based martingale approximation.
-Denote the probabilities of the vertices by $p_{j}$ (which correspond to $\mathbb{P}(U\in A_{j}),\leavevmode\nobreak\ A_{j}\in\pi_{k}$) and the locations by $\ell_{j}$.
+Denote the probabilities of the vertices by $p_{j}$ (which correspond to $\mathbb{P}(U\in A_{j}),\penalty 10000\ A_{j}\in\pi_{k}$) and the locations by $\ell_{j}$.
 It follows that
 $\mathbb{E}[(U-M_{k})^{2}]=\sum_{j=1}^{2^{k}}p_{j}\mathbb{E}[(U-\ell_{j})^{2}\mid U\in A_{j}]=:\sum_{j=1}^{2^{k}}p_{j}d_{j}^{2}.$
-Note that $\sum_{j}p_{j}=1$ and $\sum_{j}d_{j}\leqslant\sup U-\inf U\leqslant 1$. It follows analogously as the Hölder’s inequality argument in (41) that
+Note that $\sum_{j}p_{j}=1$ and $\sum_{j}d_{j}\leqslant\sup U-\inf U\leqslant 1$. It follows analogously to the Hölder’s inequality argument in (69) that
 $\mathbb{E}[(U-M_{k})^{2}]$ $\leqslant\max_{j\in[2^{k}]}(p_{j}d_{j}^{2})^{2/3}.$
@@
 Lemma B.1.
-Let $\delta>0$. Then for any joint distribution $(X,Y)$ with $X\geqslant\mathbb{E}[Y]$, it holds that
+Let $\delta>0$. Then for any joint distribution $(X,Y)$ with $X\geqslant\mathbb{E}[Y]$ or $X\leqslant\mathbb{E}[Y]$, it holds that
 $\mathrm{Var}(Y)\leqslant(1+\delta)\,\mathbb{E}[(X-Y)^{2}]+\frac{(1+\delta)^{3}}{\delta^{3}}(\sup\operatorname{supp}X-\inf\operatorname{supp}X)^{2}.$ (43)
 Proof.
-Let $C=(1+\delta)^{3}/\delta^{3}$. Without loss of generality, we assume that $\mathbb{E}[Y]=0$. Equivalent to (43) is
+By symmetry, we assume that $X\geqslant\mathbb{E}[Y]$. Let $C=(1+\delta)^{3}/\delta^{3}$. Without loss of generality, we assume that $\mathbb{E}[Y]=0$. Equivalent to (43) is
 $(1+\delta)\,\mathbb{E}[X^{2}]-2(1+\delta)\,\mathbb{E}[XY]+\delta\mathbb{E}[Y^{2}]+C(\sup\operatorname{supp}X-\inf\operatorname{supp}X)^{2}\geqslant 0.$
@@
 This shows (44) and thus proves (43). ∎
+Lemma B.2 (Oscillation growth under cyclic axis-aligned refinement).
+Let $\{\pi_{k}\}_{k\geqslant 0}$ be the sequence of partitions generated by the cyclic splitting schedule, and let $k=dq+r$ with $q\in\mathbb{N}_{0}$ and $0\leqslant r<d$. Suppose that
+$g(\mathbf{x})=\sum_{i=1}^{d}g_{i}(x_{i})$
+is an additive representation with each $g_{i}$ of bounded variation. For a cell $A=\prod_{i=1}^{d}I_{i}(A)\in\pi_{k}$, define
+$\mathrm{osc}_{A}(g):=\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x}).$
+Then
+$\sum_{A\in\pi_{k}}\mathrm{osc}_{A}(g)\leqslant 2^{\,k-\lfloor k/d\rfloor}\sum_{i=1}^{d}\|g_{i}\|_{\mathrm{TV}}.$
+Consequently, after taking the infimum over additive representations of $g$,
+$\sum_{A\in\pi_{k}}\mathrm{osc}_{A}(g)\leqslant 2^{\,k-\lfloor k/d\rfloor}\|g\|_{\mathrm{TV}}.$
+Proof.
+Define
+$\mathrm{osc}_{I}(g_{i}):=\sup_{u\in I}g_{i}(u)-\inf_{u\in I}g_{i}(u).$
+For every cell $A=\prod_{i=1}^{d}I_{i}(A)$, additivity gives
+$\mathrm{osc}_{A}(g)\leqslant\sum_{i=1}^{d}\mathrm{osc}_{I_{i}(A)}(g_{i}).$
+Hence it suffices to control, for each coordinate $i\in[d]$,
+$S_{i}(k):=\sum_{A\in\pi_{k}}\mathrm{osc}_{I_{i}(A)}(g_{i}).$
+At level $k$, the cyclic rule splits along coordinate
+$j_{k}:=1+(k\bmod d).$
+If a cell $A$ is not split, then its contribution to every $S_{i}$ is unchanged. If a cell $A$ is split into $A^{L}$ and $A^{R}$ along coordinate $j_{k}$, then for every $i\neq j_{k}$,
+$I_{i}(A^{L})=I_{i}(A^{R})=I_{i}(A),$
+so the contribution to $S_{i}$ is duplicated:
+$\mathrm{osc}_{I_{i}(A^{L})}(g_{i})+\mathrm{osc}_{I_{i}(A^{R})}(g_{i})=2\,\mathrm{osc}_{I_{i}(A)}(g_{i}).$
+For the split coordinate $j_{k}$, the child intervals partition the parent interval, hence
+$\mathrm{osc}_{I_{j_{k}}(A^{L})}(g_{j_{k}})+\mathrm{osc}_{I_{j_{k}}(A^{R})}(g_{j_{k}})\leqslant\mathrm{osc}_{I_{j_{k}}(A)}(g_{j_{k}}).$
+Therefore
+$S_{j_{k}}(k+1)\leqslant S_{j_{k}}(k),\qquad S_{i}(k+1)\leqslant 2S_{i}(k)\quad\text{for }i\neq j_{k}.$
+Fix $i\in[d]$. By time $k=dq+r$, coordinate $i$ has been selected exactly $q+\mathbf{1}_{\{i\leqslant r\}}$ times, and has not been selected exactly
+$k-q-\mathbf{1}_{\{i\leqslant r\}}$
+times. Since $S_{i}$ never increases when coordinate $i$ is selected and can at most double otherwise, we obtain
+$S_{i}(k)\leqslant 2^{k-q-\mathbf{1}_{\{i\leqslant r\}}}S_{i}(0)\leqslant 2^{k-q}\|g_{i}\|_{\mathrm{TV}},$
+because $\mathrm{osc}_{I_{i}(A_{0})}(g_{i})\leqslant\|g_{i}\|_{\mathrm{TV}}$ at level 0.
+Summing over $i$ yields
+$\sum_{A\in\pi_{k}}\mathrm{osc}_{A}(g)\leqslant\sum_{i=1}^{d}S_{i}(k)\leqslant 2^{\,k-\lfloor k/d\rfloor}\sum_{i=1}^{d}\|g_{i}\|_{\mathrm{TV}}.$
+Taking the infimum over all additive representations of $g$ gives the final claim. ∎
 Proof of Theorem 3.
@@
 $\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\bigg((1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}(\left\lVert g\right\rVert_{\mathrm{TV}}\mathrm{Var}(Y))^{2/3}\bigg).$ (45)
-To this end, we need an upper bound of $\mathrm{Var}(Y)$. Fix $g\in\mathcal{G}$ and let $g_{*}:=(\sup g+\inf g)/2$. By Minkowski’s inequality,
-$\begin{split}\mathrm{Var}(Y)^{1/2}=\min_{y\in\mathbb{R}}\mathbb{E}[(Y-y)^{2}]^{1/2}\leqslant\mathbb{E}[(Y-g_{*})^{2}]^{1/2}&\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\mathbb{E}[(g(\mathbf{X})-g_{*})^{2}]^{1/2}\\ &\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\frac{1}{2}\left\lVert g\right\rVert_{\mathrm{TV}}.\end{split}$ (46)
+To this end, we need an upper bound for $\mathrm{Var}(Y)$. Fix $g\in\mathcal{G}$ and let $g^{*}:=(\sup g+\inf g)/2$. By Minkowski’s inequality,
+$\begin{split}\mathrm{Var}(Y)^{1/2}=\min_{y\in\mathbb{R}}\mathbb{E}[(Y-y)^{2}]^{1/2}\leqslant\mathbb{E}[(Y-g^{*})^{2}]^{1/2}&\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\mathbb{E}[(g(\mathbf{X})-g^{*})^{2}]^{1/2}\\ &\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\frac{1}{2}\left\lVert g\right\rVert_{\mathrm{TV}}.\end{split}$ (46)
 Squaring both sides yields that for any $\delta>0$,
@@
 By definition, $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$.
-Define $g_{j}=g|_{I_{j}}$, so that $\left\lVert g\right\rVert_{\mathrm{TV}}=\sum_{j}\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
-Define $\mathrm{Ran}(g_{j})=[\inf g_{j},\sup g_{j}]$ and recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
+Define $g_{j}=g|_{I_{j}}$ and $\mathrm{Ran}(g_{j})=[\inf g_{j},\sup g_{j}]$, and recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
 The key step is to decompose the local mean squared error $\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ into two parts for each $j$, with the help of Lemma B.1.
@@
   1. (i)
-    $y_{k,j}\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and
-    $V_{j}=\min\big{\{}(1+\delta^{-1})\,\mathbb{E}[(g(\mathbf{X})-y_{k,j})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}],\,\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\big{\}}.$
+[…]
   2. (ii)
-    $y_{k,j}\not\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and
-    $V_{j}=\min\Big{\{}\frac{(1+\delta)^{3}}{\delta^{3}}\,p_{j}(\Delta g_{j})^{2},\,\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\Big{\}}.$
+[…]
 In case (i), (48) follows immediately from the polarization identity $(a+b)^{2}\leqslant(1+\delta)a^{2}+(1+\delta^{-1})b^{2}$.
@@
 $V_{j}\leqslant(1+\delta^{-1})p_{j}(\Delta g_{j})^{2}<p_{j}(1+\delta^{-1})^{3}(\Delta g_{j})^{2},$
-as desired.
-On the other hand, suppose that $k=dq+r$ for some integer $q$ and $0\leqslant r<d$, and denote by $\pi_{k}$ the resulting partition of $\mathbb{R}^{d}$.
-Observe (by induction on $q,r$, and elementary geometry) that
+as desired. On the other hand, suppose that $k=dq+r$ for some integer $q$ and $0\leqslant r<d$, and denote by $\pi_{k}$ the resulting partition of $\mathbb{R}^{d}$. We use Lemma B.2 and observe that
 $\sum_{A\in\pi_{k}}\Big(\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\Big)\leqslant 2^{q(d-1)+r}\left\lVert g\right\rVert_{\mathrm{TV}}.$
@@
 The only difference from the proof of Theorem 3 is (52), since for non-marginally atomless measures (10) and (11) may not hold.
-Instead, by our assumptions, (10) is be replaced by
-$\max\Big{\{}\varphi_{L}(\hat{x}_{j}),\varphi_{R}(\hat{x}_{j})\Big{\}}\leqslant\frac{1}{2}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]+\frac{(\Delta Y)^{2}}{N},$
-because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $(\Delta Y)^{2}/N$. As a consequence, if we denote by
+Instead, by our assumptions, (10) is replaced by
+$\max\Big\{\varphi_{L}(\hat{x}_{j}),\varphi_{R}(\hat{x}_{j})\Big\}\leqslant\frac{1}{2}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]+\frac{(\Delta Y)^{2}}{N}.$ (56)
+To see this, let $B$ and $C$ be disjoint events with $\mathbb{P}(C)=m$, and suppose that $Y\in[a,b]$ almost surely, so $\Delta Y=b-a$.
+We let $\mu_{B}:=\mathbb{E}[Y\mid B]$.
+Since the conditional mean minimizes squared loss,
+$\mathbb{P}(B\cup C)\mathrm{Var}(Y\mid B\cup C)$ $\leqslant\mathbb{E}[(Y-\mu_{B})^{2}{\mathbbm{1}}_{B\cup C}]$
+$=\mathbb{P}(B)\mathrm{Var}(Y\mid B)+\mathbb{E}[(Y-\mu_{B})^{2}{\mathbbm{1}}_{C}]\leqslant\mathbb{P}(B)\mathrm{Var}(Y\mid B)+m(\Delta Y)^{2}.$
+Therefore, adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $(\Delta Y)^{2}/N$.
+As a consequence of (56), if we denote by
 $u_{k}:=\max_{A\in\pi_{k}}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}],$
 then $u_{k}$ satisfies the recursive inequalities
-$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{(\Delta Y)^{2}}{N},\leavevmode\nobreak\ k\geqslant 0;\quad u_{0}=\mathrm{Var}(Y).$
+$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{(\Delta Y)^{2}}{N},\penalty 10000\ k\geqslant 0;\quad u_{0}=\mathrm{Var}(Y).$
 To solve this, let $v_{k}:=u_{k}-2(\Delta Y)^{2}/N$.
@@
 $\sum_{A\in\pi_{k}}\Big(\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\Big)\leqslant\sum_{i=1}^{d}\xi_{k,i}\left\lVert g_{i}\right\rVert_{\mathrm{TV}},$ (58)
-for some (random) positive integers $\xi_{k,i},\leavevmode\nobreak\ i\in[d]$ depending on the samples from $\Sigma$.
+for some (random) positive integers $\xi_{k,i},\penalty 10000\ i\in[d]$ depending on the samples from $\Sigma$.
 See also (62) below.
@@
 Proof of Theorem 10.
-In the following, we use $S$ to denote a random variable with law $\Sigma$. We first establish an upper bound of
+In the following, we use $S$ to denote a random variable with law $\Sigma$. We first establish an upper bound for
 $\mathbb{E}[(Y-\mathbb{E}_{n}[M_{k}^{S}])^{2}].$
@@
 $\mathbb{E}[(Y-\mathbb{E}_{n}[M_{k}^{S}])^{2}]\leqslant\mathbb{E}[\mathbb{E}_{n}[\mathbb{E}[(Y-M_{k}^{S})^{2}]]].$ (59)
-Here, the outer expectation is with respect to the randomness from the empirical measure $\mathbb{E}_{n}$ (the law of $(\mathbf{X},Y)$ is fixed in the settings of Theorems 2 and 6), and the inner expectation is with respect to the joint law of $(\mathbf{X},Y,\{M_{k}^{S}\})$.
+Here, the outer expectation is with respect to the randomness from the empirical measure $\mathbb{E}_{n}$ (the law of $(\mathbf{X},Y)$ is fixed in the settings of Theorems 3 and 7), and the inner expectation is with respect to the joint law of $(\mathbf{X},Y,\{M_{k}^{S}\})$.
 In other words, we trivially bound the risk of the average by the average of the risks among the $n$ trees.
@@
 copies $\{\xi^{\prime}_{k-1,i}\}_{i\in[d]}$ and $\{\xi^{\prime\prime}_{k-1,i}\}_{i\in[d]}$, independently sample a uniform $\zeta_{k}\in[d]$, and define
-    $\xi_{k,\zeta_{k}}=\max\{\xi^{\prime}_{k-1,\zeta_{k}},\xi^{\prime\prime}_{k-1,\zeta_{k}}\};\quad\xi_{k,i}=\xi^{\prime}_{k-1,i}+\xi^{\prime\prime}_{k-1,i},\leavevmode\nobreak\ i\neq\zeta_{k}.$
+    $\xi_{k,\zeta_{k}}=\max\{\xi^{\prime}_{k-1,\zeta_{k}},\xi^{\prime\prime}_{k-1,\zeta_{k}}\};\quad\xi_{k,i}=\xi^{\prime}_{k-1,i}+\xi^{\prime\prime}_{k-1,i},\penalty 10000\ i\neq\zeta_{k}.$
 See Figure 11 for an illustration of the first few values of $\{\xi_{k,i}\}$.
@@
 By induction hypothesis, we have
-    $\sum_{j\in J_{k}}\Delta g_{j,i}^{\ell}$ $\leqslant\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ u\end{subarray}}\Delta g_{j,i}^{\ell}+\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ v\end{subarray}}\Delta g_{j,i}^{\ell}$
+    $\sum_{j\in J_{k}^{\ell}}\Delta g_{j,i}^{\ell}$ $\leqslant\sum_{\begin{subarray}{c}j\in J_{k}^{\ell}\\ j\succ u\end{subarray}}\Delta g_{j,i}^{\ell}+\sum_{\begin{subarray}{c}j\in J_{k}^{\ell}\\ j\succ v\end{subarray}}\Delta g_{j,i}^{\ell}$
     $\preceq_{\mathrm{st}}\xi^{\prime}_{k-1,i}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}+\xi^{\prime\prime}_{k-1,i}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}=\xi_{k,i}\left\lVert g_{i}\right\rVert_{\mathrm{TV}},$
     where $j\succ u$ means that the set $I_{j}$ is a descendant of $u$ in the binary subtree representation.
-  • For dimension $\zeta_{k}$, we again use the fact that the variation $\sum_{j\in J_{k}}\Delta g_{j,\zeta_{k}}^{\ell}$ is at most the sum of variations among the descendants of $u,v$, but the total variation $\left\lVert g_{\zeta_{k}}\right\rVert_{\mathrm{TV}}$ is already decomposed into two, say $\left\lVert g_{\zeta_{k}}^{(i)}\right\rVert_{\mathrm{TV}},\,i=1,2$, corresponding to $u,v$.
+  • For dimension $\zeta_{k}$, we again use the fact that the variation $\sum_{j\in J_{k}^{\ell}}\Delta g_{j,\zeta_{k}}^{\ell}$ is at most the sum of variations among the descendants of $u,v$, but the total variation $\left\lVert g_{\zeta_{k}}\right\rVert_{\mathrm{TV}}$ is already decomposed into two, say $\left\lVert g_{\zeta_{k}}^{(i)}\right\rVert_{\mathrm{TV}},\,i=1,2$, corresponding to $u,v$.
 Therefore, by inductive hypothesis,
-    $\sum_{j\in J_{k}}\Delta g_{j,\zeta_{k}}^{\ell}$ $\leqslant\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ u\end{subarray}}\Delta g_{j,\zeta_{k}}^{\ell}+\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ v\end{subarray}}\Delta g_{j,\zeta_{k}}^{\ell}$
+    $\sum_{j\in J_{k}^{\ell}}\Delta g_{j,\zeta_{k}}^{\ell}$ $\leqslant\sum_{\begin{subarray}{c}j\in J_{k}^{\ell}\\ j\succ u\end{subarray}}\Delta g_{j,\zeta_{k}}^{\ell}+\sum_{\begin{subarray}{c}j\in J_{k}^{\ell}\\ j\succ v\end{subarray}}\Delta g_{j,\zeta_{k}}^{\ell}$
     $\preceq_{\mathrm{st}}\xi^{\prime}_{k-1,{\zeta_{k}}}\left\lVert g_{\zeta_{k}}^{(1)}\right\rVert_{\mathrm{TV}}+\xi^{\prime\prime}_{k-1,{\zeta_{k}}}\left\lVert g_{\zeta_{k}}^{(2)}\right\rVert_{\mathrm{TV}}$
@@
 The term $2^{k}U^{2}/N$ is absorbed by the term $U^{4}2^{k}\log(Nd)/N$ therein.
-The other term $2^{k}U^{2}d/(Nn^{1/3})$ is carried into (B.35), and is bounded by $C2^{k}(\log N)d/(Nn^{1/3})$ under the choice of $U=\left\lVert g^{*}\right\rVert_{\infty}+2\sigma\sqrt{2\log N}$ at the end of that proof, where $C>0$ may depend on $\left\lVert g^{*}\right\rVert_{\infty}$ and $\sigma$.
+The other term $2^{k}U^{2}d/(Nn^{1/3})$ is carried into (B.35), and is bounded by $C2^{k}(\log N)d/(Nn^{1/3})$ under the choice of $U=\left\lVert g_{*}\right\rVert_{\infty}+2\sigma\sqrt{2\log N}$ at the end of that proof, where $C>0$ may depend on $\left\lVert g_{*}\right\rVert_{\infty}$ and $\sigma$.
 This concludes (26).
 ∎
+To prove Theorem 11, we first record the expected oscillation estimate that replaces the deterministic cyclic counting argument.
+Lemma B.4.
+Assume the setting of Definition 3.1, and let
+$g(\mathbf{x})=\sum_{j\in S}g_{j}(x_{j})\in\mathcal{G}_{S}.$
+Then for every $k\geqslant 0$,
+$\mathbb{E}\left[\sum_{A\in\pi_{k}^{\Xi}}\left(\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\right)\right]\leqslant 2^{k}e^{-\rho k/2}\sum_{j\in S}\|g_{j}\|_{\mathrm{TV}}.$
+Consequently,
+$\mathbb{E}\left[\sum_{A\in\pi_{k}^{\Xi}}\left(\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\right)\right]\leqslant 2^{k}e^{-\rho k/2}\|g\|_{\mathrm{TV}}.$
+Proof.
+Recall the notation from Lemma B.2. For $j\in S$, define
+$T_{j}^{\Xi}(k):=\sum_{A\in\pi_{k}^{\Xi}}\mathrm{osc}_{I_{j}(A)}(g_{j}).$
+As in the proof of Lemma B.2,
+$\sup_{A}g-\inf_{A}g\leqslant\sum_{j\in S}\mathrm{osc}_{I_{j}(A)}(g_{j}),$
+so it suffices to bound $\mathbb{E}[T_{j}^{\Xi}(k)]$.
+Fix $j\in S$.
+Conditional on $\pi_{k}^{\Xi}$, each splittable cell $A\in\pi_{k}^{\Xi}$ is split along coordinate $j$ with conditional probability at least $\rho$.
+If $A$ is split along coordinate $j$, then the sum of the oscillations of $g_{j}$ over the two children is at most $\mathrm{osc}_{I_{j}(A)}(g_{j})$.
+If $A$ is split along a coordinate different from $j$, then the $j$th interval is duplicated, so the sum of the oscillations over the two children is exactly $2\mathrm{osc}_{I_{j}(A)}(g_{j})$.
+Therefore,
+$\mathbb{E}\!\left[T_{j}^{\Xi}(k+1)\,\middle|\,\pi_{k}^{\Xi}\right]\leqslant\bigl(\rho\cdot 1+(1-\rho)\cdot 2\bigr)T_{j}^{\Xi}(k)=(2-\rho)T_{j}^{\Xi}(k).$
+Taking expectations and iterating,
+$\mathbb{E}[T_{j}^{\Xi}(k)]\leqslant(2-\rho)^{k}T_{j}^{\Xi}(0).$
+Since $2-\rho=2(1-\rho/2)\leqslant 2e^{-\rho/2}$ and $T_{j}^{\Xi}(0)\leqslant\|g_{j}\|_{\mathrm{TV}}$, we obtain
+$\mathbb{E}[T_{j}^{\Xi}(k)]\leqslant 2^{k}e^{-\rho k/2}\|g_{j}\|_{\mathrm{TV}}.$
+Summing over $j\in S$ proves the first claim, and taking the infimum over additive representations proves the second. ∎
+Proof of Theorem 11.
+Fix a realization of the randomization $\Xi$.
+Conditional on $\Xi$, the tree is deterministic, and the part of the proof of Theorem 3 that leads to (45) goes through verbatim, with a suitable adaptation of the upper bound (54) for the oscillation term.
+More precisely, for every $g\in\mathcal{G}_{S}$, the conditional predictor $M_{k}^{\Xi}$ satisfies
+$\mathbb{E}\!\left[(Y-M_{k}^{\Xi})^{2}\,\middle|\,\Xi\right]\leqslant\,$ $(1+\delta)\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+(1+\delta^{-1})(\mathrm{Var}(Y)\Theta_{k}(\Xi;g))^{2/3},$ (63)
+where
+$\Theta_{k}(\Xi;g):=2^{-k}\sum_{A\in\pi_{k}^{\Xi}}\left(\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\right).$
+Taking expectations over $\Xi$ and using Jensen’s inequality for the concave map $x\mapsto x^{2/3}$ yields
+$\mathbb{E}\!\left[\Theta_{k}(\Xi;g)^{2/3}\right]\leqslant\left(\mathbb{E}[\Theta_{k}(\Xi;g)]\right)^{2/3}.$
+By Lemma B.4,
+$\mathbb{E}[\Theta_{k}(\Xi;g)]\leqslant 2^{-k}\cdot 2^{k}e^{-\rho k/2}\|g\|_{\mathrm{TV}}=e^{-\rho k/2}\|g\|_{\mathrm{TV}}.$
+Hence
+$\mathbb{E}\!\left[\Theta_{k}(\Xi;g)^{2/3}\right]\leqslant e^{-\rho k/3}\|g\|_{\mathrm{TV}}^{2/3}.$
+Substituting this estimate into (63) and then following step I of the proof of Theorem 3 gives (27). ∎
 B.3 Proofs of results in Section 4
@@
 Then the function $\varphi_{L}$ is non-decreasing and continuous, and $\varphi_{R}$ is non-increasing and continuous.
-Moreover, under the (cyclic) MinimaxSplit construction,
+Moreover, denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ from the cyclic MinimaxSplit construction.
+Then
 $\max_{j\in J_{k}}\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\leqslant 2^{-k}\mathbb{E}[\log(1+e^{-YM_{0}})]=2^{-k}H(Y).$
@@
 Proof.
-We may assume that $\sup\operatorname{supp}X-\inf\operatorname{supp}X$, otherwise the claim is trivial.
+We may assume that $\sup\operatorname{supp}X-\inf\operatorname{supp}X<\infty$, otherwise the claim is trivial.
 Given the marginal law of $X$, the right-hand side of (64) is minimized if $X$ and $Y$ are comonotonic.
+Moreover, with $\sup\operatorname{supp}X$ and $\inf\operatorname{supp}X$ fixed, the right-hand side of (64) is minimized when we have the equivalence of events $\{X=\sup\operatorname{supp}X\}=\{Y=1\}$ and $\{X=\inf\operatorname{supp}X\}=\{Y=-1\}$.
 This is achieved by setting, with $a=\inf\operatorname{supp}X\leqslant\sup\operatorname{supp}X=b$, that $p=\mathbb{P}(X=b,Y=1)=1-\mathbb{P}(X=a,Y=-1)$, for some $p\in[0,1]$.
 It remains to show the deterministic inequality
-$\kappa_{a,b}(p):=h(p)+(p\log(1+e^{-b})+(1-p)\log(1+e^{a}))+\frac{1}{4}|b-a|\geqslant 0.$
+$\kappa_{a,b}(p):=-h(p)+(p\log(1+e^{-b})+(1-p)\log(1+e^{a}))+\frac{1}{4}|b-a|\geqslant 0.$
 Note that $\kappa_{a,b}$ is convex in $p$ and $\kappa_{a,b}^{\prime}(p)=0$ if and only if
@@
 Note that $M_{k}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}=y_{k,j}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}$, where $y_{k,j}=\log(\mathbb{P}(Y=1\mid\mathbf{X}\in I_{j})/\mathbb{P}(Y=-1\mid\mathbf{X}\in I_{j}))$.
-Define $g_{j}=g|_{I_{j}}$, so that $\left\lVert g\right\rVert_{\mathrm{TV}}=\sum_{j}\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
-Recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
+Define $g_{j}=g|_{I_{j}}$ and recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
 By applying Lemma B.6 to the joint distribution $(g(\mathbf{X}),Y)\mid\mathbf{X}\in I_{j}$, decompose
@@
 where $U_{j}=\mathbb{E}[\log(1+e^{-Yg(\mathbf{X})}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and
-$V_{j}=\min\Big{\{}\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}],\frac{p_{j}\Delta g_{j}}{4}\Big{\}}.$
-We have by construction and (64) that
+$V_{j}=\min\Big\{\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}],\frac{p_{j}\Delta g_{j}}{4}\Big\},$
+where $p_{j}:=\mathbb{P}(\mathbf{X}\in I_{j})$. We have by construction and (66) that
 $\mathbb{E}[\log(1+e^{-YM_{k}})]=\sum_{j\in J_{k}}\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\leqslant\sum_{j\in J_{k}}U_{j}+\sum_{j\in J_{k}}V_{j}.$ (67)
@@
 $\max\Big\{\varphi_{L}(\hat{x}_{j}),\varphi_{R}(\hat{x}_{j})\Big\}\leqslant\frac{1}{2}\mathbb{P}(\mathbf{X}\in A)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A))+\frac{\log N}{N},$
-because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $\log N/N$ (which can be seen from the extreme case with a sample $1$ added to samples of 1-1).
-As a consequence, if we denote by
+because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most
+$\frac{1}{N}\,\log N+\Big(1-\frac{1}{N}\Big)\log\Big(\frac{1}{1-\frac{1}{N}}\Big)\leqslant\frac{1+\log N}{N},$
+which can be seen from the extreme case with a sample $1$ added to samples of 1-1. As a consequence, if we denote by
 $u_{k}:=\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A)),$
 then $u_{k}$ satisfies the recursive inequalities
-$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{\log N}{N},\leavevmode\nobreak\ k\geqslant 0;\quad u_{0}=H(Y).$
+$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{\log(eN)}{N},\penalty 10000\ k\geqslant 0;\quad u_{0}=H(Y).$
 Solving this yields
-$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A))=u_{k}\leqslant 2^{-k}H(Y)+\frac{2\log N}{N}.$ (70)
-It is also easy to check that (70) is also satisfied for the case $u_{0}<2\log N/N$, that is, (70) holds in general. Replacing (68) by (70) leads to
-$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\bigg{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\Big{(}\left\lVert g\right\rVert_{\mathrm{TV}}(1+2^{k+1}\frac{\log N}{N})\Big{)}^{1/2}\bigg{)}.$
+$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A))=u_{k}\leqslant 2^{-k}H(Y)+\frac{2\log(eN)}{N}.$ (72)
+It is also easy to check that (72) is also satisfied for the case $u_{0}<2\log(eN)/N$, that is, (72) holds in general. Replacing (70) by (72) leads to
+$\mathbb{E}[\log(1+e^{-YM_{k}})]\leqslant\inf_{g\in\mathcal{G}}\bigg(\mathbb{E}[\log(1+e^{-Yg(\mathbf{X})})]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\Big(\left\lVert g\right\rVert_{\mathrm{TV}}(1+2^{k+1}\frac{\log(eN)}{N})\Big)^{1/2}\bigg).$
 Applying concavity of the function $x\mapsto x^{1/2}$ for $x>0$ and that
-$\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}\Big{(}2^{k+1}\frac{\log N}{N}\Big{)}^{1/2}\leqslant\frac{\left\lVert g\right\rVert_{\mathrm{TV}}}{2}+\frac{2^{k}\log N}{N},$
+$\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}\Big(2^{k+1}\frac{\log N}{N}\Big)^{1/2}\leqslant\frac{\left\lVert g\right\rVert_{\mathrm{TV}}}{2}+\frac{2^{k}\log(eN)}{N},$
 we obtain (32). ∎
+Proof of Theorem 14.
+The proof follows in exactly the same way as Theorem 4.3 of Klusowski and Tian (2024) in the C4.5 case, as the arguments therein do not depend on how the tree is constructed, as long as the splitting rule is axis-aligned.
+∎
 Appendix C Additional example for anti-symmetry-breaking preference
-In this appendix, we provide an empirical version of Example 8, showing that the ASBP persists even when samples are involved and the symmetry is not perfect.
+In this appendix, we provide an empirical version of Example 9, showing that the ASBP persists even when samples are involved (meaning that the symmetry is not perfect under the empirical measure).
 Example 25.
-Define $f(\mathbf{x})=x_{1}+|x_{2}|,\leavevmode\nobreak\ \mathbf{x}=(x_{1},x_{2})\in[-1,1]^{2}$, which is Lipschitz and bounded.
+Define $f(\mathbf{x})=x_{1}+|x_{2}|,\penalty 10000\ \mathbf{x}=(x_{1},x_{2})\in[-1,1]^{2}$, which is Lipschitz and bounded.
 Let $\mathbb{P}_{*}$ denote the law of $\mathbf{X}$ in Example 9 and $\mathbb{P}_{n}$ denote the empirical law of $\{(\mathbf{X}_{i},Y_{i})\}_{1\leqslant i\leqslant n}$, where each $\mathbf{X}_{i}$ is sampled from $\mathbb{P}_{*}$ and $Y_{i}=f(\mathbf{X}_{i})$.
 Denote by $\mathcal{R}$ the set of all axis-aligned rectangles in $1,1]^{2}$.
-It follows from standard arguments in van der Vaart and Wellner (2013) that the class $\{f{\mathbbm{1}}_{R}\}_{R\in\mathcal{R}}$ is a Donsker class, so with $m_{R}=\mathbb{E}^{\mathbb{P}_{*}}[f(\mathbf{X})\mid\mathbf{X}\in R]$, we have $\sup_{R\in\mathcal{R}}|\mathbb{E}^{\mathbb{P}_{n}}[f(\mathbf{X})\mid\mathbf{X}\in R]-m_{R}|=O_{\mathbb{P}}(1/\sqrt{n})$.
+Since $\left\lVert f\right\rVert_{\infty}\leqslant 2$ and the set of axis-aligned rectangles forms a VC class, the class
+$\mathcal{H}:=\{(f-m)^{2}\mathbf{1}_{R}:\ R\in\mathcal{R},\ m\in[-2,2]\}$
+is a bounded Donsker class. For
+$T_{R}:=\inf_{m\in[-2,2]}\mathbb{E}^{\mathbb{P}_{*}}[(f(\mathbf{X})-m)^{2}\mathbf{1}_{\{\mathbf{X}\in R\}}],\qquad V_{R}:=\inf_{m\in[-2,2]}\mathbb{E}^{\mathbb{P}_{n}}[(f(\mathbf{X})-m)^{2}\mathbf{1}_{\{\mathbf{X}\in R\}}],$
+we therefore have
+$\sup_{R\in\mathcal{R}}|V_{R}-T_{R}|\leqslant\sup_{h\in\mathcal{H}}|\mathbb{E}^{\mathbb{P}_{*}}[h]-\mathbb{E}^{\mathbb{P}_{n}}[h]|=O_{\mathbb{P}}\Big(\frac{1}{\sqrt{n}}\Big).$
 Here, we use the notation $\xi_{n}=O_{\mathbb{P}}(\alpha_{n})$ if the sequence $\{\xi_{n}/\alpha_{n}\}$ is tight.
-Let $g_{R}(\mathbf{x})=(f(\mathbf{x})-m_{R})^{2}$ for $R\in\mathcal{R}$.
-The class $\{g_{R}{\mathbbm{1}}_{R}\}_{R\in\mathcal{R}}$ is also Donsker.
-For $R\in\mathcal{R}$, let $V_{R}$ denote the empirical risk on $R$, i.e.,
-$V_{R}=\frac{1}{n}\sum_{i=1}^{n}((f(\mathbf{X}_{i})-\mathbb{E}^{\mathbb{P}_{n}}[f(\mathbf{X})\mid\mathbf{X}\in R])^{2}{\mathbbm{1}}_{\mathbf{X}_{i}\in R}).$
-It follows from the Donsker property and triangle inequality that
+Moreover, for each $R\in\mathcal{R}$, if $M_{R}:=\sum_{i=1}^{n}\mathbf{1}_{\{\mathbf{X}_{i}\in R\}}$, then conditional on $M_{R}=m\geqslant 1$,
+$\mathbb{E}[V_{R}\mid M_{R}=m]=\frac{m-1}{n}\mathrm{Var}(f(\mathbf{X})\mid\mathbf{X}\in R),$
+while $V_{R}=0$ when $m=0$. Hence
+$0\leqslant T_{R}-\mathbb{E}[V_{R}]=\frac{\mathbb{P}(M_{R}\geqslant 1)}{n}\mathrm{Var}(f(\mathbf{X})\mid\mathbf{X}\in R)\leqslant\frac{\|f\|_{\infty}^{2}}{n}\leqslant\frac{4}{n}.$
+By the triangle inequality,
 $\sup_{R\in\mathcal{R}}|V_{R}-\mathbb{E}[V_{R}]|=O_{\mathbb{P}}\Big(\frac{1}{\sqrt{n}}\Big).$ (73)
@@
 $R_{a,b;u}^{\leftarrow}$ $=[a,u]\times[-1,1],\quad R_{a,b;u}^{\rightarrow}=[u,b]\times[-1,1],\quad u\in[a,b].$
-Also denote by $m(p)=\mathbb{E}[(M-1){\mathbbm{1}}_{\{M\geqslant 1\}}/M]$ where $M\buildrel\mathrm{law}\over{\sim}\mathrm{Bin}(n,p)$.
+Also denote by $m(p)=\mathbb{E}[(M-1){\mathbbm{1}}_{\{M\geqslant 1\}}/M]$ where $M\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Bin}(n,p)$.
 A standard computation (by first conditioning on the number of samples in the rectangle) yields that
@@
 In the following, the constants $C$ (depending on $\varepsilon$ only) may vary from line to line.
-Note that m(p)=1O(1/(np))m(p)=1-O(1/(np)) jointly as $np\to\infty$ (see e.g., Corollary 1 of Wang et al.
-(2008)).
+Note that m(p)=1O(1/(np))m(p)=1-O(1/(np)) jointly as $np\to\infty$ (see e.g., Corollary 1 of Wang and others (2008)).
 The following results hold if $b-a\geqslant C/n^{1/6}$ for $n$ large enough.
@@
 As a consequence, the consistency of the MinimaxSplit algorithm is not always guaranteed in the large-sample regime, unlike VarianceSplit (Theorem 4.3 of Klusowski and Tian (2024)) and cyclic MinimaxSplit (Theorem 8).
-Appendix D Further numerics
-D.1 Performance metrics
+Appendix D Additional synthetic experiments for the classification setting
+In this appendix, we complement Section 4 with additional synthetic experiments that isolate the behavior of the classification splitting rules in one dimension.
+Throughout, we consider $Y\in\{\pm 1\}$ and write
+$\eta(x)\coloneqq\mathbb{P}(Y=1\mid X=x),\qquad g_{*}(x)\coloneqq\log\Big(\frac{\eta(x)}{1-\eta(x)}\Big).$
+We compare three axis-aligned tree constructions: (i) entropy-CART, which minimizes the weighted sum of child entropies; (ii) Gini-CART, which minimizes the weighted sum of child Gini impurities; [^10] The criterion $h(p)=-p\log p-(1-p)\log(1-p)$ was used for entropy-CART, while Gini-CART uses the Gini impurity criterion 2p(1p)2p(1-p).
+and (iii) the entropy-based MinimaxSplit rule from Section 4, which minimizes the maximum of the two weighted child entropies.
+Since $d=1$ in all experiments below, MinimaxSplit and cyclic MinimaxSplit coincide.
+Unless otherwise stated, each tree is grown with minimum leaf size $8$, training sample size $250$, and depth budget $k\in\{1,\dots,8\}$.
+D.1 Piecewise signal: predictive risk as a function of depth
+We first consider a piecewise-constant Bayes log-odds,
+$g_{*}(x)=\begin{cases}2,&0\leqslant x<0.30,\\ -1.5,&0.30\leqslant x<0.60,\\ 1,&0.60\leqslant x\leqslant 1,\end{cases}\qquad X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1),$
+so that $\eta(x)=e^{g_{*}(x)}/(1+e^{g_{*}(x)})$ is piecewise constant with two change-points.
+The left panel of Figure 12 reports the expected log-loss, which is the empirical analogue of the surrogate risk in (30) and (32).
+The middle panel reports the corresponding misclassification error.
+Two features of Figure 12 are worth emphasizing.
+First, entropy-CART and Gini-CART achieve the best shallow-tree performance: at depth $k=2$ they attain an expected log-loss of around $0.52$, whereas MinimaxSplit is still near $0.54$.
+This is consistent with the fact that the CART rules are more aggressive in optimizing the immediate impurity decrease.
+Second, once the depth exceeds the first few levels, the ranking reverses.
+From depth $k=3$ onward, MinimaxSplit has the smallest expected log-loss, and the gap becomes substantial by depths $k=5,6,7,8$.
+In particular, the MinimaxSplit curve stabilizes near $0.541$, while the two CART baselines drift upward toward $0.589$.
+Thus, on this piecewise signal, the worst-child criterion trades some initial greediness for a substantially more stable partition at moderate and large depths.
+The middle panel shows that the advantage is more pronounced for cross-entropy than for the 011 loss.
+This is also in line with Section 4: the primary theoretical control is on $\mathbb{E}[\log(1+e^{-YM_{k}})]$, and the excess classification risk is obtained only after the calibration step of Theorem 14.
+Empirically, the misclassification curves are close after depth $k=3$, whereas the cross-entropy curves remain clearly separated.
+The right panel shows that MinimaxSplit reaches a larger number of terminal cells by depth $5$, but, as we explain next, these leaves are much more evenly populated.
+Figure 12: Synthetic one-dimensional classification experiment with piecewise Bayes log-odds.
+Left: expected log-loss versus depth.
+Middle: test misclassification versus depth.
+Right: average number of leaves versus depth.
+In dimension $d=1$, the entropy-based MinimaxSplit rule coincides with cyclic MinimaxSplit.
+The main empirical advantage appears in the surrogate risk: after a short shallow-tree disadvantage, MinimaxSplit attains a lower and markedly more stable expected log-loss than entropy-CART and Gini-CART.
+D.2 Leaf-balance as the geometric mechanism
+Figure 13 makes the geometric mechanism behind the previous risk comparison explicit.
+The left panel plots the mean number of samples per leaf on a logarithmic scale, and the right panel plots the standard deviation of the leaf sizes, again on a logarithmic scale.
+The mean leaf sizes are broadly comparable once the depth is moderate, but the dispersion is dramatically different.
+For example, at depth $k=4$ the standard deviation of the leaf sizes is about $4.9$ for MinimaxSplit, compared with roughly $13$ for the two CART baselines; at depth $k=6$ the corresponding values are about $3.3$ versus about 6.5.
+This is the empirical signature of the worst-child criterion.
+Entropy-CART and Gini-CART reduce the average impurity, and therefore can continue to benefit from highly unbalanced splits as long as one child becomes pure enough.
+By contrast, MinimaxSplit penalizes any split that leaves one child both large and highly impure.
+The resulting trees are therefore more regular in the sense that the sample sizes of the terminal cells remain much closer to one another.
+Viewed together with Figure 12, Figure 13 suggests that the deeper-tree log-loss advantage of MinimaxSplit is not merely a numerical accident: it is explained by a clear difference in partition geometry.
+Figure 13: Leaf-size statistics for the one-dimensional piecewise classification problem.
+The left panel shows the mean number of samples per leaf, and the right panel shows the standard deviation of the leaf sizes.
+MinimaxSplit produces terminal cells with much smaller leaf-size dispersion, which is the classification analogue of avoiding highly unbalanced end-cuts.
+D.3 Continuous versus atomic covariates
+Finally, Figure 14 compares the same piecewise signal in two sampling regimes.
+In the left panel, $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$ is marginally atomless, corresponding to the setting of Theorem 12.
+In the right panel, $X$ is sampled from a uniform grid of $41$ equally spaced points in $[0,1]$, which places us in the purely atomic setting of Theorem 13.
+The plotted quantity is again the expected log-loss.
+The qualitative ordering is the same in both regimes.
+In each case, the CART baselines improve quickly at the first few levels and then deteriorate as the tree continues to refine, whereas the MinimaxSplit curve becomes nearly flat once it reaches its best value.
+The advantage of MinimaxSplit is somewhat smaller in the atomic experiment than in the atomless one, but it remains visible throughout the deeper levels.
+This is consistent with the form of (32), where exponential decay in the atomless case is accompanied by an additional discrete-resolution term.
+Empirically, the atomic curves flatten earlier, which is natural because once the grid resolution and the minimum-leaf constraint become active, all three procedures have limited room for further refinement.
+Overall, Figure 14 shows that the classification analogue of MinimaxSplit is not tied to the marginally atomless setting: its advantage in cross-entropy persists when the covariate distribution is discrete, although the attainable gain is moderated by the finite support.
+Figure 14: Expected log-loss for the one-dimensional piecewise classification problem under continuous and atomic covariates.
+Left: $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$.
+Right: $X$ sampled from a uniform grid on $[0,1]$.
+The same qualitative pattern appears in both settings: entropy-CART and Gini-CART improve fastest at very small depths, while MinimaxSplit is more stable and attains the best deep-tree surrogate risk.
+Appendix E Further numerics
+E.1 Performance metrics
 Mean Squared Error (MSE).
@@
 In our experiments the images are normalized to $[0,1]$, hence $L=1$, and we use Gaussian weights $w_{uv}$ and mean aggregation as in skimage.metrics.structural_similarity.
-D.2 Higher-dimensional input domain
+E.2 Two-dimensional input domain
+For this experiment, we use a synthetic dataset generated from a complex, non-linear function of two features, $\mathbf{x}=(x_{1},x_{2})$.
+The true function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ used for generating the target values is defined as:
+$\begin{split}f(\mathbf{x})&=\left(2-2.1u_{1}^{2}+\frac{u_{1}^{4}}{3}\right)u_{1}^{2}+u_{1}u_{2}+\left(-4+4u_{2}^{2}\right)u_{2}^{2};\\ &u_{1}=3(x_{1}-0.5),\\ &u_{2}=3(x_{2}-0.5),\end{split}$ (76)
+where $x_{1},x_{2}$ are independently and uniformly sampled from $[0,1]$.
+This function incorporates both quadratic and higher-order polynomial terms, creating a complex surface with multiple peaks and valleys.
+The visualizations of the prediction surfaces, as shown in Figures 15 and 16, further illustrate the improved accuracy and smoother transitions in the predicted values when using the proposed methods.
+In addition to decision trees fitted with MSE ($L^{2}$) loss with these three kinds of splitting criteria, we also provide trees fitted with $L^{1}$ loss for completeness.
+Figure 15: Predictions and true function values of (76), visualized using heatmaps.
+Top Row: The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models.
+The next three subplots depict the predictions of the variance-based decision tree, minimax decision tree, and cyclic minimax decision tree (depth $k=10$), respectively.
+Each plot is annotated with the corresponding MSE to quantitatively assess the accuracy.
+Bottom Row: The final four subplots visualize the predictions from four different random forest models (using Algorithm 1 of Appendix F), each built with varying combinations of the three error methods.
+Again, each plot is labeled with the MSE to facilitate direct comparison.
+Figure 16: The predictions and true function values are visualized using heatmaps, allowing for a clear comparison of how well each model approximates the underlying function (76).
+The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models.
+The next three subplots depict the predictions of the $L^{1}$ VarianceSplit, $L^{1}$ MinimaxSplit, $L^{2}$ VarianceSplit, and $L^{2}$ MinimaxSplit.
+Top Row: maximum depth 2; Middle Row: maximum depth 6; Bottom Row: maximum depth 10.
+In this example (as well as in Sections 5.3 and 3 below), we also incorporate an $L^{1}$ variation of the VarianceSplit and MinimaxSplit algorithms, where in the splitting rule, we replace the variances in (1), (4), and (5) by the $L^{1}$ distance from the mean.
+For instance, the $L^{1}$ variation of (1) is
+$\begin{split}(j,x_{j})&=\operatorname*{arg\,min}_{(j,\,x_{j})}\Big(\mathbb{E}\big[|Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}]|{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}\big]\\ &\hskip 56.9055pt+\mathbb{E}\big[|Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}]|{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}\big]\Big).\end{split}$ (77)
+The optimal $L^{1}$ split point may not be unique, in which case we break ties arbitrarily.
+E.3 Higher-dimensional input domain
 Consider a dimension $d\geqslant 1$ that is a multiple of $4$. The Powell function on $0,1]^{d}$,
@@
 The quartic terms in the function introduce sharp curvatures in the response surface, necessitating sophisticated splitting strategies to approximate these non-linear relationships accurately.
-Sample Size Dimension Sklearn VarianceSplit MinimaxSplit
+Sample Size Dimension Scikit-learn VarianceSplit MinimaxSplit
 10 4 $4.48\times 10^{2}$ $\boxed{2.01\times 10^{2}}$ $2.57\times 10^{2}$
@@
 The MinimaxSplit method often demonstrates superior performance in these high-dimensional, data-scarce scenarios.
-For instance, with 10 samples and 256 dimensions, the MinimaxSplit approach achieves an MSE of 8.91 × 104, outperforming both scikit-learn (1.07 × 105) and the VarianceSplit method (9.58 × 104).
-Appendix E Pseudocode for algorithms
+For instance, with 10 samples and 256 dimensions, the MinimaxSplit approach achieves an MSE of 1.14 × 105, outperforming both scikit-learn (1.24 × 105) and the VarianceSplit method (1.35 × 105).
+Appendix F Pseudocode for algorithms
 Input: Node $A_{t}$ at depth $k$ with sample $D_{t}=\{(x_{i},y_{i}):x_{i}\in A_{t}\}$; feature set $d]=\{1,\dots,d\}$ and $m_{try}=d$ by default
@@
 1
+20.15em
 VarianceSplit (CART)
@@
 4 $\mathrm{Crit}(j,s)\leftarrow P_{t}(A^{L}_{j,s})\,\mathrm{Var}(Y\mid A^{L}_{j,s})+P_{t}(A^{R}_{j,s})\,\mathrm{Var}(Y\mid A^{R}_{j,s})$
-4
+5
 […]
+70.35em
 MinimaxSplit
@@
 9 $\mathrm{Crit}(j,s)\leftarrow\max\big\{P_{t}(A^{L}_{j,s})\,\mathrm{Var}(Y\mid A^{L}_{j,s}),\ P_{t}(A^{R}_{j,s})\,\mathrm{Var}(Y\mid A^{R}_{j,s})\big\}$
-8
+10
 […]
+120.35em
 Cyclic MinimaxSplit
@@
 14 $\mathrm{Crit}(j_{k},s)\leftarrow\max\big\{P_{t}(A^{L}_{j_{k},s})\,\mathrm{Var}(Y\mid A^{L}_{j_{k},s}),\ P_{t}(A^{R}_{j_{k},s})\,\mathrm{Var}(Y\mid A^{R}_{j_{k},s})\big\}$
-12
+15
 $s^{\star}\leftarrow\arg\min_{s\in S_{j_{k}}(A_{t})}\mathrm{Crit}(j_{k},s)$;  split $A_{t}\to A^{L}_{j_{k},s^{\star}},\,A^{R}_{j_{k},s^{\star}}$.
@@
 Figure 17: Comparison of node-level split selection in two random forest variants.
-Left: a random-dimension random forest selects a random feature uniformly for each node and choose the best split over the chosen feature.
-Right: a regular RF node samples $m_{\mathrm{try}}$ coordinates uniformly from the $d$ features and chooses the best split over all features according to the chosen impurity/criterion (the highlighted feature illustrates the winner).
+Left: a random-dimension random forest selects a random feature uniformly for each node and chooses the best split over the chosen feature.
+Right: a regular RF node samples $m_{\mathrm{try}}$ coordinates uniformly from the $d$ features and chooses the best split over all sampled features according to the chosen impurity/criterion (the highlighted feature illustrates the winner).
 Input: Sample $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{N}$, number of trees $B\in\mathbb{N}$, maximum depth $K\in\mathbb{N}$, minimum leaf size $n_{\min}\in\mathbb{N}$, per-node feature budget $m_{\mathrm{try}}$, split rule $\mathrm{SplitRule}$.